A Complete Bibliography of Publications in IEEE MultiMedia

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 September 2024 Version 1.46

Title word cross-reference

2 [CBWD11, DMA^{+}12]. $26.85 [Swi96]. $29.95 [Hat96, Wil96]. 3 [ATS^{+}13, Ano02y, Ano14o, AHR^{+}13, BYZZ13, BCLH14, CCB^{+}98, CBWD11, CHWB11, CV14, DKRS04, GD14, Gou98b, Gou98c, Gou99b, Gro96b, Gro98b, Hor96d, IGB^{+}08, JGN^{+}13, KTY00, KWG10, LTD05, LD06, MLM^{+}07, MTG97, MTS^{+}06, NYNI99, UTZH12, XPZ^{+}13]. $34.95 [Bay97]. $39.95 [Gou98a, Tee98b]. $39.99 [Gou99d]. 4 [PAS^{+}13]. $40 [Ma96]. $42.95 [Pet96]. $44.95 [Gou00, San00a, Tee99]. $48.00 [Bha00]. $49.50 [Hoc95]. $49.95 [GSW97, Gou99c, Gou99d]. $50 [Ma96]. $54.95 [Str00]. $54.99 [Ver98]. $6 [GSW97]. $6.99 [Gou99e]. $64.95 [Hoc96]. $68 [MS95]. $69.00 [SB97]. $69.95 [Ara99b, Gol97c]. $83.00 [Str99a]. $84.95 [Kra00]. $95.00 [Str99b]. $99.95 [Car96, GSW97]. $995 [Gou98b, Gou98c]. ^{3} [AAHA20a, AAHA20b]. ^{∘} [CSC^{+}20, KAB^{+}23, CBC^{+}20].

0-12-208875-1 [Dav98a]. 0-12-691398-8 [Bro96]. 0-13-190992-4 [Str96]. 0-13-231374-X [Gol97c]. 0-13-272295-X [Bay97]. 0-201-70314-9 [San00a]. 0-201-84760-4 [Swi96]. 0-306- [Str99b]. 0-387-94211-4 [Hoc95]. 0-387-94441-9 [Hat96]. 0-440-22423-3 [Gou99e]. 0-471-12253-X [Pet96]. 0-471-14159-3 [Wil96]. 0-471-17805-5 [Tee98b]. 0-471-18359-8 [Ver98]. 0-471-25346-4 [Gou99d]. 0-7923-9432-1 [MS95]. 0-7923-9604-9 [GSW97]. 0-87930-481-2 [Gou98a]. 0-89006-870-4 [Str99a]. 0-89006-884-4 [SB97]. 0-89186-6876-4 [Ma96].

1 [Abo99a, Ano02k, Bir94a, KL07, PEB09c, Sto06]. 1-55860-346-8 [Hoc96]. 1-55860-533-9 [Ara99b]. 1-55860-584-3 [Str00]. 1-55860-598-3 [Gou00]. 1-56881-087-3 [Bha00]. 1-85233-228-X [Kra00]. 1-886801-65-7 [Tee99]. 1-886801-71-1 [Gou99c]. 1-886801-78-9 [Gou99d]. 10K [Ano17-27]. 13 [HZK^{+}23]. 13-D [HZK^{+}23]. 1394 [Tho00]. 16 [IMK^{+}16]. 19 [Che21d]. 1990s [Bla94].

2 [Abo99b, BCL00, Bir94b, Bul02, GKL^{+}98, LSL24, SJ06, Tee99, Wil98]. 2-D [LSL24]. 2.0 [Bol07a, Bul02]. 2/e [Tee99]. 2004 [JJO^{+}04]. 2005 [SJM06]. 2006 [Tit06]. 2011 [FJ12]. 2012 [JIW13]. 2013 [OK14]. 2016 [LSG^{+}17, SZ17]. 2017 [Ano16b, BEG^{+}18, Kel17]. 2019 [Ano18f, Ano18m, Ano19t]. 2020 [Ano20t, Ano20u]. 2021.indd [Ano23t]. 2023.indd [Ano23u]. 21 [BVH^{+}03, DB05, KLKW09, RDD09, TNJ05, TLS04, Vet04, Wan04, ZRD23]. 21st [Ano22-43, Ano23-46, Ano14q, HLT^{+}15, OB00]. 2D [ZZC^{+}21].

3 [CGW^{+}23, KP24, LPSZ08, LLZ22, LY19, SD07b, SZS^{+}20a, SZS^{+}20b]. 3-D [CGW^{+}23, KP24, LLZ22, LY19, SZS^{+}20a, SZS^{+}20b]. 3.0 [Hol10]. 324M [SJ04]. 360-Degree [TCM18, YZSY21]. 3D [Gou98b, Gou98c, Gou99d, KWYK22, NKZ^{+}22]. 3D-BCNN-Based [KWYK22]. 3DTV [SOS13]. 3G [SJ04]. 3G-324M [SJ04]. 3Learning [AAHA20a, AAHA20b].

4 [Bal98, BCL99, BCL00, DKRS04, NAK^{+}22, PFAHA03, SS01, SYJC06, XDL04]. 4.0 [Bal98]. 45772-5 [Str99b].

5 [EMM^{+}98, FCM^{+}22]. 5.50

[Gou98b, Gou98c]. 5G [Ano17c, Ano17d, Ano17r, SSEKE19, YWWW19, ZRW^{+}19].

7 [Ano02k, Ano02j, BCGU10, NL99a,

NL99b, TLS04].

802.11 [And98]. 8K [GYL^{+}23].

97 [GY97].

= [HN07].

Abandoned [AMM21]. Absence [SJM06]. Abstract [SCY^{+}16]. Academia [SS11]. Academic [GSW97, MS95]. Accelerating [Ano24i, TCPD10]. Accelerator [Car96]. Acceptability [AFKN95]. Access [DTG^{+}08, Jai06, Jai97d, MPM^{+}17, Rei94b, SDEFK09, SH10, SBB^{+}13, UTZH12, Vet04]. Accessibility [DG08]. Accessible [AVW08, GLN^{+}08, LPSZ08, NSS01, PMGCCPRM11]. Accessing [EL06, Oom11]. Accompaniment [PFMMF23]. Account [Ros03]. Accuracy [TGB^{+}22]. Achievements [BVH^{+}03]. Achieving [HWD10]. ACM [COCB14, JJO^{+}04, SJM06, Swi96, COCB14, FJ12, Kel17, RJ14]. Acoustic [GWD^{+}22, SM15]. Acquisition [DRB09]. Acronyms [Met00]. Acting [CCM^{+}04]. Action [Eff05, FFW^{+}20, Gou98a, WZL12, WYZ^{+}16]. Active [Sch94, TWY^{+}22]. Active3D [CNN02]. Active3D-Build [CNN02]. ActivePresenter [GSW97]. Activities [AHR^{+}13, Bry00, Kah96, YMON08]. Activity [MNKI07, VKV^{+}12]. Ad [Ano18h, Ano18t, Ano18u, Ano18v, Ano18-30, Ano18-29, Ano18-35, Ano18-33, Ano18-34, GQ10]. Adaptation [AAHA20a, AAHA20b, CDR^{+}10, Gec97, ISSZ08, JM17, LTC^{+}20, Vet04, ZLZ06]. Adaptations [BBMC02]. Adapting [CRD^{+}11]. Adaptive [BQLN^{+}18, CBG09, DTF17, Dav94, DDV03, GWC17, GRP21, GT11, HYLK21, LLP^{+}16, LWN^{+}21, LWW^{+}22, LW22, LLZ03, MPG07, QZ14, TH05, VW03, VJS^{+}23, YZSY21, YWT^{+}24, ZF24]. Adaptively [LYL^{+}15]. Add [Meh99, Nac10a]. Add-in [Meh99]. Addiction [HLADCR^{+}21]. Adding [AYK08, BQLN^{+}18, Gro96a, RCG^{+}99]. Addison [San00a, Swi96]. Addison-Wesley [San00a]. Addresses [Rei96a]. Adjacent [ZZZ^{+}21]. Adults [MYBP^{+}22]. Advanced [PPDH94]. Advances [APCH23, Ano01d, GG23, RMPK07, SZS^{+}20a, SZS^{+}20b]. Advancing [Ano16c, MB00, ZKWH17b]. Advantage [MD00]. Adversarial [CYL22, GLZW20, LWL^{+}20, LWW^{+}22, WLL^{+}21, ZSZ^{+}20]. Advertisement [Ano13e, Ano13i, Ano13h, Ano14d, Ano14i, Ano14j, Ano14k, Ano14l, Ano14p, Ano14o, Ano14r, Ano14q, Ano14w, Ano15f, Ano15h, Ano15n, Ano15r, Ano15s, Ano15u, Ano15y, Ano16e, Ano16g, Ano16j, Ano16k, Ano16p, Ano16t, Ano16r, Ano16u, Ano16v, Ano16z, Ano16y, Ano16-28, Ano16-27, Ano17e, Ano17f, Ano17g]. Advertising [Ded95, WY07]. Advice [Kol96, Rei94a]. Advisory [RDP99]. Aerial [DSGG21, LYGT19]. Aeronautical [HS99]. Aesthetic [GA20a, GA20b]. Aesthetics [Ada03, DV01, DV03, Nac03a, ZYNS16]. Affect [AAHA20a, AAHA20b, CVDL05, Ing13, KAB^{+}23]. Affect-Driven [AAHA20a, AAHA20b]. Affectionate [Sch06]. Affective [BB03, FSC08, SCY^{+}16, VVS21, ZXHS21]. Affective-Motion [VVS21]. AffectiveNet [VVS21]. Affinity [ZG09]. Affinity-Propagation [ZG09]. Against [Kan20]. Age [Che00, DP95, GBL97, LLT10, XSS14, SS00, San00b]. Agenda [FS01]. Agent [TIS96]. Agents [ASC96, CC98, CL10, GB98, RWS^{+}94, LH97]. Aggregation [FZXL24, NPG13, PLL^{+}11]. Agreements [MB99]. Ahead [Pan06, Rui14c]. AI [AAE23, Ano22o, DLW^{+}19, Jai16, KP24, PFMMF23, ZK18]. AI-Oriented [DLW^{+}19]. Aid [DNM^{+}10]. Aids [BN01]. aims [EMM^{+}98]. Air [GS95, ZYJ^{+}13]. Airports [Ano02i]. al [Ver98]. Album [GL03, LT11]. Alesso [San00a]. Algebra [WDG95]. Algorithm [HZK^{+}23, HLT23, LLL17a, LLLH18, LSL24, YWT^{+}07, YH16, ZAKNY23]. Algorithms [DWN01, DDV03, GL03, LCSC11, LLL^{+}17b, LLH22, iWD15]. Alignment [PGHK13]. All-in-One [Gol08b]. Allocation [BP96, GQ10, HMS17, MXH21]. AlloSphere [AKMHP09]. along [Bra05]. Alzheimer [DGT^{+}22, APS^{+}18, DNM^{+}10]. Ambient [Aar04, Bak06, Bak07]. American [RBK^{+}00]. Ames [Wil96]. among [DH01]. AN.general hHalf jz.indd [Ano23b]. Analyses [CZ96]. Analysis [AF18, APS^{+}18, Ano02f, AHR^{+}13, Bed05, BHS13, CNI^{+}04, Che21a, Che21b, CEG^{+}13, DZS^{+}02, DD04, DLB^{+}19, Gol03b, GAL^{+}18, LHD^{+}20, LLY14, LTS^{+}07, LXH^{+}22, MM14, MBX14, MTH^{+}20, NG11, Par12, SGK^{+}07, Sla11b, Sto06, SJ06, SD07b, SD07a, SSW20, WGY^{+}21, YZSY21, ZMZB11, ZZC^{+}21]. Analytics [Ano16g, Ano16f, Ano17f, PVK^{+}16, WFZ^{+}20]. Analyzing [DOD22, SW09, Ste95]. Anatomy [TDZ^{+}12]. Ancestral [DGR13]. Anchor [HCLK23, OWZ^{+}20]. Anchor-Based [OWZ^{+}20]. Anchor-Free [HCLK23]. Ancient [ROAS00]. Ancillary [Ano02y]. Andrea [Wil96]. Andrew [Gou00]. Android [WLZL20]. Angle [GLB^{+}21]. Angular [LCCL21]. Animal [VWA^{+}07]. Animated [RWS^{+}94]. Animation [DV96, Gou99c, Gou99d, TN98]. Animations [VD00]. Annals [Ano22h, Ano22i, Ano22j, Ano23j, Ano23l, Ano23k].

Annotating [FNHB05]. Annotation [ABCB02, BBDS10, DGLW07, KTP06, LT11, NSS01, PMGCCPRM11, SBL00, SZ00, VW03, YMON08, YNC09, KC00]. Annotation-Retrieval [DGLW07]. Annotations [AFO^{+}20, SvOPS06]. Annotea [HS08]. Annual [Ano95a, Ano96a, Ano97, Ano98a, Ano99a, Ano00a, Ano01a, Ano02a, Ano03a, Ano04a, Ano05a, Ano06a, Ano07a, Ano08a, Ano09a, Ano11a, Ano13a]. Anomaly [ASL^{+}24, DSGG21, RT06]. Anonymous [WXL^{+}17]. Answer [GS94]. Answering [HWL^{+}12, LMH^{+}10, LTB23, TML^{+}14]. Antinoise [WXL^{+}17]. Antispoofing [CYW21]. Anyone [Bre99, God08]. Anyway [GS94, Jai94d, Rei98d, Dav97, Pac99]. AOMedia [CMW^{+}23]. Aotearoa [Ran08]. AP [Bro96, Dav98a]. Apnea [TWY^{+}22]. App [Gol04d, Rei98a, Wea00]. Apple [GO96]. Applet [SdOG98]. Applet-Based [SdOG98]. Application [BB02, CCLS03, DPC05, GM01a, HCAMJ17, JYC08, KMF^{+}01, LJJT12, MLM^{+}07, OR02, SKK^{+}12, SDEFK09, WZ08, ZZC^{+}21, ZRW^{+}19, Bra97b]. Application-Layer [JYC08]. Application-Specific [HCAMJ17, WZ08]. Applications [AO15, AFO^{+}20, Ano16e, Ano16d, Ano16h, Ano16i, Ano16j, Ano17e, Ano17c, Ano17d, BYZZ13, CZY20, CRD^{+}11, CSJ^{+}18, CH02, CPCT03, CCPC11, CEG^{+}13, DZS^{+}02, EMS96, For02, GCP01, Has02, HRvL96b, Hua97, IMMP99, IGB^{+}08, JKP12, JN04, JXS^{+}17, JW17, LL08, LCG^{+}94, MTH^{+}20, MTS^{+}06, MG96, New95b, NXKS13, PCMT16, RHP^{+}06, RT06, SW09, SERL01, SSR96, SC15, SMdS^{+}21, Str96, Ude02, iWD15, WJ07, WNZ10, XW15, ZMvD^{+}08, ZZLG21, Chi95, Vin98, Ano22k, Ano22l, Ano22m, Ano23n, Ano23o]. Applying [KNC01]. Appointed [Gro97a]. Approach [Ano20j, BMDG20, Bro02, CAJ04, CSZL20, CW09, GTZ^{+}00, Han17, LWL^{+}17, LZDC19, LLL^{+}17b, MKYH03, MKK^{+}15, OSS04, PRC00, RDP99, SdOG98, SSW20, TLJ^{+}19, VRG18, WLZL20, ZYS^{+}10]. Approaches [CY13, GD14, LL08, WNZ10, YNC09, YJC07]. April [Bra05, Rei99b]. AR-Immersive [PVAMM10]. AR/VR [SSEKE19]. Arbitrary [HLW^{+}19, LNM^{+}24]. Arcade [Gri96]. Archaeological [ABC^{+}08]. Architectural [AG00]. Architecture [ALS22, Ano21g, Ano21h, Ano21i, Ano23m, Bir94a, Bir94b, CC98, FS01, PMGCCPRM11, PFMMF23, Pri99, TR01, WHTC24, WSS02, XDL04, Cam97]. Architectures [CRD^{+}11, GCGR11, LRDTP13, RR94, STL18, ZLZ06]. Archival [GS06]. Archive [HKM^{+}10]. Archived [ZMZB11]. Archives [AN97, EL06, Gro97b]. Arden [CRB^{+}08]. Arizona [Gol94]. Art [Ano01b, Bed05, Bre04, Bro05, BKCR13, CV16, Cle08, DJBF07, Del01a, Fal07, FSC08, Hig08, Ing13, JJO^{+}04, Jai97e, Khu07, MM03b, Mir03, MN07, SJM06, Sto06, SJ06, SD07b, SD07a, UTZH12, VRSW08, Wal03, RP97]. Art-Science [BKCR13]. Artech [SB97, Str99a]. Artech-House [Str99a]. Artful [Abo99a, Abo99b, And08, Ano01b, Bre99, Bur09a, Del01a, Fal07, Fle06, Hal01, HR00, HWV^{+}09, Jam07, Khu07, Mal00, McG06, Neu00, OB00, Pac99, Sto06, SJ06, SD07b, Wei08]. Arthroplasty [TGC^{+}06]. Article [Che18a, Rui17a]. Articles [Ano22r, Ano22s, Ano22p, Ano22q, Ano21f]. Artificial [Ano23a, Bra05, Che19b, IGR^{+}20, KTB^{+}22, RWS^{+}94, Rui17c]. Artist [LT11]. Artistic [AKMHP09, Gol02c, LLEK23]. Artists [GA20a, GA20b, HR00]. Arts [Kel16, WSR^{+}95, HR00]. Artwork [BCK^{+}05]. Artworks [BBC00]. Asia [Sha95]. Asian [GO96]. Ask [Rei94a]. Aspects [YGN18]. Assessing [CH02]. Assessment [Ano17r, Fer18, HYC^{+}19, HG15, JCM^{+}23, LXH^{+}22, MSF16, QPZ^{+}21, TLJ^{+}19, ZGL^{+}14]. Assets [Che22a].

Assisted [GCCGW17, HWZ^{+}23, YKHI94]. Assistive [DG08, PCMT16]. Assurance [SSSS95]. Asymmetrically [FZH^{+}21]. Asynchronous [Ros03]. Athletic [PPP^{+}11]. ATM [FGL^{+}01, GKL^{+}98, RC95, WO97]. ATM-Based [RC95]. Attack [QGZF23]. Attacks [LJ05]. attendee [GSW97]. Attention [CYW23, JCM^{+}23, MNKI07, MB16, PXL10, SHH09, WCQ^{+}22, XZL21, ZWH^{+}21]. Attention-Ranking [SHH09]. Attentive [WLL^{+}21]. attract [MB99]. Attracts [MB99]. Attribute [WPC^{+}20, WLN^{+}21, WTW23]. Attribute-Guided [WPC^{+}20]. Attributed [SMK^{+}10]. Auction [YCZ^{+}12]. auctions [Bal99a, MB99]. Audience [Cle08, HS17]. Audiences [WZBKB17, HR00]. Audio [CDD^{+}00, CNYL21, EL06, GPA20a, GPA20b, Goo03, GRF^{+}16, Gou15, GS95, GCK12, Kri03a, LKGR94, MKYH03, Pan95, PGHK13, PT15, Qua13, TN98, VWA^{+}07, VTJL^{+}08, WNY09, WBKW96, WHW^{+}22]. Audio-Video [MKYH03]. Audio-Visual [GPA20a, GPA20b]. Audiovisual [CCW07, DS15, Oom11]. Auditory [LG05, MJSE15, MOK^{+}15, PKL15]. Augmented [Che21c, CD14, HK17, Lin17, Rui17c, STL18, SZS^{+}20a, SZS^{+}20b, ZDS19]. Augmenter [LCCL21]. Augmenting [CHWB11, DAL^{+}16]. Aula [PVAMM10]. Authentication [GQFZ23, GCK12, ISSZ08, YKW01]. Authoring [GLR95, GSW97, KCRO10, MM03a, MMC01, NAV04, Nac14, SBS97, SK03, SMdS^{+}21, VKS99, Sus99]. Authorization [CPCT03]. Authors [Ano94c, Ano96b, Kol96]. Authorship [SSS^{+}19]. Autism [cSC15]. Autoblocking [LLLH18, YH16]. Autoencoders [LWW^{+}22]. Automata [BCF04, DMA^{+}12]. Automated [KZ06, SJS07, SZ00, WCQ^{+}22]. Automatic [Mad06]. Automatically [YW14, Yeh16]. Automating [Pic16]. Automotive [Gug03]. Autonomous [Che19d, DSGG21]. aVCSR [YWT^{+}24]. Averted [Gro95c]. Awaiting [Gol04d]. Award [Ano16q, Ano16-28, Ano17n, Ano17-27, Ano19e, Che19a, Ano16b, Ano18-29, Ano23-29]. Awards [Ano19d, Ano21j, Ano21k, Ano22a, Ano23s, Ano20-30]. Aware [AD24, CDR^{+}10, DTF17, FOW^{+}22, FWL22, GTH^{+}13, HGHW16, HYC^{+}19, HMS17, HTL96, HZSH23, HWL^{+}14, HLW^{+}19, LSM^{+}21, LLZ22, MBM21, SHG^{+}07, WGY^{+}21, WTW23, ZGL^{+}14, RLJ^{+}23]. Awareness [GTZ^{+}00, WDZ^{+}17, ZWH^{+}21].

B [Ano16b]. Back

[Ano09b, Ano18a, GO96, Nac14, SA07b]. Back-End [SA07b]. Background [LH23, LTC^{+}20]. Backpack [DB05]. Backward [RAE16]. Backward-Compatible [RAE16]. Bag [WH11]. Bag-of-Words [WH11]. Balanced [GT11, WLL^{+}21]. Ball [RR05]. Ballet [MKK^{+}15]. Balloon [Smi10c]. BAMMF [Liu16]. Bandwidth [Ano01n, HMS17, HZSH23, Mos03]. Bandwidth-Aware [HZSH23]. Banking [Gro95c, KGG96]. Baochun [ZKWH17a]. Based [ASL^{+}24, AGR^{+}21, ASLB20, Ano02f, Ano16c, Ano17r, AAHA20a, AAHA20b, BP96, BPT06, BD03, CWW^{+}19, CH02, CWZ^{+}05, CL06, Che22a, CNYL21, CDWT10, CNN02, Cur02, DTF17, DFG^{+}14, Dje02, DGT^{+}22, EFS01, FZH^{+}21, FCL^{+}19, GD14, GPA20a, GPA20b, GZP^{+}24, GR17, GRP21, GWD^{+}22, HCLK23, Hay98, Hir97, HG23, Hor96d, HHLW20, HXXS18, HRH^{+}21, HZK^{+}23, JNK06, JGN^{+}13, KTB^{+}22, KGS^{+}02, Kim96, KD17, KGU07, LTD05, LVL09, LLLH18, LZH^{+}23, LSKP09, LTM03, LCM09, LLD^{+}16, LLL^{+}17b, LLH22, LZD^{+}22, LMS^{+}22, LNM^{+}24, LRDTP13, LWH16, LY19, hMLDM98, MTXH19, MNKI07, MXH21, MMSB15, Mou06, MTS^{+}06, NKZ^{+}22, OWZ^{+}20, PRC00, PZKV11, PXL10, PFMMF23, PFAHA03, QPZ^{+}21, QJC^{+}18, QGZF23, QSL^{+}19, RSR^{+}01, RC95, Rou99, SdOG98, SZ94, SGS01, SMK^{+}10, SRA08, SLW98, SSS^{+}19, VCBM19, WHM15, WCS^{+}21, WEV03, WBKW96, XZL^{+}16, XCZY14, YMON08, YHFC14]. Based [YH16, YKHI94, YKW01, ZAKNY23, ZJC^{+}20, ZMFB21, ZLL^{+}20, ZF24, ZSZ^{+}20, dLA01, Bra98, KC00, WJ99, KWYK22]. Bases [MdRCS15]. Basic [Dav98a]. Basics [Tee98b, Ver98]. basis [Chi95]. Battle [Ano01r, Gro95c]. Battles [Gro96a, MB00]. BCNN [KWYK22]. Be [Ano16s, Ano18p, Aut11, Kol96, HR00]. Beagle [Bur09b]. Beat [Bol08, Bur09b, Cle08, DJBF07, FSC08, Hig08, MN07, SJM06, Tit06, Tit07, VRSW08]. Beats [Kri03b]. Beautiful [DV01]. Behavior [AF18, APS^{+}18, Ano16i, Ano16j, HS17, TJBBFB15, VWA^{+}07, XPZ^{+}13]. Behaviors [HXXS18]. Behind [OB00]. being [Bre99]. Bell [Gro96b]. Belur [Ma96]. Ben [Ara99b, Dav98a]. Benchmarking [LSG^{+}17]. Benchmarks [Gri96, OLIG04]. Berlin [Kra00]. Bermed [KO94]. Best [Ano17-27, Che18a, Rui15a, Rui17a, Jef99, Ano18w, Ano18x, Ano19x]. Beta [ASL^{+}24]. Better [Ano16v, Ano17q, Jef99]. better-than-best-effort [Jef99]. Between [HS08, KRB19, MTD^{+}11, WZL^{+}21]. Beyond [DLW^{+}19, Jai03b, Kri03a, LBZL22, Rui15b, SKSH08, SMH^{+}17, HK17]. Bias [LTB23, Dav98b]. Bidirectional [MMSB15]. Big [Ano16g, Ano16f, Ano17f, Ano18c, Ano24c, Hat96, MM14, Rui14a, SZW^{+}19, TCS^{+}15, VGH^{+}18, WFZ^{+}20, WB16, ZCWH15]. Bilinear [ZLT^{+}12]. Bill [Gou99d]. BilVideo [BCGU10, DSUG03]. BilVideo-7 [BCGU10]. Bimodal [DS15]. Binarization [LLL^{+}17b]. Binocular [FZXL24]. Bioelasticity [NMK^{+}06]. Biofeedback [MOK^{+}15]. Biological [AGR07]. Biometric [Ran14]. Biometrics [AHN^{+}18, Gol03b, KZ06]. Biopsy [TDZ^{+}12]. Biped [Abo99a, Abo99b]. Bird [VRSW08]. Bit [FHK08b]. Bitrate [GWC17, WHW^{+}22, ZCH^{+}19]. Bits [LLL17a, Dav97]. Blend [Nac09]. Blind [GLN^{+}08, JCM^{+}23, MOK^{+}15]. Blobs [MJ02, WX14]. Block [Ma96]. Blockchain [Che22a, DWX23, XXZ^{+}23]. Blockchain-Based [Che22a]. Blockchain-Empowered [XXZ^{+}23]. Blockchains [ZRD23]. Blocks [Ano01f]. Blogging [PP05]. Bluetooth [MD00]. Board [Ano18v, Gro99b, Gro00c, WZL^{+}24, Ano21x, Ano21y, Ano21z, Ano22-35, Ano22-36, Ano22-37, Ano22-38]. Boardroom [Gri96]. bodies [RFM95]. Body [Bur09a, ZWL^{+}20]. BOF [GHD^{+}14]. Book [Bro96, Gou99a, Tee98a, Gou99c]. Books [Bal01, Car98, Gou98a, LPSZ08, WSI^{+}05, Sch00]. Boosting [ZLT^{+}12]. Borko [GSW97]. Boundaries [JJO^{+}04]. Boundary [SZW^{+}19]. Box [Gol08b, Dav99]. Boxes [Meh99]. Brain [Ano02t]. Brazil [FCA08]. Breathing [ROAS00]. Brews [Gro96b]. Bridge [WO97]. Bridging [DGLW07, DV03, HKR10, LLJW15, Pan02]. Briefs [Bal99b, GL97b]. Bringing [Che00, DB00]. Broad [Hoc96]. Broadband [Kri03b, Rat00, San00a]. Broadcast [CHWB11, ROW^{+}97, YS09]. Broadcasting [BBG^{+}96, CHK^{+}09, FCA08, HFdW11, LZH^{+}07, FWI98, FWI99]. Broker [NS95]. Brook [Tee99]. Brothers [Hat96]. Browse [Car96]. Browser [Car96, Gro96a, Sch14]. Browsers [Car98, Rei97d]. Browsing [Rei96d, SJ00, SH08, WKS^{+}03, YNC09, NM98]. BTC [QJC^{+}18]. Budding [ZK18]. Buffering [SA07b, WS02]. Build [CNN02]. Building [AFO^{+}20, ABH^{+}95, Ano01f, Ano01h, CLSS97, CT01, Hat96, Hor96d, LLZ22, MPL00]. Bullets [Tit03b]. Burgeoning [XCZH19]. Business [Bal98, DK04, GY97, GL97b, MB99, TCPD10]. Businesses [MB99]. Button [Hal94]. buyers [MB99].

C [Gol97c, Ver98]. CABAC [HCAMJ17]. Cache [HCAMJ17]. Caching [CWW^{+}19, CWZ^{+}05, SSEKE19, ZCH^{+}19]. CAD [Jai96b]. CADW [QGZF23]. CAE [WHW^{+}22]. Calibration [SSLS14]. Calif [Ara99b, Gou00, Str00]. Call [Ano13c, Ano13d, Ano14d, Ano14a, Ano14b, Ano14c, Ano14h, Ano15e, Ano15b, Ano15c, Ano15d, Ano15o, Ano15p, Ano15r, Ano15q, Ano16b, Ano16c, Ano16e, Ano16d, Ano16g, Ano16f, Ano16h, Ano16i, Ano16j, Ano16p, Ano16q, Ano17e, Ano17f, Ano17c, Ano17d, Ano17g, Ano17n, Ano18e, Ano19e, Ano19d, Ano19h, Ano19g, Ano19f, Ano20f, Ano20g, Ano20h, Ano20d, Ano20e, Ano20s, Ano21j, Ano22r, Ano22s, Ano22p, Ano22q, Ano22t, Ano22u, Ano22v, Ano23s, Ano23-30, Ano24o, Ano24p, FGR12, MBPM98]. call-detail [MBPM98]. Call-for-Major-Awards-2023-full-ml.indd [Ano23s]. Camera [DFG^{+}14, LKBE08, VCBM19, VAAI24]. Camera-Based [DFG^{+}14, VCBM19]. Cameras [Hay98, KRB19, TCM18]. Campus [FHL04]. Can [LT11, Sch06, Tai02, And98]. Capabilities [BQLN^{+}18]. Capacity [LLY^{+}18]. Capella [GSW97]. Captioning [AD24, STL18, WB08]. Captions [BC09]. Capture [CKRW00, GS06, SKWB22]. Capturing [CFR^{+}10, DOD22, RC06]. Car [Ano02u, MW11]. Card [Ara99b, CMC^{+}20]. Cardiomorphologies [Khu07]. Cards [Meh99]. Care [APS^{+}18, BEG^{+}18, BM22, Hua97, DB97, Kim95]. Career [Ano16u, Ano19u, Ano19v, Ano19w, Ano20k, Ano21l, Ano24i, Ano24q]. Cascading [THJG13]. Case [BVT^{+}23, Gou98a, PEB09a]. Casual [NYNI99]. Catalog [PHS03]. Categorizing [HB04]. Cats [Smi13a]. CBIR [MMS13].

CBMI [IMK^{+}16]. CD [Gou99c, San00a, Tee99, Gol97c, Neu00]. CD-ROM [Gou99c, San00a, Gol97c]. CD-ROMs [Neu00]. CDN [DWX23, FRM^{+}07]. CDN-Supported [FRM^{+}07]. Celebrity [Zim03]. Cell [Ano17r, Tit06]. Cellular [BCF04, DMA^{+}12, ZCH^{+}19]. Center [Ano16s, Ano18p, Ano24q, CLSY12, MPL00, PVAMM10]. Centered [Ano01h, Jai06, Ovi96, PCMT16]. Centers [ZCZ20]. Centric [RGH^{+}21, Sch14, SdOG98, ZMvD^{+}08]. Century [Ano22-43, Ano23-46, Ano14q, OB00]. CFP [Ano22v]. CG&A [Ano14h, Ano23-31]. CGA [Ano23t]. CGA-general-half-2021.indd [Ano23t]. CGAN [QGZF23]. CGAN-Based [QGZF23]. Chain [RDD09, ZRD23]. chalk [SBN97]. Challenge [FJ12, Gro95c, JIW13, OK14, Smi11b, Chi97, SS11]. Challenges [Bol15b, CY13, DK04, GD14, GB98, JW17, PAS^{+}13, PMV^{+}10, SSG19, SAB^{+}20, Woj98, YJC07, ZZLG21]. Chance [GB98, Jai07]. Change [Gro98a]. Changes [Ano09d, MB00, Rui17b, TJBBFB15]. Changing [Coo01, Tho00]. Channel [LNM^{+}24, Smi11d]. Chaotic [HZK^{+}23, LFX^{+}18, LMS^{+}22]. Character [XPZ^{+}13]. Characteristic [ZZC^{+}21]. Characteristics [CIL15, LH23]. Characterizing [VWA^{+}07, WKY^{+}15]. Characters [Bre04]. Charge [Tai02]. Charles [Gou99c, Gou99d, Pet96, Tee99]. Checking [Gro95c, LLD^{+}16]. Cheese [Gol06c]. Chest [JCM^{+}23]. CHI [Bol17]. Chief [Jai95a, Jai95d, Jai95b, Jai95c, Jai96b, Jai96c, Jai96a, Jai96d, Pan07b]. Child [GGX^{+}17]. Child-Safety [GGX^{+}17]. ChildGuard [GGX^{+}17]. Children [AdAFM95, CRL^{+}03, DP95, RHP^{+}06]. China [LZH^{+}07, XCZH19]. Chinese [GO96]. Chip [Gro96b, HZSH23]. Chips [GO96]. Choose [Ano16u]. Choropleth [ZSN^{+}05]. Christian [ZKWH17a]. Chronicles [Jai03d]. Cinema [MD00, PVAMM10, Vai95]. Cinemagraphs [Yeh16]. Cinematics [Ano01b]. Cinematography [DFG^{+}14, Dav98a, Dav98a]. Circle [Nac14]. Circuit [WZL^{+}24]. Circulating [Ran08]. CiSE [Ano21b]. Cities [RBK^{+}00]. Citizen [Swi96]. City [BVA^{+}15, DLW^{+}19, RBK^{+}00]. CitySensing [BVA^{+}15]. Civil [CNN02]. Civilizing [Swi96]. Claim [GO96]. Class [WLL^{+}21]. Class-Balanced [WLL^{+}21]. Classification [AK22, AHR^{+}13, GSOI10, GL20, HLT23, LWH^{+}18, MXH21, WBKW96]. Classifier [FGC^{+}14]. Classroom [Gol97c, PEB09a, PEB09b, PEB09c]. ClaviNet [LCL21]. Cleansing [ZYS^{+}10]. Clearing [LS00]. ClearType [LS00]. Click [SS12]. Clickers [Jai97b]. Clicking [Smi10a]. Client [BDPN94, GS01, Jai05b, SA07b, XDL04]. Client-Server [BDPN94]. Climate [Gou15]. Clips [DWZ^{+}08, HPH05]. Clones [VD00]. Close [ZLS14]. Closed [BC09, SW04]. Closed-Loop [SW04]. Closing [Nac14]. Clothes [ZWL^{+}20]. Clothing [Ano01e, ZLL^{+}20]. Cloud [Ano14i, Ano15o, Ano15p, Ano15r, Ano15q, Ano16p, BM22, GWL^{+}22, GRP21, JW17, KUM^{+}19, LLS11, PLL^{+}11, PFMMF23]. Cloud-Mobile [LLS11]. Clouds [KP24]. Clues [PXL10, XWJ16]. Cluster [PZKV11]. Cluster-Based [PZKV11]. Clustering [Fer18, OWZ^{+}20, WS02, YW14, ZCZ20, ZG09]. Clusters [CKN^{+}11]. CNN [ZMFB21]. CNN-Based [ZMFB21]. Co [HS08]. Co-Annotea [HS08]. Codecs [RM94]. Coding [AGR^{+}21, AA05, Che04, CMW^{+}23, CJ12, Goo03, GTH^{+}13, GRP21, HG23, HZSH23, Kal06, KL07, LZZ^{+}21, Ma96, MZZ^{+}16, QZ14, Qua13, WFI07, WYZ^{+}16, WHW^{+}22, ZF24, Ma96].

Coding-Based [HG23]. Coffee

[CSC^{+}20, CBC^{+}20]. Cognition [Kel17]. Cold [TYW23]. Cold-Start [TYW23]. Collaboration [Cur02, TY00, VRSW08]. Collaborative [AB94, CT01, FRM^{+}07, GT08, LL08, MDK10, Mur15, NN08, Oom11, VJS^{+}23, WYZ^{+}16, YGM^{+}16, Sus99]. Collections [DRB09, HCH12, KSPW07, PZKV11, SH10]. Collision [XNS^{+}13]. Collocation [ZLL^{+}20]. Color [Ano94a, BBMC02, FGD01, LLP00, Nac10b, OB00, VWS95]. Colors [GA20a, GA20b, MMS13]. Columns [Gro00b]. Combinatorial [PRC00]. Combined [GLB^{+}21]. Combining [NL12]. come [Dav95]. Comedy [RP97]. Comes [Gro96b, GBL97]. comfort [AK99]. Coming [DSB96]. Commerce [GS95, Has02, Lit94, TNJ05, CPCT03]. Commercial [DWZ^{+}08, ZYXW11]. Common [Bol20, Che00, WJ07]. Communicating [CVDL05]. Communication [AdAFM95, Bir94a, Bir94b, CR04, CNI^{+}04, He04, MTD^{+}11, SB97, SREH04, Str99a, Stü95, SY05b, TH05, VJS^{+}23, Yan15, Chu95, SAH^{+}99]. Communications [Ano01n, Ano17c, Ano17d, DL22, MG99, PM16, Str99b, Tho00, ZRW^{+}19, Bra97a, Bra97b, Ens98, Lug99, Str99b]. Communities [Jai98a, MB99, ZLS14]. Community [BKCR13, LMH^{+}10, MP17, MHL10, New10, RWS^{+}94, Rui17b, WNZ10]. Community-Contributed [LMH^{+}10, MHL10, New10, WNZ10]. CommunityPVR [HG09]. Compact [DLB^{+}19, NPG13]. Companion [PGHK13, Gou99c]. Company [Abo99a, GOVI15, Swi96, Abo99b]. Comparative [RGH^{+}21]. Comparisons [Bul02]. Compatibility [Nac10c]. Compatible [BCGU10, RAE16, Hof94]. Competition [GL97b, Sch14]. Complementary [Chu02]. Complete [Car96, TCP04]. Complex [AMM21, CSZL20, Fer18, Hav00, HZK^{+}23, LSL24, VKV^{+}12, ZZC^{+}21]. Complexity [Bro02, GM01a, LSK^{+}22]. Compliance [LLD^{+}16]. Component [LZZ^{+}21]. Composing [DKRS04]. Compositing [Hay98]. Composition [WEV03, WDG95]. Comprehensive [TR01, WXY^{+}22]. Compressed [CDWT10, FZH^{+}21, ISSZ08, LMS^{+}22, WLZL12, XL22, YWT^{+}24]. Compressed-Domain [ISSZ08, WLZL12]. Compressing [TL97]. Compression [ASLB20, BD03, GYL^{+}23, GHD^{+}14, Hoc96, HHLW20, KD17, LHD^{+}20, Ma96, NAK^{+}22, Pan95, PMZ13, ZMFB21, FWI99, Hoc95, Hoc96]. Compression-Encryption-Hiding [BD03]. Compression-Then-Encryption-Based [ASLB20]. Compressive [YWT^{+}24]. COMPSAC [Ano18m, Ano18f]. Computational [Ada03, DV01, DV03, Gol03b, SCY^{+}16]. Computer [Ano13g, Ano14j, Ano15s, Ano16t, Ano16s, Ano16q, Ano16r, Ano16u, Ano17m, Ano17n, Ano17u, Ano18q, Ano18n, Ano18p, Ano18o, Ano19r, Ano19p, Ano19q, Ano20q, Ano20r, Ano20v, Ano20w, Ano20-27, Ano20-30, Ano20x, Ano20-28, Ano20y, Ano20-29, Ano20z, Ano21g, Ano21h, Ano21i, Ano21q, Ano21r, Ano21v, Ano21s, Ano21x, Ano21y, Ano21t, Ano21w, Ano21u, Ano21z, Ano22t, Ano22u, Ano22k, Ano22l, Ano22m, Ano22-28, Ano22-35, Ano22-29, Ano22-36, Ano22-32, Ano22-31, Ano22-30, Ano22-37, Ano22-33, Ano22-34, Ano22-38, Ano23m, Ano23n, Ano23o, Ano23-32, Ano23-33, Ano23-35, Ano23-36, Ano23-30, Ano23-38, Ano23-34, Ano23-31, Ano23-37, Ano23q, Ano24r, Ano24o, Ano24t, Ano24q, Ano24s, Ano24p, Ano24n, Ano24b, AAHA20a, AAHA20b, BH07, Dus00, EH05, Fal07]. Computer [FHK08b, Gou99d, Gou00, GCCGW17, Hol10, Jai16, Ma96, Sto06, SJ06, SD07b, SD07a, TN98, WSR^{+}95, YGM^{+}16, RM99, Ano18r, Ano18s, Ano19s, Ano23c]. Computer-Assisted [GCCGW17]. Computer-Based [AAHA20a, AAHA20b]. Computer-Navigated [BH07]. Computer-Supported [Dus00, YGM^{+}16]. Computers [GSW97, Ano19h, Ano20f, Ano20g, Ano20h, Ano24d, Ano24e]. Computing [Ano02c, Ano02e, Ano14i, Ano15p, Ano20i, Ano20s, Ano22w, Ano23w, Ano23v, Ano23x, Ano23p, Ano24j, Ano24a, BBCN12, Bol15a, Che21b, CJ12, CZCJ16, Del00, FBB^{+}05a, GH15, ISDH14, Jai94a, Jai96a, KM95, KMB97, KWYK22, LH18, MLJ06, MTH^{+}20, Nac00, Nac04, OO04, PCMT16, PEB09b, PEB09c, PPL08, RMPK07, Rui14b, Sch06, SB97, SSEKE19, TCS16, TATS94, Wea00, ZMvD^{+}08, ZXHS21, ZCWH15, SS00, Ano15o, Ano15r, Ano15q, Ano16p, Ano17b, Ano18d, Ano19a, Ano19b, Ano19c, Ano20a, Ano21a, Ano21b, Ano22r, Ano22s, Ano22b, Ano22c, Ano22d, Ano22h, Ano22i, Ano22j, Ano23d, Ano23e, Ano23g, Ano23f, Ano23j, Ano23l, Ano23k, Ano23r, Ano24g, Ano24h]. ComputingEdge [Ano20b, Ano20c, Ano21f, Ano21c, Ano21d, Ano21e, Ano22e, Ano22f, Ano22g, Ano23h, Ano23i]. Concept [LCSC11, NST^{+}06, SH10]. Concepts [Dav98a, Smi14, EMM^{+}98]. Conceptual [GCP01]. Concerns [Bal99b]. Concurrent [Emo09]. Condition [DHTF98]. Conducting [SGVT04]. Conference [Ano23a, Bol17, COCB14, CKRW00, HLT^{+}15, Kel17, RC06, RJ14, Reb95, JP13]. Conferences [Ano15f, GBH15, KAHHM96]. Conferencing [For02]. Confidence [GL20]. Configurable [Cou99]. Conformance [FF05]. Congress [Ano19t]. Connect [Ano17t, Ano19i, Ano19j]. Connected [Ano15u, Bov01]. Connection [Ano14r]. Conservation [KBB04]. Considerations [Pai13]. Consistency [LWZ^{+}18, TGB^{+}22, ZZZ^{+}21]. Consistent [Kas04]. Consortium [Hos98]. Constrained [HHLW20, ZLT^{+}12]. Constraint [GL20]. Constructed [YW14]. Constructing [CRB^{+}08, HOK00, KRN97, SS01, dLA01]. Construction [Ano02y]. Consumer [CCW07, Tes05, Dia99]. Consumers [SMF06]. Consumption [AAE23, Ded95, FDM^{+}15, KMJ17, LGT17]. Contact [ABPR^{+}19, McG04, Neu00]. Contact-Expressive [McG04]. Contact-Free [ABPR^{+}19]. Content [Ang03, Ano02f, Ano02k, Bal98, BCF04, BDL^{+}09, CDR^{+}10, CRD^{+}11, CASC10, Chu02, DYG^{+}00, DZS^{+}02, DD04, Dje02, EFS01, Fei09, FS04, FLZ17, Gol02b, Gol02c, GTH^{+}13, Gro99a, GR17, HKM^{+}10, HNDP17, HS17, KMJ17, LTM03, LRDTP13, LHX11, MM03a, McK05, MTS^{+}06, Nac00, NSS01, OSS04, PRC00, PMZ13, PP05, Rui15b, SYN^{+}18, SGK^{+}07, SMdS^{+}21, Sin04, Sla11b, SC97, SZ94, SMF06, SY05a, Sub06, SMH^{+}17, WKS^{+}03, WTW23, WBKW96, XSS14, YKHI94, YKW01, Pur97, CLSY12]. Content-Aware [GTH^{+}13, WTW23]. Content-Based [Ano02f, Dje02, GR17, MTS^{+}06, PRC00, SZ94, WBKW96, YKHI94, YKW01]. Content-Driven [SMdS^{+}21]. Contents [Ano14s, Ano14t, Ano16w, Ano16x, Ano17v, Ano17w, Ano17x, Ano17y, Ano18-36, Ano18-37, Ano18-38, Ano18-39, Ano19-30, Ano19-31, Ano19-32, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano21-32, Ano21-33, Ano21-34, Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano23-47, Ano23-48, Ano23-50, Ano23-49, Ano24w, Ano24x, ZH02, Ano13m, Ano14u, Ano14v, Ano15v, Ano15w, Ano15x, Ano19-33, Ano21-31]. Context [Ano02m, CDR^{+}10, CFR^{+}10, DD04, FOW^{+}22, Jai00d, MBM21, QZ14, SHG^{+}07, VBL^{+}14]. Context- [FOW^{+}22]. Context-Adaptive [QZ14]. Context-Aware [CDR^{+}10, MBM21, SHG^{+}07]. Context-Tagged [VBL^{+}14]. Contextual [WHM15]. Continuator [Ano02s]. Continue [Che22d, MB00]. Continuing [GR17]. Continuous [RR05, TR01]. Contract [RDBRD15]. Contracts [ZRD23]. Contradiction [Nac03a]. Contrary [CWY^{+}22]. Contrast [HT06, SS19]. Contrast-Sensitive [HT06]. Contrastive [GQFZ23]. Contributed [LMH^{+}10, MHL10, New10, WNZ10]. Contributions [Rui16a]. Control [Bak07, CMW^{+}23, EP06, FGR12, FRM^{+}07, LZZ^{+}18, MNKI07, MSW18]. Controlled [KMS16, SY05b]. Controller [DTF17]. Controlling [GPLR95]. Converge [San00a]. Convergence [El 18, Jai02b, LLS11, Meh99, MD00, XCZH19, Jai99c]. Converging [GL97b]. Conversational [SJ04]. Conversations [HRH^{+}21]. Conversion [CK96, MMSB15, ZLL^{+}23]. Converting [CBWD11]. Convexification [JLD07]. Convolution [GZP^{+}24, LCCL21]. Convolutional [FFW^{+}20, FOW^{+}22, FWL22, HZY^{+}23, LSK^{+}22, RLW^{+}21, VAAI24]. Cookbook [Tee99]. Cool [GBL97]. Cooperative [Dus00, YWWW19, KO94]. Coordinated [EMG97]. Coordinating [DWX23]. Coping [SSG19]. Copy [LCM^{+}12, LLY^{+}11, THJG13]. Copy-Detection [THJG13]. Copy-Location [LLY^{+}11]. Copyright [Bal98, DN00, Gol98, Gro95c, Gro98b, MB00]. Copywrong [DN00]. Core [Ano15y, Ano16-27]. Corner [Dav96a]. Cornerstone [Pet96]. Corporate [Ano02l]. Corporation [Car96]. Correlated [Gol04c]. Correlation [CL06, ZSZ^{+}20]. Corridor [GBL97]. Cosegmentation [LYGT19]. Cost [BP96, CZ96, DT94, DSGG21, GWC17, JWW^{+}14]. Cost-Based [BP96]. Cost-Effective [DT94]. Cost-Efficient [GWC17]. Couch [Jai97b]. Could [Bol07a, KAB^{+}23]. Counts [Nac00]. Coupling [ZJC^{+}20]. Courseware [MZSH01]. Cover [Ano09b, Ano16l, Ano16m, Ano17h, Ano17i, Ano17j, Ano18a, Ano18j, Ano18k, Ano18l, Ano19k, Ano19l, Ano19m, Ano19n, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano22x, Ano22y, Ano22z, Ano22-27, Ano23y, Ano23z, Ano23-28, Ano23-27, Ano24k, Ano24l, LT11, Ano13f, Ano14e, Ano14f, Ano15k, Ano15i, Ano15j, Ano18i, Ano19o, Ano21m, Ano21n, Ano21o, Ano21p]. Coverage [Bro96, Gro00a, PPP^{+}11]. Covered [Ano20-27, Ano20-28, Ano20-29, Ano21v, Ano21w, Ano22-32, Ano22-33, Ano23-32, Ano23-33, Ano23-34, Ano24s]. Covers [Ano14g, Ano15l, Ano16n, Ano16o, Ano17k]. COVID [Che21d]. COVID-19 [Che21d]. Create [MM03a, Yeh16]. Created [MB99]. Creating [DH01, Gri96, HH18, Ike02, Jam07, RSR^{+}01]. Creation [KTY00, McK05]. Creative [DHTF98, Jai96b, WZBKB17]. Creativity [Kel17, Pan07a]. Creator [CHCH22]. Creature [Gou99d, Gou99b]. Credit [Gro96b]. Crime [Gro96a, IGR^{+}20]. Critical [Kel14, Mak05]. Cropping [LL14]. Cross [CHSC22, CGW^{+}23, LWZ^{+}18, LW00, MLHK17, TYW23, VRSW08, WFZ^{+}20, WCQ^{+}22, ZDZ15, ZSZ^{+}22, ZZS^{+}24, ZYS^{+}10, ZSZ^{+}20, ZZS^{+}22]. Cross-Disciplinary [VRSW08]. Cross-Domain [CHSC22, CGW^{+}23, TYW23]. Cross-Media [LWZ^{+}18]. Cross-Modal [MLHK17, WFZ^{+}20, WCQ^{+}22, ZSZ^{+}22, ZZS^{+}24, ZYS^{+}10, ZSZ^{+}20, ZZS^{+}22]. Cross-Organizational [LW00]. Cross-Platform [ZDZ15]. Crossmodal [VRG18]. Crowcroft [Str00]. Crowdsensing [LJSP17]. Crowdsourced [SMH^{+}17]. Crowdsourcing [New10, WDZ^{+}17]. Cryptanalysis [LFX^{+}18]. Cryptanalyzing [LLL17a, LLLH18, LSL24, LLH22]. Cryptographic [DWN01]. Cryptography [Hat96]. CS [Ano14w, Ano16z, Ano16y, Ano18g, Ano18h, Ano23u]. CS-generic-full-km-2023.indd [Ano23u]. CSCW [GCM94]. CTU [LZZ^{+}18]. CTU-Level [LZZ^{+}18]. CUBE [MPL00]. Cues [DAL^{+}16, LTC^{+}20, LG05]. Cultural [CD14, Dav98b]. Culture [Bro05, Jai06, Kel15, Mak05]. Curating [Mur15]. Curious [Dav98b]. Curiously [Meh04]. Currency [Mos03]. Current [CIL15, GCK12, LLY14, VDL^{+}22]. curve [Dav98b]. Customers [GBL97]. Customizable [MTW97]. Customization [SY05a]. Cut [YWT^{+}07]. Cutout [YWT^{+}07]. Cyber [Ano16e, Ano16d, Ano17e, Gol97c, JXS^{+}17, WDZ^{+}17]. Cyber-Enabled [Ano16e, Ano16d, Ano17e, JXS^{+}17]. Cybernetic [Ano01b]. CyberPilot [Car96]. Cybersecurity [Ano14p, Ano15t, Ano16e, Ano16d, Ano17e, JXS^{+}17, Tit03b]. Cyberspace [GS95, Swi96]. Cycle [KBD^{+}05].

D [Ara99b, Str99b, ATS^{+}13, Ano02y, Ano14o, AHR^{+}13, BYZZ13, BCLH14, CCB^{+}98, CBWD11, CHWB11, CV14, CGW^{+}23, DMA^{+}12, DKRS04, GD14, Gou98b, Gou98c, Gou99b, Gro96b, Gro98b, Hor96d, HZK^{+}23, JGN^{+}13, KP24, KTY00, KWG10, LTD05, LD06, LSL24, LLZ22, LY19, MLM^{+}07, MTG97, MTS^{+}06, NYNI99, SZS^{+}20a, SZS^{+}20b, UTZH12, XPZ^{+}13]. D&I [Ano22-31, Ano24r]. D2D [ZCH^{+}19]. D2D-Enabled [ZCH^{+}19]. Daconta [Ver98]. Dagstuhl [HK17]. Daisy [LPSZ08]. Dance [Abo99a, CYL22, MDK10, Abo99b]. Dancers [Abo99a, Abo99b]. Dancing [Bro02, Nau98]. Dante [RP97]. Dares [FJ12]. Dark [Kri03a]. Dasarathy [Ma96]. DASH [Sod11]. Data [Ano16g, Ano16f, Ano17f, Ano18f, AAE23, BVA^{+}15, BLBC22, BDPN94, CCB^{+}98, CKN^{+}11, Che18c, Che21a, Che21b, Che22a, FSC08, For02, Gri96, GT11, Has02, Hoc96, HHLW20, HCH12, HUU^{+}23, ISDH14, Jai00a, Jai01b, Jai22, Kan20, KUM^{+}19, KTB^{+}22, LJSP17, LWL^{+}17, LYL^{+}15, LTS^{+}07, Ma96, MLHK17, MYBP^{+}22, Meh99, MM14, MTH^{+}20, MNK^{+}07, Pra24, RSS09, Rui14a, SBKK^{+}10, SXHH16, SKWB22, SZW^{+}19, TCS^{+}15, VGH^{+}18, WFZ^{+}20, WCS^{+}21, WB16, WNZ10, ZXHS21, ZCWH15, Ano18c, Ano24c]. Data-Based [KTB^{+}22]. Data-Driven [LYL^{+}15, WNZ10]. Data-Hiding [GT11]. Database [AN97, DSUG03, GSW97, GV07, Hor96d, KGU07, LJJT12, LCM09, MTW97, NTB97, Sak94, MDK97]. Databases [BCEH03, CG14, CFR^{+}10, Par98, SJ00, YKHI94, ZCSSM98]. dataLOFT [BM22]. Dataset [AFO^{+}20, AMM21, AAHA20a, AAHA20b, DOD22, GYL^{+}23, VKV^{+}12]. DataWeb [MTW97]. David [Bha00, MS95, Wil96]. Day [Smi11c]. Days [Ano02b]. DBMSs [PS97]. DCT [NAK^{+}22]. Dead [GR17, Rei98b, XSS14]. Death [Gro96a]. Debate [Hat96]. Decade [Bol07b]. Deception [ABPR^{+}19, KHA^{+}21]. Decision [FS04, SBKK^{+}10]. Decision-Making [SBKK^{+}10]. Decoding [HCAMJ17, HHLW20]. Deep [ALS22, Ano20j, BMDG20, Che19c, Che22b, DGT^{+}22, FFW^{+}20, GRP21, HG15, Jai03c, JCM^{+}23, LYGT19, LZDC19, LH18, Miy17, Pra23a, QPZ^{+}21, QGZF23, Rui14b, SSW20, TWY^{+}22, WFZ^{+}20, XL22, YTL^{+}24, ZZZ^{+}21, ZZS^{+}22]. Defect [WZL^{+}24]. Defenestration [DSB96]. Defined [WHTC24, ZKWH17b]. Defining [GS95, Par99, Pur98]. Definition [CMC^{+}20, Gro96b, YA14, YHX15]. Deformation [CSZL20]. Degree [TCM18, YZSY21]. Dehazing [LWN^{+}21]. Deitel [Gol97c]. déjà [Rei96b]. Delay [GHBC94, LFX^{+}18, LLH22].

Delay-Sensitive [GHBC94]. Delight [Zim03]. Deliver [HG09]. Delivering [Mil98, New94, SSG19]. Delivers [Hor96d, MB99]. Delivery [BDL^{+}09, CWZ^{+}05, FLZ17, HNDP17, LPL^{+}13, PAS^{+}13, PFAHA03, SYN^{+}18, SW04, SMH^{+}17, YA14, MDK97]. Dell [Gou99e]. Demand [BP96, BD03, BB04, DWX23, FHL04, HG09, Hor96d, LTS17, LL97, LV94, hMLDM98, SRW24, Kri99]. Democracy [GA15]. Democratization [GA15]. Demonstration [HS99]. Demystifying [Goo03]. Denoiser [XL22]. Denoising [FWL22]. DenseNet [WHW^{+}22]. Density [LLEK23]. Department [Che18a, Rui17a]. Departments [Ano09c, Fot06a]. Deployment [HNDP17, Jai06, JTCC09]. Depression [LXH^{+}22]. Depth [CV14, FZXL24, HLW16, LLEK23, SOS13, Wil96, Ano21f]. Depth-Guided [FZXL24]. Deregulation [GS95]. DeRose [MS95]. Description [Ano02k, Ano02j, KNC01, MJ02, WFI07]. Descriptions [MM03a]. Descriptors [DLB^{+}19, NPG13]. Design [AB94, BB02, Bla94, Bol08, BM98, FBB05b, HZSH23, KCRO10, Mak05, MMC01, Pai13, Par98, SD97, Ude02, XNS^{+}13, YGM^{+}16, Kog98, SAH^{+}99]. Designing [Gou15, MFG03, Reb95]. Designs [SH08]. Desire [NvOH05, vONH04]. Desktop [DB00]. Destruction [XLT^{+}21]. Detachable [ZF24]. detail [MBPM98]. Details [LWL^{+}20]. Detecting [FGD01, GLZW20, RT06, SNK99]. Detection [ASL^{+}24, ABPR^{+}19, AMM21, CCPC11, CWY^{+}22, DSGG21, HG23, HRH^{+}21, KHA^{+}21, KD17, LCL23, LW22, LCM^{+}12, LLY^{+}11, LLD^{+}16, LMS^{+}22, Mad06, MYBP^{+}22, MAHT98, NAK^{+}22, NL12, PZKV11, RLJ^{+}23, SS19, THJG13, TGB^{+}22, WZL^{+}24, WLZL20, WZL^{+}21, YWT^{+}24, ZDZ15, ZZL^{+}22]. Detectors [NL12]. Developing [EMS96]. Development [Agu96, BL01, GM01a, MZSH01]. Developments [GCK12]. Device [GCP01, GHD^{+}14, MOK^{+}15]. Device-Independent [GCP01]. Devices [Ano20j, BMDG20, CYW23, CCS16, HFdW11, HLB^{+}10, HRCK08, McG04, PFMMF23, SH08, ZDS19, ZOR^{+}09]. Devices-Interface [SH08]. DHNet [NAK^{+}22]. Diabetes [KTB^{+}22]. Diagnosis [DGT^{+}22, Par12]. DIBR [LZD^{+}22]. Dictionary [FGC^{+}14, HG23]. Diego [ZKWH17a]. Difference [BGC^{+}15]. Differences [ABPR^{+}19]. Different [CDD^{+}00, LHH^{+}24, LCL21]. Differentiated [Met00]. Diffuse [Bry00]. Diffusion [GT11]. Digesting [DWZ^{+}08]. Digital [Ano01j, Ano01r, Bal98, Bal99b, BDS^{+}09, Bir94a, Bir94b, Bol08, Bol20, BB04, Bro05, Bro02, CWH^{+}17, CDR^{+}10, Che22a, Che00, COH00, Cle08, DMA^{+}12, Dav98a, Dav02, Dim99, El 18, FCA08, FS01, Fei09, GQFZ23, GB98, Gol02b, Gol08a, GO96, GL97b, HT08, JJO^{+}04, Jai95a, Jai01c, LZH^{+}07, LLT10, MB99, MB00, MD00, MLJ06, Mil98, MTG97, Nac00, Nac04, Nau98, Pan02, PMR^{+}04, Pet96, PBM^{+}07, Pur97, Sch00, Shn02, SMF06, Vet04, Wan04, WNY09, WFRZ21, XXZ^{+}23, Zhe03, ZOR^{+}09, Bra98, Tob95]. Digital-Audio [WNY09]. Digital-Video [Ano01j]. Digital-Watermarking [GB98]. Digitally [FS01]. Dilemma [GL97b]. DIME [BB03]. Dimension [BQLN^{+}18, Gri96]. Dimensional [Bha00, Hav00, KCKC00, LLY^{+}18]. Dimensionality [GTQG16]. Diminishing [Sye01]. DiOS [BLA^{+}23]. Directed [FFW^{+}20]. Direction [MB00, WTW23]. Directional [SS19]. Directions [Hig08, LLY14]. Director [SSS^{+}19]. Dirichlet [MXH21]. Disability [MMRM08]. Disambiguation [PZW^{+}16]. Disaster [Che18d, GLZW20, Rei98a]. Disaster-Related [GLZW20]. Disciplinary [VRSW08]. Discipline [DH01]. Disciplines [DH01, Pan07a]. Disciplining [Gon00]. Discovering [JK19, YGN18, ZYXW11]. Discovery [AKMHP09, MHL10]. Discrete [MLHK17]. Discriminant [LTS^{+}07]. Discriminative [LWH^{+}18]. Discriminators [LTC^{+}20]. Disease [DGT^{+}22, DNM^{+}10]. disk [Car96]. Disks [GHBC94]. Disparity [JWW^{+}14]. Displaced [BBD^{+}00]. Display [Ano01d, AR05, EMG97, Foo06, SZS^{+}20a, SZS^{+}20b]. Displaying [JGN^{+}13]. Displays [BBD^{+}00, Pla00]. Dissociative [Meh04]. Distance [Ste01, Sye01, Tob95, Bra98]. Distant [FGL^{+}01]. Distillation [ZSZ^{+}22, ZZS^{+}24]. Distorted [QPZ^{+}21]. Distortion [WHW^{+}22]. Distortions [WLZL12]. Distraction [RLJ^{+}23]. Distraction-aware [RLJ^{+}23]. Distributed [Ano01j, BBCN12, BP96, BR96, CLD02, CT01, DH97, EMS96, For02, Gec97, GPLR95, HZK^{+}23, JN04, KOCG99, LMS09, MG96, PFMMF23, QZ14, SB97, SGK^{+}07, VKvBG95, WKC^{+}15, ZCSSM98, ZSM02, ZCC^{+}13, GHR00, Vin98]. Distributing [JGN^{+}13]. Distribution [Ano16c, LBZL22, OO04, ZKWH17a, ZKWH17b, HHI^{+}95]. DIVE [Hag96]. Diver [PMR^{+}04]. Diverse [CGW^{+}23]. Diversifying [ZG09]. Diversity [Ano23w, Ano23v, Ano23x, GM01a, Kun01, OWZ^{+}20, YGN18]. Diversity-Induced [YGN18]. Divide [HKR10, Pan02]. Dividing [MB99]. Divine [RP97]. Divx [Meh99]. DMB [SKK^{+}12]. Do [CSC^{+}20, CBC^{+}20, Gol04a, Tho00]. Dobson [Kra00]. Document [LW00]. Documentaries [ZMZB11]. Documentation [Hav00]. Documenting [KTP06]. Documents [DLH01, HB04, Jai98b, Jai98d, KE96, LM02, dLA01]. Does [Ada03, Bol09, CRL^{+}03, SA07a, Sla11b]. Doesn’t [RBK^{+}00]. Doing [Jai03a, NN08].

Domain [AS20, CHSC22, CGW^{+}23, ISSZ08, LTC^{+}20, QZ14, Rui16c, TYW23, WLZL12, YWT^{+}24]. Domains [CYW21]. Dominic [Gou98a]. Don’t [Gol06c, SS12, Ben96]. Doomed [BGPQ21]. Dot [ZCC^{+}13]. Double [NAK^{+}22]. Download [YS09]. Drawing [LLEK23]. DreamWorks [Gro95a]. Drive [Ano23w, Ano23v, Ano23x]. Driven [Ano18f, AAHA20a, AAHA20b, LYL^{+}15, OSS04, SMdS^{+}21, WHTC24, WNZ10]. Drives [Meh99]. Driving [Che19d]. DRM [TSDD08]. DSL [Gro98b]. DSMGN [YXH^{+}23]. DTT [Meh99]. DTV [Meh99]. Dual [KHY04, SLZ^{+}22, YXH^{+}23]. Dual-Supervised [YXH^{+}23]. Dually [LWW^{+}22]. Dubois [Bur09a]. Dueling [Meh99]. Dunhuang [LLP00]. Duplicate [CY13, LJZ^{+}13, NXKS13]. Durand [MS95]. DVD [Meh99, Par99, Tay99]. DVD-Video [Tay99]. Dynamic [BDS^{+}09, CYW23, CW09, Dav94, HNDP17, JM17, ROB00, STH08]. Dynamics [GCCGW17].

E-Books [Bal01, Sch00]. E-Commerce [Has02, TNJ05, CPCT03]. E-Health [MG12]. E-mail [Rei98a]. E-Revolution [Hal01]. E-Services [KGS^{+}02]. E-Speak [KGS^{+}02]. e-Video [San00a]. E.A.T. [Ano02b]. each [Gou99e]. Ear [TKBR^{+}07]. Early [Ano02b, GRG21, WXY^{+}22]. Early-Stage [GRG21]. Ears [FBB05b, Gol03a]. Eastman [Gou99e]. ECC [AS20]. ECG [HHLW20]. Ecological [MJSE15]. Ecosystem [SJ09]. Ecosystems [BLBC22]. eCubeLand [HUU^{+}23]. Ed [Bro96]. Edge [Ano22w, Ano24j, Che21b, HWZ^{+}23, LLEK23, MB00, RLJ^{+}23, SSEKE19, WLZ23, Ano20i, Ano19a, Ano19b, Ano19c, Ano20a, Ano23d, Ano23p, Ano24a]. Edge-Assisted [HWZ^{+}23]. edited [Gou98a, Gou99e, Hoc95]. Editing [ALK^{+}96, Dav03, LCCL21, WTW23, YCZ^{+}12]. Editor [Hat96, LS00, San00b, PCS^{+}15, And01, Ang03, BYZZ13, Del00, Dim07, Dje02, Dus00, Eff98, Gib98, Jai95a, Jai95d, Jai95b, Jai95c, Jai96b, Jai96c, Jai96a, Jai96d, Kla96, Lug99, Pan07b, Sin99, Sin04, Smi05, Sye01]. Editor-In-Chief [Jai95a, Jai95d, Jai95b, Jai95c, Jai96b, Jai96c, Jai96a, Jai96d]. Editorial [BLM^{+}22, JYS^{+}14]. Editors [LSZM00, BB04, CR04, CKAI09, DHH15, DG08, DV03, FGS09, GS06, GM01b, GM01a, GB96, GT08, HH05, Hua97, Moe97, NTB97, PK06, ROAS00, RMPK07, ZYNS16, LLP^{+}16]. eds [Kra00]. Education [AVW08, Ano21k, CBG09, FGL^{+}01, FHK08b, Jai97a, hMLDM98, SBL00, Sye01]. Education-on-Demand [hMLDM98]. Educational [DP95, FHK08a, Kel16, PVK^{+}16]. EEG [HB05]. Effect [LLJW15, Zha12]. Effective [CSZL20, DT94, DGT^{+}22, WdJ09]. Effects [ALK^{+}96, DFG^{+}14, Eff05, MJSE15, PKL15]. Efficiency [CMW^{+}23, HZSH23]. Efficient [ALS22, AAE23, CAJ04, CBWD11, CG14, EFPS16, FGSS08, FDM^{+}15, GWC17, GHD^{+}14, HLT23, LCM^{+}12, LSK^{+}22, XZYY13, YNC09]. Effort [MMC01, Jef99]. EFX [ALK^{+}96]. EGGAN [TSM21]. EIA [KMF^{+}01]. EIC [Ano09d, Ano10a, Gol02a, Gol03c, Gol03d, Gol03e, Gol04a, Gol04c, Gol04b, Gol04d, Gol06a, Gro97a, Gro98a, Gro99a, Gro99b, Gro00b, Gro00a, Gro00c, Gro01, Pan06, Pan07a, Pan08]. Elderly [PFMMF23]. Electrocardiography [LLLH18, YH16]. Electromagnetic [Bak07]. Electronic [Ano94b, CG95, Ded95, Jai03d, WKS^{+}03, WSI^{+}05, MDK97, Reb95]. electronics [Dia99]. Element [LLS11]. Email [TR01]. Embedded [ZJC^{+}20]. Embodied [Ing13]. Embodying [Ano01b]. Embracing [CO12, Che21a]. emerge [RFM95]. Emergency [CAJ04]. Emerging [AB12, Add00, DLB^{+}19, GCGR11, HRvL96a, HRvL96b, LPL^{+}13, MTH^{+}20, SC15, TCS16].

Emoji [LSM^{+}21]. Emotion [FOW^{+}22, GPA20a, GPA20b, GWD^{+}22, HYC^{+}19, HRH^{+}21, HYLK21, MBM21, Pic16, TJBBFB15, WCS^{+}21, ZYNS16]. Emotion-Aware [HYC^{+}19]. Emotional [CC98, GH15, XWJ16]. Emotions [SSW20, WEV03]. Empirical [DOD22]. Empowered [WLZ23, XXZ^{+}23]. Empowerment [KRB19]. Enable [DRB09]. Enabled [Ano16e, Ano16d, Ano17e, CDR^{+}10, FS01, JXS^{+}17, SZH^{+}08, TGC^{+}06, ZCH^{+}19]. Enabling [BBZB07, DWX23, GTH^{+}13, TNJ05, Vet04, Wan04]. Ename [PCKS00]. Encoding [FCL^{+}19, ZRD23]. Encounters [Kel16]. Encryption [ASLB20, AS20, BD03, Hat96, HZK^{+}23, ISSZ08, KWYK22, LLL17a, LFX^{+}18, LLLH18, LSL24, LLH22, PMZ13, YH16, ZZC^{+}21]. Encyclopedia [AK99]. End [HYLK21, SA07b, ZLL^{+}20]. End-to-End [HYLK21, ZLL^{+}20]. Ending [Hal94]. Endurance [GRG21]. Energy [FDM^{+}15, FLZ17, HZK^{+}23]. Energy-Efficient [FDM^{+}15]. Engagement [dG04]. Engaging [Bre04, RL06]. Engineering [Ano14q, CNN02, DH01, GM01b, GM01a, GS01, KJKS01, KNC01, SERL01, UTZH12, Ano21a, Ano21b, Ano22b, Ano22c, Ano22d, Ano23e, Ano23g, Ano23f]. Engines [NC17]. Enhance [ADV05, BDL^{+}09, VTJL^{+}08]. Enhanced [Ano16g, Ano16f, Ano17f, CLD02, GWL^{+}22, MSR^{+}15, PVK^{+}16, SY05b, ZZS^{+}22]. Enhancement [CHSC22, GLLW24, LMS^{+}22]. Enhancing [CWW^{+}19, DFG^{+}14, MBPM98, WHTC24, WH11, YZSY21, dG04]. Enjoys [JP13]. Enlightenment [McG06]. Enough [FHK08b, TGB^{+}22]. Enriched [BDS^{+}09, LGT17, MPM^{+}17, SH10]. Ensemble [PZW^{+}16, VAAI24]. Enters [DP95]. Entertainment [MPL00, MD00, ZCL^{+}04]. Entities [AD24, Cle08]. Entry [DYG^{+}00]. Environment [AB12, ALK^{+}96, AB94, BHLP16, Cur02, FGL^{+}01, KOCG99, LG05, OO04, ROB00, TY00, DF99, WGH94]. Environmental [MQC^{+}10]. Environments [Dav94, DSB96, God08, HOY99, Jai96a, Kra00, KG05, MZ97, RDLT06, SK03, STH08, VKV^{+}12, ZSM02, dG04, ABH^{+}95, Kra00]. ePub [Ano13e]. Equipment [MPL00]. Era [Ano01r, Gol02a, Miy17, VGH^{+}18]. Erica [Tee99]. Erratum [Ano20j, Jai97c, ZZS^{+}24]. Error [GT11]. Escaping [Ser05]. Escritoire [AR05]. Essence [GM01a]. Establishing [Rui15a]. Esteem [HLADCR^{+}21]. Estimate [SXHH16]. Estimating [MMC01]. Estimation [CDWT10, FZXL24, LZZ^{+}18, NGA23, PM16, SSLS14]. Ethernet [WO97]. Ethical [SAB^{+}20]. Ethnographic [GS94]. Europe [Meh99]. European [Tat96]. Europeans [Gro95a]. Evaluate [ZDS19]. Evaluating [BBMC02, LLL^{+}17b, MFG03, PS97, RM94, Str99a, WSR^{+}95]. Evaluation [LSG^{+}17, LDZS18, MG12, RGH^{+}21, SWdR^{+}08, Str99a]. Event [CLD02, Jai99a, LTM03, MXH21, PZKV11, TNJ05, WGY^{+}21, WJ07, WLPN11, ZDZ15]. Event-Aware [WGY^{+}21]. Event-Based [LTM03]. Events [Ano95b, Ano95c, Ano96c, Ano96d, Ano98c, Ano98d, Ano99b, Ano99c, Ano99d, Ano99e, Ano00b, Ano00c, Ano01o, Ano01p, Ano01q, Ano02v, Ano02w, Ano02x, Ano03g, Ano03h, Ano04h, Ano04i, Ano05e, Ano05f, Ano06e, Ano06f, Ano06g, Ano06h, Ano07g, Ano07h, Ano07i, Ano08e, Ano08f, Ano08g, Ano08h, Ano09h, Ano09i, Ano09j, Ano09k, Ano10f, Ano10g, Ano10h, Ano10i, Ano11g, Ano11h, Ano11i, FNHB05, PPP^{+}11, Smi13d, TML^{+}14]. Ever [Tai02]. Every [Smi11c, TH01]. Everything [FHK08b, NL99a, NL99b]. Evidence [GA20a, GA20b]. Evolution [EL00, MM03b, Sch00, Stü95]. Evolve [GB98]. Evolving [Ano20k, Ano21l]. Example [CDWT10, LWH16, Bay97]. Example-Based [CDWT10, LWH16]. Examples [Bul02]. Exceptional [Rui16a]. Exchange [Col94, Rei94c]. Execution [SK03]. Exemplars [LYL^{+}15]. Exhibition [Ano02r, Zhe00, JJO^{+}04, SJM06]. Exist [RBK^{+}00]. Expanding [Gol03c]. Experience [Ano04e, Bak06, CR04, CBWD11, Cle08, CSC^{+}20, CBC^{+}20, Coo01, CD14, GS06, Jai00a, Jai01a, Jai01c, Jai03c, MST^{+}06, PM16, PT15, Ste01, SY05b, ZYJ^{+}13, Fri95, SS00]. Experienced [WSR^{+}95]. Experiences [ADV05, COCB14, DFG^{+}14, Hol10, Jai02a, KJKS01, McG06, OGM^{+}17, PVK^{+}16, Rad05, SHG^{+}07, VTJL^{+}08]. Experiencing [ABC^{+}08, Tit07]. Experiential [Ano02c, WTL^{+}14]. Experiment [LH23]. Experiments [Bea95, Gol02b, WKC^{+}15]. Expertise [LEGFD08]. Experts [Ano15y, Ano16-27, Rui16c]. Explained [Car98, Hat96, MBE95]. Explaining [GSW97]. Explicit [GRP21]. Exploiting [GLLW24, XWJ16, XLL^{+}22, Zim03]. Exploration [AKMHP09, Gri96]. Explorations [Cam97]. Explore [Ano21l]. Explored [Gou00]. Explorer [Kla96]. Exploring [CL06, HWV^{+}09, KSPW07]. Express [Smi11a]. Expression [CNI^{+}04, LTS^{+}07, RLW^{+}21, RDBRD15, SLZ^{+}22, TSM21, TLJ^{+}19, Wan04, XLT^{+}21, ZZLG21]. Expressive [CDD^{+}00, MP17, McG04, ST16]. Expressiveness [CVDL05, HN07]. Extend [Bal98]. Extended [BLA^{+}23, FGR12, HLW16]. Extending [Cha96, MB00]. Extensible [TCPD10]. Extensions [YA14]. Extraction [AGR07, HLW^{+}19]. Extractor [CYL22]. Extreme [CLSY12, JM17]. Extreme-Dynamic-Range [JM17]. Eye [KMS16, NL12, Rei96c]. Eye-Controlled [KMS16].

Fabrication [May96]. Facade [LLZ22]. Face [Ano02i, BDD^{+}12, Bha00, CYW21, CGW^{+}23, CEG^{+}13, JKP12, NL12, VD00, Bha00]. Face-Detection [NL12]. Faces [FGS09, GB98, Gro95c, Gro00b, SNK99, WN94, YW14]. Facets [Smi13b]. Facial [CNI^{+}04, RLW^{+}21, TLJ^{+}19, XLT^{+}21, ZZLG21]. Facilitate [MTD^{+}11]. Facilitating [KGS^{+}02]. Factory [DF99]. Factual [LBZL22]. Fad [Bal99a]. Fail [BGPQ21]. Failure [RT06]. Fair [Dav02]. Faithfulness [Nac07]. Fall [LMS^{+}22, Sch06, ZZL^{+}22]. Family [Bol09, RAE16, WLZL20]. Farewell [Gro01]. Fashion [LLY14, SM18]. Fast [PMV^{+}22, XWJ16]. Fax [PHG94]. Fax-MIME [PHG94]. Feature [BCLH14, GZP^{+}24, GLB^{+}21, LBZL22, LZD^{+}22, LMS^{+}22, VVS21, WPC^{+}20, XZL21, ZLL^{+}20]. Feature-Guided [XZL21]. Features [BC09, CYW21, CMW^{+}23, FCL^{+}19, GSOI10, SSS^{+}19, YTL^{+}24]. Federated [BM22]. Feedback [MJSE15, PKL15, PT15, RR05, SM15, WMS05]. feeding [DB97]. Fi [Gou99e]. Fictionalization [Gor05]. Fictive [Mir03]. Field [HWZ^{+}23, LCCL21, Yu17]. Fields [CCS16]. Fights [IGR^{+}20]. Film [GCM94, SRA08, SSS^{+}19, ZMZB11]. Film-Grain [SRA08]. Films [JGN^{+}13]. Filter [LSK^{+}22, MZZ^{+}16, SRA08]. Filtering [BBD^{+}00, BB03, FGSS08, HLW16, LCSC11, Pla00]. Filters [ZGL^{+}14]. Finally [Jai02b]. Finance [ZK18]. Finch [Meh04]. Finding [Ano02d, BN01, DV01, MM14, PBLB12]. Fine [TSM21]. Fine-Grained [TSM21]. Finger [KZ06, ZYJ^{+}13]. Fingerprint [CNYL21, Gro98b, KHY04]. Fingerprinted [FS04]. Fingerprints [LCM^{+}12]. Fingers [FBB05b]. Finite [KWYK22]. First [Ano16-28, Ano17-27, Car96, GAL^{+}18, LFX^{+}18, LDZS18, LLH22, Rei99b]. First-Impression [GAL^{+}18]. First-Order [LFX^{+}18, LLH22]. Fish [TKBR^{+}07, VRSW08]. Fish-Bird [VRSW08]. Fisher [Hoc95]. Fit [Ada03, Ano02d, Ano16u, Ano16v, Ano17q]. Fixed [Ano01n]. Flashes [Rei99b]. flaws [Ano17l]. FLeak [WCQ^{+}22]. FLeak-Seg [WCQ^{+}22]. Fleming [Gou99d]. Flexibility [RCG^{+}99]. Flexible [EP06, KJKS01, TIS96]. Flow [GLB^{+}21, LW00, SS19, WXL^{+}17]. flows [Jef99]. Fluckiger [Str96]. Fluorescein [WCQ^{+}22]. Focus [Ano15g, Ano15h, Ano16k, GO96]. Fog [ZAKNY23]. Fog-Based [ZAKNY23]. Folk [Ano02e, LSKP09]. FoodLog [AO15]. Footage [WKY^{+}15]. Forecasters [SRW24]. Foreground [LTC^{+}20]. Foreground/Background [LTC^{+}20]. Forensics [BEUW12, JKP12, LJJT12, VAAI24]. Forest [DYG^{+}00]. Forges [Rui14c]. Forget [NN15]. Forging [ZLS14]. Formalization [RDBRD15]. Format [DTG^{+}08, DYT^{+}09, SKK^{+}12, SDEFK09]. Formats [CV14, DPC05]. Formerly [Gou99c]. Forms [Hor96d]. Forum [MB99, Rei94a, Rei94b, GO96]. Forward [AF18, Ano09d, Che18b, Che22d, Pan08]. forwarding [Jef99]. foundation [Bra97a]. Foundational [Tia06]. Four [KAHHM96]. Fractal [Hoc95]. Fractals [Meh99, Hoc95]. Frail [MYBP^{+}22]. Frame [ALS22, AFKN95, MMSB15, ZF24]. Frame-Skipping [ALS22]. Framework [AB03, AMM21, BLBC22, BDL^{+}09, CHK^{+}09, FNHB05, GTQG16, GWD^{+}22, HRvL96a, Hol10, HG15, HRH^{+}21, HLW^{+}19, KNC01, Lar03, LCM09, SBKK^{+}10, SMK^{+}10, SLZ^{+}22, THJG13, WXY^{+}22, ZWL^{+}20, ZLL^{+}20, GHR00]. Frameworks [MG96]. Francisco [Ara99b, Gou00, Str00]. Francois [Str96]. Free [ABPR^{+}19, HCLK23, SZS^{+}20a, SZS^{+}20b]. FreeD [Hor96d]. Freeman [Gou98a].

FreeWalk [NYNI99]. Frequencies [HS99]. Frequency [CYW21]. Fresco [LLP00]. friendly [SAH^{+}99]. Front [Ano14g, Ano15l, Ano16l, Ano16n, Ano16m, Ano16o, Ano17h, Ano17k, Ano17i, Ano17j, Ano18i, Ano18j, Ano18k, Ano18l, Ano19l, Ano19m, Ano19n, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano22x, Ano22y, Ano22z, Ano22-27, Ano23y, Ano23z, Ano23-28, Ano23-27, Ano24k, Ano24l, Ano13f, Ano14e, Ano14f, Ano15k, Ano15i, Ano15j, Ano19o, Ano21m, Ano21n, Ano21o, Ano21p]. Frontier [FCM^{+}22, Foo06, GH15, NTB97]. Fruition [BHLP16]. full [Ano23u, Ano23s]. Fully [LZDC19]. Function [FLZ17, GRP21, JW17, ZKWH17b]. Functions [Hor96d]. Fund [Ano22-31, Ano24r]. Fundus [WCQ^{+}22]. Furht [GSW97]. Fuse [CYW21]. Fusing [BVA^{+}15, SXHH16]. Fusion [AGR^{+}21, Ano17g, BCLH14, CWY^{+}22, DGT^{+}22, GZP^{+}24, GWD^{+}22, MBM21, MPST18, Pan07a, PZW^{+}16, SRW24, SLZ^{+}22, VRG18, WXY^{+}22, WTW23, YXH^{+}23, ZLL^{+}20]. Fusion-Based [DGT^{+}22]. Future [AB12, AA05, Ano02u, Bak07, Bal01, Bol20, BHS13, Che21d, GG23, GI11, GCK12, Ike02, Juh11, LVL09, LZZ^{+}21, LLY14, McG04, Nac00, Nac04, Pan07b, PKN^{+}04, Sto14, TSDD08, VDL^{+}22, Vai95, WZ10, ZKWH17a]. Fuzzy [ZAKNY23]. Fuzzy-Based [ZAKNY23].

G [Bha00, Kra00, MS95, IGB^{+}08, PAS^{+}13]. Gadgets [Nac03b]. Gaile [Bha00]. Gaining [GL97b]. Gains [Meh99]. Gait [LXH^{+}22, MMS13]. Gallery [Rei99a, Cam97, RP97]. galore [Gol94]. Game [ABC^{+}08, DB00, Hol10, Nac01, Smi10b]. Games [DOD22, Gou98a, Mal00, PEB09a, PEB09b, PEB09c, Wea00]. Gaming [Ano20j, BMDG20, GY97, Kri03b, Ser05]. GAN [LMS^{+}22, ZJC^{+}20]. GAN-Based [LMS^{+}22]. Gap [DGLW07, DV03, Smi12a]. Garment [CHCH22]. Garments [CHCH22]. Gary [Str99a]. Gated [CZY20, LTC^{+}20]. Gateway [PHG94]. Gateways [LW00]. Gaussian [XL22]. Gaze [GTZ^{+}00, SSLS14]. Gender [ABPR^{+}19]. Gene [LTS^{+}07]. General [Bro96, Car98, Ano22v, Ano23t]. Generalized [CYW21, LBZL22, MSW18]. Generate [LCL21]. Generated [CLSY12, XWJ16]. Generating [BCF04, CWH^{+}17, CYL22, XZL^{+}16]. Generation [AYK08, CV14, GV07, GHD^{+}14, Hir97, HWZ^{+}23, Hub04, LBZL22, MZZ^{+}16, MTG97, SvdM11, SH08, VBL^{+}14, XCZH19, YXH^{+}23, Vin98]. Generations [GI11]. Generative [CYL22, GL20, KP24, LWL^{+}20, WLL^{+}21]. Generic [Ano02k, Ano23u]. Genetic [GL03, ZAKNY23]. GeoDec [SBKK^{+}10]. Geodesics [LTD05]. Geodesics-Based [LTD05]. Geographic [CCB^{+}98, New10, WJ99]. Geographical [Kra96]. Geolocalization [ZHC^{+}20]. Geolocation [HWL^{+}14]. Geolocation-Aware [HWL^{+}14]. Geometric [KWG10, LJ05, LZD^{+}22, WLZL12]. Geosocial [JYS^{+}14]. Geospatial [SBKK^{+}10]. Geras [Gro95a]. Gestural [Que96]. Gesture [LSKP09, PKL15, PPL08, iWD15]. Gesture-Based [LSKP09]. Get [Ano15m, Ano15n, Ano20q, Ano20r, Ano24m, Dav99, HNM03, Pla96]. Gets [GO96]. Getting [Bal98, MD00]. Getty [Wil98]. Giblin [Bha00]. GISs [WSS02]. Give [GB98]. Gives [Bro96]. Glasgow [EL00]. glass [RM99, PHS03]. Glasses [SZS^{+}20a, SZS^{+}20b]. Glasses-Free [SZS^{+}20a, SZS^{+}20b]. Glassner [Gou00]. Global [GWL^{+}22, TY00]. Globalizing [Jai96c]. Glove [KCRO10]. Go [Ano16z, Ano14w, Ano16y, MB00, HOK00]. Goals [BVH^{+}03]. Goes [Ano02l, ROW^{+}97, Rui15b]. Going [HLNS17]. Golden [GA20a, GA20b]. Gone [Smi12b]. Good [Kol96, SRW24]. Goodman [Str99b]. Gordon [Bha00]. Got [Ano17l]. Governed [CRD^{+}11]. GPS [Dia99]. Gradient [AGR^{+}21, SS19]. Gradient-Based [AGR^{+}21]. Grail [Ano02f]. Grain [SRA08]. Grained [TSM21]. Grand [FJ12, JIW13, OK14, SS11]. Granular [WMS05]. Graph [ASL^{+}24, FFW^{+}20, FOW^{+}22, HZY^{+}23, QSL^{+}19, RLW^{+}21, SMK^{+}10, XCZY14, YKW01]. Graph-Based [XCZY14]. Graphic [GO96]. Graphical [FNM98]. Graphics [Ano22k, Ano22l, Ano22m, Ano23n, Ano23o, Gou00, Gro96a, Gro98b, Qui03, TN98, WSR^{+}95, Gou99a, RM99, Gou00]. graphing [Wil95]. Gray [FGD01]. Gray-Scale [FGD01]. Grayscale [PMV^{+}22]. Great [Smi11b]. Green [FDM^{+}15]. Grid [DSGG21, LVL09]. Grid-Based [LVL09]. Grids [BBCN12]. Groping [Rei96d]. Grosky [Gro97a]. Ground [Meh99, Pla96]. Group [LSZM00, Low95]. Grouping [LNM^{+}24]. Grouping-Based [LNM^{+}24]. Grow [Del01a]. Growing [GOVI15, LTS17]. Grows [Gro96b, GY97]. GSM [MB99]. Guest [JYS^{+}14, And01, Ang03, BYZZ13, BB04, CR04, CKAI09, DHH15, Del00, DG08, Dim07, Dje02, DV03, Dus00, Eff98, FGS09, GS06, Gib98, GM01b, GM01a, GB96, GT08, HH05, Hua97, LLP^{+}16, Lug99, Moe97, NTB97, PK06, PCS^{+}15, ROAS00, RMPK07, Sin99, Sin04, Smi05, Sye01, ZYNS16]. GUI [Gol07b]. Guidance [BBZB07]. Guide [Hoc96, MS95, Str96, Swi96, Wil95]. Guided [FZXL24, HGHW16, HG15, HLW16, JCM^{+}23, LJZ^{+}13, LHH^{+}24, LNM^{+}24, WPC^{+}20, XZL21]. Guidelines [CM01]. Guides [CCLS03]. Guitar [Fer18]. Guojun [SB97].

H [San00a]. H.264 [Kal06]. Habits [MZSH01]. half [Ano23t]. Halftone [GL10, GT11]. Hall [Gol97c, Str96]. Hallinan [Bha00]. Hand [Ano15f, PKL15, PPL08]. Hand-Gesture [PKL15, PPL08]. Handbook [Gou99c, Bro96, Gou99c]. Handheld [SH08, VTJL^{+}08]. Handley [Str00]. Handling [MAHT98]. Handwriting [AYK08]. Happy [Gol06a]. Haptic [CHK^{+}09, DFG^{+}14, JNK06, MNK^{+}07, NMK^{+}06, PK06, PBM^{+}07, SZH^{+}08, TGC^{+}06, WO06]. Haptic-Enabled [SZH^{+}08, TGC^{+}06]. Haptics [MTB06]. Hard [XSS14, ZCZ20]. Hardware [Hor97d]. Harmonizing [DTT^{+}13]. Harvesting [Emo09]. Harvey [Gol97c]. Hash [Gou99c, XNS^{+}13]. Hashing [GQFZ23, LZ16, LWZ^{+}18, MLHK17, YTL^{+}24]. HCFGs [YHFC14]. HCI [FBB05b, OGM^{+}17]. HDR [FCM^{+}22]. HDTV [Gro96b, Meh99]. Head [DOD22, KAB^{+}23]. Headers [GC96]. Heads [XGN06, Tab97]. Health [Bal99b, BMO18, BEG^{+}18, BM22, GO96, Hua97, Jai22, MYBP^{+}22, MG12, Kim95, BM22]. health-care [Kim95]. Health-X [BM22]. Healthcare [AO15, ASLB20, Dim07, KMB97, WdJ09, ZZL^{+}22]. Hearing [AdAFM95, MTD^{+}11, PPL08, SS01]. Hearing-Impaired [AdAFM95, MTD^{+}11]. Heights [Ano15y, Ano16-27]. Help [McG06, Rei97d]. Helps [Pla96]. Herding [Smi13a]. Here [Jai02b, SS12]. Heritage [Add00, AG00, CD14, FBB^{+}05a, Hav00, Oom11, PCKS00, ROAS00, SO00, Zhe00]. Heterogeneous [Ano01j, Ban97, HZY^{+}23, HLY10, SXHH16, XCZH19]. HEVC [HCAMJ17, LZZ^{+}18, MTXH19, YA14]. hHalf [Ano22v]. hHoriz [Ano23-45]. Hiding [BD03, GL10, GT11]. Hierarchical [BR96, KS06, LYGT19, ZZS^{+}22]. Hierarchy [JYXT11]. Higginbottom [Str99a]. High [AGR07, Ano01n, CWW^{+}19, CYW21, CMW^{+}23, CYW23, CW09, CMC^{+}20, FGC^{+}14, Gro96b, HS99, HZSH23, LLY^{+}18, NL12, RSR^{+}01, SGK^{+}07, Ser05, Vai95, YA14, YHX15]. High- [CYW21]. High-Bandwidth [Ano01n]. High-Dynamic-Range [CYW23, CW09]. High-Efficiency [HZSH23]. High-Performance [NL12, SGK^{+}07]. High-Quality [FGC^{+}14]. High-Resolution [Vai95]. High-Speed [CWW^{+}19]. High-Throughput [AGR07]. Highlights [TCP04]. Highway [Swi96]. Hip [TGC^{+}06]. Histograms [NAK^{+}22]. Historic [HOK00]. History [Ano22h, Ano22i, Ano22j, Ano23j, Ano23l, Ano23k, Car98, HLP^{+}14, Smi11c, ZZLG21]. HIT [Cam97]. Hits [MD00]. Hoc [GQ10]. Hoffman [Hat96]. Holodeck [Foo06]. Holographic [CCS16, ZDS19]. HoloLens [LDZS18]. Holoscopic [ATS^{+}13]. Holy [Ano02f]. Home [AYK06, JNK06, LTM03, MPL00, PKN^{+}04, Kim95]. Home-Based [JNK06]. Honeymoon [Gol07a]. Honeywell [HP96]. Hongjian [GSW97]. Honoring [Ano20s]. Hopes [GO96]. Horizon [Rui17b]. Horus [SGK^{+}07]. Hospitality [JP13]. HOST [Ano20t, Ano20u]. hot [Ano21f]. Hotels [Tit03a]. House [Ano14d, Ano14i, Ano14j, Ano14l, Ano14o, Ano14r, Ano14q, Ano14w, Ano15r, Ano16e, Ano16g, Ano16j, Ano16k, Ano16p, Ano16t, Ano16-28, Ano17e, Ano17f, Ano17g, LD06, SB97, Str99a, Ano15n, Ano16z]. HTML [Bal98, Gro95c, MB00]. HTTP [BQLN^{+}18, DTF17]. Human [Ano01h, Ano16i, Ano16j, AAE23, Bha00, DAL^{+}16, EH05, EYT^{+}06, GB98, Jai06, KWG10, KG05, LMCP18, MMS13, Shn02, TL97, iWD15, ZWL^{+}20, ZZL^{+}22]. Human-Centered [Jai06]. Human-Computer [EH05]. Human-Face [Bha00]. Human-Scale [KG05]. Human-Vehicle [EYT^{+}06]. Humans [Zen16]. Humphrey [Ano23-29]. Hybrid [CYW23, DWX23, LTS^{+}07, MYBP^{+}22, YNC09]. Hyperlinking [VRG18]. Hypermedia [CM01, CRL^{+}03, DF99, GLR95, hMLDM98, MS95, MdRCS15, TY00, GHR00, Low95]. Hyperspace [Rei97a]. Hyperspectral [ASL^{+}24]. HyTime [MS95, New95a, New95b]. HyVIS [Low95].

Ian [Str00]. ICME [IMK^{+}16]. ICMR [RJ14]. Ideas [CWH^{+}17, Pan08]. Identification [CNYL21, LHX11, WA16, WLZL20]. Identifying [WN94]. Identity [AHN^{+}18]. IEEE [Ano95a, Ano96a, Ano96b, Ano97, Ano98a, Ano99a, Ano00a, Ano17u, Ma96, And98, Ano94a, Ano94c, Ano03a, Ano03f, Ano04g, Ano05a, Ano06a, Ano06d, Ano07f, Ano08d, Ano11f, Ano13g, Ano14i, Ano14h, Ano14j, Ano14k, Ano15o, Ano15p, Ano15r, Ano15q, Ano15s, Ano16p, Ano16t, Ano16s, Ano16q, Ano16r, Ano16u, Ano17m, Ano17n, Ano18m, Ano18q, Ano18n, Ano18p, Ano18o, Ano18r, Ano18s, Ano19r, Ano19p, Ano19q, Ano19s, Ano19t, Ano20v, Ano20w, Ano20-27, Ano20-30, Ano20x, Ano20-28, Ano20y, Ano20-29, Ano20z, Ano20-31, Ano21q, Ano21r, Ano21v, Ano21s, Ano21x, Ano21y, Ano21t, Ano21w, Ano21u, Ano21z, Ano22t, Ano22u, Ano22-28, Ano22-35, Ano22-29, Ano22-36, Ano22-32, Ano22-31, Ano22-30, Ano22-37, Ano22-33, Ano22-34, Ano22-38, Ano23a, Ano23-32]. IEEE [Ano23-33, Ano23-35, Ano23-36, Ano23-30, Ano23-38, Ano23-34, Ano23-31, Ano23-37, Ano23-39, Ano24r, Ano24o, Ano24t, Ano24q, Ano24s, Ano24p, Ano24n, Gol03c, Gro97a, Rui14c, Rui15a, SZW^{+}19, Tho00, Ano17b, Ano18b, Ano18c, Ano18d, Ano19h, Ano20f, Ano20g, Ano20h, Ano20q, Ano20r, Ano21g, Ano21h, Ano21i, Ano22r, Ano22s, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22o, Ano22n, Ano23j, Ano23l, Ano23k, Ano23m, Ano23p, Ano23n, Ano23o, Ano23q, Ano23r, Ano24m, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Che18b]. IEEE OA OnePageAD.indd [Ano23-40]. if [Kol96]. II [Bra97b, FWI99, HRvL96b, New95b]. Illnesses [AC18]. Illumination [GLLW24]. Illusion [VTJL^{+}08]. Image [AD24, Ano20j, AK22, AP00, BMDG20, BN24, BBC00, BB03, Bla94, CZY20, CG14, CHCH22, DMA^{+}12, DGLW07, DGR13, DTT^{+}13, FWL22, GSW97, GLLW24, Hay98, Hir97, Hoc95, HG15, HZK^{+}23, HTH^{+}22, JK19, JYXT11, JCM^{+}23, Kas04, KWYK22, LJJT12, LL14, LLP00, LJZ^{+}13, LLL17a, LFX^{+}18, LLLH18, LWH^{+}18, LWN^{+}21, LSL24, LHH^{+}24, LJ05, LCM^{+}12, LLH22, LWH16, LH18, LCCL21, Ma96, MM14, NPG13, NC17, Par98, PPDH94, QJC^{+}18, QGZF23, Rui14a, Sak94, SJ00, STL18, SZ00, Sto06, SJ06, SD07b, SD07a, TH01, VCBM19, WLL^{+}21, WXY^{+}22, XZYY13, YNC09, YWT^{+}07, YCZ^{+}12, YXH^{+}23, YH16, Yeh16, YGN18, ZGL^{+}14, ZZC^{+}21, ZG09, ZH02, Wil95]. Image-Based [Hir97]. Image-Editing [YCZ^{+}12]. Image-Relevant [AD24]. Image-Scrambling [LLL17a]. Image-to-Image [CHCH22, HTH^{+}22]. Imagery [BDD^{+}12, BCK^{+}05]. Images [CW09, CDWT10, EBG98, FZH^{+}21, FGD01, GQFZ23, Gol94, GBL97, GL10, GT11, HLADCR^{+}21, Hor96d, HB04, LWL^{+}20, LT11, MSF16, PMV^{+}22, Ran08, TL97, VBL^{+}14, WFRZ21, YH13, ZYS^{+}10, ZWL^{+}20]. Imaging [AGR07, BYZZ13, DSGG21, GSW97, Jai95a, LD06, Shn02, WH97]. Immersive [AB12, APCH23, ATS^{+}13, AKMHP09, DSB96, FCM^{+}22, HOY99, HWZ^{+}23, Jai97d, KCRO10, LCL23, Moe97, PVAMM10, SYN^{+}18, VDL^{+}22, WLZ23, WSI^{+}05, ZCC^{+}13]. IMMPS [SD97]. Impact [Ano01f, Bak06, Bak07, CRB^{+}08, Che19a, DN00, EL06, Fei09, FHK08b, God08, KNLZ07, Nac00, Nac01, Nac07, NJ06, PEB09c, Pic16, RL06, Sch06, Ste01, WB16, ZMvD^{+}08]. Impacts [PT15]. Impaired [AdAFM95, MNK^{+}07, MTD^{+}11, PPL08, RHP^{+}06, SS01, Dim99]. Implementation [CAJ04, NG11, SZS^{+}20a, SZS^{+}20b, YZSY21, EMM^{+}98]. Implications [Kun01]. Implicit [GRP21, WCS^{+}21]. Importance [RDLT06, SRA08]. Impression [GAL^{+}18]. Improved [HZY^{+}23, JWW^{+}14, NGA23, WLWW21]. Improvement [LFX^{+}18]. Improvements [IGB^{+}08]. Improving [SA07b]. IMTV [HHI^{+}95]. In-Kernel [LKLH13]. In-Loop [MZZ^{+}16]. In-Vehicle [GRF^{+}16]. Incidental [XZL^{+}16]. includes [San00a]. Inclusion [Ano23w, Ano23v, Ano23x]. Incomplete [GL20, SXHH16]. Incongruity [WZL^{+}21]. Increasing [WDZ^{+}17]. Incremental [HLT23, ZCZ20]. Independent [GCP01]. Index [Ano95a, Ano96a, Ano97, Ano98a, Ano99a, Ano00a, Ano01a, Ano02a, Ano03a, Ano04a, Ano05a, Ano06a, Ano07a, Ano08a, Ano09a, Ano11a, Ano13a]. Indexes [Dav96b]. Indexing [BCGU10, Dje02, Gro97b, KAHHM96, LM02, SZ94]. Individualized [JNK06]. Individuals [MOK^{+}15]. Induced [YGN18]. Induction [WKY^{+}15]. Industrial [Miy17, VKV^{+}12]. Industry [FJ12, KSW12, KLKW09, SS11, Sti95, Ano21f]. Inference [XCZY14]. Infobahns [Sha95]. Information [Ano94c, Ano96b, Ano13g, Ano18g, Ano18q, Ano22-34, Ano23-35, Ano23-36, Ano23-38, Ano23-37, Ano24t, Ara99b, BCL00, Bry00, Car98, CHWB11, CHSC22, Che18d, CDP99, Gol04b, Gri96, Gro94, GL97b, GWD^{+}22, Jai97d, Jai00a, KLKW09, Kem95, Kra96, Liu94, MDK10, New10, NKZ^{+}22, PBLB12, Rei94c, Swi96, Tee98b, Tee98a, Tia06, WJ99, WHM15, WLZL20, XLL^{+}22, Ara99a, AK99, DH97, Kim95, Low95, NM98]. Information-Based [NKZ^{+}22].

Informationitis [Jai00b]. Informative [Wal03]. Informed [Swi96]. Infoscopes [Jai95b]. Infrared [YXH^{+}23, ZSZ^{+}22, ZLL^{+}23, ZZS^{+}24, ZCC^{+}13]. Infrastructure [KNLZ07, Liu94, Mil98]. Infrastructures [LVL09, Rad05]. Initial [CGM97]. Initiation [FGR12]. Initiative [LSG^{+}17]. Initiatives [Bal01]. Inmarsat [FTCS99]. Inner [Khu07]. Innovation [Ano16-28, Oom11, VGH^{+}18]. Innovations [Tit06]. Innovative [Gou98a]. Inpainting [LZH^{+}23]. Input [AYK08, MW11]. InSense [BPT06]. Insertion [LLJW15, LSM^{+}21, WY07]. Insights [BMO18, Kun01]. Inspection [CM01]. Inspired [PCMT16]. Instagram [HLADCR^{+}21]. Instance [LW22, MSW18, QPZ^{+}21]. Institute [PEB09a]. instruction [SBN97]. Instrument [FNM98, Pai13]. Instruments [CWH^{+}17]. Integer [LLY^{+}18]. Integrated [He04, HLN98, SJ00]. Integrating [cSC15, Dia99, GAL^{+}18, HS17, HOY99, Kra96, LKGR94, Pri99, Tho00, TCP04, WJ99, ZCSSM98]. Integration [Ano17g, BG96, BKCR13, GS01, Has02, Hos98, MPST18, WLZL20]. Integrity [SKWB22]. Intel [GO96, GL97b]. Intellectual [BHLP16, GS95]. Intelligence [Aar04, Ano18f, Ano23a, BEUW12, BBCN12, Che19b, IGR^{+}20, MKK^{+}15, Rui17c, WFZ^{+}20, WLZ23]. Intelligence-Empowered [WLZ23]. Intelligent [AdAFM95, Ano22o, Ano22n, AYK08, GZC09, HUU^{+}23, KMM^{+}10, KRB19, LHD^{+}20, LCCL21, SMdS^{+}21, XCZH19]. Intent [AYK06]. Intentions [CDD^{+}00, YWT^{+}07]. Inter [HZY^{+}23]. Inter-Relational [HZY^{+}23]. Interaction [Ano13d, Cha06, EYT^{+}06, FSC08, GB96, Ing13, Kra00, KMS16, MJSE15, Mak05, MTS^{+}06, Nac10b, NMK^{+}06, NGA23, PVK^{+}16, Que96, WZ10, YGM^{+}16].

Interactions [Ano16i, Ano16j, HWV^{+}09, HXXS18, Kra00, WDZ^{+}17]. Interactive [AAA^{+}09, BM98, Bre04, CVDL05, CCM^{+}04, COCB14, CK96, CCPC11, CBG09, COH00, Cle08, Cre05, DHH15, DDV03, FBB^{+}05a, FBB05b, Gou15, Gou98a, Gro95a, Hag96, HH05, HFK95, HRCK08, Hor96d, HW98, JNK06, JJO^{+}04, Jai97d, KMJ17, LL97, LV94, MSW18, Mal00, Mil98, PMR^{+}04, PD00, RGH^{+}21, SM15, SJM06, SHH09, VW03, WX23, Wal03, ZSN^{+}05, ABH^{+}95, Reb95]. Interactivity [Bak06, Gou98a, HN07]. Interchange [New95a, New95b]. Interdisciplinary [HH18]. Interest [BPT06, JCM^{+}23, YWT^{+}24]. Interest-Based [BPT06]. Interesting [Yeh16]. Interface [Ano01n, Ano18t, Ano18u, Bla94, Bro02, Del02, Gou15, Kim96, KGU07, Lar03, LSKP09, SREH04, SH08, TIS96, TGC^{+}06, Ano19i, Ano19j]. Interfaces [ADV05, Gol06c, Gug03, HKN04, KMS16, Ovi96, PK06, PKL15]. Interleaved [LCCL21, MdRCS15]. Interlinking [RSS09]. International [COCB14, HLT^{+}15, RJ14]. Internet [Coo01, MB00, San00a, Tee98b, Bra98, Bra97a, Bra97b, Cas96, CNI^{+}04, CWZ^{+}05, CMC^{+}20, El 20, FWI98, FWI99, GS95, Gro95c, GO96, Hor96d, LPL^{+}13, Lit94, LLZ03, ML18, Meh99, MN07, PKN^{+}04, PFAHA03, Sad20, San00a, San00b, Sod11, SSEKE19, Tee98a, ZAKNY23]. Internetworked [Str00]. Internetworking [Str00]. Interoperability [FF05, GC96, Gro95c, KLKW09, SS08]. Interoperable [TH05, Wan04, WSS02]. Interpolation [LWW^{+}22]. Interpretability [GG23]. Intersection [CV16]. Intervention [cSC15]. Interview [ZKWH17a]. Intranet [Gro96a]. Intraprediction [AGR^{+}21]. Introduction [And01, Ang03, BB04, CR04, CKAI09, Del00, DG08, Dim07, Dje02, DV03, Dus00, Eff98, FGS09, FHK08a, GS06, Gib98, GM01b, GM01a, GB96, GT08, HH05, Hoc96, Hos98, Hua97, Lug99, Moe97, NTB97, PK06, Pet96, ROAS00, RMPK07, Sin99, Sin04, Smi05, Sye01, ZXHS21, BYZZ13, DHH15, LLP^{+}16, PCS^{+}15, ZYNS16]. Intrusive [Bak06]. Invariant [GWL^{+}22, LZD^{+}22]. Invasive [LRCY07]. IoT [BLBC22, PFMMF23, ZZL^{+}22]. IP [FGL^{+}01, Gro98b, MB99]. IP/ATM [FGL^{+}01]. IPng [Bra97a]. IPTV [Jai05a]. ISBN [Ara99b, Bay97, Bha00, Bro96, Dav98a, Gol97c, GSW97, Gou98a, Gou99c, Gou99d, Gou99e, Gou00, Hat96, Hoc95, Hoc96, Kra00, Ma96, MS95, Pet96, San00a, SB97, Str96, Str99a, Str99b, Str00, Swi96, Tee98b, Tee99, Ver98, Wil96]. Island [JTCC09]. ISM [SZ17]. ISMA [FF05]. Isn’t [Jai03c]. ISO [Liu94]. Isochronous [WO97]. Isotropic [AK22]. Issue [Ano13d, BLM^{+}22, ZXHS21, WZ10]. Issues [JTCC09, LJSP17, PBM^{+}07, Sub06, ZMvD^{+}08, EMM^{+}98]. IT-general-CFP-hHalf [Ano22v]. Item [TYW23, Vet04]. Iterative [QJC^{+}18]. ITU [Gro96b].

J [Bay97, Hat96, MS95, Str99b, Tee99]. Jabber [SA07a]. Jackson [Bay97]. Japan [Cas96, Kah96]. Japanese [Sti95]. Java [Bay97, Ver98, Bay97, FGR12, Gro96b, Hor96d, Ver98, dOHS^{+}03]. Java-Based [Hor96d]. JavaScript [Tee99, Ver98, Tee99]. JavuNetwork [Kam99]. Jeff [Gou99c]. Jingle [SA07a]. Job [Ano15g, Ano15h, Ano16k, Ano18w, Ano18x, Ano19x]. Jobs [Ano18v, Ano21x, Ano21y, Ano21z, Ano22-35, Ano22-36, Ano22-37, Ano22-38]. Jock [Ara99b]. John [Gou99d, Pet96, Tee98b, Ver98, Wil96]. Joint [AS20, PMZ13, TML^{+}14]. Jointly [FGC^{+}14]. Jon [Str00]. Journal [Ano20q, Ano20r, Ano23q, Ano24b]. Journalism [DB05]. Journey [Khu07, Pan06, Yu17]. JPEG [AAE23, DTT^{+}13, EFPS16, HLNS17, RAE16].

JPSearch [DTT^{+}13]. Judge [LT11]. Jumping [PT15]. Just [Bal99a, Dav96a, OO04, Pac99, Smi13b]. Just-in-Time [OO04].

K-DIME [BB03]. Kalman [SRA08]. Karaoke [SLW98]. Kaufmann [Ara99b, Gou00, Hoc96, Str00]. Keep [Ano19u, Ano19v, Ano19w]. Keeping [Bay97, Rei96c]. Kernel [LKLH13]. Key [HG23, LHD^{+}20]. Key-Point [LHD^{+}20]. Keyboard [GO96]. Keyframe [PXL10]. Keyframe-Based [PXL10]. Keywords [ZH02]. Khalid [Hoc96]. Killer [Gol04d, Rei98a]. Kills [Bak07]. Kind [Kel16]. Kinect [MMS13, Zha12, ZYJ^{+}13]. King [Gol98]. Kinka [ROB00]. Kiosks [Gol98]. Kitsou [Bur09a]. Klewel [GOVI15]. Kluwer [GSW97, MS95]. km [Ano23u]. Know [Gol04a, NL99a, NL99b, Ben96]. Knowledge [AGR07, AD24, DWX23, FOW^{+}22, GLLW24, Jai01a, MHL10, NMK^{+}06, SY05b, YKHI94, Ano17s, Ano18-31, Ano18-32]. Knowledge-Assisted [YKHI94]. Knowledge-Aware [AD24, FOW^{+}22]. Known [Gou99c]. Kodak [Gou99e]. Kooks [Foo05]. Korea [Kri03b]. ks.indd [Ano23-45]. Kuji [ROB00]. Kulit [GI11].

L [Bay97, Bha00, Wil96]. Lab [Agu96, Pic16, Cam97]. Label [GL20, PTM11, WKC^{+}13]. Labeling [HXXS18]. Labels [GL20, ZSYW22]. Laboratory [KM95, RSR^{+}01, HP96]. Labs [Gro96b]. Lacey [Kra00]. Lag [LSL24, ZZC^{+}21]. Lag-Complex [LSL24, ZZC^{+}21]. LAN [And98, SA07b]. Lance [Hat96]. Landmark [MBX14, PZKV11]. Landscapes [KSPW07]. Landscaping [WZ10]. Language [Ano17g, AAA^{+}09, CL06, GAL^{+}18, KGU07, LTB23, MPST18, Ovi96, RDBRD15, TC01, Wan04, WLWW21, Car98, Hos98]. Language-Based [KGU07]. Languages [Cha96, JN04, RFM95]. Large [AMM21, BBZB07, CY13, DLH01, DGR13, DLW^{+}19, HCH12, JYS^{+}14, LL14, LLY^{+}11, LY19, NST^{+}06, PPP^{+}11, SWW11, VCBM19, XNS^{+}13, YHS11, YJC07, ZHC^{+}20, ZCSSM98, ZXHS21]. Large-Scale [BBZB07, CY13, DLH01, DGR13, DLW^{+}19, HCH12, JYS^{+}14, LL14, LLY^{+}11, LY19, NST^{+}06, SWW11, YHS11, YJC07, ZHC^{+}20, ZCSSM98, ZXHS21]. Late [WXY^{+}22]. Latency [CGM97]. Latent [CWY^{+}22, JK19, LJSZ14, MXH21, TSM21, TYW23, WTW23, YGN18]. Later [Car96]. Latest [SZ17]. Latin [RBK^{+}00]. Launched [Ano19-29]. Law [Foo05]. Layer [JYC08]. Layers [Tho00]. Layout [BBMC02]. Layouts [GL03, KS06]. LCEVC [FCM^{+}22]. LDA [JK19]. Leadership [Ano17u]. Leading [Ano15y, Ano16-27]. Leakage [WCQ^{+}22]. Leaps [Gri96]. Learnable [CYL22]. Learned [Gor05]. Learner [SC16]. Learners [SC16]. Learning [ALS22, Ano16g, Ano16f, Ano17f, ABC^{+}08, AAHA20a, AAHA20b, BBDS10, BVT^{+}23, BC09, Cha06, CL06, Che19c, CYW21, Che22b, CMW^{+}23, CGW^{+}23, CYW23, CT01, DGT^{+}22, EFS01, FGC^{+}14, FZH^{+}21, FHL04, GG23, GQFZ23, GPA20a, GPA20b, GWL^{+}22, GRP21, HYC^{+}19, HG23, HZY^{+}23, HRH^{+}21, HTH^{+}22, HLT23, HYLK21, Kel15, LLP^{+}16, LWH^{+}18, LWW^{+}22, LBZL22, LTS^{+}07, Miy17, QPZ^{+}21, Sch94, SK00, Smi13c, Ste01, TSM21, TWY^{+}22, VVS21, Ver98, WPC^{+}20, WCQ^{+}22, WH11, WKC^{+}13, XLT^{+}21, XZL^{+}16, YGM^{+}16, YH13, ZHC^{+}20, ZSYW22, ZZZ^{+}21, ZSZ^{+}20, Bra98, Dav98b, Tob95]. Learning-Based [FZH^{+}21, GPA20a, GPA20b, GRP21, ZSZ^{+}20]. Lecture [LKBE08, LCM09]. Leeuw [Dav98a]. Legacy [GS01]. Legal [KRB19, SAB^{+}20]. less [Ano15m, Ano15n]. Lessons [Gor05, Smi13c]. Let [Jai15, MSR^{+}15, Nac09]. Letter [LS00, San00b]. Letters [Ano18r, Ano18s, Ano19s, WSR^{+}95, Ano21g, Ano21h, Ano21i, Ano23m]. Level [IGR^{+}20, LZZ^{+}18, MB99, AK99]. Levels [LHH^{+}24]. Leveraging [Ano20j, BMDG20, CJ12]. Lexical [GWD^{+}22]. LFI [LCCL21]. LFI-Augmenter [LCCL21]. Li [ZKWH17a]. Libraries [BDS^{+}09, KS06]. Library [Ano14w, Ano16y, Ano18h, COH00, Ano16z]. Lies [Ano01r]. Life [BPT06, Bol08, Bra05, CRB^{+}08, Coo01, Dim07, Gro96a, KBD^{+}05, MZSH01, ROAS00, Dav99, HR00]. Life-Changing [Coo01]. Life-Long [MZSH01]. Lifeblood [Jai22]. Lifecycle [Gol04b]. Lifestyle [HT08]. Light [GLLW24, HWZ^{+}23, LCCL21, Yu17]. Light-Field [Yu17]. Like [Gol03d]. LIMAN [NKZ^{+}22]. Limited [Kol96]. Line [Gro95a, GY97, LLEK23, Rei94c, RWS^{+}94]. Linear [JLD07]. Link [Gro95a]. Links [HLADCR^{+}21, Hor96d]. Lisa [BBG^{+}96]. list [Ma96]. Little [FHK08b]. Live [BBG^{+}96, CHWB11, HOY99, Jai97e, NG11, WEV03, WZBKB17, XSS14]. Living [ROB00, Tit07]. Load [XZL^{+}16]. Load-Based [XZL^{+}16]. Local [GSOI10, GWL^{+}22, HHI^{+}95, JWW^{+}14, NKZ^{+}22]. Localization [GSOI10, THJG13, WZL12]. Location [GHD^{+}14, LLY^{+}11, SLW98, XCZY14]. Location-Based [SLW98]. Locative [Hig08, Wei08]. Logbook [VBL^{+}14]. Logging [BDD^{+}12, BPT06]. Logic [WKY^{+}15]. Logistic [LSL24, ZZC^{+}21]. Long [FS01, HG09, LZH^{+}23, MZSH01, SSW20, XSS14]. Long-Term [FS01, SSW20]. Look [LDZS18, Pra23a, WZL^{+}24]. Looking [Ano09d, Ano18w, Ano18x, Ano19x, Pan07b, Pan08, RM99]. Looks [HK17]. Loop [MZZ^{+}16, SW04]. Lopez [ZKWH17a]. Loss [GRP21, HYLK21, JYC08, ZHC^{+}20, ZZZ^{+}21]. Lossless [LHD^{+}20]. Lost [Tit06, Rei99c]. Love [Sch06]. Low [CYW21, DSGG21, GLLW24, LSK^{+}22, PMV^{+}22, Ser05, WHW^{+}22]. Low-Complexity [LSK^{+}22]. Low-Cost [DSGG21]. Low-Frequency [CYW21]. Low-Light [GLLW24]. Low-Resolution [PMV^{+}22]. LSB [FGD01]. Lu [SB97]. Lucent [MB99]. Lurking [EH05].

M [ABC^{+}08]. M-Learning [ABC^{+}08].

Machine

[AAE23, CRB^{+}08, GG23, MKK^{+}15, Sch06]. Machines [Kan20, Kel15]. Mackinlay [Ara99b]. Made [Smi11c]. Magazine [Ano22o]. Magic [SKSH08]. Magnitude [GLB^{+}21]. Mail [CG95, Rei98a, Hor96d]. Mailing [HLN98]. Major [Ano16q, Ano17n, Ano19d, Ano21j, Ano23s, Sub06]. Majority [WCS^{+}21]. Make [Gol98, MPL00]. Maker [Car96]. Making [Ano02g, BGC^{+}15, DT94, GLN^{+}08, MS95, NSS01, Neu00, SBKK^{+}10]. Malaysia [GO96, GBL97]. Malware [WLZL20]. Manageable [DWX23]. Management [AN97, Ano01j, BMR99, BR96, CDR^{+}10, CKAI09, Che18d, Che21d, Che22a, DTT^{+}13, DSUG03, DLW^{+}19, Gol04b, HLY10, ISDH14, LRDTP13, LY19, MLJ06, Wan04]. Management-Enabled [CDR^{+}10]. Managing [BHLP16, For02, GM01a, Gro99a, Hor96d, Kra00, LMS09]. Manga [AFO^{+}20, HLW^{+}19]. Manga109 [AFO^{+}20]. Manifold [GL20]. Manifolds [ZG09]. Manipulating [YHX15]. Manipulation [TSM21]. Manual [YNC09]. Manuals [DLH01]. Many [FGS09, LKLH13, Smi14]. Map [Car96, CV14, HLW16, LSL24, SMK^{+}10, ZZC^{+}21]. Map-Search [SMK^{+}10]. Mapping [Gro96a, HGHW16, MSW18]. Maps [BN01, COH00, Hor96d, ZSN^{+}05]. Mark [Gol98, Str00]. Market [Gro95a, WSR^{+}95, Gro96b]. Markets [Meh99]. Markov [SvdM11, iWD15, WKY^{+}15]. Mars [Gol07a]. Mask [YXH^{+}23]. Mass [Bha00, Gou99c, Gou99d, RL06, San00a, Str99a]. Mass-Scale [RL06]. Masses [TCM18, Wea00, Tay99]. Master [Gou99c, Sto06, SJ06, SD07b]. Mastering [Gou98b, Gou98c]. Masterpiece [MTS^{+}06]. Masterworks [Gou98a]. Masthead [Ano18y, Ano18z, Ano18-27, Ano18-28, Ano19y, Ano19z, Ano19-27, Ano19-28, Ano20-32, Ano20-33, Ano20-34, Ano20-35, Ano20-36, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano23-41, Ano23-42, Ano23-44, Ano23-43, Ano24u, Ano24v]. Matching [FCL^{+}19, JKP12, JWW^{+}14, LJZ^{+}13, SMK^{+}10, XZL21, XZYY13]. Math [AVW08]. Matter [Gol07b, Sla11b]. Matters [Ano14l]. Max [PP05]. Maze [Jai95d, Rei98c]. MBone [Cas94]. McClellan [Bay97]. Me [KS12]. Meaning [DV01, Foo05]. Measurement [LMCP18]. Measuring [OR02]. Mechanism [LLY^{+}18]. Mechanisms [CYW23, ZWH^{+}21]. Med [Hof94]. Media [Abo99a, Abo99b, Ada03, And08, Ano98b, Ano01b, Ano01f, Ano01g, Ano01h, Ano01i, Ano02f, Ano02h, Ano02y, Ara99b, Ara99a, Bak06, Bak07, Bay97, BBCN12, Bha00, BMO18, Bov01, Bre99, Bro96, Bur09a, Bur09b, CZY20, Car96, CRB^{+}08, CL10, Cha06, CGM97, CWZ^{+}05, CEG^{+}13, Dav98a, Dav02, Del01a, DP95, DN00, DV01, DV03, DK04, DAC05, Fal07, Fei09, Fle06, Foo06, FRM^{+}07, FHK08b, GA15, God08, Gol95, Gol96, Gol97a, Gol97b, Gol97c, GSW97, GB98, Gol04c, Gol06b, Gol06c, Gol07a, Gol07b, Gol08b, Gol08a, GV07, Gou98a, Gou98b, Gou98c, Gou99c, Gou99a, Gou99d, Gou99e, Gou99b, Gou00, GTH^{+}13, Hal01, HR00, HKM^{+}10, Hat96, HM95, HWV^{+}09, Hig08, Hoc95, Hoc96, HLP^{+}14, HKN04, HS08, Ing13, Jai97b, Jai97c, Jai97d, Jai97e, Jai97a, Jai98a, Jai98b, Jai98c]. Media [Jai98d, Jai99c, Jai99a, Jai99b, Jai00a, Jai00b, Jai00c, Jai00d, Jai01a, Jai01b, Jai01c, Jai02a, Jai07, Jam07, JW17, Juh11, KNLZ07, KZ06, Khu07, Kol96, Kra00, LLP^{+}16, LWZ^{+}18, Luo04, Ma96, MSW18, Mal00, MST^{+}06, MPG07, McG06, Meh98, MS95, Mir03, Nac00, Nac01, Nac04, Nac07, Nac10b, Neu00, NJ06, OB00, Pac99, PVK^{+}16, Par98, Pet96, PEB09c, RL06, RDD09, RDBRD15, Rou99, SSG19, San00a, Sch06, Sch14, SB97, She99, Sin99, Smi12c, SBB^{+}13, Ste01, Sto06, SJ06, SD07b, Str96, Str99a, Str99b, Str00, Swi96, Tee98b, Tee99, Tia06, TCPD10, Tit02, TSDD08, TR01, Ver98, WZL12, WFZ^{+}20, WLN^{+}21, WLZ23, Wei08, WSR^{+}95, Wil95, Wil96, WO06, WTL^{+}14, XCZY14, XCZH19, YJC07, ZMvD^{+}08, ZRD23, Pur97]. MediaEval [LSG^{+}17]. Mediascape [Rou99]. Mediascapes [SHG^{+}07]. MediaStation [LKGR94]. Mediating [MP17]. Medical [BLBC22, LHH^{+}24, LH18, Par12, WH97, Hof94]. Medici [Bre99]. Medicine [HZY^{+}23, LH18, QSL^{+}19, KO94]. Medusa [WGH94]. Meet [Bol07a, JW17, ZSYW22]. Meeting [CKRW00, PBLB12, Shn02, BHS13]. Meetings [NYNI99]. Meets [Ano17g, Che22b, MPST18, PKN^{+}04, SS11, SM18]. Mega [JS13]. Member [Gro00c]. Members [Gro99b, Pra23b, Ma96]. Membership [Ano13h, Ano14l, Ano16v, Ano17q, Ano18-35, Ano17s, Ano18-31, Ano18-32]. Memory [CG14, CWY^{+}22, DD04, DAL^{+}16, HCLK23, MNKI07]. Memory-Based [MNKI07]. Memory-Efficient [CG14]. Mental [AC18]. Menu [PKL15]. Mercuri [HP96]. Mercury [Hor96d]. Merge [LL97]. Merging [Jai96d]. MERL [ABH^{+}95]. Merwin [Ano17u, Ano18-29]. Mesh [LLY^{+}18, LY19, MMSB15, XLL^{+}22]. Mesh-Based [MMSB15]. Meshes [LTD05]. Message [Ano09d, Ano10a, Del02, Gol02a, Gol03c, Gol03d, Gol03e, Gol04a, Gol04c, Gol04b, Gol04d, Gol06a, Gro98a, Gro99a, Gro99b, Gro00b, Gro00a, Gro00c, Jai95a, Jai95d, Jai95b, Jai95c, Jai96b, Jai96c, Jai96a, Jai96d, Pan06, Pan07a, Pan08]. Messaging [KG97]. Meta [Nac00, Nac04]. Metadata [Bul04, CKAI09, FDM^{+}15, KBD^{+}05, LMS09, NvOH05, SS06, SBB^{+}13, SGS01, Tes05, vONH04]. Metalanguage [TKBR^{+}07]. Metamorphosis [OKN^{+}99]. Metaphor [AP00, CT01]. Metaverse [BLA^{+}23, Che22c, DL22, FCM^{+}22, XXZ^{+}23]. Method [Ano17r, CWH^{+}17, CCW07, FWL22, LY19, NGA23, WNY09]. Methodology [BL01, MG12]. Methods [HKC07, MTH^{+}20]. Metric [GPA20a, GPA20b, WH11]. Metrics [MMC01, MM14]. MHEC [EMM^{+}98]. MHEC-5 [EMM^{+}98]. MHEG [Col94, MBE95]. MHEGAM [KG97]. Michael [Gou99e, Ver98]. Michelle [Tee98b]. Micro [BBCN12, CHH98, JS13]. Micro-University [CHH98]. Microarray [LTS^{+}07]. Microcosm [Hal01, GHR00]. Microelectronics [May96]. Microexpression [GLB^{+}21, SLZ^{+}22, VVS21]. Microsoft [Tee98b, Gro95a, Kla96, Rei96a, Tee98a, Zha12]. Microtuning [Mou06]. Middleware [Cou99, TCPD10, WSS02]. MIDI [LHA06, Mou06]. Might [Ano02u]. Migrating [Nac03b]. Milano [Gou98a]. Millennium [BCL00, BCL99]. Miller [Gou98a, Swi96]. Million [SMH^{+}17]. MIME [PHG94]. Minding [Smi12a]. Minimal [EL06]. Minimal-Impact [EL06]. Minimally [LRCY07]. Mining [BHS13, HS17, LEGFD08, SWW11, WCS^{+}21, WLPN11, WKC^{+}13, YHS11]. MIPR [SZW^{+}19]. Mixed [BVT^{+}23, CCM^{+}04, HFdW11, HS08, LRCY07, Rad05]. Mixed-Media [HS08]. Mixed-Reality [HFdW11, LRCY07, Rad05]. Mixing [MKYH03]. Mixture [WHW^{+}22]. ml.indd [Ano23s]. MMAC [ZXHS21]. MMD [GSW97]. MMHealth [BEG^{+}18]. MMT [LPL^{+}13]. MMX [GL97b]. MobiDENK [KBB04]. Mobile [Ano01i, BBZB07, CG14, CFR^{+}10, DNM^{+}10, GCGR11, GHD^{+}14, Hak15, HFdW11, HWV^{+}09, HLB^{+}10, HRCK08, Hub04, ISDH14, Juh11, KBB04, LLS11, Lug99, Nac03b, OO04, PM16, ROW^{+}97, SKK^{+}12, STH08, Str99b, SY05a, Sub06, SSEKE19, SH08, Tit07, VBL^{+}14, WZL12, WDZ^{+}17, Wei08, WZ10, WKC^{+}13, WTL^{+}14, YHFC14, ZOR^{+}09, Dia99]. Mobility [HLY10, Jay03, Meh99]. Mobilized [Aut11]. Modal [MLHK17, WFZ^{+}20, WCQ^{+}22, ZSZ^{+}22, ZLL^{+}23, ZZS^{+}24, ZYS^{+}10, ZSZ^{+}20, ZZS^{+}22]. Modalities [WZL^{+}21]. Modality [MTD^{+}11]. Model [Ded95, GL20, HFK95, KE96, LZZ^{+}18, Meh99, MBM21, OSS04, TYW23, TWY^{+}22, WZL^{+}24, WJ07, ZLL^{+}23, ZLT^{+}12, ZDS19, EL00]. Model-Driven [OSS04]. Modeling [AYK06, Ang03, Car98, CL10, CPCT03, CEG^{+}13, GCP01, HUU^{+}23, KCKC00, LTM03, MBX14, Ovi96, QZ14, SCY^{+}16, SC16, SSW20, WZL^{+}21, ZYNS16, KC00, HLT^{+}15]. Models [Dav96b, KWG10, SvdM11, UTZH12, WH11, YWT^{+}07]. Modern [Bol09, CJ12]. Modes [Mur15]. Modification [LW22, SS19]. Modulation [ST16]. Molecular [Hal01]. Molecules [QSL^{+}19]. Momentous [Che20]. Mona [BBG^{+}96]. Money [GS95]. Monitoring [AC18, BCD^{+}02, GGX^{+}17, MQC^{+}10, TWY^{+}22]. Monroe [Tee99]. Monument [KBB04]. Moon [WSR^{+}95]. Moratorium [Bul04]. Moreland [Wil96]. Morgan [Ara99b, Gou00, Hoc96, Str00]. Morphing [CDD^{+}00]. Motion [CSZL20, KTP06, VTJL^{+}08, VVS21]. Motions [KAB^{+}23]. Motivate [MZSH01]. Motivator [PEB09b, PEB09c]. Mouse [Jai97b]. Mouth [WDZ^{+}17]. Move [Gol06c]. Movement [DOD22, Eff05]. Movements [dG04]. Movie [DFG^{+}14, KSW12, VW03]. Movies [YW14]. Moving [Che18b, Che22d, VGH^{+}18]. MP3 [MB99]. MPEG [Ano02k, Ano02j, Bal98, BCGU10, BCL99, BCL00, BNZ05, BVH^{+}03, Chi95, DKRS04, DPC05, DTG^{+}08, DYT^{+}09, DB05, DLB^{+}19, DAC05, FCM^{+}22, GCGR11, GKL^{+}98, HKM^{+}10, KLKW09, LPL^{+}13, NL99a, NL99b, NAK^{+}22, Pan95, PFAHA03, QH05, Qua13, RDD09, SS01, SDEFK09, SYJC06, Smi05, Sod11, TCPD10, TNJ05, TLS04, VDL^{+}22, Vet04, Wan04, Wil98, XDL04, ZRD23]. MPEG-2 [GKL^{+}98, Wil98]. MPEG-21 [BVH^{+}03, DB05, KLKW09, RDD09, TNJ05, TLS04, Vet04, Wan04, ZRD23]. MPEG-4 [Bal98, BCL99, BCL00, DKRS04, NAK^{+}22, PFAHA03, SS01, SYJC06, XDL04]. MPEG-5 [FCM^{+}22]. MPEG-7 [Ano02k, Ano02j, NL99a, NL99b, TLS04]. MPEG-7-Compatible [BCGU10]. MPEG-A [DPC05, HKM^{+}10]. MPEG/Audio [Pan95]. MRI [XL22]. MSN [Gro95c]. MTAC [GY97]. Much [Bol09]. Mulsemedia [CSC^{+}20, CBC^{+}20, SMdS^{+}21]. Multi [Ano15y, Ano16-27, MSW18, WTW23, ZCH^{+}19]. Multi-Attribute [WTW23]. Multi-Bitrate [ZCH^{+}19]. Multi-Core [Ano15y, Ano16-27]. Multi-Instance [MSW18]. Multiattention [NKZ^{+}22]. Multicast [JYC08, JTCC09, LLZ03, YS09]. Multichannel [WFRZ21]. Multicross [LZZ^{+}21]. Multidevice [KMM^{+}10]. Multidimensional [WX23]. Multidomain [CHCH22]. Multifaceted [SC16]. Multigraph [ZZS^{+}22]. Multilabel [GL20, ZSYW22]. Multilevel [AB97]. MultiMedia [HLT^{+}15, Aar04, APCH23, AN97, AB97, Agu96, ALS22, AO15, AFO^{+}20, AdAFM95, AKMHP09, Ang03, Ano94c, Ano01f, Ano01i, Ano01j, Ano01n, Ano01r, Ano02d, Ano02l, Ano02k, Ano02m, Ano02i, Ano16c, Ano16e, Ano16d, Ano16g, Ano16f, Ano16h, Ano16i, Ano16j, Ano17e, Ano17f, Ano17c, Ano17d, Ano17g, AB94, ABC^{+}08, BYZZ13, BL01, BEUW12, BCL00, Bea95, Ben96, BGC^{+}15, BP96, Bol06, Bol07a, Bol07b, Bol09, Bol15a, Bol15b, Bol17, BMO18, BEG^{+}18, BCF04, BHS13, BB04, Bra98, BBMC02, BDPN94, Cam97, CR04, CDD^{+}00, CKN^{+}11, CAJ04, CRD^{+}11, CG95, CO12, CSJ^{+}18, CLD02, Cha06, Cha96, CHH98, CKAI09, Che04, CL06, CCPC11, Che18d, Che18c, Che19b, Che19d, Che19c, Che21a, Che21b, Che21d, Che21c, Che22a, Che22b, Che22c, CBG09, Che00, cSC15, CKRW00, Col94, CM01, Cou99]. Multimedia [CZCJ16, DLH01, Del00, Del01b, DF99, Dia99, DPC05, DYG^{+}00, DD04, Dim07, DS15, DAL^{+}16, Dje02, DTG^{+}08, DL22, DNM^{+}10, DDV03, Dus00, Eff98, EMS96, EFS01, El 18, El 20, Emo09, Ens98, EP06, FGL^{+}01, FS04, For02, FTCS99, FHK08a, FHK08b, FJ12, FLZ17, Fur94, FWI98, FWI99, GL97a, Gec97, GHBC94, GT08, Gol03a, Gol97c, GSW97, Gol98, Gol02c, Gol03b, Gol03e, Gol04c, Gol04b, Gol06b, Gol06c, GV07, Gon00, Gou98a, Gro94, GO96, Gro96b, GBL97, GY97, GZC09, GR17, Gug03, GPLR95, GH15, GRG21, Hak15, Han15, Han17, HPH05, He04, HTL96, HRvL96a, HRvL96b, HLN98, HS17, HLADCR^{+}21, HFK95, HRCK08, Hol10, HS99, HWL^{+}12, HH18, Hor96d, Hos98, HP96, Hua97, HW98, HCH12, IMMP99, Ike02, ISDH14, JJO^{+}04, Jai06, Jai94a, Jai94b, Jai94c, Jai94d]. Multimedia [Jai98b, Jai99b, Jai01b, Jai03d, Jai03a, Jai05b, Jai07, JYS^{+}14, JN04, JXS^{+}17, Kam99, Kan20, KMF^{+}01, KGG96, KE96, Kel14, Kem95, KMM^{+}10, KG97, KTB^{+}22, Kim95, KCRO10, Kog98, Kol96, KBD^{+}05, Kra96, Kri03a, Kri99, KBB04, KGU07, KMS16, LMS09, LSG^{+}17, LCL23, LPSZ08, LM02, LD06, LVL09, LJSP17, LWL^{+}17, LGT17, LPL^{+}13, LEGFD08, LCG^{+}94, LZ16, LDZS18, LXH^{+}22, LH97, MTB06, Mak05, MBPM98, MLM^{+}07, MG99, MPL00, May96, McG06, MHL10, MTH^{+}20, MPST18, MMRM08, MDK97, MTS^{+}06, MSR^{+}15, MG96, MZSH01, NAV04, NvOH05, NST^{+}06, New94, New95a, New95b, NTB97, OO04, OLIG04, Pan02, PK06, PCMT16, PMZ13, PMGCCPRM11, PCS^{+}15, PHG94, PS97, PD00, Pla96, PBLB12, PFAHA03, Pra23a, Pra24, RL06, RCG^{+}99, RR94, Reb95, ROW^{+}97, Rei98b, RBK^{+}00, RC95, RP97, RC06, RLZ13]. Multimedia [RJ14, RMPK07, Rui14b, Rui15b, Rui16b, Rui17b, SKK^{+}12, Sad20, SA07a, Sch94, SK00, SB97, SJM06, Sha04a, SSR96, SD97, SAH^{+}99, SZ17, SC15, SAB^{+}20, SKWB22, Sla11a, Sla11b, Smo94, SWdR^{+}08, Sod11, SM18, SvOPS06, SHG^{+}07, Sti95, Str96, Str99b, Str00, Stü95, SZW^{+}19, SY05b, Sun05, SH08, TLJ^{+}19, Tat96, TWY^{+}22, TKBR^{+}07, Tho00, TCS^{+}15, TCS16, TH05, Tit03a, Tob95, TNJ05, Vai95, VKS99, Vet04, VWS95, VKvBG95, WX23, WY07, WJ07, WSR^{+}95, WZ10, Wil98, WH97, WSS^{+}12, WO97, XSS14, XW15, XCZH19, YHS11, Yan15, YHFC14, YS09, YKHI94, ZAKNY23, ZLZ06, ZLS14, Zen16, ZSM02, ZYNS16, ZXHS21, ZRW^{+}19, ZCWH15, ZKWH17a, ZKWH17b, Zim03, dOHS^{+}03, vONH04, ABH^{+}95, And98, BCL99, Bra97a, Bra97b, CLSS97, Chi95, Chi97, Dav99, DH97, Fri95, Gol94, Hof94, Jef99, Lug99]. multimedia [NM98, Pac99, Pur98, SBN97, Tay99, WGH94, Ano94a, Ano03a, Ano03f, Ano04g, Ano05a, Ano06a, Ano06d, Ano07a, Ano07f, Ano08d, Ano11f, Gol03c, Gro97a, Rui14c, Rui15a, JP13, JIW13, OK14, SS11, Ste95, Str00, Ano95a, Ano96a, Ano96b, Ano97, Ano98a, Ano99a, Ano00a, Che18b]. Multimedia-Enhanced [SY05b]. Multimedia-friendly [SAH^{+}99]. Multimedia-Hard [XSS14]. Multimedia-on-Demand [BP96]. Multimodal [ABPR^{+}19, APS^{+}18, BCD^{+}02, BG96, BH07, BCLH14, CVDL05, CCM^{+}04, Che21a, CWY^{+}22, DGT^{+}22, EYT^{+}06, FOW^{+}22, GLZW20, GPA20a, GPA20b, GB96, GWD^{+}22, HXXS18, HYLK21, KHA^{+}21, MPM^{+}17, MYBP^{+}22, MXH21, MBX14, MBM21, MW11, Ovi96, PZW^{+}16, RHP^{+}06, SKWB22, SRW24, TYW23, TGC^{+}06, VRG18, WCS^{+}21, WGY^{+}21, WZL^{+}21, ZXHS21]. Multimodality [dG04]. Multiparty [VJS^{+}23]. Multiple [Ano02g, BQLN^{+}18, CYL22, HKN04, HS08, KUM^{+}19, NAK^{+}22, QPZ^{+}21, SSG19, Tit03a, WB08, WJ99, WFI07, WFRZ21, WZL^{+}24, WN94, WLZL20, ZWH^{+}21]. Multiple-Description [WFI07]. Multiple-Source [BQLN^{+}18]. Multiple-Strategy [WZL^{+}24]. Multiplicative [MBM21]. Multipoint [Bea95, YWWW19]. Multiprocessor [HZSH23]. Multipurpose [Nac03b]. Multiscale [CZY20, LCM^{+}12, RLW^{+}21]. Multiscreen [HWL^{+}14]. Multisensor [SXHH16]. Multisensory [Bol15a, CR04, OGM^{+}17, ZCL^{+}04]. Multispectral [MLM^{+}07]. Multistream [WXL^{+}17]. Multitask [FGC^{+}14]. Multitone [GL10]. MultiTube [Bol07a]. Multiuse [MP17]. Multiuser [Hag96, ABH^{+}95]. Multiview [FWL22, HUU^{+}23, LTB23, LWZ^{+}18, OWZ^{+}20, WHM15, WYZ^{+}16, XLL^{+}22, ZSYW22]. Mumford [Bha00]. Museum [BBZB07, Wil98]. Music [BHLP16, BNZ05, Eff98, GCCGW17, HKN04, KSPW07, LH23, LT11, LGT17, LCL21, LCM09, LLJW15, Mad06, MPM^{+}17, MAHT98, PRC00, RSS09, Sch00, SvdM11, SWW11, Smo94, WZ08, WZBKB17]. Music-Related [RSS09]. Musical [BM98, CWH^{+}17, FNM98, Fer18, KOCG99, LCL21, Pai13, ZCL^{+}04]. Musicians [Ano02s, MB99]. MusiKalscope [FNM98]. Musings [Rei97a]. Must [Nac05]. MVIIDroid [WLZL20]. My [Foo06, Gol06c, Jai05a, NN08]. myComputer [Ano13i]. myCS [Ano17o, Ano17p, Ano18-30]. Mystic [WSR^{+}95]. Myths [Gol03e].

N [Str99a]. NAB [Tit06]. Nadeau [Wil96]. Name [SNK99]. Name-It [SNK99]. Naming [PTM11, SNK99]. Napster [Tai02]. Narrowband [New94]. NASA [Gri96]. Natick [Bha00]. Natura [PVAMM10]. Natural [ADV05, AYK08, Jai96a, KGU07, MSF16]. Naturalistic [LH23]. Navigated [BH07]. Navigating [SBS97]. Navigation [CCB^{+}98, HLB^{+}10, LG05, WLWW21]. Navigational [CH02, Jai97d]. Navigator [WLWW21, Kla96]. Near [CY13, NXKS13]. Near-Duplicate [CY13, NXKS13]. NEAT [RDP99]. Need [Ano20k, Ano21l, Bol09, TGB^{+}22]. Needed [Gro95c, Jai97d]. Needle [MM14]. Needs [Rei97d, Shn02]. Negative [HTH^{+}22]. Neglected [GH15]. Neighborhood [GRP21]. Nerve [MD00]. Nested [XZYY13]. Nested-SIFT [XZYY13]. Net [CP05, GB98, GS95, Gro95c, Gro96b, Gro98b, Hor96d, MB99, Gro96b, MD00]. NetCarta [Car96]. NetMedia [ZSM02]. Netscape [Kla96]. NetSci [Gou99e]. NetSci-Fi [Gou99e]. NetStudy [Gou99e]. NetView [ZCSSM98]. NetVLAD [CZY20]. Network [ASL^{+}24, AK22, Bed05, BDPN94, BGPQ21, CHSC22, CWY^{+}22, CYL22, DGT^{+}22, FLZ17, FOW^{+}22, FZXL24, GLZW20, HCLK23, Hof94, HZY^{+}23, JW17, LWN^{+}21, LBZL22, LSK^{+}22, NKZ^{+}22, PM16, Sha04a, SdOG98, Stü95, SMH^{+}17, VAAI24, WFI07, WPC^{+}20, WLL^{+}21, WGY^{+}21, WHW^{+}22, XZL21, YXH^{+}23, YW14, ZSYW22, ZKWH17b, Gro95a, MB00]. Network-Centric [SdOG98]. Networked [MZ97, MN07, Str96, WH97, WGH94]. Networking [Ano18-33, Ano18-34, Bro96, DRB09, GTH^{+}13, LZ16, WHTC24, WO97, ZKWH17b, Ban97]. Networking-Driven [WHTC24]. Networks [ALS22, Ano17r, BMR99, Fer18, FFW^{+}20, FWL22, GQ10, HNDP17, HLY10, Hub04, LWL^{+}20, MYBP^{+}22, Mou06, New94, PAS^{+}13, RL06, RLW^{+}21, Rat00, ROW^{+}97, Rui14b, SWW11, Str99a, WXL^{+}17, WX14, XGN06, XCZH19, ZCH^{+}19, CLSS97, Str99a]. Neural [ASL^{+}24, FFW^{+}20, FWL22, GLZW20, LSK^{+}22, MYBP^{+}22, PM16, Rui14b, STL18, VAAI24]. Neutral [ST16]. Newer [Rei94b]. Newlywed [Smi10b]. News [AD24, Ano01c, Ano01d, Ano01e, Bal98, Bal99a, Bal99b, Bal01, Del02, GB98, Gol98, Gri96, GS95, Gro95a, Gro95b, Gro95c, GO96, Gro96a, Gro96b, GBL97, GY97, Gro97c, GL97b, Gro98b, Meh99, MB99, MB00, MD00, PTM11, Rei99b, SNK99, Sch00, Smi11d, Wea00, WA16, Ben96]. Newspapers [Gro95a]. Next [AYK08, CV14, FCM^{+}22, GV07, Hub04, MZZ^{+}16, SO00, SH08, XCZH19, Vin98, CLSY12]. Next-Generation [AYK08, CV14, GV07, Hub04, MZZ^{+}16, Vin98]. NFV [Ano17r]. NFV-Based [Ano16c, Ano17r]. Niche [Pet96]. Night [MD00]. Nightmare [KRB19]. Nights [MN07]. Nixon [Kra00]. No [Gro95c, QPZ^{+}21, ZGL^{+}14]. No-Reference [QPZ^{+}21, ZGL^{+}14]. Nodes [SMH^{+}17]. Noise [CHSC22, GT11, SRA08]. Noise-Balanced [GT11]. Noise-Robust [CHSC22]. Noisy [HTH^{+}22]. Nomadic [HWV^{+}09]. Nominations [Ano16b, Ano16q, Ano17n, Ano19e, Ano19d, Ano20s, Ano21j, Ano21k]. Non [QJC^{+}18]. Non-uniform [QJC^{+}18]. Nondirect [ZJC^{+}20]. Nonlinear [MLHK17, MPG07]. Nonlocal [MZZ^{+}16]. Nonparametric [MSF16]. Nonrigid [CSZL20]. Nonuniform [QPZ^{+}21]. Nonverbal [AF18]. Normal [GRG21, LBZL22]. Normalization [CYL22]. Norwood [Str99a]. Notebook [Gou00]. Notes [Bal98, GB98, Gol98, Gro95a, GY97, GL97b, Meh99, MB00, MD00]. Notification [MQC^{+}10]. Novel [AVW08, BLBC22, CW09, GTQG16, HG23, LCG^{+}94, Rad05, WKY^{+}15]. NT [Car96]. Nuage [EHH09]. NUS [CLSY12]. NUS-Tsinghua [CLSY12].

Object [AMM21, CLD02, Col94, FGC^{+}14, GD14, GS01, GSOI10, KTY00, KD17, NvOH05, PS97, PFAHA03, RLJ^{+}23, SS19, TGB^{+}22, WS02, WHW^{+}22, XXZ^{+}23, ZLT^{+}12, ZYXW11, vONH04]. Object-Based [PFAHA03]. Object-Detection-Based [KD17]. Object-Oriented [GS01]. Objective [JJ14]. Objectives [TT11]. Objects [HS08, LLY^{+}18, LYGT19, RBF03]. Obscure [NvOH05, vONH04]. Obsessives [Foo05]. Obstructive [TWY^{+}22]. Oculus [BVT^{+}23]. ODP [GL97a]. ODP-RM [GL97a]. Off [Gol96, Pla96]. Off-the-Shelf [Gol96]. Offered [Hor96d]. Offers [Bol15a]. office [Chu95]. Olympics [Gro96a]. On-Demand [FHL04, HG09]. On-Device [GHD^{+}14]. On-Line [Gro95a, Rei94c, RWS^{+}94]. On-the-Shelf [Gol95]. On-Water [SM15]. Once [WZL^{+}24]. One [Ano17s, Ano18-31, Ano18-32, Gol08b, LTD05, LKLH13, MM03b, Ano21f]. One-to-Many [LKLH13]. One-to-One [LTD05]. Online [Bal99a, COCB14, CBG09, FGC^{+}14, Jam07, LJSZ14, MPM^{+}17, MB99, PLL^{+}11, PHS03, SSLS14, TC01, WX23, Wea00, Yan15, YCZ^{+}12]. Only [DP95, Jai03c, WZL^{+}24]. onto [GS95, Rei99a]. Ontologies [BBDS10, BDS^{+}09]. Ontology [FNHB05, NST^{+}06, RDD09]. Open [Ano19u, Ano19v, Ano19w, Ano20q, Ano20r, Ano23q, Ano24b, Bir94a, Bir94b, CO12, SDEFK09, SC16, Smi11d, TSDD08, VGH^{+}18, WSS02, CLSS97, GHR00, WZBKB17]. Opening [Col94, MD00]. Operating [BLA^{+}23, Ste95]. Operation [LSZM00]. Opinion [Smi12d]. Opportunities [Ano20k, Ano21l, Ano24i, KMB97, Pan08, PAS^{+}13, Tit07]. Optical [GLB^{+}21, TDZ^{+}12]. Optimal [GQ10, Pla00, QJC^{+}18, WCS^{+}21]. Optimization [HNDP17, JYXT11, KUM^{+}19, LJSZ14, NGA23, XNS^{+}13]. Optimized [HLW^{+}19]. Optimizing [WX23, XW15]. Options [Ano16v, Ano17q, Ano19u, Ano19v, Ano19w, Ano21l]. Oracle [iWD15, Gro95a]. Oracles [NC17]. Orchestra [SGVT04]. Orchestral [HB05, PGHK13]. Order [LFX^{+}18, LLH22]. Organ [OB00]. Organizational [LW00]. Organizations [LMCP18]. Oriented [DLW^{+}19, GS01, HH18, TLJ^{+}19]. Other [Gou98a]. Our [Ano19-29, Bla94, Bol20, ROAS00]. Outlook [VDL^{+}22]. Output [Gro96b]. over-rainbow-podcast-hHoriz-ks.indd [Ano23-45]. Overhauled [Gro96b]. Overlay [WFI07]. Overload [MDK10]. Overloading [Jai98d]. Overview [Ano02j, Fur94, GAB^{+}98, SZW^{+}19, HRvL96a]. overviews [Ano21f]. Own [MPL00]. Owners [SMF06]. Ownership [SMF06].

P [Kra00]. P2P [Del01b, DWX23, YJC07]. P2P-CDN [DWX23]. Page [BBMC02, GL03]. pages [Bur09b, Swi96]. Paintings [GA20a, GA20b, SCY^{+}16]. Palm [Ano15f]. panacea [Sus99]. Pancreas [KTB^{+}22]. Pandemics [Che21d]. Panel [HLW^{+}19]. Panorama [Che04]. Panoramas [Zhe03]. Paper [Ano17-27, Che18a, Rui17a]. Papers [Ano13d, Ano14d, Ano14a, Ano14b, Ano14h, Ano15e, Ano15c, Ano15d, Ano15o, Ano15p, Ano15r, Ano15q, Ano16c, Ano16e, Ano16d, Ano16g, Ano16f, Ano16h, Ano16i, Ano16j, Ano16p, Ano17e, Ano17f, Ano17c, Ano17d, Ano17g, Ano18e, Ano19h, Ano19g, Ano19f, Ano20f, Ano20g, Ano20h, Ano20d, Ano20e, Ano22t, Ano22u, Ano22v, Ano23-30, Ano24o, Ano24p, Rui15a, Ano13c, Ano14c, Ano15b]. Paradigm [CASC10, CYW23, ML18, PCMT16, WX14]. Paradigms [KAHHM96, MG96, Par98]. Parallel [CJ12, Lee98, SYJC06, SGK^{+}07]. Parallel-Horus [SGK^{+}07]. Parameter [HLW^{+}19, LZZ^{+}18]. Parameter-Optimized [HLW^{+}19]. Parameterization [LTD05]. Parameterized [LZDC19]. Parametric [HB05]. Paries [Gou99c]. Paris [Bra05]. Parkinson [APS^{+}18]. Parks [ABC^{+}08]. Parsing [LLZ22, TML^{+}14]. Part [Ano02k, Ano02j, HH18, LWH^{+}18, NL99a, NL99b, PEB09b, ZJC^{+}20, vONH04, Abo99a, Abo99b, BCL99, BCL00, Bir94a, Bir94b, Bul02, HRvL96a, HRvL96b, NvOH05, New95a, New95b, PEB09c, Sto06, SJ06, SD07b]. Part-Based [ZJC^{+}20]. Partial [LJZ^{+}13]. Partial-Duplicate [LJZ^{+}13]. Participation [Bol17, WZBKB17]. Partition [ZF24]. Partition-Based [ZF24]. Partitioning [Kas04]. Partners [Gro95a]. Partnets [Gri96]. Passing [Che22d]. Passthrough [BVT^{+}23]. Past [ABC^{+}08, ROAS00]. Patch [LNM^{+}24]. Patches [HTH^{+}22]. Path [AF18, CBWD11]. Patient [AS20]. Patients [APS^{+}18]. Patriots [Dav02]. Pattern [ZCC^{+}13]. Patterned [BBD^{+}00, Pla00]. Patterns [Bha00, BM98, LLD^{+}16]. Paul [Gol97c]. Paved [GBL97]. Payload [MG99]. PAZ [GRF^{+}16]. PC [Bro96, Car96, Gou99a, GO96, Jai02b]. PCs [Jai99c, Rei97c]. PDAs [Rei98d]. pDisVPL [LWH^{+}18]. Peak [CNYL21]. Peak-Based [CNYL21]. Pedestrian [LW22]. Peer [JTCC09, NG11, SWW11, Wea00]. Peer-to-Peer [JTCC09, NG11, SWW11, Wea00]. Pen [Kim96, ZMvD^{+}08]. Pen-Based [Kim96]. Pen-Centric [ZMvD^{+}08]. People [Jai96b, Jai07, LEGFD08, MTD^{+}11, PTM11, Wea00]. Perception [Eff05, LH23, MBM21, Mou06, TJBBFB15, YZSY21, ZYNS16]. Perception-Based [Mou06]. Perceptual [GQFZ23, PBM^{+}07, WX23, XW15, YTL^{+}24]. Performance [HH18, Ing13, IGB^{+}08, LGT17, MDK10, MM14, MN07, NL12, PGHK13, SA07b, SGK^{+}07, SK03, Str99a]. Performance-Oriented [HH18]. Performances [MPM^{+}17, WZBKB17]. Performing [Jam07]. perils [Sha95]. Person [EYT^{+}06, LZDC19, NN08, PCMT16, ZJC^{+}20]. Person-Centered [PCMT16]. Personal [AR05, BMO18, BEG^{+}18, EL06, GS06, Ros03]. Personality [GAL^{+}18]. Personalization [Ang03, LVL09]. Personalized [BM22, GRF^{+}16, KTP06, LTM03, NC17, PPP^{+}11, RR94, WKS^{+}03, ZSYW22]. Personalizing [TLS04]. Perspective [Aar04, Bol15a, BGPQ21, DL22, FLZ17, He04, Liu94, Ma96, MM03b, Tia06]. Perspectives [GO96, HLP^{+}14]. Pervasive [Ano22r, Ano22s, FBB^{+}05a, GZC09, SS00]. Peter [Bha00, San00a]. Peters [Bha00]. Petri [CP05]. Petrovsky [Tee98b]. PGAN [ZJC^{+}20]. Phenomenon [EH05]. Philip [PHS03]. Philips [CNYL21]. Phone [NN08, Tit06]. Phones [BBZB07, MB99, Wei08]. Phoning [GL97b]. Photo [Jai07, KS06, LTM03, PZKV11, VBL^{+}14]. Photographs [AP00, Hor96d]. Photography [Sto14]. PhotoPlethysmoGraphy [PMV^{+}22]. Photorealism [Gou99d]. Photos [CCW07, NN15, New10, Tes05]. Phrase [LHA06]. Phylogeny [DGR13]. Physical [HKR10, MTS^{+}06, WDZ^{+}17]. Physics [RSR^{+}01]. Piano [BVT^{+}23]. Pics [Hof94]. Pictorial [BCK^{+}05]. Pictorially [BDS^{+}09]. Picture [CFR^{+}10]. Pictures [Ano02r, JLD07, Ano94a]. Picturing [Fle06]. Pieces [Fer18]. Pioneer [MM03b].

Piracy [MD00, San00b]. Pirates [Dav02]. Pitch [CNYL21]. Pitch-Shift [CNYL21]. Pivot [MKYH03]. Pixel [LLL17a]. Place [WLN^{+}21]. Places [MP17]. Plan [Gro95a]. Planning [LLY^{+}18, TGC^{+}06, XPZ^{+}13]. Planogram [LLD^{+}16]. Plans [CAJ04]. Platform [BKCR13, CDR^{+}10, Cou99, CNN02, HLADCR^{+}21, Jam07, LMCP18, VGH^{+}18, ZDZ15]. Play [BGPQ21, Nac01, NJ06, Wea00]. Playback [KE96]. Players [SH08]. Playing [Ano02s]. Playout [CK96]. Playspaces [MP17]. Pleno [EFPS16]. Plenum [Str99b]. PML [RCG^{+}99]. Podcast [Ano22-43, Ano23-46, Ano23-45]. Point [GWL^{+}22, GRP21, HG23, KP24, LHD^{+}20]. Pointing [KZ06]. Poised [Bal99b, Gro98b]. Policy [Hat96, Swi96]. Political [KHA^{+}21]. Pollution [Jai94b]. Pool [DJBF07]. Pooling [LLL^{+}17b]. Pooling-Based [LLL^{+}17b]. Poor [Smi12c]. Popular [Mad06, SW09]. Portals [DYG^{+}00, Jai99b, MB00, Sch00]. Pose [CYL22]. Positioning [BH07]. Possibilities [EL00]. Post [MTXH19]. Post-Processing [MTXH19]. Posterity [BDD^{+}12]. Postgraduate [LXH^{+}22]. Postprocessing [ZMFB21]. Posture [JLD07]. Potatoes [Jai97b]. Power [Del01b, DSGG21, Swi96]. Poynton [Pet96]. pp [Ara99b, Bay97, Bha00, Dav98a, GSW97, Gou99c, Gou99d, Gou00, Hoc95, Hoc96, Kra00, Ma96, MS95, Pet96, San00a, SB97, Str96, Str99a, Str99b, Str00]. PP8K [GYL^{+}23]. Practical [CG95, WSR^{+}95]. Practice [Ano02s, HH18]. Practices [Tes05]. Precision [KWYK22, LH18, Sla11a]. Precision-Recall [Sla11a]. Prediction [FZH^{+}21, LZZ^{+}21, LZDC19, WCS^{+}21, WLN^{+}21]. Preferences [BC09]. Prefetching [CWW^{+}19, HCAMJ17]. Premo [HRvL96a, HRvL96b]. Prentice [Gol97c, Str96]. Prentice-Hall [Str96].

Prepare [Ano17t]. Preproduction

[GCM94]. Presence [SJM06]. Presence/Absence [SJM06]. Presentation [CCW07, HRvL96a, HRvL96b, Hol10, KMM^{+}10, PCKS00, SD97, Fri95]. Presentations [GSW97, HW98, RCG^{+}99, RC06, ZSM02]. Presents [GL97b]. Preservation [HKM^{+}10, LLP00]. Preserved [Ano20j, BMDG20]. Preserving [AK22, GI11, LWL^{+}17, LMS^{+}22, WH11, XXZ^{+}23]. Press [Bay97, Ma96, Swi96]. Press/Addison [Swi96]. Preventive [LH18]. Price [DTF17]. Price-Based [DTF17]. primacy [Pur97]. Primary [LYGT19]. Primer [Wil96]. Principles [New95b]. Print [Gro96b]. Printed [WZL^{+}24]. Printing [Ano14o, LLT10]. Prior [SY05b]. Priors [LLEK23]. Privacy [Ano22-43, Ano23-46, AK22, Bal99b, BGC^{+}15, Kan20, LJSP17, LWL^{+}17, LMS^{+}22, MB00, XXZ^{+}23, Ano14k, Ano18b, Ano24m, Ano24f]. Privacy-Preserving [AK22, LWL^{+}17, LMS^{+}22, XXZ^{+}23]. Pro [Car96]. Probabilistic [LWH^{+}18, WMS05]. Problem [GO96, Jai07]. Problems [Gro98b, Smi12b, XSS14]. proceedings [Reb95]. Process [GPLR95, HHLW20, RLZ13, Ude02]. Processes [ZRD23]. Processing [BBC00, CKN^{+}11, Che04, Dim07, DYT^{+}09, GYL^{+}23, GSW97, ISSZ08, KUM^{+}19, MTXH19, VWA^{+}07, WTL^{+}14, Wil95]. Produce [Zim03]. Producers [CASC10]. Producing [San00a]. Product [DLH01, GO96, SRW24]. Production [CEG^{+}13, LGT17, Sub06, Kam99]. Products [Ano01k, Ano01l, Ano01m, Ano02n, Ano02o, Ano02p, Ano02q, Ano03b, Ano03c, Ano03d, Ano03e, Ano04b, Ano04c, Ano04d, Ano05b, Ano05c, Ano06b, Ano06c, Ano07b, Ano07c, Ano07d, Ano07e, Ano08b, Ano08c, Ano09c, Ano09e, Ano09f, Ano09g, Ano10b, Ano10c, Ano10d, Ano10e, Ano11b, Ano11c, Ano11d, Ano11e, Ano12a, Ano12b, Ano12c, Ano13j, Ano13k, Ano13l, Ano14m, Ano14n, Fot06a, Fot06b, Fot08a, Fot08b, Fot12, Hor95a, Hor95b, Hor95c, Hor95d, Hor96d, Hor96a, Hor96b, Hor96c, Hor97d, Hor97a, Hor97b, Hor97c, Hor98a, Hor98b, Hor98c, Hor98d, Hor99, Shu99a, Shu99b, Shu99c, Shu00a, Shu00b, Shu00c, Shu00d, Shu01, Gou98a]. Professional [Bro96, Dav98a]. Profiles [XZL^{+}16]. Profiling [YCZ^{+}12]. Profit [Tit02]. Program [BP96, Gou99c, JJO^{+}04]. programmable [CLSS97]. Programmers [Ver98]. Programming [DV96, Gro95a, HMS17, JLD07, McK05, Gou99a]. Progressive [ML18, YWT^{+}07, YS09]. Project [Agu96, ABH^{+}95, AK99, Ban97, CLSS97, DJBF07, DH97, EMS96, Ens98, FWI98, FWI99, HP96, Jef99, KC00, KM95, Low95, NM98, Sha95, SSR96, SAH^{+}99, SBL00, Tab97, Vin98, BBG^{+}96, Bry00, KO94, PMR^{+}04]. Projected [AR05, BCK^{+}05]. Projection [LJSZ14, WKC^{+}13]. Projections [VWS95]. Projects [HH18, Ens98]. Prominent [GA20a, GA20b]. Promise [Bal01, Bed05]. promises [Sha95]. Promotes [GO96]. Propagation [LSZM00, LZH^{+}23, PTM11, ZG09]. Propagation-Based [LZH^{+}23]. Property [BHLP16, GS95]. Proportions [GA20a, GA20b]. Proposals [Ano19g, Ano20d, Ano20e]. Proposed [Bal98]. Prosodic [Kim96]. Prospects [GG23, LV94]. Protecting [Gro98b]. Protection [Bal98, Gro95c, LL14, Ran14]. Protocol [LL97, SJ04, FGR12]. Protocols [GMV01, Bra97a, Bra97b]. Prototypes [CWH^{+}17]. Prototyping [GRG21, TIS96]. Provides [Hor96d]. Providing [AAE23]. Provisioning [LVL09]. Proximity [SREH04]. Proxy [CWZ^{+}05]. Pseudo [WKC^{+}13]. Pseudolabels [LHH^{+}24].

Psychological [SSW20]. PTR [Gol97c]. Public [Ing13, MQC^{+}10, RL06]. Publications [Bry00]. Published [Ano20q, Ano20r, Ano24m]. Publishers [Ara99b, GSW97, Gou00, Hoc96, MS95, Str00]. Publishing [GSW97, Gou99e, MTW97, Str99b, Swi96]. Pursuit [LJSZ14]. Push [Bol07b, Gol98, GBL97]. Push-y [GBL97]. Pushing [Gro97c, SZW^{+}19]. Put [RLZ13]. Putting [Jai98c].

Q&A [AHN^{+}18]. QoE

[CWW^{+}19, HYC^{+}19, HMS17, TLJ^{+}19, YZSY21, ZDS19]. QoE-Aware [HMS17]. QoE-Oriented [TLJ^{+}19]. QOS [NS95, VKvBG95, BDL^{+}09, IMMP99, JN04, Sha04a]. Qualities [SCY^{+}16]. Quality [Ano04e, Ano17r, BCEH03, CH02, CDWT10, FGC^{+}14, HG15, HHLW20, JCM^{+}23, KAB^{+}23, MSF16, MFG03, OR02, PM16, QPZ^{+}21, SW09, SSR96, SSSS95, WX23, WHTC24, XW15, XWJ16, YGM^{+}16, ZGL^{+}14, ZYNS16]. Quality-Aware [ZGL^{+}14]. Quality-Constrained [HHLW20]. Quality-of-Service [SSR96]. Quantitative [LLL^{+}17b]. Quantization [GRP21]. Quantum [Ano20-31, Ano23-39, ZAKNY23]. Queries [TML^{+}14]. Query [DTG^{+}08, DYT^{+}09, SBKK^{+}10]. Querying [AP00, KGU07, LMS09, NC17, SJ00, WX14]. Quest [BVT^{+}23, GC96]. Question [HWL^{+}12, LMH^{+}10, LTB23]. Quintuplet [ZHC^{+}20].

R [Bay97, Wil96, ZKWH17a]. Radio

[GB98, YWWW19]. Rainbow [Ano22-43, Ano23-46, Ano23-45]. Ramakrishna [Ano16b]. Random [And08]. Range [CYW23, CW09, JM17]. Rank [FCL^{+}19]. Rank-Based [FCL^{+}19]. Ranking [LCSC11, SHH09]. RanKloud [CKN^{+}11]. Rapid [BB02]. Rate [AFKN95, CMW^{+}23, GQ10, LZZ^{+}18, MMSB15]. Ratings [HS17]. Ratio [FZH^{+}21]. Rau [Ano16b]. Ray [JCM^{+}23]. Raychaudhuri [Str99b]. Rays [KWG10]. Reaches [MB99]. reaching [AK99]. Reader [Ano04f, Ano05d, Sha04b, SD07a, Sub06]. Reading [Ara99a, HLW^{+}19, MZSH01, San00a]. Readings [Ara99b]. Ready [Bal98, MD00, Rei94c]. Real [CNI^{+}04, DSGG21, Foo05, FZXL24, GS95, GLLW24, HB05, HZK^{+}23, Jai95c, Jai96d, Jai07, JM17, KRN97, LSZM00, LLY^{+}11, MD00, MQC^{+}10, Nac10a, Pic16, PGHK13, RDLT06, SYN^{+}18, SWdR^{+}08, SSLS14, VJS^{+}23, WB08, WLZL12, WHTC24, XGN06, XZL21, XW15]. Real-Time [CNI^{+}04, FZXL24, HZK^{+}23, JM17, LSZM00, LLY^{+}11, MQC^{+}10, PGHK13, SYN^{+}18, SWdR^{+}08, SSLS14, VJS^{+}23, WB08, WLZL12, WHTC24, XGN06, XZL21, XW15]. Real-World [Pic16]. Realistic [Ike02, SJS07, VD00]. Reality [BVT^{+}23, BLA^{+}23, Car98, CCM^{+}04, Che21c, EFPS16, Gol06b, Gou98b, Gou98c, Gro95a, HFdW11, HK17, HOK00, KRN97, LRCY07, Lin17, MTB06, OB00, PCKS00, Rad05, SZS^{+}20a, SZS^{+}20b, TCM18, VJS^{+}23, YWWW19, Yu17, ZDS19]. Really [CRL^{+}03, Gol04a]. Rear [CKRW00]. Reassure [MB99]. Recall [Sla11a]. Receivers [WFRZ21, Dia99]. Recognition [Bal01, Bha00, BCLH14, CHSC22, Che18a, Che20, EYT^{+}06, FFW^{+}20, FOW^{+}22, GPA20a, GPA20b, GHD^{+}14, GWD^{+}22, HYLK21, MNKI07, MMS13, NKZ^{+}22, Pic16, RLW^{+}21, SLZ^{+}22, SSS^{+}19, TLJ^{+}19, VVS21, VKV^{+}12, WB08, WYZ^{+}16, XLT^{+}21, ZLT^{+}12, ZZLG21]. Recognized [LL14]. Recognizing [JLD07, Rui16a]. Recommendation [PLL^{+}11, TYW23, VBL^{+}14, WHM15]. Recommended [Car98]. Recompression [ZF24]. Reconstructing [KWG10, LD06, Smi13d]. Reconstruction [CCS16, LLZ22, RBK^{+}00, XLT^{+}21, XL22]. Record [AS20]. Recording [Emo09, LKBE08]. Records [PBLB12]. Recovery [JYC08, QJC^{+}18, Zhe00]. Recreational [Gou00]. Rectification [LZD^{+}22]. Recurring [LLD^{+}16]. Red [Gol03d, Smi10c]. Redefining [KSW12, San00b]. Reducing [CGM97]. Reduction [GTQG16, LTB23, WHW^{+}22]. Reference [QPZ^{+}21, ZGL^{+}14, ZF24]. Refinement [JWW^{+}14, ZSYW22]. Reflecting [KMJ17]. Reflective [HN07]. Refocusing [HPH05]. Regenerative [MG99]. Region [CYL22, JCM^{+}23, LWN^{+}21, YWT^{+}24]. Region-of-Interest [YWT^{+}24]. Region-of-Interest-Guided [JCM^{+}23]. Regions [Yeh16]. Regularization [GL20, OWZ^{+}20]. Regularizing [LWW^{+}22]. Rehabilitative [PCMT16]. Rehearsal [KOCG99]. Reidentification [GZP^{+}24, LZDC19, WPC^{+}20, ZJC^{+}20, ZWH^{+}21]. Reinforcement [Che22b, HRH^{+}21, ZZZ^{+}21]. Reinvention [GR17]. Related [GLZW20, RSS09]. Relation [ZSYW22]. Relational [HZY^{+}23]. Relations [ASC96]. Relationship [HXXS18, WCS^{+}21, ZLS14]. Relationships [DGR13, HS08]. Relay [LKLH13]. Relevance [Han15]. Relevant [AD24]. Reliable [WHTC24]. Relic [KCKC00]. Remarks [Ano04f, Ano05d, Sha04b, SD07a, Sub06]. Remember [Gol98, Nac05, NN15]. Remote [Bol17, BGPQ21, PMV^{+}22, RDP99, RGH^{+}21, UTZH12, WZL12, WO06, Kam99]. Removal [SRA08]. Rendering [LCL23, LLEK23, MNK^{+}07]. Replacement [MTD^{+}11]. Report [Cas96, Kah96]. Reporting [TNJ05, MBPM98]. Reports [Agu96, ABH^{+}95, AK99, Ban97, CLSS97, DH97, EMS96, Ens98, FWI98, FWI99, HP96, IMK^{+}16, Jef99, KC00, KM95, NM98, Sha95, SSR96, SAH^{+}99, SBL00, Tab97, Vin98, Low95]. Repositories [LLY^{+}11]. Repository [XNS^{+}13]. Representation [BNZ05, CP05, EFPS16, GWL^{+}22, HZY^{+}23, LWW^{+}22, ZZZ^{+}21, ZSZ^{+}20, ZZS^{+}22]. Representations [MPM^{+}17]. Representing [CTCC99, FNHB05]. Repurposing [AYK06, OSS04, PMR^{+}04, Sin04]. Requirements [AN97, SW04]. Rerank [YH13]. Reranking [HKC07, TT11]. Research [Che21d, Che22c, DHTF98, FS01, GOVI15, Han17, HPH05, Juh11, SZ17, SC15, SAB^{+}20, TCS16, UTZH12]. reservation [Bra97b]. Residence [XCZY14]. Residual [FFW^{+}20, MTXH19]. Residual-Based [MTXH19]. Resistance [WLZL12]. Resists [LL14]. Resolution [PMV^{+}22, Ser05, Vai95, Sch00]. Resource [BMR99, Bra97b, KUM^{+}19, KNC01, SW04, Ano21f]. Resources [Ano01j, Ano16u, Bal01, DN00, Emo09, MD00, Met00, Nac00, Nac01, PMGCCPRM11, SGS01, Tat96, Tho00]. Response [Che21d]. Responsive [AB97, AB03, KG05]. Restoration [BBC00, KCKC00, LLP00]. Restructuring [RT01]. Resume [Ano19u, Ano19v, Ano19w, Ano21l]. Retargeting [MB16]. Retinex [LNM^{+}24]. Retinex-Guided [LNM^{+}24]. Retransmissions [NG11]. Retrieval [AP00, BBDS10, BCGU10, BCLH14, CY13, CZY20, CDP99, DGLW07, DZS^{+}02, Dje02, DTG^{+}08, EBG98, GD14, GS06, GV07, GR17, Jai07, JKP12, JP13, LJJT12, LLP00, LJZ^{+}13, LSKP09, LTM03, LCSC11, NPG13, OLIG04, PZW^{+}16, QSL^{+}19, RJ14, Sak94, Sch14, SZ94, SWdR^{+}08, Tia06, TDZ^{+}12, VCBM19, WXY^{+}22, WBKW96, WSS^{+}12, WKC^{+}13, XZYY13, YHS11, YHFC14, YKHI94, YGN18, ZG09, ZH02, ZSZ^{+}20, ZZS^{+}22]. Retrievals [WN94]. retrieve [NM98]. Retrieved [LYL^{+}15]. Reusable [EFS01].

Reuse [Bol07b, SERL01]. Reveal [Bol07b]. Reversal [CHSC22]. Reversible [ZLL^{+}23]. Review [Ara99a, Gou99a, Gou99b, Tee98a]. Reviewers [Ano03f, Ano04g, Ano06d, Ano07f, Ano08d, Ano11f, Ano13b, Ano15a, Ano16a, Ano17a]. Reviewing [GSW97, Gou99b, Gou00, Tee98a]. Reviews [Ano98b, Ano01g, Ano01h, Ano02h, Ara99b, Bay97, Bha00, Bro96, Car96, Cha06, Dav98a, DP95, Gol96, Gol97a, Gol97b, Gol97c, GSW97, GV07, Gou98a, Gou98b, Gou98c, Gou99c, Gou99a, Gou99d, Gou99e, Gou00, Hat96, Hoc96, KZ06, Kol96, Kra00, Luo04, Ma96, Meh98, MS95, Par98, Pet96, San00a, SB97, She99, Str96, Str99a, Str99b, Str00, Swi96, Tee98b, Tee99, Tia06, Ver98, WSR^{+}95, Wil95, Wil96, Gol95, Hoc95]. Revisited [Gro95c]. Revisiting [SD07a]. Reviving [CYW23]. Revolution [Aut11, Hal01, Jai97a]. ReWeb [RT01]. Rhythm [CCPC11, LMCP18]. Rhythmic [MJSE15]. Ribcage [KWG10]. Rich [DAC05, Smi12c]. Richard [Ano17u, Ano17u]. Ride [GO96]. Right [Han17]. Rights [CDR^{+}10, KLKW09, MLJ06, Wan04]. Rights-Information [KLKW09]. Rise [Yan15]. Risky [MYBP^{+}22]. Rival [Gro96b]. Rivals [Gro95c]. River [Gou99c, Gou99d, Tee99]. RMI [PKN^{+}04]. Road [Juh11, LWL^{+}20, WKS^{+}03]. Roadmap [Oom11]. Robotic [Ano01b, SK03]. Robotics [Miy17]. Robust [CHSC22, CNYL21, FWL22, LJSZ14, QGZF23, WNY09]. Rochester [PEB09a]. Rock [Ano14p, Ano14o, Ano15t]. Rockland [Gou99c, Gou99d]. Role [DYG^{+}00, Smo94]. Rolling [RR05]. ROM [Gou99c, San00a, Gol97c]. Romance [ZK18]. ROMs [Neu00]. Room [CKRW00, Tit03a, Tit07]. Rotation [GWL^{+}22]. Rotation-Invariant [GWL^{+}22]. Roundup [SS06, Wea00]. Route [Zhe03].

Rowing [SM15]. Rule [BBDS10, WKY^{+}15]. Rumor [CWY^{+}22]. Rx [Bal99b].

S [Kra00, Mal00]. S5 [BN24]. Sadun [Tee99]. Safe [GL20]. Safety [GGX^{+}17, Kah96, MQC^{+}10]. Saliency [HG15, JK19, LJZ^{+}13]. Saliency-Guided [HG15, LJZ^{+}13]. Saliency-Weighted [JK19]. Salient [RLJ^{+}23, SS19]. Sampling [SRA08]. Sampling-Based [SRA08]. San [Ara99b, Gou00, Str00]. Sarcasm [WZL^{+}21]. Satellite [BMR99, Bra98, IMMP99, MG99, PD00, Pri99, Lug99]. Satellite-based [Bra98]. Satisfied [FZH^{+}21]. Sayood [Hoc96]. Scalable [BBCN12, CKN^{+}11, GTH^{+}13, JYXT11, JTCC09, LKLH13, MPG07, Qui03, THJG13, WKC^{+}13, YA14]. Scale [BBZB07, CY13, CYL22, DLH01, DGR13, DLW^{+}19, FGD01, HGHW16, HCH12, JYS^{+}14, KG05, LL14, LLY^{+}11, LY19, MJ02, NST^{+}06, NPG13, NXKS13, RL06, SWW11, Sla11b, YHS11, YJC07, ZHC^{+}20, ZCSSM98, ZXHS21]. Scale-Aware [HGHW16]. Scale-Space [MJ02]. Scales [VCBM19]. SCAN [BD03]. SCAN-Based [BD03]. Scanner [MLM^{+}07]. Scenario [DB05]. Scenarios [AMM21, BBCN12, HFK95, KMM^{+}10, SJS07, VKS99]. Scene [Ano20j, BMDG20, BN24, JK19, LW22, LYL^{+}15, LY19, YMON08]. Scene-Adaptive [LW22]. Scenes [DKRS04, Ike02, KRN97, LY19]. Schedule [KE96]. Scheduler [ZAKNY23]. Scheme [BMR99, HZSH23, HT06]. Schemes [SYJC06]. Scholarship [Ano17u, Ano17u]. School [RSR^{+}01]. Schoolchildren [Kim96]. Science [AVW08, Ano21a, Ano21b, Ano22b, Ano22c, Ano22d, Ano23e, Ano23g, Ano23f, BKCR13, Che18c, FHK08b, Gol03e, Gou15]. Sciences [Dim07]. Scientific [AKMHP09, GBH15]. Scientists [WSR^{+}95]. Scope [Gol03c]. Score [PGHK13]. Scores [Kri03a]. Scrambling [DMA^{+}12, LLL17a]. Screen [HLW^{+}19, Jai03c, LLS11, SKSH08, CMC^{+}20]. Screen-Aware [HLW^{+}19]. Screens [GLN^{+}08]. Scripting [AB97, RFM95, SSSS95]. SDN [Ano16c]. SDN/NFV [Ano16c]. SDN/NFV-Based [Ano16c]. SeamlessDesign [KTY00]. Search [AHN^{+}18, Ano15g, Ano15h, Ano16k, CG14, CLSY12, DGLW07, DTT^{+}13, Foo05, GCGR11, Han15, HKC07, JYXT11, LTS17, LEGFD08, LHX11, ML18, MTS^{+}06, NXKS13, RGH^{+}21, Rui14a, SHH09, SS08, SMK^{+}10, WDG95, WBKW96, XNS^{+}13]. Searching [Jai00c, Jai00d, SC97]. Section [FHK08a]. Secure [ASLB20, DWX23]. Secured [CRD^{+}11]. Securing [PMZ13]. Security [AHN^{+}18, AS20, Ano02i, Ano14k, Ano18b, Ano22-43, Ano23-46, Bal01, BEUW12, BLBC22, Gol03b, Gro98b, LJSP17, Pra24, Sun05, ZLZ06]. See [HNM03]. Seek [Gro95a]. Seeker [FSC08]. Seeking [Dav94, Dav95]. Seeks [Ano23c, Gri96]. Seg [WCQ^{+}22]. Segment [CWW^{+}19, CWZ^{+}05]. Segment-Based [CWW^{+}19, CWZ^{+}05]. Segmentation [DS15, LHH^{+}24, PMV^{+}22, WCQ^{+}22]. Seine [Bra05]. Selecting [Yeh16]. Selection [PRC00, PKL15]. Selective [LWL^{+}17]. Self [HLADCR^{+}21, JJ14, LL14, VTJL^{+}08, ZSZ^{+}22, ZZS^{+}24]. Self-Esteem [HLADCR^{+}21]. Self-Motion [VTJL^{+}08]. Self-Recognized [LL14]. Self-Supervised [ZSZ^{+}22, ZZS^{+}24]. Sellers [YCZ^{+}12]. Semantic [ABCB02, CKAI09, DYT^{+}09, DV03, LM02, LWZ^{+}18, NSS01, PMGCCPRM11, TDZ^{+}12, WLN^{+}21, WTW23, XWJ16, YHFC14, ZLL^{+}20, ZSZ^{+}20, ZWH^{+}21, ZZS^{+}22, SvOPS06]. Semantics [CTCC99, CDP99, DRB09, FGS09, Jai94c, KNC01, MXH21, PPDH94, SJ09, WH11]. Semantics-Based [MXH21]. Semantics-Preserving [WH11]. Semi [GTQG16, SZ00]. Semi-Automated [SZ00]. Semi-Supervised [GTQG16]. Seminar [HK17]. Semiotic [CL10]. Semisupervised [LW22, LHH^{+}24]. Sense [Ano02g, HWL^{+}14, RDLT06, RLZ13]. Sensed [CZCJ16]. Senses [ABC^{+}08]. SenseWeb [KNLZ07]. Sensing [AC18, APS^{+}18, Ano16i, Ano16j, BN24, JM17, KNLZ07, LMS^{+}22, SOS13, SSW20, XL22, YWT^{+}24]. Sensitive [GHBC94, HT06]. Sensor [Zha12, ZYJ^{+}13]. Sensorial [SSG19]. Sensors [Gro98b, Jai98c]. Sensory [Jai98d]. Sentence [WXY^{+}22]. Sentiment [LSM^{+}21, WGY^{+}21]. Sentiment-Aware [LSM^{+}21]. Sequence [HZK^{+}23, LHD^{+}20, LSM^{+}21, WHM15]. Sequences [KTP06]. Series [KMJ17, Liu16]. Serious [AC18]. Server [BDPN94, CKN^{+}11, DT94, Hor96d, Tee98b, Vai95, SAH^{+}99]. Servers [BCD^{+}02, CZ96, CGM97, Lee98, RC95, SYJC06]. Serves [Gro95a]. Service [Ano20-30, BCEH03, CLD02, Gro95a, HG09, HS99, LVL09, MB99, PFMMF23, PD00, SSR96, XCZH19, YS09, Jef99]. Service-Level [MB99]. Services [Ano19t, BM22, BDL^{+}09, DWX23, DAC05, FGL^{+}01, Gro95a, HRCK08, KGS^{+}02, KJKS01, MG12, Met00, Mil98, MG96, NG11, PFMMF23, RDP99, SKK^{+}12, SW04]. Session [FGR12]. Set [Meh99, Woj98, Kog98, GB98]. Set-top [Meh99]. Sets [CCB^{+}98]. SGI [Gro95a]. SGML [New95a, New95b]. SGML/HyTime [New95a, New95b]. Shanghai [Ran08]. Shape [BCLH14, CGW^{+}23, NKZ^{+}22]. Shaping [Bol20]. Share [Bol07b, MSR^{+}15]. Shared [DS15, KNLZ07]. Sharing [BDPN94, HLADCR^{+}21, QJC^{+}18]. Shastra [AB94]. Shelf [Ano98b, Gol95, Gol96, Gol97a, Gol97b, Gol97c, GSW97, Hat96, Meh98, She99]. Shift [CNYL21, SBN97]. Shirakawa [HOK00]. Shirakawa-go [HOK00].

Shneiderman [Ara99b]. Shonan [BHS13]. Short [HPH05, LZH^{+}23]. Short-Long-Term [LZH^{+}23]. Shot [LWN^{+}21, LBZL22, SSS^{+}19]. Show&Tell [SZ00]. Showdown [Sch14]. Shrinking [GBL97]. Sickness [DOD22]. Side [Kri03a]. SIFT [XZYY13]. Sighted [MOK^{+}15]. Sigmoidal [HMS17]. Sign [AAA^{+}09]. Signal [Dim07, VWA^{+}07]. Signals [Ano16h, Ano16i, Ano16j, BMO18, CSJ^{+}18, GH15, JM17, RLZ13]. Signature [MB99]. signposts [Dav95]. Signs [Bra05]. SignTutor [AAA^{+}09]. Silicon [GBL97, Liu16]. Silver [Tit03b]. Similarity [EBG98, QSL^{+}19]. Simpler [Rei94b]. Simulation [Fei09, LRCY07, SZH^{+}08, WO06, XPZ^{+}13, dLA01]. Simulation-Based [dLA01]. Single [Bry00, LWN^{+}21]. Site [CT01, Gro95a, MFG03, Rei96d, Rei96c, Rei96a]. Sites [Bal99b, Gou98b, Gou98c, RT01]. Situations [MYBP^{+}22]. Size [BN24]. Sketch [BN24]. Sketch-to-Image [BN24]. Skills [AdAFM95, Ano20k, Ano21l]. Skin [PMV^{+}22]. Skipping [ALS22]. Sleep [TWY^{+}22]. Slideshow [CCW07]. Small [Ano17r, MB99, WB16]. SmallBlue [LEGFD08]. Smart [ASLB20, Ano01e, Ano20j, BMDG20, CCS16, DLW^{+}19, GL97b, Kra00, Sto14, XCZH19, ZRD23]. Smarter [Ano18f, Dav96a]. Smartphone [HLADCR^{+}21]. Smell [CSC^{+}20, CBC^{+}20]. SMIL [Bul02, CP05]. Smoliar [GSW97]. Soccer [SH10]. Social [Ano18-33, Ano18-34, CZY20, Chu95, CZCJ16, DRB09, GA15, GH15, HKR10, HWL^{+}14, HXXS18, Juh11, KRB19, LLP^{+}16, Nac10c, PVK^{+}16, PCS^{+}15, Rei98a, RLZ13, SC16, VJS^{+}23, WDZ^{+}17, WFZ^{+}20, WLN^{+}21, XCZY14, YMON08, YW14, ZDZ15]. Social-Sensed [CZCJ16]. Socially [MSR^{+}15, SH10]. Societal [Bol15b]. Society [Ano17u, Ano20-30, Ano21x, Ano21y, Ano21z, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ma96, Ano13g, Ano14j, Ano15s, Ano16t, Ano16s, Ano16q, Ano16r, Ano16u, Ano17m, Ano17n, Ano18q, Ano18n, Ano18p, Ano18o, Ano18r, Ano18s, Ano19r, Ano19p, Ano19q, Ano19s, Ano20v, Ano20w, Ano20-27, Ano20x, Ano20-28, Ano20y, Ano20-29, Ano20z, Ano21q, Ano21r, Ano21v, Ano21s, Ano21t, Ano21w, Ano21u, Ano22t, Ano22u, Ano22-28, Ano22-29, Ano22-32, Ano22-31, Ano22-30, Ano22-33, Ano22-34, Ano23-32, Ano23-33, Ano23-35, Ano23-36, Ano23-30, Ano23-38, Ano23-34, Ano23-31, Ano23-37, Ano24r, Ano24o, Ano24t, Ano24q, Ano24s, Ano24p, Ano24n, Ano20q, Ano20r, Ano23q, Ano24b]. Soft [FS04]. Soft-Decision [FS04]. Software [Ano14q, CO12, DP95, GTZ^{+}00, GSW97, Gou98b, Gou98c, Hor97d, JGN^{+}13, Mak05, Mur15, WHTC24, ZKWH17b, Wil95]. Software-Based [JGN^{+}13]. Software-Defined [WHTC24, ZKWH17b]. Solos [Fer18]. Solution [Bal98, GO96, Hor96d, JGN^{+}13, Kri99]. Solutions [Gol95, Gol96, Jai00c, ZLZ06]. Solve [Gro98b, Jai07]. Somatics [Gol07b]. Someone [MB99]. Song [LSKP09]. Sonic [HWV^{+}09, PT15, RR05]. Sonification [Ano13d, DHH15, Eff05, FBB05b, HH05, HB05, SM15, TJBBFB15, WMS05, ZSN^{+}05]. Sons [Gou99d, Pet96, Tee98b, Ver98, Wil96]. Sorting [Gou99e]. Sought [Gro95c, Rei99c]. sought-in [Rei99c]. Sound [And08, CCS16, Gol03d, HWV^{+}09, HLT23, LLJW15, dG04]. Sound-Effect [LLJW15]. Sounding [RBF03]. Soundscape [WEV03]. Source [BQLN^{+}18, CO12, VAAI24]. Sourcebook [Tee98b, Wil96]. Sources [CGW^{+}23]. Sourcing [Bry00]. South [Kri03b]. Sovereign [BM22]. SP [LHA06]. SP-MIDI [LHA06]. Space [HCLK23, MJ02, MKYH03, NYNI99, Rou99, SBS97, TSM21, XPZ^{+}13, dG04]. Space-Time [HCLK23]. Spaces [CTCC99, Fri95, Sin99]. Sparse [HG23, WYZ^{+}16, WKC^{+}13]. Sparsity [ZLT^{+}12]. Sparsity- [ZLT^{+}12]. Spatial [Kem95, LCCL21, NM98, WS02, WHW^{+}22, XZL21, ZWH^{+}21]. Spatial-Angular [LCCL21]. Spatially [HGHW16]. Spatio [MBX14]. Spatio-Temporal [MBX14]. SPC [GSW97]. Speak [KGS^{+}02]. Speaker [WA16, WLWW21]. Speakers [WB08]. Special [Ano13d, BLM^{+}22, Cas96, FHK08a, Kah96, WZ10, ZXHS21]. Specific [HCAMJ17, WZ08]. Specification [BCEH03, HRvL96b, JN04]. Specified [Sha04a]. Specifying [VKS99]. Spectrum [Bak07]. Specular [LCL23]. Speech [CHSC22, HTL96, Lar03, MAHT98, PPL08, Qua13, ST16, WB08]. Speech-Aware [HTL96]. Speed [CWW^{+}19]. Spencer [Meh04]. Spies [Ano01r]. Split [FFW^{+}20, LL97]. Spoken [Gol03a, Ovi96]. Sports [ABCB02, CHWB11, HFdW11, SYN^{+}18, TCP04, WY07, WKY^{+}15]. Spotting [GLB^{+}21]. Spread [Gro98b]. Springer [Hat96, Hoc95, Kra00]. Springer-Verlag [Hat96, Hoc95, Kra00]. Sprint [Gro95a]. Stack [MM14]. Stage [FS01, GRG21, SLW98]. Stakes [GO96]. Standard [Ano02k, BCL00, CYW23, DTT^{+}13, DLB^{+}19, DAC05, FDM^{+}15, GCGR11, Gro95c, HRvL96a, HRvL96b, LPSZ08, LPL^{+}13, Pla96, Sod11, BCL99, GL97a, AAE23, Kal06, KL07, VDL^{+}22]. Standard-Dynamic-Range [CYW23]. Standardization [PMV^{+}10, Ran14, Chi97]. Standardized [KZ06, RDBRD15]. Standardizing [GC96, Gug03]. Standards [And98, Ano01f, Ano01n, Bal01, BCL99, BCL00, BN01, Bra97a, Bra97b, Bry00, Bul01, Car98, Chi95, Chi97, CJ12, DTG^{+}08, DLW^{+}19, EMM^{+}98, FCA08, GL97a, GC96, GS95, Gro95c, GO96, Gro96b, HRvL96a, HRvL96b, Hos98, Kal06, KL07, KLKW09, LZH^{+}07, Liu94, LRDTP13, Met00, MMRM08, NST^{+}06, New95a, New95b, Par99, Pla96, RAE16, RFM95, SA07a, Sha04a, SS06, SS08, SWdR^{+}08, SvOPS06, SGS01, Sus99, Tay99, TKBR^{+}07, Tho00]. Standards-Based [LRDTP13]. StarGAN [CHCH22]. Stars [Ano14p, Ano14o, Ano15t]. Start [TYW23]. Startup [LTS17]. Startups [GRG21, Jai16, Yan15]. State [Gol94, ZZZ^{+}21]. Static [Nac07]. Station [MNKI07]. Statistical [GSOI10]. Status [VDL^{+}22]. Stay [Ano14r, Ano15u]. Stealing [LHA06]. Steer [Ano02u]. Steganography [CW09, FGD01, WFRZ21]. Stephen [GSW97]. Stepping [Ano01e]. Stepwise [ZWL^{+}20]. Stereo [FCL^{+}19, JWW^{+}14, XZL21, XLL^{+}22]. Stereoscopic [FZH^{+}21]. Steven [MS95, Swi96]. Still [Dav95, Gol04d, HLNS17]. Stock [Rei97b]. stop [Ano21f]. Storage [BR96, Gri96, Kam99]. Stories [JS13]. Story [Bro02, Dav94, GS94, KS12, MSR^{+}15]. Storytelling [BVA^{+}15, CCM^{+}04, Dav96a, MSR^{+}15, PCS^{+}15]. strata [KC00]. strata-based [KC00]. Strategies [Ano15y, Ano16-27, Kas04, TT11]. Strategy [WZL^{+}24, WKY^{+}15]. Stream [CK96, FGSS08, SYJC06]. Streaming [Ano02t, BQLN^{+}18, BGPQ21, CWZ^{+}05, DTF17, DWX23, EP06, FRM^{+}07, GWC17, Gro97c, HLY10, JTCC09, LKLH13, MPG07, Pra23a, Sod11, Tit02, WFI07, WHTC24, Wea00, XDL04, YZSY21, YJC07, ZSM02]. Streams [DWZ^{+}08, GKL^{+}98, KUM^{+}19]. Streets [GBL97]. Stress [Pic16]. Strong [HLNS17]. Structural [LWL^{+}20]. Structuralizing [Jai01b]. Structure [CSZL20, Mad06, XLL^{+}22]. Structured [EMG97, LZDC19, TATS94]. Structures [CH02, Hav00, WLPN11, New95a]. Stuart [Ara99b]. Student [Ano17u, Ano18-35, LXH^{+}22]. Studies [Gou98a, KMF^{+}01]. Studio [BBG^{+}96, Gro95a]. Studios [Gib98, GAB^{+}98]. Study [BVT^{+}23, DOD22, PEB09a, PEB09b]. Studying [LH23, PEB09c]. Stumbles [GS95]. Sturgeon [Foo05]. Style [CHCH22, Gol97c, LWL^{+}20, LNM^{+}24, LWH16]. Styles [LCL21, YCZ^{+}12]. Stylization [Ano20j, BMDG20]. Subject [Ano13a, Gou99e]. submissions [Ano94b]. Subscribe [Ano21b]. Subscription [Ano13e]. Subspace [LCSC11, LJSZ14, OWZ^{+}20]. Subway [MNKI07]. Success [Bal99b]. Successful [Ano01h, Gou98a, Str96]. Successive [JLD07]. Suggestions [NN08]. Summarization [TCP04, XWJ16]. Summary [PXL10]. Sun [Gro96b]. SunSoft [Bay97]. Super [GBL97]. Superimposing [BCK^{+}05]. Superscape [Gou98b, Gou98c]. Supervised [GTQG16, MXH21, TYW23, YXH^{+}23, ZSZ^{+}22, ZZS^{+}24]. Support [BBZB07, CK96, CV14, Sha04a, SSR96, Kim95]. Supported [Dus00, FRM^{+}07, YGM^{+}16]. Supporting [GWC17, KO94, Rad05, Vin98]. Supports [Agu96]. Suppositional [XLL^{+}22]. Surface [TJBBFB15]. Surgeries [LRCY07]. Surgery [BH07, SZH^{+}08, WO06]. Surgical [WSI^{+}05]. Surround [HWV^{+}09, QH05]. Surround-Sound [HWV^{+}09]. Surroundings [HKR10]. Surveillance [BDD^{+}12, KD17, NGA23, SJS07, ZZL^{+}22]. Survey [BL01, CIL15, JN04, SOS13, VKvBG95, ZZLG21]. Sustainable [Agu96, Ano17b, Ano18d, Ano23r, Ano24g, Ano24h]. Swap [LNM^{+}24]. Symbolic [BNZ05]. Symmetrically [FZH^{+}21]. Symmetry [LLZ22]. Symmetry-Aware [LLZ22]. Symphonic [GCCGW17]. Symphony [WZBKB17]. Synchronization [ASC96, CL10, CL06, HW98, MdRCS15, SYJC06, WO06]. Synchronized [Hos98]. Synchronizing [TN98]. Syncing [DS15]. Synopsis [NGA23]. Synthesis [BN24, CNI^{+}04, CGW^{+}23, WLL^{+}21, WCS^{+}21, WMS05, RM99]. Synthetic [WEV03]. System [ATS^{+}13, AdAFM95, ASLB20, AAA^{+}09, BCGU10, BB03, BLA^{+}23, BR96, DSUG03, EP06, Fle06, GGX^{+}17, Hag96, HFdW11, He04, HLN98, HHLW20, HZSH23, HZK^{+}23, HS08, KTP06, KG97, KCRO10, KHY04, LFX^{+}18, LSKP09, LLH22, NG11, NL12, PKN^{+}04, QSL^{+}19, SD97, SHH09, SH10, SZ00, SLW98, TWY^{+}22, TC01, TDZ^{+}12, VJS^{+}23, XCZH19, YHFC14, ZCC^{+}13, Kim95, MDK97, Ste95]. Systems [Ano01h, Ano01n, Ano16c, Ano16h, Ano16i, Ano16j, Bir94a, Bir94b, BP96, BM98, BGPQ21, CVDL05, CDD^{+}00, CSJ^{+}18, Del00, Dim07, DTG^{+}08, Eff98, EFS01, Fur94, Gec97, Gro94, Gro96a, GZC09, IMMP99, Ike02, Jai94c, Kem95, KD17, Kra96, KG05, LJ05, MSW18, NTB97, PK06, Pri99, RGH^{+}21, SWdR^{+}08, WJ99, Wei08, ZCL^{+}04, DH97, Low95, GSW97, SB97, Ano22o, Ano22n].

Table [Ano13m, Ano14s, Ano14t, Ano14u, Ano14v, Ano15v, Ano15w, Ano15x, Ano16w, Ano16x, Ano17v, Ano17w, Ano17x, Ano17y, Ano18-36, Ano18-37, Ano18-38, Ano18-39, Ano19-30, Ano19-31, Ano19-32, Ano19-33, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano23-47, Ano23-48, Ano23-50, Ano23-49, Ano24w, Ano24x]. Tables [Hor96d]. Tackle [Jai16]. Tactile [El 20, KCRO10, Sad20, SSEKE19]. Tag [NN08, NC17, PLL^{+}11]. Tag-Cloud [PLL^{+}11]. Tagged [PZKV11, VBL^{+}14, ZYS^{+}10]. Tagging [GT08, HS08, LL08, LSM^{+}21, NC17, TYW23, YNC09]. Tail [HG09]. Taiwan [GBL97]. Take [Ano14w, Ano15y, Ano16z, Ano16y, Ano16-27]. Takes [Bol15b, LTS17, WSR^{+}95]. Taking [Pra23a, Rei97b]. Tale [CSC^{+}20, CBC^{+}20]. Talking [Rei97c, Tab97, XGN06].

Tampering [QJC^{+}18]. Tapping [GLN^{+}08, TJBBFB15]. Target [WZL12]. Targeted [Ded95]. Task [CCLS03]. Tasks [BH07, XZL^{+}16]. Tattoo [LJJT12]. Taxonomy [HM95, JN04, MZ97]. Teaching [AdAFM95, BGC^{+}15, Cur02, Gol97c, HH18, McK05]. Team [LKBE08, Pra23b]. Tech [Ano18w, Ano18x, Ano19x, Bal99b, Bal01, Bol08, Gri96, Mal00, Meh99, MB00, MD00, Tit06, Tit07]. TechIgnite [Ano17z]. Technical [LDZS18, MG12, PMV^{+}10, Pet96, Ano21f]. Technique [ASLB20, LL14]. Techniques [AG00, BYZZ13, BL01, LLY14, NXKS13, YKW01]. Techno [Bre99]. technological [Chi95]. Technologies [AC18, AVW08, Ano02g, CRD^{+}11, DG08, DLW^{+}19, El 18, GCGR11, Goo03, Juh11, LGT17, San00a, SK00, Shn02, MBPM98]. Technology [Ano01d, Ano02m, Ano16g, Ano16f, Ano17f, Ano17c, Ano17d, Ben96, Bro05, CV16, DNM^{+}10, Gro96a, GBL97, GL97b, Hat96, KGG96, Kel16, LTS17, MPL00, McG06, OB00, WdJ09, Woj98, ZRW^{+}19, PEB09a, Str96]. Telecollaboration [SdOG98, dOHS^{+}03]. Telecom [GBL97]. Telecommunicant [VD00]. Telecommunication [Bea95, EMS96]. Telecommunications [AB12, Gro96b, PPDH94]. Teleconferencing [ZCC^{+}13]. Teleimmersive [AHR^{+}13]. Telematics [Ano02u]. Telemedicine [Gro95b, PD00]. Telementoring [SZH^{+}08]. Telephony [MB99, Nac03b, SJ04]. Telepresence [Bol17, Jai95c, Moe97]. Telerehabilitation [JNK06]. Television [AB03, COCB14, Cre05, FCA08, LZH^{+}07, ZOR^{+}09]. Tell [Gro00c, KS12]. Template [Ran14]. Temple [ROB00]. Temporal [ASC96, HFK95, MBX14, SRW24]. Tensor [NPG13]. Term [FS01, LZH^{+}23, SSW20]. Terrestrial [LZH^{+}07, Pri99]. Terrorism [IGR^{+}20]. Test [Gro95a]. Testbed [DSGG21]. Tested [Gro96a]. Testing [BB02, RHP^{+}06]. Texas [JP13]. Text [Ano20j, BMDG20, GL20, Mal00, TML^{+}14, VWS95, WLL^{+}21]. Textbooks [Gol08a]. Textual [MLHK17]. Textural [LWH16]. Texture [FWL22, Par12, YTL^{+}24]. Texture-Aware [FWL22]. Thanks [Ano03f, Ano04g, Ano06d, Ano07f, Ano08d, Ano11f]. Theater [FS01, PKN^{+}04]. Theatre [WKC^{+}15]. Their [Gol03a, Gol98, HR00]. Them [RLZ13]. Thematic [ZYXW11]. Theme [MBX14]. Theory [Ano17c, Ano17d, SZS^{+}20a, SZS^{+}20b, ZRW^{+}19, Hoc95]. Therapy [HKN04]. Thermal [ZSZ^{+}22, ZLL^{+}23, ZZS^{+}24]. These [Rei98d]. Thieves [Ano01r]. Things [Smi10a, Dav95, CMC^{+}20, ML18, ZAKNY23]. Think [Gro00c, Ara99a, Ara99b]. thinking [Dav99]. Third [BCL00, HWD10, Kel16, LLS11, BCL99]. Those [DNM^{+}10]. Three [Bha00, Hav00, KCKC00, LLY^{+}18, GSW97]. three-attendee [GSW97]. Three-Dimensional [Bha00, Hav00, KCKC00, LLY^{+}18]. Threefold [VKV^{+}12]. Thrive [SC16]. Throughput [AGR07]. Thule [LD06]. Tiled [ZWL^{+}20]. Tiling [CCW07]. Time [Bul04, CNI^{+}04, CHSC22, FZXL24, Gro98a, HCLK23, HB05, HZK^{+}23, JM17, LSZM00, LFX^{+}18, LLY^{+}11, LLH22, MdRCS15, MQC^{+}10, OO04, PGHK13, SYN^{+}18, SBS97, SWdR^{+}08, SSLS14, VJS^{+}23, WB08, WLZL12, WHTC24, XGN06, XZL21, XW15, Dim99]. Time-Delay [LFX^{+}18, LLH22]. Time-Reversal [CHSC22]. Timed [CP05]. Timeshared [Kri99]. Timmerer [ZKWH17a]. Titel [Bro96]. TNG [GHR00]. Today [Ano19u, Ano19v, Ano19w, MW11, Ano21b, Ano21l]. Tomorrow [MW11]. Tool [AO15, DKRS04, Hor96d, Rui14b, GSW97]. Toolkit [CWH^{+}17, Gou99d]. Tools [AG00, Ano02j, Dav96a, Gol02c, GS95, NAV04, PPDH94, San00a, SS01, Zim03]. ToothPic [VCBM19]. Top [MB99, Meh99]. Topic [BN01, CWY^{+}22, TYW23, XZL^{+}16]. Topic-Based [XZL^{+}16]. Topics [JK19, Ano21f]. Torch [Che22d]. Torque [Bak07]. Touch [GLN^{+}08, McG04, RDLT06]. Touches [MB99]. Touching [GLN^{+}08]. Tourism [WGY^{+}21]. Tracing [DGR13, FS04, WXL^{+}17]. Tracker [HCLK23]. Tracking [CHWB11, FGC^{+}14, LJSZ14, MJ02, VD00, WZL12, ZSZ^{+}22, ZLL^{+}23, ZZS^{+}24]. Trademark [EBG98]. Traders [SMF06]. Trading [XXZ^{+}23]. Traffic [EP06, HMS17, Juh11, Kah96, Meh99, SXHH16, YHX15]. Trails [Gol06a]. Train [Ano02y, CWW^{+}19]. Training [DLH01, May96, MKK^{+}15, SM15, WSI^{+}05, ZGL^{+}14, Hof94]. Trampoline [PT15]. Transactions [Ano17b, Ano18c, Ano18d, Ano19h, Ano20f, Ano20g, Ano20h, Ano23r, Ano24m, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, GS95, Gro95c]. Transcendent [McG06]. Transcoding [LHA06, NSS01, STH08]. Transcription [WZ08]. Transfer [CYL22, HYC^{+}19, JYXT11, LWL^{+}20, LNM^{+}24, LWH16]. Transferring [NMK^{+}06, XL22]. Transformed [Jai02a]. Transformer [GZP^{+}24, GWD^{+}22]. Transformers [SRW24]. Transforming [DLH01]. Transitioning [GL97b]. Translation [CHCH22, HTH^{+}22, TC01, ZWL^{+}20]. Translational [LLZ22]. Transmission [GQ10, GKL^{+}98, SYJC06, WHTC24, YWWW19, ZAKNY23]. Transmitting [TL97]. Transport [XCZH19, Bra97b]. Trend [Meh99]. Trends [Add00, GCK12, MB99]. Trials [Gro95a]. Triangle [LTD05]. Triangular [LY19]. Triggers [Miy17]. True [Jai05b]. Trumpet [Gol03d]. Truncation [Ma96]. Tsinghua [CLSY12]. tu [Rei96b]. TUI [Gol07b]. Turbulence [MN07]. Turmoil [Gro95c].

Tutorial [Lee98, Pan95]. Tutoring [AAA^{+}09, CRL^{+}03, LLP^{+}16, TC01, WZ08]. Tutors [AYK08]. TV [DWZ^{+}08, GCM94, Gro96b, GL97b, HG09, HS17, HRCK08, HWL^{+}14, Jai99c, Jai02b, MD00, Mil98, Tit07, WA16]. TV-Net [MD00]. TVX [COCB14]. Tweets [GLZW20]. Twins [El 18]. Two [AVW08, Bha00, LWN^{+}21, LLH22]. Two- [Bha00]. Two-Shot [LWN^{+}21]. Type [KTB^{+}22]. Tyranny [Hal94].

Ubiquitous [Hak15, KMM^{+}10, KRB19, LLP^{+}16, MST^{+}06, MLJ06, Pan02, TCS16, WZ10]. Ubiquity [HWD10]. UHD [GYL^{+}23, JGN^{+}13]. Ultra [CMC^{+}20, YA14, YHX15]. Ultra-High [CMC^{+}20, YHX15]. Ultra-High-Definition [YA14]. UltraViolet [KSW12]. UMTS [Pri99]. Uncertainty [LHH^{+}24]. Uncertainty-Guided [LHH^{+}24]. uncharted [Bur09b]. Uncluttering [Rei98c]. Undeaf [Gol03a]. Underestimated [EH05]. Underpinned [LSL24]. Understanding [AA05, Ano16i, Ano16j, GCCGW17, JK19, LYL^{+}15, MS95, RT01, Rui16b, SB97, Str96, Sun05, TML^{+}14, Zen16]. Unencumbered [Que96]. Unified [LMCP18, Qua13, SBB^{+}13]. uniform [QJC^{+}18]. Unifying [DTG^{+}08, ZH02]. Unique [TH01]. Universal [SLZ^{+}22, TKBR^{+}07, Vet04]. University [CHH98, Gol94]. Unlikely [Gro95a, Rei99b]. unlimited [GSW97, Ano17s, Ano18-31, Ano18-32]. Unmanned [DSGG21]. Unpaired [HTH^{+}22]. Unscented [SRA08]. Unsupervised [GQFZ23, WA16, XLL^{+}22]. Unveiled [Rui17a]. Up-Conversion [MMSB15]. Upcoming [Ano95b, Ano95c, Ano96c, Ano96d, Ano98c, Ano98d, Ano99b, Ano99c, Ano99d, Ano99e, Ano00b, Ano00c, Ano01o, Ano01p, Ano01q, Ano02v, Ano02w, Ano02x, Ano03g, Ano03h, Ano04h, Ano04i, Ano05e, Ano05f, Ano06e, Ano06f, Ano06g, Ano06h, Ano07g, Ano07h, Ano07i, Ano08e, Ano08f, Ano08g, Ano08h, Ano09h, Ano09i, Ano09j, Ano09k, Ano10f, Ano10g, Ano10h, Ano10i, Ano11g, Ano11h, Ano11i]. Update [Gol98, Sti95, Tat96]. Upgrade [GL97b]. Upload [Ano19u, Ano19v, Ano19w, Ano21l]. Uploading [IGB^{+}08]. Upsampling [HLW16, XZL21]. UpStage [Jam07]. Urban [EP06, LWL^{+}20, MTH^{+}20, SXHH16]. Urge [GB98]. Usability [BB02, CM01, RHP^{+}06]. Usage [BL01]. Use [CAJ04, Dav02, HLADCR^{+}21, SKWB22, And98]. Usefulness [Han15]. User [CLSY12, FZH^{+}21, Kim96, LH23, MS95, Ovi96, PK06, PT15, RGH^{+}21, Sch14, SS00, SY05b, TIS96, WLN^{+}21, WSS^{+}12, XWJ16, YWT^{+}07]. User-Centered [Ovi96]. User-Centric [RGH^{+}21, Sch14]. User-Controlled [SY05b]. User-Generated [CLSY12, XWJ16]. Users [GLN^{+}08, NC17, Rei94a, Tai02, XCZY14, DB97]. Ushering [Gol02a]. Using [ALS22, Ano02y, AK22, Ara99b, BBDS10, BC09, BDL^{+}09, CWW^{+}19, CNI^{+}04, CYL22, CCS16, CHCH22, DMA^{+}12, DRB09, DWN01, DSGG21, GL03, GT11, HKM^{+}10, HMS17, Hoc95, HTH^{+}22, LD06, LLY^{+}18, LWL^{+}20, LT11, LLJW15, LCM^{+}12, LHA06, May96, MMS13, MTD^{+}11, NPG13, New95a, New95b, PMZ13, Par12, PXL10, PEB09b, PEB09c, PGHK13, RDBRD15, SK00, SC16, SS19, SY05b, TLS04, VTJL^{+}08, VWS95, VBL^{+}14, WB08, Wei08, Wil98, WN94, XGN06, YCZ^{+}12, YWT^{+}24, YW14, ZAKNY23, ZYJ^{+}13, Ara99a, Bra98, KC00, MBPM98, SS01]. Utility [DTF17]. Utility-Aware [DTF17]. Utopia [Rei99c].

V [Ma96]. vain [Rei99c]. Validation [DYT^{+}09]. Valley [Liu16]. Value [Ano02i, RDD09, ZRD23]. Variable [iWD15]. Various [Sha04a]. VBR [CZ96]. VC [KL07]. VC-1 [KL07]. vCache [GWC17]. VEC [LW00]. Vector [MKYH03, Qui03, SS19]. Vehicle [EYT^{+}06, GZP^{+}24, GRF^{+}16, WPC^{+}20, ZWH^{+}21]. Vehicles [DSGG21]. Vehicular [ALS22, WX14]. Vendors [GB98]. Verification [GMV01]. VERL [FNHB05]. Verlag [Hat96, Hoc95, Kra00]. Version [Gou98b, Gou98c, GSW97]. versus [Meh99]. Vert [EHH09]. VEs [Hag96]. Via [GLZW20, HYC^{+}19, LWN^{+}21, LSM^{+}21, BN24, CZY20, CH02, FTCS99, LJZ^{+}13, LMS^{+}22, Mil98, NAK^{+}22, TWY^{+}22, WCQ^{+}22, WHW^{+}22, XZL^{+}16]. Viability [MG12]. Video [AGR^{+}21, AYK06, ATS^{+}13, AA05, Ano01j, Ano17r, AFKN95, BDD^{+}12, BBDS10, BCGU10, BDS^{+}09, BCEH03, BD03, BC09, CY13, CBWD11, CWW^{+}19, COCB14, CZ96, CIL15, CK96, CMW^{+}23, CJ12, COH00, CMC^{+}20, Dav03, DRB09, DZS^{+}02, DWX23, DT94, DSUG03, DLW^{+}19, DLB^{+}19, EMG97, EP06, FGSS08, FDM^{+}15, FCM^{+}22, FNHB05, GYL^{+}23, GSW97, GQ10, GKL^{+}98, Gro97b, Gro97c, HFdW11, HYC^{+}19, HMS17, HOY99, Hor96d, HZSH23, HLY10, HUU^{+}23, HYLK21, ISSZ08, IGB^{+}08, Jay03, Kal06, KL07, KC00, KAHHM96, KD17, KGU07, LKGR94, Lee98, LZZ^{+}21, LZH^{+}23, LL97, LCSC11, LHD^{+}20, LV94, LLZ03, LLY^{+}11, LMS^{+}22, hMLDM98, MZZ^{+}16, MJ02, MAHT98, MTG97, MKYH03, NG11, NGA23, PAS^{+}13, PMV^{+}10, PLL^{+}11, PP05, PMR^{+}04, PXL10, PPDH94, QPZ^{+}21, QZ14, RM94, RSR^{+}01, RGH^{+}21, San00a, SBS97, Sch14, SGK^{+}07, SREH04, STH08, STL18, SHH09]. Video [SW04, SJ04, SH10, SZ94, SH08, SMH^{+}17, Tay99, THJG13, TCP04, TATS94, TLS04, TML^{+}14, VDL^{+}22, Vai95, VWA^{+}07, VAAI24, VRG18, WLZL12, WHM15, WHTC24, Wea00, WDG95, WKY^{+}15, WNZ10, WKC^{+}13, YMON08, YZSY21, YWT^{+}24, YA14, YHX15, ZMFB21, ZF24, ZCH^{+}19, Bra98, Dim99, Kam99, Kri99, Tob95, Wil98, FWI99, Pet96]. Video-Based [hMLDM98, RSR^{+}01]. Video-Card [CMC^{+}20]. Video-Content [DZS^{+}02]. Video-on-Demand [DWX23, LL97, LV94, Kri99]. Video-Stream [FGSS08]. Videoconferences [Emo09]. Videoconferencing [GTZ^{+}00, Gro96b]. Videodisk [Gro95c]. VideOlympics [SWdR^{+}08]. Videophone [SS01, XGN06]. Videos [ABCB02, CYW23, CYL22, HLADCR^{+}21, KAB^{+}23, LMH^{+}10, LYGT19, PTM11, SNK99, WLPN11, XWJ16, ZYXW11]. View [Bol06, CKRW00, GD14, MTG97, WLZL20]. View-Based [GD14]. Viewing [Hor96d, KAB^{+}23]. Viewpoint [HWZ^{+}23, WHM15]. Viewport [YZSY21, ZCC^{+}13]. Viewport-Adaptive [YZSY21]. Views [Ano01r, Bol06, Bol07a, Bol07b, Bol09, Chu95, Coo01, Dav95, Dav96b, Dav96a, DB97, Dav97, Dav98b, DHTF98, Dav99, DSB96, Dim99, DYG^{+}00, Fri95, HN07, HT08, Hor96d, Jai06, Ker00, MTB06, Nau98, Pan02, Pur97, Pur98, Ran08, Rat00, RM99, SKSH08, SBN97, SS00, Sti95, SMF06, Tat96, Tit03a, ZSYW22]. Villages [HOK00]. Viper [AB03]. Virtual [Abo99a, Add00, Ano02s, BBC00, BBG^{+}96, CTCC99, Che21c, DV96, FBB^{+}05a, Gib98, GAB^{+}98, God08, GB98, Gou98b, Gou98c, GS95, Gro95a, Hay98, HK17, HNDP17, HOK00, Hir97, HKR10, HWZ^{+}23, Jai96d, KRN97, KCKC00, KSPW07, Kog98, LKBE08, LG05, MZ97, MTB06, MPL00, MTG97, MKK^{+}15, NYNI99, OKN^{+}99, OB00, PCKS00, ROAS00, ROB00, RDLT06, SGVT04, Smi12d, SO00, SLW98, TCM18, VJS^{+}23, Woj98, XGN06, XPZ^{+}13, YWWW19, Yu17, Zhe00, dG04, Abo99b, Cam97, Chu95, Car98, DF99]. Virtual-Reality [PCKS00, TCM18]. Virtualization [FLZ17, JW17, ZKWH17b].

Virtualized [AG00, KRN97, LLS11]. Virus [AK99]. Visible [HT06, TL97, YXH^{+}23]. Vision [And08, Ano01i, Ano17g, Ara99b, Fal07, Foo06, Gol06b, Gol06c, Gol07a, Gol07b, Gol08b, Gol08a, GZP^{+}24, GAL^{+}18, Jai16, Jai97b, Jai97c, Jai97d, Jai97e, Jai97a, Jai98a, Jai98b, Jai98c, Jai99c, Jai99a, Jai99b, Jai00a, Jai00b, Jai00c, Jai00d, Jai01a, Jai01b, Jai01c, Jai02a, Jai07, MPST18, Sak94, Sto06, SJ06, SD07b, WLWW21, Ara99a, RM99]. Vision-and-Language [WLWW21]. Visions [Ano01r, Bol06, Bol07a, Bol07b, Bol09, BHS13, Chu95, Coo01, CD14, Dav95, Dav96b, Dav96a, DB97, Dav97, Dav98b, DHTF98, Dav99, DSB96, Dim99, DYG^{+}00, Fri95, HN07, HT08, Jai06, Jai98d, Ker00, MTB06, Nau98, Pan02, Pur97, Pur98, Ran08, Rat00, RM99, SKSH08, SBN97, SS00, Sti95, SMF06, Tat96]. Visitor [ADV05]. Visitors [PVAMM10]. VistaMail [HLN98]. Visual [AF18, Ano16-28, AAE23, BVA^{+}15, BC09, Cha96, CPCT03, CCPC11, CG14, CDP99, DV96, EFPS16, GPA20a, GPA20b, GCGR11, GHD^{+}14, HLB^{+}10, HKC07, HLP^{+}14, KUM^{+}19, KM95, KMB97, LJZ^{+}13, LTS17, LWH^{+}18, LTB23, LJSZ14, LHX11, MLHK17, MB16, MNK^{+}07, NPG13, PXL10, SJS07, ST16, Smi14, SSW20, Tia06, TT11, XNS^{+}13, XPZ^{+}13, ZHC^{+}20, ZCSSM98, ZZL^{+}22, ZG09, ZH02, Low95, Jai15]. Visual-Textual [MLHK17]. Visualization [Ara99b, Fei09, Hal01, MPM^{+}17, Ara99a]. Visualize [SBKK^{+}10]. Visualizing [FSC08]. Visually [MNK^{+}07, MTD^{+}11, RHP^{+}06, SC97]. Vocabulary [JYXT11]. VOD [BR96, SA07b]. Voice [Bal01, Gro95c, Meh99, MB99, Ros03, SW09, Sch00]. Voices [GL97b]. VoiceXML [Lar03]. VoIP [SW09]. Vol [Ano05a, Ano06a, Ano07a]. Volume [Ano95a, Ano96a, Ano97, Ano98a, Ano99a, Ano00a, Ano03a]. Volunteer [Ano20-30]. Volunteered [New10]. voyage [Bur09b]. VR [AP00, DB00, DOD22, Gri96, MTS^{+}06, SSEKE19]. VR2Gather [VJS^{+}23]. VRML [Wil96, Car98, EL00, Wil96]. vs [Kla96]. vu [Rei96b]. Vulnerable [GL97b].

W [GSW97]. W3C [Lar03]. Wakeman [Str00]. Walking [MJSE15, MMS13]. Wall [CMC^{+}20]. Wallets [Bal99b]. Want [GL97b, Jai05a]. Wanted [NL99a, NL99b]. WarpClothingOut [ZWL^{+}20]. Was [Gol98]. Watch [Ano15y, Ano16-27, Ano22o, Gro96a]. Water [SM15]. Watermark [GMV01, KHY04, QJC^{+}18, TH01]. Watermark-Fingerprint [KHY04]. Watermarking [ASLB20, AS20, DWN01, GB98, GO96, HT06, Kun01, LLY^{+}18, LJ05, LZD^{+}22, PBM^{+}07, QGZF23, WNY09, WLZL12, WXL^{+}17]. Watermarks [GL10]. Wave [GO96, HWD10]. Wavelet [ASL^{+}24, AS20, GSOI10, HHLW20, QZ14]. Wavelet-Based [HHLW20]. Wavelet-Domain [QZ14]. Way [MZZ^{+}16, Rei94b, Tho00]. Wayang [GI11]. Weakly [ZYS^{+}10]. Wearable [MOK^{+}15]. Weave [Jai15]. Weaves [Gro95a]. Web [Hos98, Ano01f, Bal98, Bal01, BCD^{+}02, BL01, BB02, BN01, Bol07a, BBMC02, BDL^{+}09, CY13, CH02, CL06, CNN02, Cur02, DH01, DYG^{+}00, DN00, EFS01, GM01b, GM01a, GS01, GCP01, Gou98b, Gou98c, Gou99e, GS95, Gro95a, GO96, Gro96a, GL97b, Gro98b, Has02, Hol10, Hor96d, HB04, Jai96c, Jai98c, Jai99a, Jai15, Ker00, KGS^{+}02, KJKS01, KNC01, LMH^{+}10, LCM09, LH97, May96, MM03b, MB00, MD00, MMC01, Met00, MFG03, MTW97, MMRM08, Nac00, Nac01, NvOH05, NSS01, NPG13, NXKS13, OR02, PMV^{+}10, RSS09, Rei96d, Rei96c, Rei96a, Rei97b, Rei97d, Rei98c, Rei99a, RT01, RT06, Rou99, SERL01, Sla11b, SC97, SvOPS06, SBB^{+}13, SGS01, Tat96, Ude02, WJ99, WLPN11, YMON08, YH13, dLA01, vONH04]. Web-Based [CH02, CL06, CNN02, Cur02, EFS01, KGS^{+}02, LCM09, Rou99, SGS01, WJ99]. Web-Scale [NPG13, NXKS13, Sla11b]. Web3D [LY19]. Web Insight [Rei98b, Rei98a, Rei98c, Rei98d, Rei99a, Rei99c, Rei99b]. Web Sight [Rei96d, Rei96c, Rei96a, Rei96b, Rei97a, Rei97c, Rei97b, Rei97d]. Webcast [GOVI15]. Webcasting [GSW97]. Webcasts [WY07]. WebDAV [Sus99]. WebMap [Car96]. Webmaster [Gou98b, Gou98c]. WebQEM [OR02]. Website [Ano19-29]. Webvertising [GY97]. Week [Ano23-39, Ano20-31]. Weighted [JK19, LCSC11]. Weightless [Bur09a]. Welcome [Gro99b, Nac00, Pra23b]. Wesley [San00a, Swi96]. Whalebone [LD06]. Where [Ada03, Bol07a, Del01a, Foo06, New10, WSS^{+}12, Ano16u]. Wherever [Ano16z, Ano14w, Ano16y]. Whiteboard [Che00]. Whither [Gro96b]. Who [Rei97d, SSS^{+}19]. Whose [Dav97, GS94]. Wide [GS95, Hos98, Jai95d]. Widely [Gro98b]. Wider [Gro95a]. Wiley [Gou99d, Pet96, Tee98b, Ver98, Wil96]. Will [Aut11, Tit02]. Windows [Car96, DSB96, Jai03b]. Wireless [And98, Ano01n, FHL04, GQ10, KD17, Sch00, Sha04a, Wea00, XGN06, Ker00]. Without [HHLW20, DSB96]. Woes [Sch00]. Wolff [Gou99e]. Women [Ano20s]. Word [Gol03a, WDZ^{+}17, XZL^{+}16]. Word-Learning [XZL^{+}16]. Words [WH11]. Work [Ano01j, ABC^{+}08, Ben96, Bol17, Bra98, BBMC02, Cam97, CDD^{+}00, Che00, CM01, CRL^{+}03, Del01b, DF99, Dia99, Dus00, Emo09, EP06, FGL^{+}01, HTL96, HRCK08, Kam99, KGG96, Kim95, Kog98, Kri99, KGU07, LD06, LEGFD08, LH97, MBPM98, MPL00, May96, Meh04, MS95, MDK97, MTS^{+}06, MZSH01, Reb95, RP97, RC06, RLZ13, SHG^{+}07, SH08, Tob95, Vai95, WY07, Wil98]. workable [Kri99]. Workers [Ano02y]. Workflow [VKV^{+}12]. Working [Rui16c, TIS96]. Works [PHS03]. Workshop [BEG^{+}18, Gou99d, Gou99b, Gou99d, KMJ17]. World [Ano15y, Ano16-28, Ano17-27, Ano18f, Ano19t, Gol03a, GLLW24, Hir97, Jai95d, Ker00, Kri03b, Pic16, Ser05, Ano16-27, GS95, Hos98]. World-Wide [Jai95d]. Worlds [DV96, KRN97, MPL00]. Would [Kol96, HR00]. Would-Be [Kol96]. Writing [MM03a, ZYJ^{+}13]. Wrong [Sla11a]. WSICC [KMJ17].

X [Bay97, Gol97c, Kra00, Pet96, BM22, JCM^{+}23, KWG10]. X-Ray [JCM^{+}23]. X-Rays [KWG10]. X.400 [CG95]. xDSL [MB99]. XML [Bal98, BN01, MB99]. XSEAD [BKCR13]. xt [Mal00, RAE16].

Yawn [Gro95a]. Years [Che20]. Yield [Tit02]. York [Gou99d, Str99b]. You-Only-Look-Once [WZL^{+}24]. Yourself [Smi11a]. YouTube [CIL15]. Yuille [Bha00]. Yuval [Hoc95].

Zero [DWX23, LBZL22, LZD^{+}22]. Zero-Shot [LBZL22]. Zero-Watermarking [LZD^{+}22]. Zhang [GSW97]. Zones [GRF^{+}16]. ZoneTag [NN08].

References

Altunbasak:2005:UFV

[AA05] Yucel Altunbasak and Hasan Ates. Understanding the future of video coding. IEEE MultiMedia, 12(1):88, 87, January/March 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                                        http://csdl.computer.

org/comp/mags/mu/2005/
01/u1088.pdf.

Aran:2009:SIS

[AAA^{+}09] Oya Aran, Ismail Ari, Lale Akarun, Bülent Sankur, Alexandre Benoit, Alice Caplier, Pavel Campr, Ana Huerta Carrillo, and François-Xavier Fanard. SignTutor: An interactive system for sign language tutoring. IEEE Multi-Media, 16(1):81–93, January/March 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ascenso:2023:JAS

[AAE23] João Ascenso, Elena Alshina, and Touradj Ebrahimi. The JPEG AI Standard: Providing efficient human and machine visual data consumption. IEEE Multi-Media, 30(1):100–111, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Athanasiadis:2020:DADa

[AAHA20a] C. Athanasiadis, M. Amestoy, E. Hortal, and S. Asteriadis. e^{3}3Learning: a dataset for affect-driven adaptation of computerbased learning. IEEE MultiMedia, 27(1):49–60, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Athanasiadis:2020:DADb

[AAHA20b] Christos Athanasiadis, Matteo Amestoy, Enrique Hortal, and Stylianos Asteriadis. e^{3}3Learning: a dataset for affect-driven adaptation of computerbased learning. IEEE MultiMedia, 27(1):49–60, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Aarts:2004:AIM

[Aar04] Emile Aarts. Ambient intelligence: a multimedia perspective. IEEE Multi-Media, 11(1):12–19, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
01/u1012abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2004/01/u1012.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/01/
u1012.pdf.

Anupam:1994:SMC

[AB94] Vinod Anupam and Chandrajit L. Bajaj. Shastra: Multimedia collaborative design environment. IEEE MultiMedia, 1(2):39– 49, Summer 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Agamanolis:1997:MSR

[AB97] Stefan Agamanolis and V. Michael Bove, Jr. Multilevel scripting for responsive multimedia. IEEE MultiMedia, 4(4):40–50, October/December 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1997/pdf/
u4040.pdf;   http://www.
computer.org/multimedia/
mu1997/u4040abs.htm.

Agamanolis:2003:VFR

[AB03] Stefan Agamanolis and V. Michael Bove, Jr. Viper: a framework for responsive television. IEEE Multi-Media, 10(3):88–98, July/ September 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
03/u3088abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/03/u3088.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/
u3088.pdf.

Abbasi:2012:IEE

[AB12] Ameer Abbasi and Uthman Baroudi. Immersive environment: An emerging future of telecommunications. IEEE Multi-Media, 19(1):80, 86, January/March 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ardito:2008:MWE

[ABC^{+}08] Carmelo Ardito, Paolo Buono, Maria F. Costabile, Rosa Lanzilotti, Thomas Pederson, and Antonio Piccinno. Multimedia at work: Experiencing the past through the senses: An M-learning game at archaeological parks. IEEE MultiMedia, 15(4):76–81, October/December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Assfalg:2002:SAS

[ABCB02] Jürgen Assfalg, Marco Bertini, Carlo Colombo, and Alberto Del Bimbo. Semantic annotation of sports videos. IEEE MultiMedia, 9(2):52–60, April/June 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
02/u2052abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/02/u2052.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/02/
u2052.pdf.

Anderson:1995:PRB

[ABH^{+}95] David B. Anderson, John W.

Barrus, John H. Howard, Charles Rich, Chia Shen, and Richard C. Waters. Project reports: Building multiuser interactive multimedia environments at MERL. IEEE MultiMedia, 2(4):77–82, Winter 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u40077.
pdf.

Abouaf:1999:AMBa

[Abo99a] Jeffrey Abouaf. Artful media: “Biped”: a dance with virtual and company dancers, Part 1. IEEE MultiMedia, 6(3):4–7, July/ September 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u3004.
pdf.

Abouaf:1999:AMBb

[Abo99b] Jeffrey Abouaf. Artful media: “Biped”: a dance with virtual and company dancers, Part 2. IEEE MultiMedia, 6(4):5–7, October/December 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/

                                        books/mu1999/pdf/u4005.
                                        pdf.

Abouelenien:2019:GDM

[ABPR^{+}19] M. Abouelenien, M. Burzo, V. Pérez-Rosas, R. Mihalcea, H. Sun, and B. Zhao. Gender differences in multimodal contact-free deception detection. IEEE MultiMedia, 26(3):19–30, July/ September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Abdullah:2018:STM

[AC18] S. Abdullah and T. Choudhury. Sensing technologies for monitoring serious mental illnesses. IEEE MultiMedia, 25(1):61–75, January/March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Ajankar:2024:IRE

[AD24] Sonali Ajankar and Tanima Dutta. Image-relevant entities knowledge-aware news image captioning. IEEE MultiMedia, 31(1):88–98, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Adams:2003:WDC

[Ada03]  Brett Adams.  Where

does computational media aesthetics fit? IEEE MultiMedia, 10(2):18–27, April/June 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
02/u2018abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/02/u2018.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/02/
u2018.pdf.

Alonso:1995:TCS

[AdAFM95] Fernando Alonso, Angélica de Antonio, José L. Fuertes, and César Montes. Teaching communication skills to hearing-impaired children with an intelligent multimedia system. IEEE MultiMedia, 2(4):55– 67, Winter 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u40055.
pdf;   http://www.computer.
org/multimedia/mu1995/
u4055abs.htm.

Addison:2000:ETV

[Add00] Alonzo C. Addison. Emerging trends in virtual heritage. IEEE MultiMedia, 7(2):22–25, April/ June 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2022.
pdf;   http://www.computer.

                                        org/multimedia/mu2000/
                                        u2022abs.htm.

Alisi:2005:NIE

[ADV05] Thomas M. Alisi, Alberto Del Bimbo, and Alessandro Valli. Natural interfaces to enhance visitors’ experiences. IEEE Multi-Media, 12(3):80–85, July/ September 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Akhtar:2018:VNB

[AF18] Z. Akhtar and T. H. Falk. Visual nonverbal behavior analysis: The path forward. IEEE MultiMedia, 25(2):47–60, April/June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Apteker:1995:VAF

[AFKN95] Ronnie T. Apteker, James A. Fisher, Valentin S. Kisimov, and Hanoch Neishlos. Video acceptability and frame rate. IEEE MultiMedia, 2(3):32–40, Fall 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u30032.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1995/
                                        u3032abs.htm.

Aizawa:2020:BMD

[AFO^{+}20] Kiyoharu Aizawa, Azuma Fujimoto, Atsushi Otsubo, Toru Ogawa, Yusuke Matsui, Koki Tsubota, and Hikaru Ikuta. Building a manga dataset Manga109 with annotations for multimedia applications. IEEE MultiMedia, 27(2):8–18, April/June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Addison:2000:VAH

[AG00] Alonzo C. Addison and Marco Gaiani. Virtualized architectural heritage: New tools and techniques. IEEE MultiMedia, 7(2):26– 31, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2026.
pdf;   http://www.computer.
org/multimedia/mu2000/
u2026abs.htm.

Ahmed:2007:KEH

[AGR07] Wamiq M. Ahmed, Arif Ghafoor, and J. Paul Robinson. Knowledge extraction for high-throughput biological imaging. IEEE MultiMedia, 14(4):52–62, October/December 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Abdoli:2021:GBI

[AGR^{+}21] Mohsen Abdoli, Thomas Guionnet, Mickael Raulet, Gosala Kulupana, and Saverio Blasi. Gradientbased intraprediction fusion for video coding. IEEE MultiMedia, 28(3):88–96, July/September 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

AguierreSmith:1996:PRM

[Agu96] T. G. Aguierre Smith. Project reports: Multimedia lab supports sustainable development. IEEE MultiMedia, 3(1):79–83, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u1079.
                                        pdf.

Akhtar:2018:BSI

[AHN^{+}18] Z. Akhtar, A. Hadid, M. S. Nixon, M. Tistarelli, J. Dugelay, and S. Marcel. Biometrics: In search of identity and security (Q&A). IEEE MultiMedia, 25(3):22–35, July/ September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Arefin:2013:CAT

[AHR^{+}13] Ahsan Arefin, Zixia Huang, Raoul Rivas, Shu Shi, Pengye Xia, Klara Nahrstedt, Wanmin Wu, Gregorij Kurillo, and Ruzena Bajcsy. Classification and analysis of 3D teleimmersive activities. IEEE MultiMedia, 20(1):38–48, January/March 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ashmanov:1999:PRV

[AK99] Igor Ashmanov and Natalya Kasperskaya. Project reports: The Virus Encyclopedia: reaching a new level of information comfort. IEEE Multi-Media, 6(3):81–84, July/ September 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u3081.
pdf.

AprilPyone:2022:PPI

[AK22] MaungMaung AprilPyone and Hitoshi Kiya. Privacypreserving image classification using an isotropic network. IEEE MultiMedia, 29(2):23–33, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Amatriain:2009:AIM

[AKMHP09] Xavier Amatriain, JoAnn Kuchera-Morin, Tobias Hollerer, and Stephen Travis Pope. The AlloSphere: Immersive multimedia for scientific discovery and artistic exploration. IEEE MultiMedia, 16(2):64–75, April/June 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Alpert:1996:EEE

[ALK^{+}96] Sherman R. Alpert, Mark R. Laff, W. Randall Koons, David A. Epstein, Danny Soroker, David C. Morrill, and Arthur J. Stein. The EFX editing and effects environment. IEEE MultiMedia, 3(1):15–29, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u1015.
pdf;   http://www.computer.
org/multimedia/mu1996/
u1015abs.htm.

Ahmed:2022:EMF

[ALS22] Usman Ahmed, Jerry Chun-Wei Lin, and Gautam Srivastava. Efficient multimedia frame-skipping architecture using deep learning in vehicular networks. IEEE MultiMedia, 29(2):66–73, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Amin:2021:LDN

[AMM21] Fazail Amin, Arijit Mondal, and Jimson Mathew. A large dataset with a new framework for abandoned object detection in complex scenarios. IEEE MultiMedia, 28(3):75–87, July/ September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Adjeroh:1997:MDM

[AN97] Donald A. Adjeroh and Kingsley C. Nwosu. Multimedia database management — requirements and archives. IEEE Multi-Media, 4(3):24–33, July/ September 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3024.
pdf;   http://www.computer.
org/multimedia/mu1997/
u3024abs.htm.

Andren:1998:SIW

[And98] Carl Andren. Standards:

IEEE 802.11 wireless LAN: can we use it for multimedia? IEEE Multi-Media, 5(2):84–89, April/ June 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u2084.
pdf.

Andres:2001:GEI

[And01] Frederic Andres. Guest Editor’s introduction. IEEE MultiMedia, 8(4):20–21, October/December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://dlib.computer.
                                        org/mu/books/mu2001/pdf/
                                        u4020.pdf;   http://www.
                                        computer.org/multimedia/
                                        mu2001/u4020abs.htm.

Andrews:2008:AMR

[And08] Ian Andrews. Artful media: Random sound and vision. IEEE MultiMedia, 15(2): 13–15, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Angelides:2003:GEI

[Ang03] Marios C. Angelides. Guest Editor’s introduction: Multimedia content modeling and personalization. IEEE MultiMedia, 10(4):12–15, October/December 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2003/
                                        04/u4012.pdf;   http:/
                                        /csdl.computer.org/comp/
                                        mags/mu/2003/04/u4012abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/04/
                                        u4012.htm.

Anonymous:1994:CPI

[Ano94a] Anonymous. Color pictures for IEEE MultiMedia. IEEE MultiMedia, 1(3):70–??, Autumn/Fall 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:1994:ES

[Ano94b] Anonymous. Electronic submissions. IEEE MultiMedia, 1(3):69–??, Autumn/Fall 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:1994:IMI

[Ano94c] Anonymous. IEEE multimedia information for authors. IEEE MultiMedia, 1(3):68–??, Autumn/Fall 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:1995:AII

[Ano95a]  Anonymous.  1995 an-

nual index: IEEE Multi-Media volume 2. IEEE MultiMedia, 2(4):99–103, Winter 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u40099.
pdf.

Anonymous:1995:UEa

[Ano95b] Anonymous. Upcoming events. IEEE MultiMedia, 2(1):105, Spring 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u10105.
pdf.

Anonymous:1995:UEb

[Ano95c] Anonymous. Upcoming events. IEEE MultiMedia, 2(3):73–74, Fall 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u30073.
pdf.

Anonymous:1996:AII

[Ano96a]  Anonymous.  1996 an-

nual index: IEEE Multi-Media volume 3. IEEE MultiMedia, 3(4):91–95, Winter 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u4091.
pdf.

Anonymous:1996:IMI

[Ano96b] Anonymous. IEEE MultiMedia information for authors. IEEE Multi-Media, 3(2):88–90, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2088.
pdf.

Anonymous:1996:UEa

[Ano96c] Anonymous. Upcoming events. IEEE MultiMedia, 3(1):89–90, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1996/pdf/
u1089.pdf.

Anonymous:1996:UEb

[Ano96d] Anonymous. Upcoming events. IEEE Multi-Media, 3(2):91–93, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2091.
pdf.

Anonymous:1997:AII

[Ano97] Anonymous. 1997 annual index: IEEE MultiMedia volume 4. IEEE Multi-Media, 4(4):90–95, October/December 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u4090.
pdf.

Anonymous:1998:AII

[Ano98a] Anonymous. 1998 annual index: IEEE MultiMedia volume 5. IEEE Multi-Media, 5(4):91–96, October/December 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u4091.
pdf.

Anonymous:1998:MRS

[Ano98b] Anonymous. Media reviews: On the shelf. IEEE MultiMedia, 5(3):71, July/ September 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u3069.
pdf.

Anonymous:1998:UEa

[Ano98c] Anonymous. Upcoming events. IEEE MultiMedia, 5(1):(cover 2), January/March 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u1c2.
pdf.

Anonymous:1998:UEb

[Ano98d]  Anonymous.  Upcom-

ing events. IEEE MultiMedia, 5(2):6–7, April/ June 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u2006.
pdf.

Anonymous:1999:AII

[Ano99a] Anonymous. 1999 annual index: IEEE MultiMedia volume 6. IEEE Multi-Media, 6(4):97–104, October/December 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u4097.
pdf.

Anonymous:1999:UEa

[Ano99b] Anonymous. Upcoming events. IEEE MultiMedia, 6(1):20–21, January/ March 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u1020.
pdf.

Anonymous:1999:UEb

[Ano99c]  Anonymous.  Upcom-

ing events. IEEE MultiMedia, 6(2):13, April/ June 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u2013.
pdf.

Anonymous:1999:UEc

[Ano99d]  Anonymous.  Upcom-

ing events. IEEE MultiMedia, 6(3):85, July/ September 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u3085.
pdf.

Anonymous:1999:UEd

[Ano99e] Anonymous. Upcoming events. IEEE MultiMedia, 6(4):(cover 2), October/December 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u4c2.
pdf.

Anonymous:2000:AII

[Ano00a] Anonymous. 2000 annual index: IEEE MultiMedia volume 7. IEEE Multi-Media, 7(4):83–93, October/December 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u4083.
pdf.

Anonymous:2000:UEa

[Ano00b] Anonymous. Upcoming events. IEEE MultiMedia, 7(1):75, January/ March 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u1075.
pdf.

Anonymous:2000:UEb

[Ano00c]  Anonymous.  Upcom-

ing events. IEEE MultiMedia, 7(2):93, April/ June 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2093.
pdf.

Anonymous:2001:AI

[Ano01a] Anonymous. Annual index. IEEE MultiMedia, 8(4):92–101, October/ December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u4092.
pdf;   http://www.computer.
org/multimedia/mu2001/
u4092abs.htm.

Anonymous:2001:AME

[Ano01b] Anonymous. Artful media: Embodying robotic art: Cybernetic cinematics. IEEE Multi-Media, 8(3):4–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u3004.
pdf;   pdf/u3004.pdf.

Anonymous:2001:N

[Ano01c]  Anonymous.  In the

news. IEEE MultiMedia, 8(4):16–19, October/ December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u4016.
pdf;   http://www.computer.
org/multimedia/mu2001/
u4016abs.htm.

Anonymous:2001:NAD

[Ano01d] Anonymous. In the news: Advances in display technology. IEEE Multi-Media, 8(3):16–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u3016.
pdf.

Anonymous:2001:NSS

[Ano01e] Anonymous. In the news: Stepping into smart clothing. IEEE MultiMedia, 8(2):16–??, April/ June 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u2016.
pdf.

Anonymous:2001:MIM

[Ano01f] Anonymous. Media impact: Multimedia standards: Building blocks of the Web. IEEE Multi-Media, 8(3):13–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u3013.
pdf.

Anonymous:2001:MR

[Ano01g] Anonymous. Media reviews. IEEE MultiMedia, 8(4):104–??, October/ December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u4104.
pdf;   http://www.computer.
org/multimedia/mu2001/
u4104abs.htm.

Anonymous:2001:MRB

[Ano01h] Anonymous. Media reviews: Building successful human centered systems. IEEE MultiMedia, 8(3):102–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u3102.
pdf.

Anonymous:2001:MVM

[Ano01i] Anonymous. Media vision:

Mobile multimedia. IEEE MultiMedia, 8(3):1–??, July/September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u3001.
                                        pdf.

Anonymous:2001:MWD

[Ano01j]  Anonymous.  Multime-

dia at work: Digital-video management for heterogeneous and distributed resources. IEEE Multi-Media, 8(3):96–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u3096.
                                        pdf.

Anonymous:2001:NPa

[Ano01k] Anonymous. New products. IEEE MultiMedia, 8(2):109–??, April/ June 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u2109.
                                        pdf.

Anonymous:2001:NPb

[Ano01l] Anonymous. New products. IEEE MultiMedia, 8(3):104–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u3104.
                                        pdf.

Anonymous:2001:NPc

[Ano01m] Anonymous. New products. IEEE MultiMedia, 8(4):89–91, October/ December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u4089.
pdf;   http://www.computer.
org/multimedia/mu2001/
u4089abs.htm.

Anonymous:2001:SHB

[Ano01n] Anonymous. Standards:

High-bandwidth interface for multimedia communications over fixed wireless systems. IEEE MultiMedia, 8(3):87–??, July/ September 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u3087.
pdf.

Anonymous:2001:UEa

[Ano01o] Anonymous. Upcoming events. IEEE MultiMedia, 8(1):77, January/ March 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u1077.
pdf.

Anonymous:2001:UEb

[Ano01p]  Anonymous.  Upcom-

ing events. IEEE MultiMedia, 8(3):??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u3c2.
pdf.

Anonymous:2001:UEc

[Ano01q] Anonymous. Upcoming events. IEEE MultiMedia, 8(4):102–??, October/ December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u4102.
pdf;   http://www.computer.
org/multimedia/mu2001/
u4102abs.htm.

Anonymous:2001:VVS

[Ano01r] Anonymous. Visions and views: Spies, thieves, and lies: The battle for multimedia in the digital era. IEEE Multi-Media, 8(3):8–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u3008.
pdf;   pdf/u3008.pdf.

Anonymous:2002:AI

[Ano02a]  Anonymous.  2002 an-

nual index. IEEE Multi-Media, 9(4):91–102, October/December 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
04/u4091.pdf;   http:/
/csdl.computer.org/dl/
mags/mu/2002/04/u4091.
htm.

Anonymous:2002:ED

[Ano02b] Anonymous. The early days of E.A.T. IEEE MultiMedia, 9(2):4–5, April/ June 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2002/02/u2004.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/02/
u2004.pdf.

Anonymous:2002:EC

[Ano02c]  Anonymous.  Experien-

tial computing. IEEE MultiMedia, 9(2):96–??, April/June 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2002/02/u2096.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/02/
u2096.pdf.

Anonymous:2002:FMF

[Ano02d] Anonymous. Finding multimedia’s fit. IEEE MultiMedia, 9(2):c2–1, April/ June 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2002/02/u20c2.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/02/
                                        u20c2.pdf.

Anonymous:2002:FC

[Ano02e] Anonymous. Folk computing. IEEE Multi-Media, 9(3):96–??, July/ September 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2002/03/u3096.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/03/
                                        u3096.pdf.

Anonymous:2002:HGC

[Ano02f] Anonymous. The Holy Grail of content-based media analysis. IEEE MultiMedia, 9(2):6–10, April/ June 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2002/02/u2006.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/02/
                                        u2006.pdf.

Anonymous:2002:MSM

[Ano02g] Anonymous. Making sense of multiple technologies. IEEE MultiMedia, 9(2):77, 87, April/June 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2002/02/u2077.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/02/
u2077.pdf.

Anonymous:2002:MR

[Ano02h] Anonymous. Media reviews. IEEE MultiMedia, 9(1):93–95, January/ March 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2002/pdf/u1093.
pdf;   http://www.computer.
org/multimedia/mu2002/
u1093abs.htm.

Anonymous:2002:MTF

[Ano02i] Anonymous. More than face value: Airports and multimedia security. IEEE MultiMedia, 9(2):11–13, April/June 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2002/02/u2011.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/02/
u2011.pdf.

Anonymous:2002:MOM

[Ano02j]  Anonymous.  MPEG-7:

Overview of MPEG-7 description tools, part 2.

IEEE MultiMedia, 9(3):83– 93, July/September 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/dl/mags/mu/2002/03/
                                        u3083.htm;   http://csdl.
                                        computer.org/dl/mags/
                                        mu/2002/03/u3083.pdf.

Anonymous:2002:MGM

[Ano02k]  Anonymous.  MPEG-7:

The Generic Multimedia Content Description Standard, Part 1. IEEE MultiMedia, 9(2):78–87, April/June 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2002/02/u2078.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/02/
                                        u2078.pdf.

Anonymous:2002:MGC

[Ano02l] Anonymous. Multimedia goes corporate. IEEE MultiMedia, 9(3):18–21, July/ September 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2002/03/u3018.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/03/
                                        u3018.pdf.

Anonymous:2002:MTC

[Ano02m]  Anonymous.  Multime-

dia technology in context. IEEE MultiMedia, 9(3):12–17, July/ September 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2002/03/u3012.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/03/
u3012.pdf.

Anonymous:2002:NPa

[Ano02n] Anonymous. New products. IEEE MultiMedia, 9(1):90–92, 95, January/March 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2002/pdf/u1090.
pdf;   http://www.computer.
org/multimedia/mu2002/
u1090abs.htm.

Anonymous:2002:NPb

[Ano02o] Anonymous. New products. IEEE MultiMedia, 9(2):93–95, April/ June 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2002/02/u2093.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/02/
u2093.pdf.

Anonymous:2002:NPc

[Ano02p] Anonymous. New products. IEEE MultiMedia, 9(3):c3, 94–95, July/ September 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2002/03/u3c3.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/03/
                                        u3c3.pdf.

Anonymous:2002:NPd

[Ano02q] Anonymous. New products. IEEE MultiMedia, 9(4):73–75, October/ December 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2002/04/u4073.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/04/
                                        u4073.pdf.

Anonymous:2002:PE

[Ano02r] Anonymous. Pictures at an exhibition. IEEE Multi-Media, 9(2):14–17, April/ June 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2002/02/u2014.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/02/
                                        u2014.pdf.

Anonymous:2002:PVM

[Ano02s]  Anonymous.  Playing

with virtual musicians: The continuator in practice. IEEE MultiMedia, 9(3):77–82, July/ September 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2002/03/u3077.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/03/
u3077.pdf.

Anonymous:2002:SB

[Ano02t] Anonymous. Streaming the brain. IEEE MultiMedia, 9(3):4–8, July/ September 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2002/03/u3004.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/03/
u3004.pdf.

Anonymous:2002:TMS

[Ano02u] Anonymous. Telematics might steer your car into the future. IEEE MultiMedia, 9(3):9–10, July/ September 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2002/03/u3009.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/03/
u3009.pdf.

Anonymous:2002:UEa

[Ano02v]  Anonymous.  Upcom-

ing events. IEEE MultiMedia, 9(2):c3, April/ June 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2002/02/u20c3.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/02/
                                        u20c3.pdf.

Anonymous:2002:UEb

[Ano02w] Anonymous. Upcoming events. IEEE Multi-Media, 9(3):11–??, July/ September 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2002/03/u3011.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/03/
                                        u3011.pdf.

Anonymous:2002:UEc

[Ano02x] Anonymous. Upcoming events. IEEE Multi-Media, 9(4):c3, October/ December 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2002/04/u40c3.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/04/
                                        u40c3.pdf.

Anonymous:2002:UAM

[Ano02y] Anonymous. Using 3D and ancillary media to train construction workers. IEEE MultiMedia, 9(2):88– 92, April/June 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2002/02/u2088.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/02/
u2088.pdf.

Anonymous:2003:AII

[Ano03a] Anonymous. 2003 annual index, IEEE MultiMedia volume 10. IEEE MultiMedia, 10(4):100–110, October/December 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
04/u4100.pdf;   http:/
/csdl.computer.org/dl/
mags/mu/2003/04/u4100.
htm.

Anonymous:2003:NPa

[Ano03b] Anonymous. New products. IEEE MultiMedia, 10(1):72–75, January/ March 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/01/u1072.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/01/
u1072.pdf.

Anonymous:2003:NPb

[Ano03c] Anonymous. New products. IEEE MultiMedia, 10(2):83–87, April/ June 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2003/02/u2083.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/02/
                                        u2083.pdf.

Anonymous:2003:NPc

[Ano03d] Anonymous. New products. IEEE MultiMedia, 10(3):c3, July/September 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/dl/mags/mu/2003/03/
                                        u30c3.htm;   http://csdl.
                                        computer.org/dl/mags/
                                        mu/2003/03/u30c3.pdf.

Anonymous:2003:NPd

[Ano03e] Anonymous. New products. IEEE MultiMedia, 10(4):c3, October/ December 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2003/04/u40c3.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/04/
                                        u40c3.pdf.

Anonymous:2003:TIM

[Ano03f] Anonymous. Thanks to IEEE MultiMedia’s reviewers. IEEE MultiMedia, 10(1):1–??, January/ March 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/01/u1001.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/01/
u1001.pdf.

Anonymous:2003:UEa

[Ano03g] Anonymous. Upcoming events. IEEE MultiMedia, 10(1):71–??, January/ March 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/01/u1071.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/01/
u1071.pdf.

Anonymous:2003:UEb

[Ano03h] Anonymous. Upcoming events. IEEE MultiMedia, 10(2):79–80, April/ June 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/02/u2079.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/02/
u2079.pdf.

Anonymous:2004:AI

[Ano04a]  Anonymous.  2004 an-

nual index. IEEE Multi-Media, 11(4):90–96, October/December 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2004/
                                        04/u4090.pdf;   http:/
                                        /csdl.computer.org/dl/
                                        mags/mu/2004/04/u4090.
                                        htm.

Anonymous:2004:NPa

[Ano04b] Anonymous. New products. IEEE MultiMedia, 11(1):92–93, January/ March 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2004/01/u1092.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/01/
                                        u1092.pdf.

Anonymous:2004:NPb

[Ano04c] Anonymous. New products. IEEE MultiMedia, 11(2):88, 85–87, April/ June 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2004/02/u2088.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/02/
                                        u2088.pdf.

Anonymous:2004:NPc

[Ano04d] Anonymous. New products. IEEE MultiMedia, 11(4):88–89, October/December 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
04/u4088.pdf;   http:/
/csdl.computer.org/dl/
mags/mu/2004/04/u4088.
htm.

Anonymous:2004:QE

[Ano04e] Anonymous. Quality of experience. IEEE Multi-Media, 11(1):96, 95, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
01/u1096abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2004/01/u1096.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/01/
u1096.pdf.

Anonymous:2004:RR

[Ano04f]  Anonymous.  Readers’

remarks. IEEE MultiMedia, 11(1):C3, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
01/u10c3.pdf;   http:/
/csdl.computer.org/comp/
mags/mu/2004/01/u10c3abs.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/01/
u10c3.htm.

Anonymous:2004:TIM

[Ano04g] Anonymous. Thanks to IEEE MultiMedia’s reviewers. IEEE Multi-Media, 11(1):90–91, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2004/
                                        01/u1090.pdf;   http:/
                                        /csdl.computer.org/comp/
                                        mags/mu/2004/01/u1090abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/01/
                                        u1090.htm.

Anonymous:2004:UEa

[Ano04h] Anonymous. Upcoming events. IEEE MultiMedia, 11(1):94–??, January/ March 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2004/01/u1094.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/01/
                                        u1094.pdf.

Anonymous:2004:UEb

[Ano04i] Anonymous. Upcoming events. IEEE MultiMedia, 11(4):c3, October/ December 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2004/
                                        04/u40c3.pdf;   http:/

/csdl.computer.org/dl/
mags/mu/2004/04/u40c3.
htm.

Anonymous:2005:AII

[Ano05a]  Anonymous.  2005 an-

nual index, IEEE Multi-Media, vol. 12. IEEE MultiMedia, 12(4):92–101, October/December 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2005/
04/u4092.pdf.

Anonymous:2005:NPa

[Ano05b] Anonymous. New products. IEEE MultiMedia, 12(3):93–95, July/ September 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2005/
03/u3093.pdf.

Anonymous:2005:NPb

[Ano05c] Anonymous. New products. IEEE MultiMedia, 12(4):90–91, October/December 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2005/
04/u4090.pdf.

Anonymous:2005:RR

[Ano05d]  Anonymous.  Readers’

remarks. IEEE Multi-Media, 12(4):1, October/ December 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2005/
                                        04/u4001.pdf.

Anonymous:2005:UEa

[Ano05e]  Anonymous.  Upcom-

ing events. IEEE MultiMedia, 12(1):73, January/March 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2005/
                                        01/u1073.pdf.

Anonymous:2005:UEb

[Ano05f] Anonymous. Upcoming events. IEEE MultiMedia, 12(2):87, April/June 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2005/
                                        02/u2087.pdf.

Anonymous:2006:AII

[Ano06a] Anonymous. 2006 annual index, IEEE MultiMedia vol. 13. IEEE MultiMedia, 13(4):91–96, October/ December 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2006:NPa

[Ano06b] Anonymous. New products. IEEE MultiMedia, 13(2):93–94, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2006:NPb

[Ano06c] Anonymous. New products. IEEE MultiMedia, 13 (4):85, October/December 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2006:TIM

[Ano06d] Anonymous. Thanks to IEEE MultiMedia reviewers. IEEE MultiMedia, 13 (1):94–95, January/March 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2006:UEa

[Ano06e] Anonymous. Upcoming events. IEEE MultiMedia, 13(1):93, January/March 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2006:UEb

[Ano06f] Anonymous. Upcoming events. IEEE MultiMedia, 13(2):89, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2006:UEc

[Ano06g] Anonymous. Upcoming events. IEEE MultiMedia, 13(3):85, July/September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2006:UEd

[Ano06h] Anonymous. Upcoming events. IEEE MultiMedia, 13(4):7, October/ December 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2007:AIM

[Ano07a] Anonymous. 2007 annual index, MultiMedia vol. 14. IEEE MultiMedia, 14(4): i1–i6, October/December 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2007:NPa

[Ano07b] Anonymous. New products. IEEE MultiMedia, 14 (1):96–97, January/March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2007:NPb

[Ano07c] Anonymous. New products. IEEE MultiMedia, 14(2):84–86, April/June 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2007:NPc

[Ano07d] Anonymous. New products. IEEE MultiMedia, 14(3):110, July/September 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2007:NPd

[Ano07e] Anonymous. New products. IEEE MultiMedia, 14(4):92, c3, October/ December 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2007:TIM

[Ano07f] Anonymous. Thanks to IEEE MultiMedia’s reviewers. IEEE MultiMedia, 14 (1):98–99, January/March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2007:UEa

[Ano07g] Anonymous. Upcoming events. IEEE MultiMedia, 14(1):95, January/March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2007:UEb

[Ano07h] Anonymous. Upcoming events. IEEE MultiMedia, 14(2):83, April/June 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2007:UEc

[Ano07i] Anonymous. Upcoming events. IEEE MultiMedia, 14(3):c3, July/September 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2008:AI

[Ano08a] Anonymous. Annual index. IEEE MultiMedia, 15 (4):0, October/December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2008:NPa

[Ano08b] Anonymous. New products. IEEE MultiMedia, 15 (1):96, c3, January/March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2008:NPb

[Ano08c] Anonymous. New products. IEEE MultiMedia, 15(2):83–84, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2008:TIM

[Ano08d] Anonymous. Thanks to IEEE MultiMedia’s reviewers. IEEE MultiMedia, 15 (1):92–95, January/March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2008:UEa

[Ano08e] Anonymous. Upcoming events. IEEE MultiMedia, 15(1):13, January/March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2008:UEb

[Ano08f] Anonymous. Upcoming events. IEEE MultiMedia, 15(2):71, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2008:UEc

[Ano08g] Anonymous. Upcoming events. IEEE MultiMedia, 15(3):11, July/September 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2008:UEd

[Ano08h] Anonymous. Upcoming events. IEEE MultiMedia, 15(4):96, October/ December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2009:AI

[Ano09a] Anonymous. Annual index. IEEE MultiMedia, 16 (4):0, October/December 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2009:BC

[Ano09b] Anonymous. Back cover. IEEE MultiMedia, 16(3): c4, July/September 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2009:DNP

[Ano09c] Anonymous. Departments: New products. IEEE MultiMedia, 16(2):1, April/ June 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2009:EML

[Ano09d] Anonymous. EIC’s message: Looking forward to changes in 2009. IEEE MultiMedia, 16(1):1, January/March 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2009:NPa

[Ano09e] Anonymous. New products. IEEE MultiMedia, 16(1):100–101, January/March 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2009:NPb

[Ano09f] Anonymous. New products. IEEE MultiMedia, 16(3):c3, July/September 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2009:NPc

[Ano09g] Anonymous. New products. IEEE MultiMedia, 16(4):82–83, October/ December 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2009:UEa

[Ano09h] Anonymous. Upcoming events. IEEE MultiMedia, 16(1):13, January/March 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2009:UEb

[Ano09i] Anonymous. Upcoming events. IEEE MultiMedia, 16(2):96, April/June 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2009:UEc

[Ano09j] Anonymous. Upcoming events. IEEE MultiMedia, 16(3):81, July/September 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2009:UEd

[Ano09k] Anonymous. Upcoming events. IEEE MultiMedia, 16(4):84, October/ December 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2010:EM

[Ano10a] Anonymous. EIC’s message. IEEE MultiMedia, 17(1):2–3, January/March 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2010:NPa

[Ano10b] Anonymous. New products. IEEE MultiMedia, 17(1):68, January/March 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2010:NPb

[Ano10c] Anonymous. New products. IEEE MultiMedia, 17(2):94–95, April/June 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2010:NPc

[Ano10d] Anonymous. New products. IEEE MultiMedia, 17 (3):79–80, July/September 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2010:NPd

[Ano10e] Anonymous. New products. IEEE MultiMedia, 17(4):108, October/ December 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2010:UEa

[Ano10f] Anonymous. Upcoming events. IEEE MultiMedia, 17(1):67, January/March 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2010:UEb

[Ano10g] Anonymous. Upcoming events. IEEE MultiMedia, 17(2):96, April/June 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2010:UEc

[Ano10h] Anonymous. Upcoming events. IEEE MultiMedia, 17(3):13, July/September 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2010:UEd

[Ano10i] Anonymous. Upcoming events. IEEE MultiMedia, 17(4):89, October/ December 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2011:AI

[Ano11a] Anonymous. Annual index. IEEE MultiMedia, 18 (4):0, October/December 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2011:NPa

[Ano11b] Anonymous. New products. IEEE MultiMedia, 18(1):104, January/March 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2011:NPb

[Ano11c] Anonymous. New products. IEEE MultiMedia, 18(2):88, April/June 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2011:NPc

[Ano11d] Anonymous. New products. IEEE MultiMedia, 18 (3):95–96, July/September 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2011:NPd

[Ano11e] Anonymous. New products. IEEE MultiMedia, 18(4):75–76, October/ December 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2011:TIM

[Ano11f] Anonymous. Thanks to IEEE MultiMedia’s reviewers. IEEE MultiMedia, 18 (1):96–97, January/March 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2011:UEa

[Ano11g] Anonymous. Upcoming events. IEEE MultiMedia, 18(2):89, April/June 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2011:UEb

[Ano11h] Anonymous. Upcoming events. IEEE MultiMedia, 18(3):11, July/September 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2011:UEc

[Ano11i] Anonymous. Upcoming events. IEEE MultiMedia, 18(4):68–69, October/ December 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2012:NPa

[Ano12a] Anonymous. New products. IEEE MultiMedia, 19 (1):88, 87, January/March 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2012:NPb

[Ano12b] Anonymous. New products. IEEE MultiMedia, 19 (2):80–c3, May/June 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2012:NPc

[Ano12c] Anonymous. New products. IEEE MultiMedia, 19 (4):80, October/December 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2013:ASI

[Ano13a] Anonymous. 2013 annual subject index. IEEE Multi-Media, 20(4):74–75, October/December 2013. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2013:R

[Ano13b] Anonymous. 2013 reviewers. IEEE MultiMedia, 20 (4):8, October/December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2013:CP

[Ano13c] Anonymous. Call for papers. IEEE MultiMedia, 20 (4):9, October/December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2013:CPS

[Ano13d] Anonymous. Call for papers: Special issue on interaction sonification. IEEE MultiMedia, 20(4):c3, October/December 2013. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2013:ESA

[Ano13e] Anonymous. ePub subscription [advertisement]. IEEE MultiMedia, 20(4): 49, October/December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2013:FC

[Ano13f] Anonymous. [Front cover].

IEEE MultiMedia, 20 (4):c1, October/December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2013:ICS

[Ano13g] Anonymous. IEEE Computer Society information. IEEE MultiMedia, 20 (4):73, October/December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2013:MAb

[Ano13h] Anonymous. Membership [advertisement]. IEEE MultiMedia, 20(4):c4, October/December 2013. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2013:MAa

[Ano13i] Anonymous. myComputer [advertisement]. IEEE MultiMedia, 20(4):37, October/December 2013. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2013:NPa

[Ano13j] Anonymous. New products. IEEE MultiMedia, 20(2):80, April/June 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2013:NPb

[Ano13k] Anonymous. New products. IEEE MultiMedia, 20(3):96, July/September 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2013:NPc

[Ano13l] Anonymous. New products. IEEE MultiMedia, 20 (4):96, October/December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2013:TC

[Ano13m] Anonymous. Table of contents. IEEE MultiMedia, 20(4):c2–1, October/ December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:CPa

[Ano14a] Anonymous. Call for papers. IEEE MultiMedia, 21(1):9, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:CPb

[Ano14b] Anonymous. Call for papers. IEEE MultiMedia, 21(1):c3, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:CPc

[Ano14c] Anonymous. Call for papers. IEEE MultiMedia, 21(3):c3, July/September 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:CPH

[Ano14d] Anonymous. Call for papers [house advertisement]. IEEE MultiMedia, 21(2): c4, April/June 2014. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:FCb

[Ano14e] Anonymous. [Front cover].

IEEE MultiMedia, 21(3): c1, July/September 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:FCc

[Ano14f] Anonymous. [Front cover].

IEEE MultiMedia, 21 (4):c1, October/December 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://www.computer.org/
csdl/mags/mu/2014/04/
mmu20140400c1.pdf.

Anonymous:2014:FCa

[Ano14g] Anonymous. Front covers. IEEE MultiMedia, 21(1): c1, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:ICCb

[Ano14h] Anonymous. IEEE CG&A call for papers. IEEE MultiMedia, 21(2):71, April/ June 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:ICCa

[Ano14i] Anonymous. IEEE cloud computing [house advertisement]. IEEE Multi-Media, 21(1):c4, January/ March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:ICS

[Ano14j] Anonymous. IEEE Computer Society [house advertisement]. IEEE Multi-Media, 21(1):59, January/ March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:ISP

[Ano14k] Anonymous. IEEE Security & Privacy [advertisement]. IEEE MultiMedia, 21(3):65, July/September 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:MMH

[Ano14l] Anonymous. Membership matters [house advertisement]. IEEE MultiMedia, 21(1):35, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:NPa

[Ano14m] Anonymous. New products. IEEE MultiMedia, 21(1):96, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:NPb

[Ano14n] Anonymous. New products. IEEE MultiMedia, 21(2):96, April/June 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:RSP

[Ano14o] Anonymous. Rock stars of 3D printing house advertisement. IEEE MultiMedia, 21(4):c4, October/December 2014. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu20140400c4.
pdf.

Anonymous:2014:RSC

[Ano14p] Anonymous. Rock stars of cybersecurity [advertisement]. IEEE MultiMedia, 21(3):c4, July/September 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:SEC

[Ano14q] Anonymous. Software engineering for the 21st Century [house advertisement]. IEEE MultiMedia, 21(2): c3, April/June 2014. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:SCH

[Ano14r] Anonymous. Stay connection [house advertisement]. IEEE MultiMedia, 21(1): 83, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:TCa

[Ano14s] Anonymous. Table of contents. IEEE MultiMedia, 21(1):c2–1, January/ March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:TCb

[Ano14t] Anonymous. Table of contents. IEEE MultiMedia, 21(2):c2–1, April/June 2014. CODEN IEMUE4.

ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:TCc

[Ano14u]  Anonymous.  Table of

contents. IEEE Multi-Media, 21(3):c2–1, July/ September 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2014:TCd

[Ano14v] Anonymous. Table of contents. IEEE MultiMedia, 21(4):c2–1, October/ December 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu20140400c2.
pdf.

Anonymous:2014:TCL

[Ano14w]  Anonymous.  Take the

CS library wherever you go! [house advertisement]. IEEE MultiMedia, 21(1): 47, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2015:R

[Ano15a]  Anonymous.  2014 re-

viewers. IEEE MultiMedia, 22(1):4–5, January/ March 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/

                                        mags/mu/2015/01/mmu2015010004.
                                        html.

Anonymous:2015:CPa

[Ano15b] Anonymous. Call for papers. IEEE MultiMedia, 22(1):c3, January/ March 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/01/mmu20150100c3.
                                        pdf.

Anonymous:2015:CPb

[Ano15c]  Anonymous.  Call for

papers. IEEE Multi-Media, 22(2):3, April/ June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/02/mmu2015020003.
                                        pdf.

Anonymous:2015:CPc

[Ano15d]  Anonymous.  Call for

papers. IEEE Multi-Media, 22(2):75, April/ June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/02/mmu2015020075.
                                        pdf.

Anonymous:2015:CP

[Ano15e] Anonymous. Call for papers. IEEE MultiMedia, 22(4):c3, October/ December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/04/mmu20150400c3.
pdf.

Anonymous:2015:CPY

[Ano15f] Anonymous. Conferences in the palm of your hand [Advertisement]. IEEE MultiMedia, 22(4):51, October/December 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/04/mmu2015040051.
pdf.

Anonymous:2015:FYJa

[Ano15g] Anonymous. Focus on your job search. IEEE MultiMedia, 22(2):c4, April/ June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/02/mmu20150200c4.
pdf.

Anonymous:2015:FYJb

[Ano15h] Anonymous. Focus on your job search [Advertisement]. IEEE Multi-Media, 22(4):79, October/ December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/

                                        mags/mu/2015/04/mmu2015040079.
                                        pdf.

Anonymous:2015:FCa

[Ano15i] Anonymous. [Front cover]. IEEE MultiMedia, 22(1): c1, January/March 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://www.computer.org/
                                        csdl/mags/mu/2015/01/
                                        mmu20150100c1.pdf.

Anonymous:2015:FCb

[Ano15j] Anonymous. [Front cover]. IEEE MultiMedia, 22(2): c1, April/June 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/02/mmu20150200c1.
                                        pdf.

Anonymous:2015:FC

[Ano15k] Anonymous. [Front cover]. IEEE MultiMedia, 22 (4):c1, October/December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://www.computer.org/
                                        csdl/mags/mu/2015/04/
                                        mmu20150400c1.pdf.

Anonymous:2015:FCc

[Ano15l] Anonymous. Front covers. IEEE MultiMedia, 22(3): c1, July/September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://www.computer.org/
csdl/mags/mu/2015/03/
mmu20150300c1.pdf.

Anonymous:2015:GMLa

[Ano15m] Anonymous. Get more, for less! IEEE MultiMedia, 22(2):17, April/ June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/02/mmu2015020017.
pdf.

Anonymous:2015:GMLb

[Ano15n] Anonymous. Get more, for less! House advertisement. IEEE MultiMedia, 22(3):43, July/ September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/03/mmu2015030043.
pdf.

Anonymous:2015:ICCa

[Ano15o] Anonymous. IEEE Cloud Computing call for papers. IEEE MultiMedia, 22(1):c4, January/ March 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/01/mmu20150100c4.
pdf.

Anonymous:2015:ICCb

[Ano15p] Anonymous. IEEE cloud computing call for papers. IEEE MultiMedia, 22(2): 9, April/June 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/02/mmu2015020009.
                                        pdf.

Anonymous:2015:ICCd

[Ano15q] Anonymous. IEEE Cloud Computing call for papers. IEEE MultiMedia, 22(4):39, October/ December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/04/mmu2015040039.
                                        pdf.

Anonymous:2015:ICCc

[Ano15r] Anonymous. IEEE Cloud Computing call for papers house advertisement. IEEE MultiMedia, 22(3): c4, July/September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://www.computer.org/
                                        csdl/mags/mu/2015/03/
                                        mmu20150300c4.pdf.

Anonymous:2015:ICS

[Ano15s] Anonymous. IEEE Computer Society [Advertisement]. IEEE MultiMedia, 22(4):93, October/ December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/04/mmu2015040093.
pdf.

Anonymous:2015:RSC

[Ano15t] Anonymous. Rock stars of cybersecurity. IEEE MultiMedia, 22(2):c3, April/ June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/02/mmu20150200c3.
pdf.

Anonymous:2015:SCA

[Ano15u] Anonymous. Stay connected [Advertisement]. IEEE MultiMedia, 22 (4):c4, October/December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://www.computer.org/
csdl/mags/mu/2015/04/
mmu20150400c4.pdf.

Anonymous:2015:TCa

[Ano15v] Anonymous. Table of contents. IEEE MultiMedia, 22(1):c2–1, January/ March 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/01/mmu20150100c2.
pdf.

Anonymous:2015:TCb

[Ano15w]  Anonymous.  Table of

contents. IEEE Multi-Media, 22(2):c2–1, April/ June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/02/mmu20150200c2.
                                        pdf.

Anonymous:2015:TCc

[Ano15x] Anonymous. Table of contents. IEEE MultiMedia, 22(4):c2–1, October/ December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/04/mmu20150400c2.
                                        pdf.

Anonymous:2015:WWL

[Ano15y] Anonymous. Watch the world’s leading experts take multi-core strategies to new heights [Advertisement]. IEEE Multi-Media, 22(4):29, October/ December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/04/mmu2015040029.
                                        pdf.

Anonymous:2016:R

[Ano16a]  Anonymous.  2015 re-

viewers. IEEE Multi-Media, 23(1):4–5, January/March 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
01/mmu2016010004.html.

Anonymous:2016:BRR

[Ano16b] Anonymous. 2017 B. Ramakrishna Rau Award call for nominations. IEEE MultiMedia, 23(4):c4, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
04/mmu20160400c4.pdf.

Anonymous:2016:CPAb

[Ano16c] Anonymous. Call for papers: Advancing multimedia distribution with SDN/NFV-based systems. IEEE MultiMedia, 23(2): 27, April/June 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
02/mmu2016020027.pdf.

Anonymous:2016:CPCb

[Ano16d] Anonymous. Call for papers: Cybersecurity for cyber-enabled multimedia applications. IEEE MultiMedia, 23(4):21, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        04/mmu2016040021.pdf.

Anonymous:2016:CPCa

[Ano16e] Anonymous. Call for papers: Cybersecurity for cyber-enabled multimedia applications house advertisement. IEEE MultiMedia, 23(3):43, July/ September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        03/mmu2016030043.pdf.

Anonymous:2016:CPMb

[Ano16f] Anonymous. Call for papers: Multimedia big data analytics in technology enhanced learning. IEEE MultiMedia, 23(4):c3, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        04/mmu20160400c3.pdf.

Anonymous:2016:CPMa

[Ano16g] Anonymous. Call for papers: Multimedia big data analytics in technology enhanced learning house advertisement. IEEE MultiMedia, 23(3):55, July/ September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
03/mmu2016030055.pdf.

Anonymous:2016:CPN

[Ano16h] Anonymous. Call for papers: New signals in multimedia systems and applications. IEEE MultiMedia, 23(4):31, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
04/mmu2016040031.pdf.

Anonymous:2016:CPNa

[Ano16i] Anonymous. Call for papers: New signals in multimedia systems and applications: Sensing and understanding human behavior and interactions. IEEE MultiMedia, 23(2): 53, April/June 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
02/mmu2016020053.pdf.

Anonymous:2016:CPNb

[Ano16j] Anonymous. Call for papers: New signals in multimedia systems and applications: Sensing and understanding human behavior and interactions house advertisement. IEEE MultiMedia, 23(3):c4, July/ September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        03/mmu20160300c4.pdf.

Anonymous:2016:FYJ

[Ano16k] Anonymous. Focus on your job search house advertisement. IEEE MultiMedia, 23(1):c3, January/March 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        01/mmu20160100c3.pdf.

Anonymous:2016:FCa

[Ano16l] Anonymous. Front cover. IEEE MultiMedia, 23(1): c1, January/March 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        01/mmu20160100c1.pdf.

Anonymous:2016:FCc

[Ano16m] Anonymous. Front cover. IEEE MultiMedia, 23(3): c1, July/September 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
03/mmu20160300c1.pdf.

Anonymous:2016:FCb

[Ano16n] Anonymous. Front covers. IEEE MultiMedia, 23 (2):c1, April/June 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
02/mmu20160200c1.pdf.

Anonymous:2016:FCd

[Ano16o] Anonymous. Front covers. IEEE MultiMedia, 23(4):c1, October/ December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
04/mmu20160400c1.pdf.

Anonymous:2016:ICC

[Ano16p] Anonymous. IEEE Cloud Computing call for papers house advertisement. IEEE MultiMedia, 23(1): c4, January/March 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
01/mmu20160100c4.pdf.

Anonymous:2016:ICSc

[Ano16q] Anonymous. IEEE Computer Society 2016 call for major award nominations. IEEE MultiMedia, 23(2):c4, April/June 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
02/mmu20160200c4.pdf.

Anonymous:2016:ICSd

[Ano16r] Anonymous. IEEE Computer Society [advertisement]. IEEE MultiMedia, 23(4):79, October/ December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
04/mmu2016040079.pdf.

Anonymous:2016:ICSb

[Ano16s] Anonymous. IEEE Computer Society: Be at the center of it all. IEEE MultiMedia, 23(2): c3, April/June 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
02/mmu20160200c3.pdf.

Anonymous:2016:ICSa

[Ano16t] Anonymous. IEEE Computer Society house advertisement. IEEE MultiMedia, 23(1):71, January/March 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
01/mmu2016010071.pdf.

Anonymous:2016:ICSe

[Ano16u] Anonymous. IEEE Computer Society is where you choose the resources that fit your career [Advertisement]. IEEE MultiMedia, 23(4):c2, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
04/mmu20160400c2.pdf.

Anonymous:2016:NMO

[Ano16v] Anonymous. New membership options for a better fit [Advertisement]. IEEE MultiMedia, 23(4):1, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
04/mmu2016040001.pdf.

Anonymous:2016:TCa

[Ano16w]  Anonymous.  Table of

contents. IEEE Multi-Media, 23(3):c2–1, July/ September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.

                                        org/csdl/mags/mu/2016/
                                        03/mmu20160300c2.pdf.

Anonymous:2016:TCb

[Ano16x]  Anonymous.  Table of

contents. IEEE Multi-Media, 23(4):2–3, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        04/mmu2016040002.pdf.

Anonymous:2016:TCLb

[Ano16y] Anonymous. Take the CS library wherever you go! [Advertisement]. IEEE MultiMedia, 23(4):59, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        04/mmu2016040059.pdf.

Anonymous:2016:TCLa

[Ano16z]  Anonymous.  Take the

CS Library wherever you go! House advertisement. IEEE MultiMedia, 23(1): 25, January/March 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        01/mmu2016010025.pdf.

Anonymous:2016:WWL

[Ano16-27] Anonymous. Watch the World’s leading experts take multi-core strategies to new heights [Advertisement]. IEEE MultiMedia, 23(4):67, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
04/mmu2016040067.pdf.

Anonymous:2016:WFV

[Ano16-28]  Anonymous.  World’s

first visual innovation award house advertisement. IEEE Multi-Media, 23(3):c3, July/ September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
03/mmu20160300c3.pdf.

Anonymous:2017:R

[Ano17a] Anonymous. 2016 reviewers. IEEE MultiMedia, 24 (2):7–8, April/June 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
02/mmu2017020007.html.

Anonymous:2017:ITS

[Ano17b] Anonymous. IEEE Transactions on Sustainable Computing. IEEE MultiMedia, 24(4):79, October/December 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
04/mmu2017040079.pdf.

Anonymous:2017:CPMb

[Ano17c] Anonymous. Call for papers: 5G multimedia communications: Theory, technology, and applications. IEEE MultiMedia, 24(3): 1, July/September 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
03/mmu2017030001.pdf.

Anonymous:2017:CPMc

[Ano17d] Anonymous. Call for papers: 5G multimedia communications: Theory, technology, and applications. IEEE MultiMedia, 24(4):1, October/December 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
04/mmu2017040001.pdf.

Anonymous:2017:CPC

[Ano17e] Anonymous. Call for papers: Cybersecurity for cyber-enabled multimedia applications house advertisement. IEEE MultiMedia, 24(1):c2, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
01/mmu20170100c2.pdf.

Anonymous:2017:CPMa

[Ano17f] Anonymous. Call for papers: Multimedia big data analytics in technology enhanced learning house advertisement. IEEE MultiMedia, 24(1):c3, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
01/mmu20170100c3.pdf.

Anonymous:2017:CPV

[Ano17g] Anonymous. Call for papers: Vision and language integration meets multimedia fusion house advertisement. IEEE MultiMedia, 24(1):35, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
01/mmu2017010035.pdf.

Anonymous:2017:FCa

[Ano17h] Anonymous. Front cover.

IEEE MultiMedia, 24(1): c1, January/March 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        01/mmu20170100c1.pdf.

Anonymous:2017:FCc

[Ano17i] Anonymous. Front cover. IEEE MultiMedia, 24(3): c1, July/September 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        03/mmu20170300c1.pdf.

Anonymous:2017:FCd

[Ano17j] Anonymous. Front cover. IEEE MultiMedia, 24 (4):c1, October/December 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        04/mmu20170400c1.pdf.

Anonymous:2017:FCb

[Ano17k] Anonymous. Front covers. IEEE MultiMedia, 24 (2):c1, April/June 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        02/mmu20170200c1.pdf.

Anonymous:2017:GF

[Ano17l] Anonymous. Got flaws?

IEEE MultiMedia, 24 (4):c3, October/December 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
04/mmu20170400c3.pdf.

Anonymous:2017:ICSa

[Ano17m] Anonymous. IEEE Computer Society. IEEE MultiMedia, 24(3):c3, July/ September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
03/mmu20170300c3.pdf.

Anonymous:2017:ICSb

[Ano17n] Anonymous. IEEE Computer Society 2017 call for major award nominations. IEEE MultiMedia, 24(3): c4, July/September 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
03/mmu20170300c4.pdf.

Anonymous:2017:Ma

[Ano17o] Anonymous. myCS. IEEE MultiMedia, 24(1):c4, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        01/mmu20170100c4.pdf.

Anonymous:2017:Mb

[Ano17p] Anonymous. myCS. IEEE MultiMedia, 24(4):c4, October/December 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        04/mmu20170400c4.pdf.

Anonymous:2017:NMO

[Ano17q] Anonymous. New membership options for a better fit. IEEE MultiMedia, 24(1): 1, January/March 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        01/mmu2017010001.pdf.

Anonymous:2017:NBV

[Ano17r] Anonymous. An NFVbased video quality assessment method over 5G small cell networks. IEEE MultiMedia, 24(4):68–78, October/December 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        04/mmu2017040068-abs.
                                        html.

Anonymous:2017:OMU

[Ano17s] Anonymous. One membership. Unlimited knowledge. IEEE MultiMedia, 24(4):c2, October/ December 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
04/mmu20170400c2.pdf.

Anonymous:2017:PC

[Ano17t] Anonymous. Prepare to connect. IEEE MultiMedia, 24(4):67, October/ December 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
04/mmu2017040067.pdf.

Anonymous:2017:RMS

[Ano17u]  Anonymous.  Richard

E. Merwin Scholarship; IEEE Computer Society Richard E. Merwin Student Leadership Scholarship. IEEE MultiMedia, 24(3): c2, July/September 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
03/mmu20170300c2.pdf.

Anonymous:2017:TCa

[Ano17v]  Anonymous.  Table of

contents. IEEE MultiMedia, 24(1):2–3, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
01/mmu2017010002.pdf.

Anonymous:2017:TCb

[Ano17w]  Anonymous.  Table of

contents. IEEE Multi-Media, 24(2):2–3, April/ June 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL https://

www.computer.org/csdl/
mags/mu/2017/02/mmu2017020002.
pdf.

Anonymous:2017:TCc

[Ano17x]  Anonymous.  Table of

contents. IEEE Multi-Media, 24(3):2–3, July/ September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
03/mmu2017030002.pdf.

Anonymous:2017:TCd

[Ano17y]  Anonymous.  Table of

contents. IEEE Multi-Media, 24(4):2–3, October/December 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.

org/csdl/mags/mu/2017/
04/mmu2017040002.pdf.

Anonymous:2017:T

[Ano17z] Anonymous. TechIgnite.

IEEE MultiMedia, 24(1): 47, January/March 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
01/mmu2017010047.pdf.

Anonymous:2017:WFB

[Ano17-27] Anonymous. World’s first 10K best paper award. IEEE MultiMedia, 24(2): c2, April/June 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
02/mmu20170200c2.pdf.

Anonymous:2018:BC

[Ano18a] Anonymous. Back cover.

IEEE MultiMedia, 25(2): c4, April/June 2018. CO-DEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:ISP

[Ano18b] Anonymous. IEEE Security & Privacy. IEEE MultiMedia, 25(3):119, July/ September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:ITB

[Ano18c] Anonymous. IEEE Transactions on Big Data. IEEE MultiMedia, 25(3): 116, July/September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:ITS

[Ano18d] Anonymous. IEEE Transactions on Sustainable Computing. IEEE MultiMedia, 25(3):118, July/ September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:CP

[Ano18e] Anonymous. Call for papers. IEEE MultiMedia, 25(3):117, July/September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:CDD

[Ano18f] Anonymous. COMPSAC 2019 data driven intelligence for a smarter world. IEEE MultiMedia, 25(3): 4, July/September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:CI

[Ano18g] Anonymous. CS information. IEEE MultiMedia, 25(2):c3, April/June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:CLA

[Ano18h] Anonymous. CS library ad.

IEEE MultiMedia, 25(1): 4, January/March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:FCa

[Ano18i] Anonymous. Front cover.

IEEE MultiMedia, 25(1): c1, January/March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:FCb

[Ano18j] Anonymous. Front cover.

IEEE MultiMedia, 25(2): c1, April/June 2018. CO-DEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:FCc

[Ano18k] Anonymous. Front cover.

IEEE MultiMedia, 25(3): C1, July/September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:FCd

[Ano18l] Anonymous. Front cover.

IEEE MultiMedia, 25(4): C1, October/December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:IC

[Ano18m] Anonymous. IEEE COMP-SAC 2019. IEEE Multi-Media, 25(4):C2, October/ December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:ICSb

[Ano18n] Anonymous. IEEE Computer Society. IEEE MultiMedia, 25(3):C3, July/ September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:ICSd

[Ano18o] Anonymous. IEEE Computer Society. IEEE Multi-Media, 25(4):C3, October/ December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:ICSc

[Ano18p] Anonymous. IEEE Computer Society: Be at the center of it all. IEEE MultiMedia, 25(3):C4, July/ September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:ICSa

[Ano18q] Anonymous. IEEE Computer Society information. IEEE MultiMedia, 25(1): c3, January/March 2018.

CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:ILCa

[Ano18r] Anonymous. IEEE letters of the Computer Society. IEEE MultiMedia, 25(3): 115, July/September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:ILCb

[Ano18s] Anonymous. IEEE letters of the Computer Society. IEEE MultiMedia, 25(4):C4, October/ December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:IAa

[Ano18t] Anonymous. Interface ad.

IEEE MultiMedia, 25(1): 96, January/March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:IAb

[Ano18u] Anonymous. Interface ad.

IEEE MultiMedia, 25(2): 104, April/June 2018. CO-DEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:JBA

[Ano18v] Anonymous. Jobs board ad. IEEE MultiMedia, 25(1):c4, January/March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:LBTa

[Ano18w] Anonymous. Looking for the BEST tech job for you? IEEE MultiMedia, 25(3): 114, July/September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:LBTb

[Ano18x] Anonymous. Looking for the BEST tech job for you? IEEE Multi-Media, 25(4):84, October/ December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:Mb

[Ano18y] Anonymous. [Masthead]. IEEE MultiMedia, 25(1): 3, January/March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:Mc

[Ano18z] Anonymous. [Masthead]. IEEE MultiMedia, 25(2):4, April/June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:Md

[Ano18-27] Anonymous. Masthead. IEEE MultiMedia, 25(3): 3, July/September 2018.

CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:Me

[Ano18-28] Anonymous. Masthead.

IEEE MultiMedia, 25(4):3, October/December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:MAA

[Ano18-29]  Anonymous.  Merwin

Award ad. IEEE Multi-Media, 25(2):1, April/June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:MA

[Ano18-30] Anonymous. myCS ad.

IEEE MultiMedia, 25(1): 95, January/March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:OMUa

[Ano18-31] Anonymous. One membership. Unlimited knowledge. IEEE MultiMedia, 25(3): 120, July/September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:OMUb

[Ano18-32] Anonymous. One membership. Unlimited knowledge. IEEE MultiMedia, 25(4):4, October/December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:SNAa

[Ano18-33] Anonymous. Social networking ad. IEEE MultiMedia, 25(2):c2, April/ June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:SNAb

[Ano18-34] Anonymous. Social networking ad. IEEE MultiMedia, 25(3):C2, July/ September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:SMA

[Ano18-35] Anonymous. Student membership ad. IEEE Multi-Media, 25(1):c2, January/ March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:TCa

[Ano18-36] Anonymous. Table of contents. IEEE MultiMedia, 25(1):1–2, January/ March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:TCb

[Ano18-37] Anonymous. Table of contents. IEEE MultiMedia, 25(2):2–3, April/June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:TCc

[Ano18-38]  Anonymous.  Table of

contents. IEEE Multi-Media, 25(3):1–2, July/ September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2018:TCd

[Ano18-39] Anonymous. Table of contents. IEEE MultiMedia, 25(4):1–2, October/ December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:CEa

[Ano19a] Anonymous. Computing Edge. IEEE MultiMedia, 26(2):4, April/June 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:CEb

[Ano19b] Anonymous. Computing Edge. IEEE MultiMedia, 26(3):C2, July/September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:CEc

[Ano19c] Anonymous. Computing Edge. IEEE MultiMedia, 26(4):C2, October/ December 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:CMA

[Ano19d] Anonymous. Call for 2019 major awards nominations. IEEE MultiMedia, 26(3): 92, July/September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:CAN

[Ano19e] Anonymous. Call for award nominations. IEEE Multi-Media, 26(1):C4, January/ March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:CPa

[Ano19f] Anonymous. Call for papers. IEEE MultiMedia, 26(3):C4, July/September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:CPP

[Ano19g] Anonymous. Call for papers and proposals. IEEE MultiMedia, 26(4):80, October/December 2019. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:CPI

[Ano19h] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE MultiMedia, 26(4):C4, October/December 2019. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:CIa

[Ano19i] Anonymous. Connect on interface. IEEE Multi-Media, 26(1):C2, January/ March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:CIb

[Ano19j] Anonymous. Connect on interface. IEEE Multi-Media, 26(1):59, January/ March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:C

[Ano19k]  Anonymous.  Cover 3.

IEEE MultiMedia, 26(3): C3, July/September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:FCa

[Ano19l] Anonymous. Front cover.

IEEE MultiMedia, 26(1): C1, January/March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:FCb

[Ano19m] Anonymous. Front cover. IEEE MultiMedia, 26(2): C1, April/June 2019. CO-DEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:FCc

[Ano19n] Anonymous. Front cover. IEEE MultiMedia, 26(3): C1, July/September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:FCd

[Ano19o] Anonymous. [Front cover]. IEEE MultiMedia, 26(4): C1, October/December 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:ICSa

[Ano19p] Anonymous. IEEE Computer Society. IEEE Multi-Media, 26(1):C3, January/ March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:ICSb

[Ano19q] Anonymous. IEEE Computer Society. IEEE MultiMedia, 26(2):C3, April/ June 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:ICS

[Ano19r] Anonymous. IEEE Computer Society. IEEE Multi-Media, 26(4):C3, October/ December 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:ILC

[Ano19s] Anonymous. IEEE letters of the Computer Society. IEEE MultiMedia, 26(1): 4, January/March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:IWC

[Ano19t] Anonymous. IEEE World Congress on Services 2019. IEEE MultiMedia, 26(2): C4, April/June 2019. CO-DEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:KYCa

[Ano19u] Anonymous. Keep your career options open: Upload your resume today! IEEE MultiMedia, 26(2): C2, April/June 2019. CO-DEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:KYCb

[Ano19v] Anonymous. Keep your career options open upload your resume today! IEEE MultiMedia, 26(3):4, July/ September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:KYCc

[Ano19w] Anonymous. Keep your career options open upload your resume today! IEEE MultiMedia, 26(4):4, October/December 2019. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:LBT

[Ano19x] Anonymous. Looking for the BEST tech job for you? IEEE MultiMedia, 26(1): 60, January/March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:Ma

[Ano19y] Anonymous. Masthead. IEEE MultiMedia, 26(1): 3, January/March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:Mb

[Ano19z] Anonymous. Masthead. IEEE MultiMedia, 26(2):3, April/June 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:Mc

[Ano19-27] Anonymous. Masthead. IEEE MultiMedia, 26(3):

3, July/September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:Md

[Ano19-28] Anonymous. Masthead.

IEEE MultiMedia, 26(4):3, October/December 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:ONW

[Ano19-29]  Anonymous.  Our new

website has launched! IEEE MultiMedia, 26(2): 92, April/June 2019. CO-DEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:TCa

[Ano19-30] Anonymous. Table of contents. IEEE MultiMedia, 26(1):1–2, January/ March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:TCb

[Ano19-31] Anonymous. Table of contents. IEEE MultiMedia, 26(2):1–2, April/June 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Anonymous:2019:TCc

[Ano19-32] Anonymous. Table of contents. IEEE MultiMedia, 26(3):1–2, July/September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2019:TCd

[Ano19-33] Anonymous. Table of contents. IEEE MultiMedia, 26(4):1–2, October/ December 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:CEa

[Ano20a] Anonymous. Computing Edge. IEEE MultiMedia, 27(1):C2, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:Ca

[Ano20b] Anonymous. ComputingEdge. IEEE MultiMedia, 27(3): C4, July/September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:Cb

[Ano20c] Anonymous. ComputingEdge. IEEE MultiMedia, 27(4): C4, October/December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:CPPa

[Ano20d] Anonymous. Call for papers and proposals. IEEE MultiMedia, 27(1):4, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:CPPb

[Ano20e] Anonymous. Call for papers and proposals. IEEE MultiMedia, 27(1):4, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:CPIa

[Ano20f] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE MultiMedia, 27(1):C4, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:CPIb

[Ano20g] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE MultiMedia, 27(1):C4, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:CPIc

[Ano20h] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE MultiMedia, 27(2):64, April/ June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:CEb

[Ano20i] Anonymous. Computing edge. IEEE MultiMedia, 27(1):C2, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:ELS

[Ano20j]  Anonymous.  Erratum

to leveraging smart devices for scene text preserved image stylization: a deep gaming approach. IEEE MultiMedia, 27(3): 99, July/September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:ECO

[Ano20k] Anonymous. Evolving career opportunities need your skills. IEEE MultiMedia, 27(3):4, July/ September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:FCa

[Ano20l] Anonymous. Front cover. IEEE MultiMedia, 27(1): C1, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:FCb

[Ano20m] Anonymous. Front cover. IEEE MultiMedia, 27(1): C1, January/March 2020.

CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:FCc

[Ano20n] Anonymous. Front cover.

IEEE MultiMedia, 27(2): C1, April/June 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:FCd

[Ano20o] Anonymous. Front cover.

IEEE MultiMedia, 27(3): C1, July/September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:FCe

[Ano20p] Anonymous. Front cover.

IEEE MultiMedia, 27(4): C1, October/December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:GPNa

[Ano20q] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE MultiMedia, 27(3):100, July/September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:GPNb

[Ano20r] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE MultiMedia, 27(4):4, October/ December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:HWC

[Ano20s]  Anonymous.  Honor-

ing women in computing: Call for nominations. IEEE MultiMedia, 27(2):4, April/June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:Ha

[Ano20t] Anonymous. HOST 2020. IEEE MultiMedia, 27(2): 62, April/June 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:Hb

[Ano20u] Anonymous. HOST 2020. IEEE MultiMedia, 27(4): C2, October/December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:ICSa

[Ano20v] Anonymous. IEEE Computer Society. IEEE Multi-Media, 27(1):C3, January/ March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:ICSb

[Ano20w] Anonymous. IEEE Computer Society. IEEE Multi-Media, 27(1):C3, January/ March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:ICSe

[Ano20x] Anonymous. IEEE Computer Society. IEEE MultiMedia, 27(2):C3, April/ June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:ICSg

[Ano20y] Anonymous. IEEE Computer Society. IEEE MultiMedia, 27(3):C3, July/ September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:ICSi

[Ano20z] Anonymous. IEEE Computer Society. IEEE Multi-Media, 27(4):C3, October/ December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:ICSc

[Ano20-27] Anonymous. IEEE Computer Society has you covered! IEEE MultiMedia, 27(2):C2, April/June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:ICSf

[Ano20-28] Anonymous. IEEE Computer Society has you covered! IEEE MultiMedia, 27(3):C2, July/September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:ICSh

[Ano20-29] Anonymous. IEEE Computer Society has you covered! IEEE MultiMedia, 27(4):144, October/ December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:ICSd

[Ano20-30] Anonymous. IEEE Computer Society Volunteer Service Awards. IEEE MultiMedia, 27(2):63, April/ June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:IQW

[Ano20-31] Anonymous. IEEE Quantum Week. IEEE MultiMedia, 27(2):C4, April/June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:Ma

[Ano20-32] Anonymous. Masthead. IEEE MultiMedia, 27(1):3, January/March 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:Mb

[Ano20-33] Anonymous. Masthead.

IEEE MultiMedia, 27(1):3, January/March 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:Mc

[Ano20-34] Anonymous. Masthead.

IEEE MultiMedia, 27(2):3, April/June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:Md

[Ano20-35] Anonymous. Masthead.

IEEE MultiMedia, 27(3): 3, July/September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:Me

[Ano20-36] Anonymous. Masthead.

IEEE MultiMedia, 27(4):3, October/December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:TCa

[Ano20-37] Anonymous. Table of contents. IEEE MultiMedia, 27(1):1–2, January/March 2020. CODEN IEMUE4.

ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:TCb

[Ano20-38] Anonymous. Table of contents. IEEE MultiMedia, 27(1):1–2, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:TCc

[Ano20-39] Anonymous. Table of contents. IEEE MultiMedia, 27(2):1–2, April/June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:TCd

[Ano20-40] Anonymous. Table of contents. IEEE MultiMedia, 27(3):1–2, July/September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2020:TCe

[Ano20-41] Anonymous. Table of contents. IEEE MultiMedia, 27(4):1–2, October/ December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:CSEa

[Ano21a] Anonymous. Computing in Science & Engineering. IEEE MultiMedia, 28(2): 107, April/June 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:CSEb

[Ano21b] Anonymous. Computing in Science & Engineering: Subscribe to CiSE today! IEEE Multi-Media, 28(4):4, October/ December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:Ca

[Ano21c] Anonymous. ComputingEdge.

IEEE MultiMedia, 28(2): C2, April/June 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:Cb

[Ano21d] Anonymous. ComputingEdge.

IEEE MultiMedia, 28(3): C2, July/September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:Cc

[Ano21e] Anonymous. ComputingEdge.

IEEE MultiMedia, 28(4): C2, October/December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:CYO

[Ano21f] Anonymous. ComputingEdge:

Your one stop resource for industry hot topics, technical overviews, and in depth articles. IEEE MultiMedia, 28(1):C2, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ICAa

[Ano21g] Anonymous. IEEE Computer Architecture Letters. IEEE MultiMedia, 28(1): 103, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ICAb

[Ano21h] Anonymous. IEEE Computer Architecture Letters. IEEE MultiMedia, 28(2):4, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ICAc

[Ano21i] Anonymous. IEEE Computer Architecture Letters. IEEE MultiMedia, 28(3): 128, July/September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:CMA

[Ano21j] Anonymous. Call for 2021 major awards nominations. IEEE MultiMedia, 28(2): 106, April/June 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:EAN

[Ano21k]  Anonymous.  Educa-

tion awards nominations. IEEE MultiMedia, 28(2): 108, April/June 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ECO

[Ano21l] Anonymous. Evolving career opportunities need your skills explore new options upload your resume today. IEEE MultiMedia, 28(1):4, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:FCa

[Ano21m] Anonymous. [Front cover].

IEEE MultiMedia, 28(1): C1, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:FCb

[Ano21n] Anonymous. [Front cover].

IEEE MultiMedia, 28(2): C1, April/June 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:FCc

[Ano21o] Anonymous. [Front cover].

IEEE MultiMedia, 28(3): C1, July/September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:FCd

[Ano21p] Anonymous. [Front cover]. IEEE MultiMedia, 28(4): C1, October/December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ICSa

[Ano21q] Anonymous. IEEE Computer Society. IEEE MultiMedia, 28(1):104, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ICSb

[Ano21r] Anonymous. IEEE Computer Society. IEEE Multi-Media, 28(1):C3, January/ March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ICSd

[Ano21s] Anonymous. IEEE Computer Society. IEEE MultiMedia, 28(2):C3, April/ June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ICSg

[Ano21t] Anonymous. IEEE Computer Society. IEEE MultiMedia, 28(3):C3, July/ September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ICSi

[Ano21u] Anonymous. [IEEE Computer Society]. IEEE MultiMedia, 28(4):C3, October/December 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ICSc

[Ano21v] Anonymous. IEEE Computer Society has you covered! IEEE MultiMedia, 28(1):C4, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ICSh

[Ano21w] Anonymous. IEEE Computer Society has you covered! IEEE MultiMedia, 28(3):C4, July/September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ICSe

[Ano21x] Anonymous. IEEE Computer Society Jobs Board. IEEE MultiMedia, 28(2): C4, April/June 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ICSf

[Ano21y] Anonymous. IEEE Computer Society Jobs Board. IEEE MultiMedia, 28(3): 4, July/September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:ICSj

[Ano21z] Anonymous. IEEE Computer Society Jobs Board. IEEE MultiMedia, 28(4): C4, October/December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:Ma

[Ano21-27] Anonymous. Masthead. IEEE MultiMedia, 28(1):3, January/March 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:Mb

[Ano21-28] Anonymous. Masthead. IEEE MultiMedia, 28(2):3, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:Mc

[Ano21-29] Anonymous. Masthead. IEEE MultiMedia, 28(3): 3, July/September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:Md

[Ano21-30] Anonymous. Masthead. IEEE MultiMedia, 28(4):3, October/December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:TCa

[Ano21-31] Anonymous. Table of contents. IEEE MultiMedia, 28(1):1–2, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:TCb

[Ano21-32] Anonymous. Table of contents. IEEE MultiMedia, 28(2):1–2, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:TCc

[Ano21-33] Anonymous. Table of contents. IEEE MultiMedia, 28(3):1–2, July/September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2021:TCd

[Ano21-34] Anonymous. Table of contents. IEEE MultiMedia, 28(4):1–2, October/ December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:A

[Ano22a]  Anonymous.  Awards.

IEEE MultiMedia, 29(2): 124, April/June 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:CSEa

[Ano22b] Anonymous. Computing in Science & Engineering. IEEE MultiMedia, 29(2): 22, April/June 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:CSEb

[Ano22c] Anonymous. Computing in Science & Engineering. IEEE MultiMedia, 29(3): 100, July/September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:CSEc

[Ano22d] Anonymous. Computing in Science & Engineering. IEEE MultiMedia, 29(4): 127, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:Ca

[Ano22e] Anonymous. ComputingEdge. IEEE MultiMedia, 29(1): C2, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:Cb

[Ano22f] Anonymous. ComputingEdge. IEEE MultiMedia, 29(2): C2, April/June 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:Cc

[Ano22g] Anonymous. ComputingEdge.

IEEE MultiMedia, 29(3): C2, July/September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:IAHa

[Ano22h] Anonymous. IEEE Annals of the History of Computing. IEEE MultiMedia, 29(2):7, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:IAHb

[Ano22i] Anonymous. IEEE Annals of the History of Computing. IEEE MultiMedia, 29(3):68, July/September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:IAHc

[Ano22j] Anonymous. IEEE Annals of the History of Computing. IEEE MultiMedia, 29 (4):96, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICGa

[Ano22k] Anonymous. IEEE Computer Graphics and Applications. IEEE MultiMedia, 29(2):10, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICGb

[Ano22l] Anonymous. IEEE Computer Graphics and Applications. IEEE MultiMedia, 29(3):90, July/September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICGc

[Ano22m] Anonymous. IEEE Computer Graphics and Applications. IEEE MultiMedia, 29(4):122, October/ December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:IISb

[Ano22n] Anonymous. IEEE Intelligent Systems. IEEE Multi-Media, 29(2):4, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:IISa

[Ano22o] Anonymous. IEEE Intelligent Systems magazine’s AI’s 10 to watch. IEEE MultiMedia, 29(1):4, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:CAa

[Ano22p] Anonymous. Call for articles. IEEE MultiMedia, 29(2):55, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:CAb

[Ano22q] Anonymous. Call for articles. IEEE MultiMedia, 29(3):111, July/September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:CAIa

[Ano22r] Anonymous. Call for articles: IEEE Pervasive Computing. IEEE MultiMedia, 29(2):113, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:CAIb

[Ano22s] Anonymous. Call for articles: IEEE Pervasive Computing. IEEE MultiMedia, 29(3):7, July/September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:CPIa

[Ano22t] Anonymous. Call for papers: IEEE Computer Society. IEEE MultiMedia, 29(3):37, July/September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:CPIb

[Ano22u] Anonymous. Call for papers: IEEE Computer Society. IEEE MultiMedia, 29 (4):13, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:CPIc

[Ano22v]  Anonymous.  Call for

papers: IT-general-CFP-hHalf (2). IEEE Multi-Media, 29(4):23, October/ December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:CE

[Ano22w] Anonymous. Computing edge. IEEE MultiMedia, 29 (4):C2, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:FCa

[Ano22x] Anonymous. Front cover. IEEE MultiMedia, 29(1): C1, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:FCb

[Ano22y] Anonymous. Front cover. IEEE MultiMedia, 29(2): C1, April/June 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:FCc

[Ano22z] Anonymous. Front cover. IEEE MultiMedia, 29(3): C1, July/September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:FCd

[Ano22-27] Anonymous. Front cover.

IEEE MultiMedia, 29(4): C1, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICSa

[Ano22-28] Anonymous. IEEE Computer Society. IEEE Multi-Media, 29(1):C3, January/ March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICSc

[Ano22-29] Anonymous. IEEE Computer Society. IEEE MultiMedia, 29(2):C3, April/ June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICSg

[Ano22-30] Anonymous. IEEE Computer Society. IEEE MultiMedia, 29(3):C3, July/ September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICSf

[Ano22-31] Anonymous. IEEE Computer Society D&I fund. IEEE MultiMedia, 29(3): 112, July/September 2022.

CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICSe

[Ano22-32] Anonymous. IEEE Computer Society has you covered! IEEE MultiMedia, 29(3):4, July/September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICSi

[Ano22-33] Anonymous. IEEE Computer Society has you covered! IEEE MultiMedia, 29(4):128, October/ December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICSj

[Ano22-34] Anonymous. IEEE Computer Society information. IEEE MultiMedia, 29 (4):C3, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICSb

[Ano22-35] Anonymous. IEEE Computer Society Jobs Board. IEEE MultiMedia, 29(1): C4, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICSd

[Ano22-36] Anonymous. IEEE Computer Society Jobs Board. IEEE MultiMedia, 29(2): C4, April/June 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICSh

[Ano22-37] Anonymous. IEEE Computer Society Jobs Board. IEEE MultiMedia, 29(3): C4, July/September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:ICSk

[Ano22-38] Anonymous. IEEE Computer Society Jobs Board. IEEE MultiMedia, 29(4): C4, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:Ma

[Ano22-39] Anonymous. Masthead.

IEEE MultiMedia, 29(1):3, January/March 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:Mb

[Ano22-40] Anonymous. Masthead.

IEEE MultiMedia, 29(2):3, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:Mc

[Ano22-41] Anonymous. Masthead. IEEE MultiMedia, 29(3): 3, July/September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:Md

[Ano22-42] Anonymous. Masthead. IEEE MultiMedia, 29(4):3, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:RCS

[Ano22-43] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE MultiMedia, 29(4):110, October/December 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:TCa

[Ano22-44] Anonymous. Table of contents. IEEE MultiMedia, 29(1):1–2, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:TCb

[Ano22-45] Anonymous. Table of contents. IEEE MultiMedia, 29(2):1–2, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:TCc

[Ano22-46] Anonymous. Table of contents. IEEE MultiMedia, 29(3):1–2, July/September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2022:TCd

[Ano22-47] Anonymous. Table of contents. IEEE MultiMedia, 29(4):1–2, October/ December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICAa

[Ano23a] Anonymous. 2023 IEEE conference on artificial intelligence. IEEE Multi-Media, 30(1):C2, January/ March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:GI

[Ano23b] Anonymous. AN.general hHalf jz.indd.

IEEE MultiMedia, 30(2): 80, April/June 2023. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:CS

[Ano23c]  Anonymous.  Computer

seeks 2025–2026. IEEE MultiMedia, 30(4):4, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:CE

[Ano23d] Anonymous. Computing Edge. IEEE MultiMedia, 30(4):C4, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:CSEa

[Ano23e] Anonymous. Computing in Science & Engineering. IEEE MultiMedia, 30(1):6, January/March 2023. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:CSEc

[Ano23f] Anonymous. Computing in Science & Engineering. IEEE MultiMedia, 30(3):7, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:CSEb

[Ano23g] Anonymous. Computing in Science & Engineering. IEEE MultiMedia, 30 (4):118, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:Ca

[Ano23h] Anonymous. ComputingEdge. IEEE MultiMedia, 30(1): C4, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:Cb

[Ano23i] Anonymous. ComputingEdge.

IEEE MultiMedia, 30(3): C4, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:IAHa

[Ano23j] Anonymous. IEEE Annals of the History of Computing. IEEE MultiMedia, 30(1):16, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:IAHc

[Ano23k] Anonymous. IEEE Annals of the History of Computing. IEEE MultiMedia, 30(3):36, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:IAHb

[Ano23l] Anonymous. IEEE Annals of the History of Computing. IEEE MultiMedia, 30(4):35, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICAb

[Ano23m] Anonymous. IEEE Computer Architecture Letters. IEEE MultiMedia, 30(4):119, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICGa

[Ano23n] Anonymous. IEEE Computer Graphics and Applications. IEEE MultiMedia, 30(1):111, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICGb

[Ano23o] Anonymous. IEEE Computer Graphics and Applications. IEEE MultiMedia, 30(4):90, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICE

[Ano23p] Anonymous. IEEE Computing Edge. IEEE MultiMedia, 30(2):C4, April/ June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:IOJ

[Ano23q] Anonymous. IEEE Open Journal of the Computer Society. IEEE MultiMedia, 30(4):120, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ITS

[Ano23r] Anonymous. IEEE Transactions on Sustainable Computing. IEEE MultiMedia, 30(3):140, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:CMA

[Ano23s]  Anonymous.  Call-for-

major-awards-2023-full-ml.indd. IEEE MultiMedia, 30(2):4, April/June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:CGH

[Ano23t] Anonymous. CGA-general-half-2021.indd. IEEE MultiMedia, 30(2):90, April/ June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:CGF

[Ano23u] Anonymous. CS-generic-full-km-2023.indd. IEEE MultiMedia, 30(2):C2, April/ June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:DDIb

[Ano23v] Anonymous. Drive diversity & inclusion in computing. IEEE MultiMedia, 30(4):C2, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:DDIa

[Ano23w] Anonymous. Drive diversity & inclusion in computing. IEEE MultiMedia, 30(2):91, April/June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:DDIc

[Ano23x] Anonymous. Drive diversity & inclusion in computing. IEEE MultiMedia, 30(3):C2, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:FCa

[Ano23y] Anonymous. Front cover. IEEE MultiMedia, 30(1): C1, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:FCb

[Ano23z] Anonymous. Front cover. IEEE MultiMedia, 30(2): C1, April/June 2023. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:FCd

[Ano23-27] Anonymous. Front cover. IEEE MultiMedia, 30(3): C1, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:FCc

[Ano23-28] Anonymous. Front cover. IEEE MultiMedia, 30(4): C1, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:HA

[Ano23-29] Anonymous. Humphrey Award. IEEE MultiMedia, 30(1):72, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICSe

[Ano23-30] Anonymous. IEEE Computer Society — call for papers. IEEE MultiMedia, 30(4):15, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICSh

[Ano23-31] Anonymous. IEEE Computer Society CG&A. IEEE MultiMedia, 30 (3):62, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICSa

[Ano23-32] Anonymous. IEEE Computer Society has you covered! IEEE Multi-Media, 30(1):4, January/ March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICSb

[Ano23-33] Anonymous. IEEE Computer Society has you covered! IEEE Multi-Media, 30(1):35, January/ March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICSg

[Ano23-34] Anonymous. IEEE Computer Society has you covered! IEEE MultiMedia, 30(3):4, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICSc

[Ano23-35] Anonymous. IEEE Computer Society information. IEEE MultiMedia, 30(1): C3, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICSd

[Ano23-36] Anonymous. IEEE Computer Society information. IEEE MultiMedia, 30(2): C3, April/June 2023. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICSi

[Ano23-37] Anonymous. IEEE Computer Society information. IEEE MultiMedia, 30(3):C3, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:ICSf

[Ano23-38] Anonymous. IEEE Computer Society information. IEEE MultiMedia, 30(4):C3, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:IQW

[Ano23-39] Anonymous. IEEE quantum week. IEEE MultiMedia, 30(1):112, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:II

[Ano23-40] Anonymous. IEEE OA OnePageAD.indd.

IEEE MultiMedia, 30(2): 92, April/June 2023. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:Ma

[Ano23-41] Anonymous. Masthead.

IEEE MultiMedia, 30(1):3, January/March 2023. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:Mb

[Ano23-42] Anonymous. Masthead.

IEEE MultiMedia, 30(2):3, April/June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:Md

[Ano23-43] Anonymous. Masthead. IEEE MultiMedia, 30(3):3, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:Mc

[Ano23-44] Anonymous. Masthead. IEEE MultiMedia, 30(4):3, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:RPH

[Ano23-45] Anonymous. over-rainbow-podcast-hHoriz-ks.indd. IEEE MultiMedia, 30(2):37, April/ June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:RCS

[Ano23-46] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE MultiMedia, 30(1):90, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:TCa

[Ano23-47] Anonymous. Table of contents. IEEE MultiMedia, 30(1):1–2, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:TCb

[Ano23-48] Anonymous. Table of contents. IEEE MultiMedia, 30(2):1–2, April/June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:TCd

[Ano23-49] Anonymous. Table of contents. IEEE MultiMedia, 30(3):1–2, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2023:TCc

[Ano23-50] Anonymous. Table of contents. IEEE MultiMedia, 30(4):1–2, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ICE

[Ano24a] Anonymous. IEEE Computing Edge. IEEE MultiMedia, 31(1):C4, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:IOJ

[Ano24b] Anonymous. IEEE Open Journal of the Computer Society. IEEE MultiMedia, 31(1):112, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ITB

[Ano24c] Anonymous. IEEE Transactions on Big Data. IEEE MultiMedia, 31 (1):32, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ITCa

[Ano24d] Anonymous. IEEE Transactions on Computers. IEEE MultiMedia, 31(1):4, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ITCb

[Ano24e] Anonymous. IEEE Transactions on Computers. IEEE MultiMedia, 31(2): 60, April/June 2024. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ITP

[Ano24f] Anonymous. IEEE Transactions on Privacy. IEEE MultiMedia, 31(1):109, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ITSa

[Ano24g] Anonymous. IEEE Transactions on Sustainable Computing. IEEE MultiMedia, 31(1):110, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ITSb

[Ano24h] Anonymous. IEEE Transactions on Sustainable Computing. IEEE MultiMedia, 31(2):68, April/ June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:CAO

[Ano24i] Anonymous. Career accelerating opportunities. IEEE MultiMedia, 31(2): C3, April/June 2024. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:CE

[Ano24j] Anonymous. Computing edge. IEEE MultiMedia, 31(2):C4, April/June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:FCa

[Ano24k] Anonymous. Front cover.

IEEE MultiMedia, 31(1): C1, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:FCb

[Ano24l] Anonymous. Front cover.

IEEE MultiMedia, 31(2): C1, April/June 2024. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:GPN

[Ano24m] Anonymous. Get published in the new IEEE Transactions on Privacy. IEEE MultiMedia, 31(2): C2, April/June 2024. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ICSg

[Ano24n] Anonymous. IEEE Computer Society. IEEE MultiMedia, 31(2):80, April/ June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ICSb

[Ano24o] Anonymous. IEEE Computer Society — call for papers. IEEE MultiMedia, 31(1):6, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ICSf

[Ano24p] Anonymous. IEEE Computer Society: Call for papers. IEEE MultiMedia, 31(2):25, April/June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ICSd

[Ano24q] Anonymous. IEEE Computer Society career center. IEEE MultiMedia, 31(1):C3, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ICSa

[Ano24r] Anonymous. IEEE Computer Society D&I fund. IEEE MultiMedia, 31(1): C2, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ICSe

[Ano24s] Anonymous. IEEE Computer Society has you covered! IEEE MultiMedia, 31(2):4, April/June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:ICSc

[Ano24t] Anonymous. IEEE Computer Society information. IEEE MultiMedia, 31(1):111, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:Ma

[Ano24u] Anonymous. Masthead.

IEEE MultiMedia, 31(1):3, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:Mb

[Ano24v] Anonymous. Masthead.

IEEE MultiMedia, 31(2):3, April/June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:TCa

[Ano24w] Anonymous. Table of contents. IEEE MultiMedia, 31(1):1–2, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anonymous:2024:TCb

[Ano24x] Anonymous. Table of contents. IEEE MultiMedia, 31(2):1–2, April/June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Aizawa:2015:FMT

[AO15] Kiyoharu Aizawa and Makoto Ogawa. Food-Log: Multimedia tool for healthcare applications. IEEE MultiMedia, 22(2):4– 8, April/June 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/02/mmu2015020004-
abs.html.

Assfalg:2000:QPV

[AP00] Jürgen Assfalg and Pietro Pala. Querying by photographs: a VR metaphor for image retrieval. IEEE MultiMedia, 7(1):52–59, January/March 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u1052.
pdf;   http://www.computer.
org/multimedia/mu2000/
u1052abs.htm.

Abdallah:2023:RAI

[APCH23] Maha Abdallah, Balakrishnan Prabhakaran, Wei Cai, and Cheng-Hsin Hsu. Recent advances in immersive multimedia. IEEE MultiMedia, 30(2):5–7, April/ June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Alvarez:2018:BAT

[APS^{+}18] F. Alvarez, M. Popa, V. Solachidis, G. Hernández-Peñaloza, A. Belmonte-Hernández, S. Asteriadis, N. Vretos, M. Quintana, T. Theodoridis, D. Dotti, and P. Daras. Behavior analysis through multimodal sensing for care of Parkinson’s and Alzheimer’s patients. IEEE MultiMedia, 25(1):14–25, January/March 2018. CO-DEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Ashdown:2005:EPP

[AR05] Mark Ashdown and Peter Robinson. Escritoire: a personal projected display. IEEE MultiMedia, 12 (1):34–42, January/March 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Arabnia:1999:RIV

[Ara99a] H. Arabnia. Reading in information visualization: using vision to think [media review]. IEEE MultiMedia, 6(4):93, October/ December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Arabnia:1999:MRR

[Ara99b]  Hamid Arabnia.  Me-

dia reviews: Readings In Information Visualization: Using Vision to Think by Stuart K. Card, Jock D. Mackinlay, and Ben Shneiderman (Morgan Kaufmann Publishers, San Francisco, Calif., 1999, $69.95, 703 pp., ISBN 1-55860-533-9). IEEE MultiMedia, 6(4):93, October/ December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u4091.
                                        pdf.

Anand:2020:JWE

[AS20] Ashima Anand and Amit Kumar Singh. Joint watermarking– encryption–ECC for patient record security in wavelet domain. IEEE MultiMedia, 27(3):66–75, July/September 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Al-Salqan:1996:TRS

[ASC96] Yahya Y. Al-Salqan and Carl K. Chang. Temporal relations and synchronization agents. IEEE MultiMedia, 3(2):30–39, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2030.
pdf;   http://www.computer.
org/multimedia/mu1996/
u2030abs.htm.

A:2024:HAD

[ASL^{+}24] Ruhan A, Danyao Shen, Lijing Liu, Juanjuan Yin, and Renpu Lin. Hyperspectral anomaly detection based on a beta wavelet graph neural network. IEEE MultiMedia, 31(2):69–79, April/June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anand:2020:CTE

[ASLB20] Ashima Anand, Amit Kumar Singh, Zhihan Lv, and Guarav Bhatnagar. Compression-then-encryption-based secure watermarking technique for smart healthcare system. IEEE Multi-Media, 27(4):133–143, October/December 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Aggoun:2013:IHV

[ATS^{+}13] Amar Aggoun, Emmanuel Tsekleves, Mohammad Rafiq Swash, Dimitrios Zarpalas, Anastasios Dimou, Petros Daras, Paulo Nunes, and Lui Ducla Soares. Immersive 3D holoscopic video system. IEEE MultiMedia, 20(1):28–37, January/ March 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Autry:2011:RWM

[Aut11] Christopher Autry. The revolution will be mobilized. IEEE MultiMedia, 18(2):85–87, April/June 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Adamo-Villani:2008:TNT

[AVW08] Nicoletta Adamo-Villani and Ronnie Wilbur. Two novel technologies for accessible math and science education. IEEE MultiMedia, 15(4):38–46, October/ December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Achanta:2006:MIH

[AYK06] Radhakrishna S. V. Achanta, Wei-Qi Yan, and Mohan S. Kankanhalli. Modeling intent for home video repurposing. IEEE Multi-Media, 13(1):46–55, January/March 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Anthony:2008:TNG

[AYK08] Lisa Anthony, Jie Yang, and Kenneth R. Koedinger. Toward next-generation, intelligent tutors: Adding natural handwriting input. IEEE MultiMedia, 15 (3):64–68, July/September 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Baker:2006:MII

[Bak06] Keith Baker. Media impact: Intrusive interactivity is not an ambient experience. IEEE MultiMedia, 13(2):4–7, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Baker:2007:MIT

[Bak07] Keith Baker. Media impact: Torque kills! future control of the ambient electromagnetic spectrum. IEEE MultiMedia, 14(1): 4–8, January/March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Baldwin:1998:NGR

[Bal98] Robin Baldwin. In the news: Getting ready for MPEG-4; business notes; XML, HTML 4.0 extend the Web; proposed solution for digital content copyright protection. IEEE MultiMedia, 5(1):6–7, January/March 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u1006.
pdf.

Baldwin:1999:NOA

[Bal99a] Robin Baldwin. In the news: Online auctions: Just another fad? IEEE MultiMedia, 6(3):12–13, July/September 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u3012.
pdf.

Baldwin:1999:NRH

[Bal99b] Robin Baldwin. In the news: Rx for health sites and digital wallets: Poised for success tech briefs and privacy concerns; tech briefs. IEEE MultiMedia, 6(4):8–9, October/ December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u4008.
pdf.

Baltes:2001:NVR

[Bal01] Cheryl Baltes. In the news: The voice recognition promise; Web resources; E-books future?; security standards initiatives; tech news. IEEE MultiMedia, 8(1):11–13, January/March 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u1011.
pdf.

Banerjea:1997:PRH

[Ban97] Anindo Banerjea. Project reports: Heterogeneous networking. IEEE Multi-Media, 4(2):84–87, April/ June 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2084.
pdf.

Bayle:1997:MRK

[Bay97]  Aime J. Bayle.  Me-

dia reviews: Keeping up with Java: Java by Example by J. R. Jackson and A. L. McClellan (SunSoft Press, 1997, 426 pp., $34.95, ISBN 0-13-272295-X). IEEE MultiMedia, 4(3):91, July/ September 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3091.
pdf.

Becker:2002:RAD

[BB02] Shirley Ann Becker and Anthony Berkemeyer. Rapid application design and testing of Web usability. IEEE MultiMedia, 9(4):38–46, October/ December 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
04/u4038abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/04/u4038.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/04/
u4038.pdf.

Bianchi-Berthouze:2003:KDA

[BB03] Nadia Bianchi-Berthouze. K-DIME: An affective image filtering system. IEEE MultiMedia, 10(3):103– 106, July/September 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/dl/mags/mu/2003/03/
u3103.htm;   http://csdl.

computer.org/dl/mags/
mu/2003/03/u3103.pdf.

Bourbakis:2004:GEI

[BB04] Nikolaos Bourbakis and George Bebis. Guest Editors’ introduction: Digital multimedia on demand. IEEE MultiMedia, 11(2): 14–15, April/June 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
02/u2014.pdf;   http:/
/csdl.computer.org/dl/
mags/mu/2004/02/u2014.
htm.

Barni:2000:IPV

[BBC00] Mauro Barni, Franco Bartolini, and Vito Cappellini. Image processing for virtual restoration of artworks. IEEE MultiMedia, 7(2):34–37, April/ June 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2034.
pdf;   http://www.computer.
org/multimedia/mu2000/
u2034abs.htm.

Bellini:2012:MGS

[BBCN12] Pierfrancesco Bellini, Ivan Bruno, Daniele Cenni, and Paolo Nesi. Micro grids for scalable media computing and intelligence in distributed scenarios. IEEE MultiMedia, 19(2):69–79, May/June 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Betrisey:2000:DFP

[BBD^{+}00] Claude Betrisey, James F. Blinn, Bodin Dresevic, Bill Hill, Greg Hitchcock, Bert Keely, Don P. Mitchell, John C. Platt, and Turner Whitted. Displaced filtering for patterned displays. SID International Symposium. Digest of Technical Papers, 31:296–299, 2000. CO-DEN DTPSDS. ISSN 0097-966X (print), 2154-6738 (electronic), 2154-6746.

                                        URL http://research.
                                        microsoft.com/~jplatt/
                                        cleartype/sid2000.pdf;
                                         http://research.microsoft.
                                        com/~jplatt/cleartype/
                                        sid2000.ps.gz.

Ballan:2010:VAR

[BBDS10] Lamberto Ballan, Marco Bertini, Alberto Del Bimbo, and Giuseppe Serra. Video annotation and retrieval using ontologies and rule learning. IEEE MultiMedia, 17(4):80–88, October/ December 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Blonde:1996:VSL

[BBG^{+}96] Laurent Blondé, Matthias Buck, Ricardo Galli, Wolfgang Niem, Yakup Paker, Wolfgang Schmidt, and Graham Thomas. A virtual studio for live broadcasting: The Mona Lisa Project. IEEE Multi-Media, 3(2):18–29, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2018.
pdf;   http://www.computer.
org/multimedia/mu1996/
u2018abs.htm.

Brown:2002:MWE

[BBMC02] David C. Brown, Esteban Burbano, Joel Minski, and Isabel F. Cruz. Multimedia at work: Evaluating Web page color and layout adaptations. IEEE MultiMedia, 9(1):86–89, January/March 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2002/pdf/u1086.
pdf;   http://www.computer.
org/multimedia/mu2002/
u1086abs.htm.

Bruns:2007:EMP

[BBZB07] Erich Bruns, Benjamnin Brombach, Thomas Zeidler, and Oliver Bimber. Enabling mobile phones to support large-scale museum guidance. IEEE MultiMedia, 14(2):16–25, April/June 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Brezeale:2009:LVP

[BC09] Darin Brezeale and Diane J. Cook. Learning video preferences using visual features and closed captions. IEEE Multi-Media, 16(3):39–47, July/ September 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Barra:2002:MMW

[BCD^{+}02] Maria Barra, Tania Cillo, Antonio De Santis, Umberto Ferraro Petrillo, Alberto Negro, and Vittorio Scarano. Multimodal monitoring of Web servers. IEEE MultiMedia, 9(3):32– 41, July/September 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2002/
                                        03/extras/u3032x1.mp3;
                                         http://csdl.computer.
                                        org/comp/mags/mu/2002/
                                        03/extras/u3032x2.mid;
                                         http://csdl.computer.
                                        org/comp/mags/mu/2002/
                                        03/u3032abs.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mu/2002/03/u3032.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/03/
                                        u3032.pdf.

Bertino:2003:QSS

[BCEH03] Elisa Bertino, Tiziana Catarci, Ahmed K. Elmagarmid, and Mohand-Saïd Hacid. Quality of service specification in video databases. IEEE Multi-Media, 10(4):71–81, October/December 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
04/u4071abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/04/u4071.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/04/
u4071.pdf.

Bottoni:2004:GMC

[BCF04] Paolo Bottoni, Marina Cammilli, and Stefano Faralli. Generating multimedia content with cellular automata. IEEE Multi-Media, 11(4):78–83, October/December 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/04/u4078.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/04/
u4078.pdf.

Bastan:2010:BMC

[BCGU10] Muhammet Bastan, Hayati Cam, Ugur Gudukbay, and Ozgur Ulusoy.

BilVideo-7: An MPEG-7-compatible video indexing and retrieval system. IEEE MultiMedia, 17(3):62–73, July/September 2010. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Bimber:2005:SPA

[BCK^{+}05] Oliver Bimber, Franz Coriand, Alexander Kleppe, Erich Bruns, Stefanie Zollmann, and Tobias Langlotz. Superimposing pictorial artwork with projected imagery. IEEE MultiMedia, 12(1):16–26, January/March 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Battista:1999:SMM

[BCL99] Stefano Battista, Franco Casalino, and Claudio Lande. Standards: MPEG-4: a multimedia standard for the third millennium. Part 1. IEEE MultiMedia, 6(4):74–83, October/ December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u4074.
                                        pdf.

Battista:2000:SMM

[BCL00] Stefano Battista, Franco Casalino, and Claudio Lande. Standards: MPEG-4: a multimedia standard for the third millennium, Part 2; for more information. IEEE MultiMedia, 7(1):76–84, January/March 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u1076.
pdf.

Bu:2014:MFF

[BCLH14] Shuhui Bu, Shaoguang Cheng, Zhenbao Liu, and Junwei Han. Multimodal feature fusion for 3D shape recognition and retrieval. IEEE MultiMedia, 21(4):38–46, October/ December 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu2014040038-
abs.html.

Bourbakis:2003:SBC

[BD03] Nikolaos Bourbakis and Apostolos Dollas. SCANbased compression-encryption-hiding for video on demand. IEEE MultiMedia, 10(3):79–87, July/ September 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
03/u3079abs.htm;   http:
//csdl.computer.org/dl/

                                        mags/mu/2003/03/u3079.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/03/
                                        u3079.pdf.

Bagdanov:2012:PLF

[BDD^{+}12] Andrew D. Bagdanov, Alberto Del Bimbo, Fabrizio Dini, Giuseppe Lisanti, and Iacopo Masi. Posterity logging of face imagery for video surveillance. IEEE MultiMedia, 19(4):48–59, October/December 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Buccafurri:2009:FUW

[BDL^{+}09] Francesco Buccafurri, Pasquale De Meo, Gianluca Lax, Domenico Ursino, Mariagrazia Fugini, Stefano Modafferi, Barbara Pernici, Roberto Furnari, Anna Goy, Pasquale Lops, Domenico Redavid, and Giovanni Semeraro. A framework for using Web services to enhance QoS for content delivery. IEEE MultiMedia, 16(1):26–35, January/March 2009. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Bucci:1994:SMD

[BDPN94] Giacomo Bucci, Riccardo Detti, Valdo Pasqui, and Stefano Nativi. Sharing multimedia data over a client-server network.

IEEE MultiMedia, 1(3): 44–55, Autumn/Fall 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Bertini:2009:DPE

[BDS^{+}09] Marco Bertini, Alberto Del Bimbo, Giuseppe Serra, Carlo Torniai, Rita Cucchiara, Costantino Grana, and Roberto Vezzani. Dynamic pictorially enriched ontologies for digital video libraries. IEEE MultiMedia, 16(2):42–51, April/ June 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Beadle:1995:EMM

[Bea95] H. W. Peter Beadle. Experiments in multipoint multimedia telecommunication. IEEE MultiMedia, 2(2):30– 40, Summer 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u20030.
pdf;   http://www.computer.
org/multimedia/mu1995/
u2030abs.htm.

Bednarzek:2005:APN

[Bed05] Wolfgang Bednarzek. The art and promise of network analysis. IEEE MultiMedia, 12(3):19–21, July/ September 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Boll:2018:MWM

[BEG^{+}18] S. Boll, T. Ebrahimi, C. Gurrin, L. Jalali, R. Jain, J. Meyer, and N. E. O Connor. MMHealth 2017: Workshop on multimedia for personal health and health care. IEEE MultiMedia, 25(1):7–11, January/March 2018. CO-DEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Bender:1996:MWT

[Ben96] Walter Bender. Multimedia at work: Technology and the news: what we don’t know. IEEE MultiMedia, 3(3):73–77, Fall 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u3073.
                                        pdf.

Battiato:2012:MFS

[BEUW12] Sebastiano Battiato, Sabu Emmanuel, Adrian Ulges, and Marcel Worring. Multimedia in forensics, security, and intelligence. IEEE MultiMedia, 19(1):17–19, January/March 2012. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Blattner:1996:MI

[BG96] Meera M. Blattner and Ephraim P. Glinert. Multimodal integration. IEEE MultiMedia, 3(4):14–24, Winter 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u4014.
pdf;   http://www.computer.
org/multimedia/mu1996/
u4014abs.htm.

Bernd:2015:TPM

[BGC^{+}15] Julia Bernd, Blanca Gordo, Jaeyoung Choi, Bryan Morgan, Nicholas Henderson, Serge Egelman, Daniel D. Garcia, and Gerald Friedland. Teaching privacy: Multimedia making a difference. IEEE MultiMedia, 22(1):12–19, January/March 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/01/mmu2015010012-
abs.html.

Bujari:2021:RPS

[BGPQ21] Armir Bujari, Ombretta Gaggi, Claudio E. Palazzi, and Giacomo Quadrio. Are remote play streaming systems doomed to fail? A network perspective. IEEE MultiMedia, 28(4):45–54, October/December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Brell:2007:PTM

[BH07] Melina Brell and Andreas Hein. Positioning tasks in multimodal computernavigated surgery. IEEE MultiMedia, 14(4):42–51, October/December 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Bhandarkar:2000:MRH

[Bha00] Suchendra M. Bhandarkar. Media reviews: Humanface recognition: Two- and Three-Dimensional Patterns of the Face by Peter L. Hallinan, Gaile G. Gordon, A. L. Yuille, Peter Giblin, and David Mumford (A. K. Peters, Natick, Mass., 1999, $48.00, 300 pp., ISBN 1-56881-087-3). IEEE MultiMedia, 7(1):86–87, January/ March 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u1085.
                                        pdf.

Barate:2016:MIP

[BHLP16] Adriano Barate, Goffredo Haus, Luca A. Ludovico, and Paolo Perlasca. Managing intellectual property in a music fruition environment.

IEEE MultiMedia, 23(2): 84–94, April/June 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
02/mmu2016020084-abs.
html.

Boujemaa:2013:FMA

[BHS13] Nozha Boujemaa, Alexander G. Hauptmann, and Shin’ichi Satoh. The future of multimedia analysis and mining: Visions from the Shonan Meeting. IEEE MultiMedia, 20 (2):4–12, April/June 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Birkmaier:1994:OADa

[Bir94a] Craig J. Birkmaier. An open architecture for digital communication systems, Part 1. IEEE MultiMedia, 1(3):77–83, Autumn/Fall 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Birkmaier:1994:OADb

[Bir94b] Craig J. Birkmaier. An open architecture for digital communication systems, Part 2. IEEE Multi-Media, 1(4):79–84, Winter 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Byrne:2013:XCP

[BKCR13] Daragh Byrne, Aisling Kelliher, Donna Cox, and Thanassis Rikakis. XSEAD: A community platform for art-science integration. IEEE MultiMedia, 20(4):4–7, October/ December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Barry:2001:SMW

[BL01] Chris Barry and Michael Lang. A survey of multimedia and Web development techniques and methodology usage. IEEE MultiMedia, 8(2):52–??, April/June 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u2052.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2001/
                                        u2052abs.htm.

Blattner:1994:OII

[Bla94] Meera M. Blattner. In our image: Interface design in the 1990s. IEEE Multi-Media, 1(1):25–36, Spring 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Braud:2023:DER

[BLA^{+}23] Tristan Braud, Lik-Hang Lee, Ahmad Alhilal, Carlos Bermejo Fernández, and Pan Hui. DiOS an extended reality operating system for the metaverse. IEEE MultiMedia, 30(2): 70–80, April/June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Bayari:2022:NSF

[BLBC22] Parkala Vishnu Bharadwaj Bayari, Apoorva Lakshman, Gaurav Bhatnagar, and Chiranjoy Chattopadhyay. A novel security framework for medical data in IoT ecosystems. IEEE MultiMedia, 29(2): 34–44, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Boll:2022:ESI

[BLM^{+}22] Susanne Boll, Jeannie S.

Lee, Jochen Meyer, Nitish Nag, and Noel E. O’Connor. Editorial special issue. IEEE MultiMedia, 29(1):5–6, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Borchers:1998:DPI

[BM98] Jan Borchers and Max Mühlhäuser. Design patterns for interactive musical systems. IEEE MultiMedia, 5(3):36–46, July/ September 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u3036.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1998/
                                        u3036abs.htm.

Boll:2022:HXD

[BM22] Susanne Boll and Jochen Meyer. Health-X dataLOFT: a sovereign federated cloud for personalized health care services. IEEE MultiMedia, 29(1): 136–140, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Bagi:2020:LSD

[BMDG20] Randheer Bagi, Sabyasachi Mohanty, Tanima Dutta, and Hari Prabhat Gupta. Leveraging smart devices for scene text preserved image stylization: a deep gaming approach. IEEE MultiMedia, 27(2):19–32, April/June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Boll:2018:HMM

[BMO18] S. Boll, J. Meyer, and N. E. O Connor. Health media: From multimedia signals to personal health insights. IEEE MultiMedia, 25(1):51–60, January/ March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Blefari-Melazzi:1999:RMS

[BMR99] Nicola Blefari-Melazzi and Gianluca Reali. A resource management scheme for satellite networks. IEEE MultiMedia, 6(4):54–63, October/December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1999/pdf/
u4054.pdf;   http://www.
computer.org/multimedia/
mu1999/u4054abs.htm.

Biezunski:2001:SXT

[BN01] Michel Biezunski and Steven R. Newcomb. Standards: XML topic maps: Finding aids for the Web. IEEE MultiMedia, 8(2):104–??, April/ June 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u2104.
pdf.

Baraheem:2024:SSI

[BN24] Samah S. Baraheem and Tam V. Nguyen. S5: Sketch-to-image synthesis via scene and size sensing. IEEE MultiMedia, 31 (2):7–16, April/June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Bellini:2005:SMR

[BNZ05] Pierfrancesco Bellini, Paolo Nesi, and Giorgio Zoia. Symbolic music representation in MPEG. IEEE MultiMedia, 12(4):42–49, October/December 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Boll:2006:VVV

[Bol06]  Susanne Boll.  Visions

and views: Visions and view in multimedia. IEEE MultiMedia, 13(2):15–17, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Boll:2007:VVM

[Bol07a] Susanne Boll. Visions and views: MultiTube — where Web 2.0 and multimedia could meet. IEEE MultiMedia, 14(1):9–13, January/March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Boll:2007:VVS

[Bol07b] Susanne Boll. Visions and views: Share it, reveal it, reuse it, and push multimedia into a new decade. IEEE MultiMedia, 14(4): 14–19, October/December 2007. CODEN IEMUE4.

ISSN 1070-986X (print), 1941-0166 (electronic).

Boll:2008:TBI

[Bol08] Susanne Boll. Tech beat:

Inside digital, life, design 2008. IEEE MultiMedia, 15 (2):8–12, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Boll:2009:VVH

[Bol09] Susanne Boll. Visions and views: How much multimedia does a modern family need? IEEE MultiMedia, 16(2):86–91, April/June 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Boll:2015:MCO

[Bol15a] Susanne Boll. Multimedia computing offers a multisensory perspective. IEEE MultiMedia, 22(2): 23, April/June 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/02/mmu2015020023.
html.

Boll:2015:MTS

[Bol15b] Susanne Boll. Multimedia takes on societal challenges. IEEE MultiMedia, 22(4):2–3, October/ December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/04/mmu2015040002-
abs.html.

Boll:2017:MCT

[Bol17] Susanne Boll. Multimedia at CHI: Telepresence at work for remote conference participation. IEEE MultiMedia, 24(3):5–9, July/ September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
03/mmu2017030005-abs.
html.

Boll:2020:SOC

[Bol20] Susanne Boll. Shaping our common digital future. IEEE MultiMedia, 27(3): 5–7, July/September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Bove:2001:CM

[Bov01] Michael Bove, Jr. Connected by media. IEEE MultiMedia, 8(4):13–15, October/December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu2001/pdf/
u4013.pdf;   http://www.
computer.org/multimedia/
mu2001/u4013abs.htm.

Bisdikian:1996:CBP

[BP96] Chatschik C. Bisdikian and Baiju V. Patel. Costbased program allocation for distributed multimediaon-demand systems. IEEE MultiMedia, 3(3):62–72, Fall 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u3062.
pdf;   http://www.computer.
org/multimedia/mu1996/
u3062abs.htm.

Blum:2006:IIB

[BPT06] Mark Blum, Alex (Sandy)

Pentland, and Gerhard Tröster. InSense: Interestbased life logging. IEEE MultiMedia, 13(4):40–48, October/December 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Bruneau-Queyreix:2018:AND

[BQLN^{+}18] J. Bruneau-Queyreix, M. Lacaud, D. Négru, J. M. Batalla, and E. Borcoci. Adding a new dimension to HTTP adaptive streaming through multiple-source capabilities. IEEE Multi-Media, 25(3):65–78, July/ September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Brubeck:1996:HSM

[BR96] David W. Brubeck and Lawrence A. Rowe. Hierarchical storage management in a distributed VOD system. IEEE MultiMedia, 3 (3):37–47, Fall 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u3037.
pdf;   http://www.computer.
org/multimedia/mu1996/
u3037abs.htm.

Braun:1997:SIPa

[Bra97a] Torsten Braun. Standards: Internet protocols for multimedia communications. I. IPng — the foundation of Internet protocols. IEEE MultiMedia, 4(3):85–90, July/September 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3085.
pdf.

Braun:1997:SIPb

[Bra97b] Torsten Braun. Standards: Internet protocols for multimedia communications. II. resource reservation, transport, and application protocols. IEEE Multi-Media, 4(4):74–82, October/December 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u4074.
pdf.

Brackett:1998:MWS

[Bra98] John W. Brackett. Multimedia at work: Satellitebased distance learning using digital video and the Internet. IEEE Multi-Media, 5(3):72–76, July/ September 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u3072.
pdf.

Braun:2005:APS

[Bra05] Carol-Ann Braun. April in Paris: Signs of artificial life along the Seine. IEEE MultiMedia, 12(3):14–18, July/ September 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Brew:1999:AMA

[Bre99] Kathy Brew. Artful media: Anyone up for being a Techno de Medici? IEEE MultiMedia, 6(2):9– 12, April/June 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u2009.
pdf.

Brew:2004:ECA

[Bre04] Kathy Brew. Engaging characters at art interactive. IEEE MultiMedia, 11(1):1, 4–7, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
01/u1001abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2004/01/u1001.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/01/
u1001.pdf.

Brokken:1996:MRN

[Bro96] Dirk Brokken. Media reviews: Networking book gives general coverage: PC Networking Handbook by Ed Titel (AP Professional, 1996, ISBN 0-12-691398-8). IEEE MultiMedia, 3 (3):92–93, Fall 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u3090.
pdf.

Brooks:2002:DDI

[Bro02]  Kevin Brooks.  Danc-

ing with digital interface complexity: a story approach. IEEE MultiMedia, 9(1):8–11, January/ March 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2002/pdf/u1008.
pdf;   http://www.computer.
org/multimedia/mu2002/
u1008abs.htm.

Broeckmann:2005:DCA

[Bro05] Andreas Broeckmann. Digital culture, art, and technology. IEEE MultiMedia, 12(4):9–11, October/ December 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Bryan:2000:SDP

[Bry00] Martin Bryan. Standards:

The Diffuse Project — single sourcing IT standards information; publications and activities. IEEE MultiMedia, 7(4):73–77, October/December 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u4073.
pdf.

Bulterman:2001:S

[Bul01] Dick C. A. Bulterman.

Standards. IEEE Multi-Media, 8(4):82–88, October/December 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u4082.
pdf;   http://www.computer.

                                        org/multimedia/mu2001/
                                        u4082abs.htm.

Bulterman:2002:SPE

[Bul02] Dick C. A. Bulterman. SMIL 2.0: Part 2: Examples and comparisons. IEEE MultiMedia, 9(1):74– 84, January/March 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://dlib.computer.
                                        org/mu/books/mu2002/pdf/
                                        u1074.pdf;   http://www.
                                        computer.org/multimedia/
                                        mu2002/u1074abs.htm.

Bulterman:2004:ITM

[Bul04] Dick C. A. Bulterman. Is it time for a moratorium on metadata? IEEE Multi-Media, 11(4):10–17, October/December 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2004/04/u4010.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/04/
                                        u4010.pdf.

Bureaud:2009:AMK

[Bur09a] Annick Bureaud. Artful media: Kitsou Dubois and the weightless body. IEEE MultiMedia, 16(1): 4–7, January/March 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Burrell:2009:MBU

[Bur09b]  Andrew Burrell.  Me-

dia beat: uncharted pages from a voyage of the Beagle. IEEE MultiMedia, 16 (2):9–11, April/June 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Balduini:2015:CFC

[BVA^{+}15] Marco Balduini, Emanuele Della Valle, Matteo Azzi, Roberto Larcher, Fabrizio Antonelli, and Paolo Ciuccarelli. CitySensing: Fusing city data for visual storytelling. IEEE Multi-Media, 22(3):44–53, July/ September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/03/mmu2015030044-
abs.html.

Burnett:2003:MGA

[BVH^{+}03] Ian Burnett, Rik Van de Walle, Keith Hill, Jan Bormans, and Fernando Pereira. MPEG-21: Goals and achievements. IEEE MultiMedia, 10(4):60–70, October/December 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
04/u4060abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/04/u4060.

                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/04/
                                        u4060.pdf.

Banquiero:2023:PMR

[BVT^{+}23] Mariano Banquiero, Gracia Valdeolivas, Sergio Trincado, Natasha García, and M.-Carmen Juan. Passthrough mixed reality with oculus quest 2: A case study on learning piano. IEEE MultiMedia, 30(2): 60–69, April/June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Bajcsy:2013:ITM

[BYZZ13] Ruzena Bajcsy, Ruigang Yang, Pietro Zanuttigh, and Cha Zhang. 3D imaging techniques and multimedia applications [Guest Editor’s introduction]. IEEE MultiMedia, 20 (1):14–16, January/March 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Canos:2004:MAE

[CAJ04] José H. Canós, Gustavo Alonso, and Javier Jaén. A multimedia approach to the efficient implementation and use of emergency plans. IEEE MultiMedia, 11(3):106–110, July/ September 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/03/u3106.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/03/
u3106.pdf.

Campbell:1997:MWE

[Cam97] Dace A. Campbell. Multimedia at work: Explorations into virtual architecture: a HIT Lab gallery. IEEE MultiMedia, 4(1):74–76, January/ March 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u1074.
pdf.

Carroll:1996:MRM

[Car96] Peter Carroll. Media reviews: Map first, browse later: CyberPilot Pro: Complete Browser Accelerator and NetCarta WebMap Maker for Windows & Windows NT, (PC disk, NetCarta Corporation, $99.95). IEEE MultiMedia, 3(2):85–86, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2083.
pdf.

Carey:1998:SVR

[Car98] Rikk Carey. Standards:

The Virtual Reality Modeling Language explained; general VRML information; VRML browsers; VRML history; recommended VRML books. IEEE MultiMedia, 5(3):84– 93, July/September 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://dlib.computer.
                                        org/mu/books/mu1998/pdf/
                                        u3084.pdf.

Casner:1994:YM

[Cas94] Stephen Casner. Are you on the MBone? IEEE MultiMedia, 1(2):76–79, Summer 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Casman:1996:SRI

[Cas96]  Jesse Casman.  Spe-

cial report: The Internet in Japan. IEEE MultiMedia, 3(2):10–11, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u2010.
                                        pdf.

Castro:2010:NPC

[CASC10] Helder Castro, Artur P. Alves, Carlos Serrao, and Brett Caraway. A new paradigm for content producers. IEEE MultiMedia, 17(2):90–93, April/June 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Comsa:2020:DSC

[CBC^{+}20] Ioan-Sorin Comşa, Estêvão Bissoli Saleme, Alexandra Covaci, Gebremariam Mesfin Assres, Ramona Trestian, Celso A. S. Santos, and Gheorghi Ghinea. Do I smell coffee? The tale of a 360^{∘} mulsemedia experience. IEEE MultiMedia, 27 (1):27–36, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Cheng:2009:IMA

[CBG09] Irene Cheng, Anup Basu, and Randy Goebel. Interactive multimedia for adaptive online education. IEEE MultiMedia, 16 (1):16–25, January/March 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Cao:2011:CVE

[CBWD11] Xun Cao, Alan C. Bovik, Yao Wang, and Qionghai Dai. Converting 2D video to 3D: An efficient path to a 3D experience. IEEE MultiMedia, 18(4): 12–17, October/December 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Camurri:1998:AEA

[CC98] Antonio Camurri and Alessandro Coglio. An architecture for emotional agents. IEEE MultiMedia, 5(4):24–33, October/ December 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u4024.
pdf;   http://www.computer.
org/multimedia/mu1998/
u4024abs.htm.

Camiciottoli:1998:NGD

[CCB^{+}98] Ricardo Camiciottoli, Jacopo M. Corridoni, Alberto Del Bimbo, Enrico Vicario, and Dario Lucarella. 3D navigation of geographic data sets. IEEE MultiMedia, 5(2):29– 41, April/June 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u2029.
pdf;   http://www.computer.
org/multimedia/mu1998/
u2029abs.htm.

Chen:2003:ATG

[CCLS03] Lih-Shyang Chen, Ku-Yaw Chang, Chii-Jeng Lin, and Pey-Shiaw Shieh. Application task guides. IEEE MultiMedia, 10(1):55–63, January/March 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
01/u1055abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/01/u1055.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/01/
u1055.pdf.

Cavazza:2004:MAM

[CCM^{+}04] Marc Cavazza, Fred Charles, Steven J. Mead, Olivier Martin, Xavier Marichal, and Alok Nandi. Multimodal acting in mixed reality interactive storytelling. IEEE MultiMedia, 11(3):30–39, July/ September 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/03/u3030.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/03/
u3030.pdf.

Chen:2011:VRD

[CCPC11] Trista P. Chen, Ching-Wei Chen, Phillip Popp, and Bob Coover. Visual rhythm detection and its applications in interactive multimedia. IEEE MultiMedia, 18(1):88–95, January/March 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Choi:2016:THR

[CCS16] Jung-Woo Choi, Byung Jin Cho, and Insik Shin. Toward the holographic reconstruction of sound fields using smart sound devices. IEEE MultiMedia, 23(3):64–74, July/ September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        03/mmu2016030064-abs.
                                        html.

Chu:2007:TSA

[CCW07] Wei-Ta Chu, Jun-Cheng Chen, and Ja-Ling Wu. Tiling slideshow: An audiovisual presentation method for consumer photos. IEEE MultiMedia, 14 (3):36–45, July/September 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Cucchiara:2014:VAC

[CD14] Rita Cucchiara and Alberto Del Bimbo. Visions for augmented cultural heritage experience. IEEE MultiMedia, 21(1):74–82, January/March 2014. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Canazza:2000:MWA

[CDD^{+}00] Sergio Canazza, Giovanni De Poli, Carlo Drioli, Antonio Rodà, and Alvise Vidolin. Multimedia at work: Audio morphing different expressive intentions for multimedia systems. IEEE MultiMedia, 7(3):79–83, July/September 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u3079.
pdf.

Colombo:1999:SVI

[CDP99] Carlo Colombo, Alberto Del Bimbo, and Pietro Pala. Semantics in visual information retrieval. IEEE Multi-Media, 6(3):38–53, July/ September 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u3038.
pdf;   http://www.computer.
org/multimedia/mu1999/
u3038abs.htm.

Carreras:2010:PCA

[CDR^{+}10] Anna Carreras, Jaime Delgado, Eva Rodriguez, Vitor Barbosa, Maria Teresa Andrade, Hemantha Kodikara Arachchi, Safak Dogan, and Ahmet M. Kondoz. A platform for contextaware and digital rights management-enabled content adaptation. IEEE MultiMedia, 17(2):74–89, April/June 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ci:2010:EBQ

[CDWT10] Wang Ci, Haoyuan Dong, Zhikai Wu, and Yap-Peng Tan. Example-based quality estimation for compressed images. IEEE MultiMedia, 17(3):54–61, July/ September 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Cosker:2013:AFA

[CEG^{+}13] Darren Cosker, Peter Eisert, Oliver Grau, Peter J. B. Hancock, Jonathan McKinnell, and Eng-Jon Ong. Applications of face analysis and modeling in media production. IEEE MultiMedia, 20(4):18–27, October/December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Christodoulakis:2010:PCC

[CFR^{+}10] Stavros Christodoulakis, Michalis Foukarakis, Lemonia Ragia, Hiroaki Uchiyama, and Takuya Imai. Picture context capturing for mobile databases. IEEE MultiMedia, 17(2):34–41, April/June 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Carrier:1995:PME

[CG95] Sylvain Carrier and Nicolas D. Georganas. Practical multimedia electronic mail on X.400. IEEE MultiMedia, 2(4):12–23, Winter 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u40012.
pdf;   http://www.computer.
org/multimedia/mu1995/
u4012abs.htm.

Chen:2014:MEI

[CG14] David M. Chen and Bernd Girod. Memory-efficient image databases for mobile visual search. IEEE MultiMedia, 21(1):14–23, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chang:1997:RIL

[CGM97] Edward Chang and Hector Garcia-Molina. Reducing initial latency in media servers. IEEE Multi-Media, 4(3):50–61, July/ September 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3050.
pdf;   http://www.computer.
org/multimedia/mu1997/
u3050abs.htm.

Chen:2023:LDF

[CGW^{+}23] Zhuo Chen, Tao Guan, Yuesong Wang, Yawei Luo, Luoyuan Xu, and Wenkai Liu. Learning 3-D face shape from diverse sources with cross-domain face synthesis. IEEE Multi-Media, 30(1):7–16, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chang:2002:AQW

[CH02] Wen-Kui Chang and Shing-Kai Hon. Assessing the quality of Web-based applications via navigational structures. IEEE Multi-Media, 9(3):22–30, July/ September 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
03/u3022abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/03/u3022.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/03/
u3022.pdf.

Chang:1996:EVL

[Cha96] Shi-Kuo Chang. Extending visual languages for multimedia. IEEE MultiMedia, 3(3):18–26, Fall 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/

books/mu1996/pdf/u3018.
pdf;   http://www.computer.
org/multimedia/mu1996/
u3018abs.htm.

Chandramouli:2006:MRL

[Cha06] Rajarathnam Chandramouli.

Media reviews: Learning through multimedia interaction. IEEE MultiMedia, 13(1):84–85, January/ March 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chou:2022:GSC

[CHCH22] Chien-Hsing Chou, Ping-Hsuan Han, Chia-Chun Chang, and Yi-Zeng Hsieh. Garment style creator: Using StarGAN for image-to-image translation of multidomain garments. IEEE MultiMedia, 29(1):85–93, January/March 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chery:2000:MWB

[Che00]  Yonald Chery.  Multi-

media at work: Bringing the common whiteboard into the digital age. IEEE MultiMedia, 7(2):90– 92, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2090.
pdf.

Chen:2004:PMC

[Che04] Chang Wen Chen. Panorama of multimedia coding and processing. IEEE Multi-Media, 11(3):111–112, c3, July/September 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2004/
                                        03/%0A;   http://csdl.
                                        computer.org/comp/mags/
                                        mu/2004/03/u3111.pdf;
                                         http://csdl.computer.
                                        org/dl/mags/mu/2004/03/
                                        u3111.htm;   http:/www.
                                        computer.org/multimedia/
                                        mu2004/u3111x1.htm.

Chen:2018:BPB

[Che18a] S. Chen. Best paper and best department article recognition. IEEE MultiMedia, 25(2):5–6, April/ June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Chen:2018:MIM

[Che18b] S. Chen. Moving IEEE MultiMedia forward. IEEE MultiMedia, 25(1):5–6, January/March 2018. CO-DEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Chen:2018:MDS

[Che18c] S. Chen. Multimedia for data science. IEEE Multi-Media, 25(4):5–7, October/ December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Chen:2018:MDI

[Che18d] S. Chen. Multimedia for disaster information management. IEEE Multi-Media, 25(3):5–7, July/ September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Chen:2019:IA

[Che19a] S. Chen. Impact and award. IEEE MultiMedia, 26(4):5–6, October/ December 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2019:AIN

[Che19b] S. Chen. Is artificial intelligence new to multimedia? IEEE MultiMedia, 26 (2):5–7, April/June 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Chen:2019:MDL

[Che19c] S. Chen. Multimedia deep learning. IEEE MultiMedia, 26(1):5–7, January/ March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Chen:2019:MAD

[Che19d]  S. Chen.  Multimedia

for autonomous driving. IEEE MultiMedia, 26(3): 5–8, July/September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2020:MYR

[Che20]  Shu-Ching Chen.  Mo-

mentous years of recognition. IEEE MultiMedia, 27 (4):5, October/December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2021:EMD

[Che21a] Shu-Ching Chen. Embracing multimodal data in multimedia data analysis. IEEE MultiMedia, 28(3): 5–7, July/September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2021:MDA

[Che21b] Shu-Ching Chen. Multimedia data analysis with edge computing. IEEE Multi-Media, 28(4):5–7, October/ December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2021:MVR

[Che21c] Shu-Ching Chen. Multimedia in virtual reality and augmented reality. IEEE MultiMedia, 28 (2):5–7, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2021:MRR

[Che21d] Shu-Ching Chen. Multimedia research for response and management of COVID-19 and future pandemics. IEEE MultiMedia, 28(1):5–6, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2022:DAB

[Che22a] Shu-Ching Chen. Digital assets and blockchainbased multimedia data management. IEEE MultiMedia, 29(2):5–7, April/ June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2022:MMD

[Che22b] Shu-Ching Chen. Multimedia meets deep reinforcement learning. IEEE MultiMedia, 29(3):5–7, July/ September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2022:MRT

[Che22c] Shu-Ching Chen. Multimedia research toward the metaverse. IEEE Multi-Media, 29(1):125–127, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2022:PTC

[Che22d] Shu-Ching Chen. Passing the torch continue moving forward. IEEE MultiMedia, 29(4):4, October/ December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chang:1998:MMU

[CHH98] Shi-Kuo Chang, Ehab Hassanein, and Chung-Yuan Hsieh. A multimedia micro-university. IEEE MultiMedia, 5(3):60–68, July/September 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u3060.
pdf;   http://www.computer.
org/multimedia/mu1998/
u3060abs.htm.

Chiariglione:1995:SMT

[Chi95]  Leonardo  Chiariglione.

Standards: MPEG: a technological basis for multimedia applications. IEEE MultiMedia, 2(1):85–89, Spring 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u10085.
pdf.

Chiariglione:1997:SCM

[Chi97]  Leonardo  Chiariglione.

Standards: The challenge of multimedia standardization. IEEE MultiMedia, 4(2):79–83, April/ June 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2079.
pdf.

Cha:2009:FHB

[CHK^{+}09] Jongeun Cha, Yo-Sung Ho, Yeongmi Kim, Jeha Ryu, and Ian Oakley. A framework for haptic broadcasting. IEEE MultiMedia, 16 (3):16–27, July/September 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chao:2022:TRE

[CHSC22] Fu-An Chao, Jeih-Weih Hung, Tommy Sheu, and Berlin Chen. Time-reversal enhancement network with cross-domain information for noise-robust speech recognition. IEEE Multi-Media, 29(1):114–124, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chung:1995:VVS

[Chu95] Joseph Chung. Visions and views: Social communication in a virtual office.

IEEE MultiMedia, 2(2):7– 9, Summer 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u20007.
pdf.

Chua:2002:CC

[Chu02] Tat-Seng Chua. Complementary content. IEEE MultiMedia, 9(4):82–83, October/December 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/dl/mags/mu/2002/04/
u4082.htm;   http://csdl.
computer.org/dl/mags/
mu/2002/04/u4082.pdf.

Cavallaro:2011:ALB

[CHWB11] Rick Cavallaro, Maria Hybinette, Marvin White, and Tucker Balch. Augmenting live broadcast sports with 3D tracking information. IEEE MultiMedia, 18(4):38–47, October/ December 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Che:2015:SCY

[CIL15] Xianhui Che, Barry Ip, and Ling Lin. A survey of current YouTube video characteristics. IEEE MultiMedia, 22(2):56–63, April/June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/02/mmu2015020056-
abs.html.

Choi:2012:LPC

[CJ12] Kiho Choi and Euee S.

Jang. Leveraging parallel computing in modern video coding standards. IEEE MultiMedia, 19(3):7– 11, July/September 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:1996:SCS

[CK96] Ming-Syan Chen and Dilip D.

Kandlur. Stream conversion to support interactive video playout. IEEE MultiMedia, 3(2):51–58, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2051.
pdf;   http://www.computer.
org/multimedia/mu1996/
u2051abs.htm.

Chbeir:2009:GEI

[CKAI09] Richard Chbeir, Harald Kosch, Frederic Andres, and Hiroshi Ishikawa. Guest Editors’ introduction: Multimedia metadata and semantic management. IEEE MultiMedia, 16(4): 8–11, October/December 2009. CODEN IEMUE4.

ISSN 1070-986X (print), 1941-0166 (electronic).

Candan:2011:RSM

[CKN^{+}11] K. Selcuk Candan, Jong Kim, Parth Nagarkar, Mithila Nagendra, and Renwei Yu. RanKloud: Scalable multimedia data processing in server clusters. IEEE MultiMedia, 18 (1):64–77, January/March 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chiu:2000:RRV

[CKRW00] Patrick Chiu, Ashutosh Kapuskar, Sarah Reitmeier, and Lynn Wilcox. Room with a rear view: Meeting capture in a multimedia conference room. IEEE MultiMedia, 7(4): 48–54, October/December 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://dlib.computer.
                                        org/mu/books/mu2000/pdf/
                                        u4048.pdf;   http://www.
                                        computer.org/multimedia/
                                        mu2000/u4048abs.htm.

Chen:2006:EMC

[CL06] Herng-Yow Chen and Kuo-Yu Liu. Exploring multimedia correlation and synchronization for Web-based language learning. IEEE MultiMedia, 13(4):50–61, October/December 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chan:2010:MMS

[CL10] Mangtang Chan and Kecheng Liu. Modeling media synchronization with semiotic agents. IEEE MultiMedia, 17(1):14–23, January/ March 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chambers:2002:MED

[CLD02] Desmond Chambers, Gerard Lyons, and Jim Duggan. A multimedia enhanced distributed object event service. IEEE MultiMedia, 9(3):56–71, July/ September 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
03/u3056abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/03/u3056.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/03/
u3056.pdf.

Cleland:2008:ABI

[Cle08] Kathy Cleland. Art beat:

Interactive art: Digital entities and the audience experience. IEEE MultiMedia, 15(4):4–7, October/ December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Campbell:1997:PRB

[CLSS97] Andrew T. Campbell, Aurel A. Lazar, Henning Schulzrinne, and Rolf Stadler. Project reports: Building open programmable multimedia networks. IEEE Multi-Media, 4(1):77–82, January/March 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u1077.
pdf.

Chua:2012:NNT

[CLSY12] Tat-Seng Chua, Huanbo Luan, Maosong Sun, and Shiqiang Yang. NExT: NUS-Tsinghua Center for Extreme Search of User-Generated Content. IEEE MultiMedia, 19(3):81–87, July/September 2012. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Costabile:2001:MWG

[CM01] Maria F. Costabile and Maristella Matera. Multimedia at work: Guidelines for hypermedia usability inspection. IEEE MultiMedia, 8(1):66–69, January/March 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u1066.
pdf.

Costa:2020:WSU

[CMC^{+}20] Miguel Costa, Ricardo Moreira, Jorge Cabral, José Dias, and Sandro Pinto. Wall Screen: an ultra-high definition videocard for the Internet of Things. IEEE MultiMedia, 27(3):76–87, July/ September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2023:LCF

[CMW^{+}23] Yi Chen, Yunhao Mao, Shiqi Wang, Xianguo Zhang, and Sam Kwong. Learning from coding features: High efficiency rate control for AOMedia video 1. IEEE MultiMedia, 30 (4):16–25, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chandrasiri:2004:ICU

[CNI^{+}04] Naiwala P. Chandrasiri, Takeshi Naemura, Mitsuru Ishizuka, Hiroshi Harashima, and István Barakonyi. Internet communication using real-time facial expression analysis and synthesis. IEEE MultiMedia, 11(3):20–29, July/ September 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/03/u3020.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/03/
u3020.pdf;   http://www.
hc.t.u-tokyo.ac.jp/~cds/
IEEEmultimedia/.

Cruz:2002:ABW

[CNN02] Christophe Cruz, Christophe Nicolle, and Marc Neveu. The Active3D-Build: a Web-based civil engineering platform. IEEE MultiMedia, 9(4):87–90, October/December 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
04/u4087abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/04/u4087.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/04/
u4087.pdf.

Chu:2021:PBP

[CNYL21] Renjie Chu, Baoning Niu, Shanshan Yao, and Jianquan Liu. Peak-based Philips fingerprint robust to pitch-shift for audio identification. IEEE MultiMedia, 28(1):74–82, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Cesar:2012:EOS

[CO12] Pablo Cesar and Wei Tsang Ooi. Embracing open source multimedia software. IEEE MultiMedia, 19 (2):11–15, May/June 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Cesar:2014:AIC

[COCB14] Pablo Cesar, Marianna Obrist, Elizabeth F. Churchill, and Tom Bartindale. ACM International Conference on Interactive Experiences for Television and Online Video (ACM TVX 2014). IEEE MultiMedia, 21(4):112–c3, October/December 2014. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu2014040112.
html.

Christel:2000:IMD

[COH00] Michael G. Christel, Andreas M. Olligschlaeger, and Chang Huang. Interactive maps for a digital video library. IEEE MultiMedia, 7(1):60–67, January/March 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u1060.
pdf;   http://www.computer.

                                        org/multimedia/mu2000/
                                        u1060abs.htm.

Colaitis:1994:OMO

[Col94] Françoise Colaitis. Opening up multimedia object exchange with MHEG. IEEE MultiMedia, 1(2):80– 84, Summer 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Cooper:2001:VVI

[Coo01] Martin Cooper. Visions and views: Internet: a life-changing experience. IEEE MultiMedia, 8(2):11– ??, April/June 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u2011.
                                        pdf;   pdf/u2011.pdf.

Coulson:1999:CMM

[Cou99] Geoff Coulson. A configurable multimedia middleware platform. IEEE MultiMedia, 6(1):62–76, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u1062.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1999/
                                        u1062abs.htm.

Chung:2005:TPN

[CP05] Soon M. Chung and Anil L. Pereira. Timed Petri net representation of SMIL. IEEE MultiMedia, 12 (1):64–72, January/March 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chang:2003:VAM

[CPCT03] Shi Kuo Chang, Giuseppe Polese, Maurizio Cibelli, and Roshan Thomas. Visual authorization modeling in E-commerce applications. IEEE Multi-Media, 10(1):44–54, January/March 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
01/u1044abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/01/u1044.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/01/
u1044.pdf.

Camurri:2004:GEI

[CR04] Antonio Camurri and Thanassis Rikakis. Guest Editors’ introduction: Multisensory communication and experience through multimedia. IEEE Multi-Media, 11(3):17–19, July/ September 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
03/u3017.pdf;   http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2004/03/u3017.
                                        htm.

Castronova:2008:MIC

[CRB^{+}08] Edward Castronova, Travis Ross, Mark Bell, Matthew Falk, Robert Cornell, and Matt Haselton. Media impact: Constructing Arden: Life inside the machine. IEEE MultiMedia, 15(1): 4–8, January/March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Carreras:2011:ATA

[CRD^{+}11] Anna Carreras, Eva Rodriguez, Jaime Delgado, Safak Dogan, Hemantha Kodikara Arachchi, Ahmet M. Kondoz, and Xavier Perramon. Architectures and technologies for adapting secured content in governed multimedia applications. IEEE MultiMedia, 18(4):48–61, October/December 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Creutzburg:2005:IT

[Cre05] Reiner Creutzburg. Interactive television. IEEE MultiMedia, 12(2):88–89, April/June 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.

org/comp/mags/mu/2005/
02/u2088.pdf.

Costabile:2003:DHR

[CRL^{+}03] Maria Francesca Costabile, Teresa Roselli, Rosa Lanzilotti, Paola Plantamura, and Antonella De Angeli. Does hypermedia really work for tutoring children? IEEE MultiMedia, 10(2):65–69, April/June 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/02/u2065.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/02/
u2065.pdf.

Cheung:2015:IMA

[cSC15] Sen ching Samson Cheung. Integrating multimedia into autism intervention. IEEE MultiMedia, 22(4):4–10, October/ December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/04/mmu2015040004-
abs.html.

Coma:2020:DSC

[CSC^{+}20] I. Com a, E. B. Saleme, A. Covaci, G. M. Assres, R. Trestian, C. A. S. Santos, and G. Ghinea. Do I smell coffee? The tale of a 360^{∘} mulsemedia experience. IEEE MultiMedia, 27 (1):27–36, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Cesar:2018:NSM

[CSJ^{+}18] P. Cesar, V. Singh, R. Jain, N. Sebe, and N. Oliver. New signals in multimedia systems and applications. IEEE MultiMedia, 25(1):12–13, January/ March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Chen:2020:EAN

[CSZL20] Xia Chen, Zhan-Li Sun, Zhigang Zeng, and Kin-Man Lam. An effective approach for nonrigid structure from motion with complex deformation. IEEE MultiMedia, 27 (3):88–98, July/September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Costantini:2001:CLD

[CT01] Fabien Costantini and Christian Toinard. Collaborative learning with the distributed building site metaphor. IEEE Multi-Media, 8(3):21–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u3021.

pdf;   http://www.computer.
org/multimedia/mu2001/
u3021abs.htm.

Chen:1999:RSV

[CTCC99] Chaomei Chen, Linda Thomas, Janet Cole, and Chiladda Chennawasin. Representing the semantics of virtual spaces. IEEE MultiMedia, 6(2):54–63, April/June 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u2054.
pdf;   http://www.computer.
org/multimedia/mu1999/
u2054abs.htm.

Curran:2002:WBC

[Cur02] Kevin Curran. A Webbased collaboration teaching environment. IEEE MultiMedia, 9(3):72–76, July/September 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
03/u3072abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/03/u3072.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/03/
u3072.pdf.

Chen:2014:NGF

[CV14] Ying Chen and Anthony Vetro. Next-generation 3D formats with depth map support. IEEE MultiMedia, 21(2):90–94, April/ June 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Camurri:2016:IAT

[CV16] Antonio Camurri and Gualtiero Volpe. The intersection of art and technology. IEEE Multi-Media, 23(1):10–17, January/March 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        01/mmu2016010010-abs.
                                        html.

Camurri:2005:CEA

[CVDL05] Antonio Camurri, Gualtiero Volpe, Giovanni De Poli, and Marc Leman. Communicating expressiveness and affect in multimodal interactive systems. IEEE MultiMedia, 12(1):43–53, January/March 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Cheng:2009:NAS

[CW09] Yu-Ming Cheng and Chung-Ming Wang. A novel approach to steganography in high-dynamic-range images. IEEE MultiMedia, 16 (3):70–80, July/September 2009. CODEN IEMUE4.

ISSN 1070-986X (print), 1941-0166 (electronic).

Calegario:2017:MTD

[CWH^{+}17] Filipe Calegario, Marcelo M.

Wanderley, Stephane Huot, Giordano Cabral, and Geber Ramalho. A method and toolkit for digital musical instruments: Generating ideas and prototypes. IEEE Multi-Media, 24(1):63–71, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
01/mmu2017010063-abs.
html.

Cao:2019:EVQ

[CWW^{+}19] Y. Cao, N. Wang, C. Wu, X. Zhang, and C. Suthaputchakun. Enhancing video QoE over high-speed train using segment-based prefetching and caching. IEEE MultiMedia, 26(4): 55–66, October/December 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2022:MFN

[CWY^{+}22] Jiaxin Chen, Zekai Wu, Zhenguo Yang, Haoran Xie, Fu Lee Wang, and Wenyin Liu. Multimodal fusion network with contrary latent topic memory for rumor detection.

IEEE MultiMedia, 29(1): 104–113, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2005:SBP

[CWZ^{+}05] Songqing Chen, Haining Wang, Xiaodong Zhang, Bo Shen, and Susie Wee. Segment-based proxy caching for Internet streaming media delivery. IEEE Multi-Media, 12(3):59–67, July/ September 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Cai:2013:LSN

[CY13] Yang Cai and Linjun Yang. Large-scale near-duplicate Web video retrieval: Challenges and approaches. IEEE MultiMedia, 20(2): 42–51, April/June 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Cheng:2022:GDV

[CYL22] Hsu-Yung Cheng, Chih-Chang Yu, and Chih-Lung Lin. Generating dance videos using pose transfer generative adversarial network with multiple scale region extractor and learnable region normalization. IEEE MultiMedia, 29 (1):47–54, January/March 2022. CODEN IEMUE4.

ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2021:GFA

[CYW21] Baoliang Chen, Wenhan Yang, and Shiqi Wang. Generalized face antispoofing by learning to fuse features from high- and lowfrequency domains. IEEE MultiMedia, 28(1):56–64, January/March 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2023:RSD

[CYW23] Peilin Chen, Wenhan Yang, and Shiqi Wang. Reviving standard-dynamic-range videos for highdynamic-range devices: a learning paradigm with hybrid attention mechanisms. IEEE MultiMedia, 30(3): 110–118, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chang:1996:CAV

[CZ96] Ed Chang and Avideh Zakhor. Cost analyses for VBR video servers. IEEE MultiMedia, 3(4):56– 71, Winter 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u4056.
pdf;   http://www.computer.

                                        org/multimedia/mu1996/
                                        u4056abs.htm.

Cui:2016:SSM

[CZCJ16] Peng Cui, Wenwu Zhu, Tat-Seng Chua, and Ramesh Jain. Social-sensed multimedia computing. IEEE MultiMedia, 23(1):92–96, January/March 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        01/mmu2016010092-abs.
                                        html.

Cao:2020:IRG

[CZY20] Yunyin Cao, Jian Zhang, and Jun Yu. Image retrieval via gated multiscale NetVLAD for social media applications. IEEE Multi-Media, 27(4):69–78, October/December 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Dufourd:2005:MSR

[DAC05] Jean-Claude Dufourd, Olivier Avaro, and Cyril Concolato. An MPEG standard for rich media services. IEEE MultiMedia, 12(4): 60–68, October/December 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Dingler:2016:MMC

[DAL^{+}16] Tilman Dingler, Passant El Agroudy, Huy Viet Le, Albrecht Schmidt, Evangelos Niforatos, Agon Bexheti, and Marc Langheinrich. Multimedia memory cues for augmenting human memory. IEEE MultiMedia, 23(2):4–11, April/June 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
02/mmu2016020004-abs.
html.

Davenport:1994:SDA

[Dav94] Glorianna Davenport. Seeking dynamic, adaptive story environments. IEEE MultiMedia, 1(3):9–13, Autumn/Fall 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Davenport:1995:VVS

[Dav95] Glorianna Davenport. Visions and views: Still seeking: signposts of things to come. IEEE MultiMedia, 2(3):9–13, Fall 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u30009.
pdf.

Davenport:1996:VVS

[Dav96a] G. Davenport. Visions and views: Smarter tools for storytelling: Are they just around the corner? IEEE MultiMedia, 3(1): 10–14, Spring 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u1010.
                                        pdf.

Davenport:1996:VVI

[Dav96b] Glorianna Davenport. Visions and views: Indexes are “Out,” models are “In”. IEEE MultiMedia, 3 (3):10–15, Fall 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u3010.
                                        pdf.

Davenport:1997:VVW

[Dav97] Glorianna Davenport. Visions and views: Whose bits are they, anyway? IEEE MultiMedia, 4(3):8– 11, July/September 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://dlib.computer.
                                        org/mu/books/mu1997/pdf/
                                        u3008.pdf.

Davenport:1998:MRB

[Dav98a] Glorianna Davenport. Media reviews: Basic digital cinematography concepts: Glorianna Davenport Digital Cinematography by Ben de Leeuw (AP Professional, 1998, 265 pp., ISBN 0-12-208875-1). IEEE MultiMedia, 5(2): 91, April/June 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u2090.
pdf.

Davenport:1998:VVC

[Dav98b] Glorianna Davenport. Visions and views: Curious learning, cultural bias, and the learning curve. IEEE MultiMedia, 5(2):14– 19, April/June 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u2014.
pdf.

Davenport:1999:VVG

[Dav99] Glorianna Davenport. Visions and views: Get a life: thinking outside the box [multimedia]. IEEE MultiMedia, 6(1):5–9, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u1005.
pdf.

Davis:2002:PPF

[Dav02] Marc Davis. From pirates to patriots: Fair use for digital media. IEEE MultiMedia, 9(4):4–7, October/December 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
04/u4004abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/04/u4004.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/04/
u4004.pdf.

Davis:2003:EVE

[Dav03]  Marc Davis.  Editing

out video editing. IEEE MultiMedia, 10(2):54–64, April/June 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
02/u2054abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/02/u2054.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/02/
u2054.pdf.

Davenport:1997:VVC

[DB97] Glorianna Davenport and Brian Bradley. Visions and views: The care and feeding of users. IEEE MultiMedia, 4(1):8–11, January/March 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/

books/mu1997/pdf/u1008.
pdf.

DeLeon:2000:BVD

[DB00] Victor DeLeon and Robert Berry, Jr. Bringing VR to the desktop: Are you game? IEEE Multi-Media, 7(2):68–72, April/ June 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2068.
pdf;   http://www.computer.
org/multimedia/mu2000/
u2068abs.htm.

Drury:2005:MBJ

[DB05] Gerrard Drury and Ian Burnett. MPEG-21 in a backpack journalism scenario. IEEE MultiMedia, 12(4):24–32, October/ December 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Dimitrova:2004:CMM

[DD04] Nevenka Dimitrova and Nevenka Dimitrova. Context and memory in multimedia content analysis. IEEE MultiMedia, 11(3):7–11, July/ September 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/03/u3007.

htm;   http://csdl.computer.
org/dl/mags/mu/2004/03/
u3007.pdf.

Doulamis:2003:AAI

[DDV03] Nikolaos D. Doulamis, Anastasios D. Doulamis, and Theodora A. Varvarigou. Adaptive algorithms for interactive multimedia. IEEE MultiMedia, 10(4):38–47, October/December 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
04/u4038abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/04/u4038.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/04/
u4038.pdf.

Dedrick:1995:CMT

[Ded95] Rick Dedrick. A consumption model for targeted electronic advertising. IEEE MultiMedia, 2(2):41–49, Summer 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u20041.
pdf;   http://www.computer.
org/multimedia/mu1995/
u2041abs.htm.

DelBimbo:2000:GEI

[Del00] Alberto Del Bimbo. Guest Editor’s introduction: Multimedia computing and systems. IEEE Multi-Media, 7(1):18–21, January/March 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u1018.
pdf.

Delaney:2001:AMA

[Del01a] Ben Delaney. Artful media: Art is where you grow it. IEEE Multi-Media, 8(1):4–7, January/ March 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u1004.
pdf.

Delaney:2001:MWP

[Del01b] Ben Delaney. Multimedia at work: The power of P2P. IEEE MultiMedia, 8(2):100–??, April/ June 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u2100.
pdf;   pdf/u2100.pdf.

Delaney:2002:NIM

[Del02]  Ben Delaney.  In the

news: The interface is the message. IEEE MultiMedia, 9(1):5–7, January/March 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2002/pdf/u1005.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2002/
                                        u1005abs.htm.

DiNapoli:1999:MWE

[DF99] Claudia Di Napoli and Mario Mango Furnari. Multimedia at work: An environment for a Virtual Hypermedia Factory. IEEE MultiMedia, 6(2):80– 83, April/June 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u2080.
                                        pdf.

Danieau:2014:THC

[DFG^{+}14] Fabien Danieau, Julien Fleureau, Philippe Guillotel, Nicolas Mollet, Marc Christie, and Anatole Lecuyer. Toward haptic cinematography: Enhancing movie experiences with camera-based haptic effects. IEEE MultiMedia, 21(2):11–21, April/June 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

deGotzen:2004:EEM

[dG04] Amalia de Götzen. Enhancing engagement in multimodality environments by sound movements in a virtual space. IEEE MultiMedia, 11(2):4– 8, April/June 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/02/u2004.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/02/
u2004.pdf.

Dick:2008:AAT

[DG08] Wayne Dick and Forouzan Golshani. Accessibility and assistive technologies: Guest Editors’ introduction: Accessibility and assistive technologies. IEEE MultiMedia, 15(4):22–26, October/December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Datta:2007:TBA

[DGLW07] Ritendra Datta, Weina Ge, Jia Li, and James Z. Wang. Toward bridging the annotation-retrieval gap in image search. IEEE Multi-Media, 14(3):24–35, July/ September 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Dias:2013:LSI

[DGR13] Zanoni Dias, Siome Goldenstein, and Anderson Rocha. Large-scale image phylogeny: Tracing image ancestral relationships. IEEE MultiMedia, 20(3):58–70, July/ September 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Dwivedi:2022:MFB

[DGT^{+}22] Shubham Dwivedi, Tripti Goel, M. Tanveer, R. Murugan, and Rahul Sharma. Multimodal fusion-based deep learning network for effective diagnosis of alzheimer’s disease. IEEE MultiMedia, 29(2):45–55, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

DeRoure:1997:PRD

[DH97] David De Roure and Wendy Hall. Project reports: Distributed multimedia information systems. IEEE MultiMedia, 4(4):68–73, October/ December 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1997/pdf/u4068.
                                        pdf.

Deshpande:2001:WEC

[DH01] Yogesh Deshpande and Steve Hansen. Web engineering: Creating a discipline among disciplines. IEEE MultiMedia, 8(2):82– ??, April/June 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u2082.
pdf;   http://www.computer.
org/multimedia/mu2001/
u2082abs.htm.

Degara:2015:ISG

[DHH15] Norberto Degara, Andy Hunt, and Thomas Hermann. Interactive sonification [Guest Editors’ introduction]. IEEE Multi-Media, 22(1):20–23, January/March 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/01/mmu2015010020-
abs.html.

Davenport:1998:VVF

[DHTF98] Glorianna Davenport, Lars Erik Holmquist, Maureen Thomas, and Future of Fun Workshop Participants. Visions and views: Fun: a condition of creative research. IEEE Multi-Media, 5(3):10–15, July/ September 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u3010.
pdf.

Diaz:1999:MWI

[Dia99] Melvin Diaz. Multimedia at work: Integrating GPS receivers into consumer mobile electronics. IEEE MultiMedia, 6(4):88–90, October/ December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u4088.
                                        pdf.

Dimitrova:1999:VVD

[Dim99] Nevenka Dimitrova. Visions and views: Digital video for the time impaired. IEEE Multi-Media, 6(2):14–17, April/ June 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u2014.
                                        pdf.

Dimitrova:2007:GEI

[Dim07] Nevenka Dimitrova. Guest Editor’s introduction: Multimedia signal processing and systems in healthcare and life sciences. IEEE MultiMedia, 14(4):20–21, October/December 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

DeLys:2007:ABP

[DJBF07] Sherre DeLys, John Jacobs, Brogan Bunt, and Marius Foley. Art beat:

The pool project. IEEE MultiMedia, 14(4):c2, 1, 4, October/December 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Djeraba:2002:GEI

[Dje02] Chabane Djeraba. Guest Editor’s introduction: Contentbased multimedia indexing and retrieval. IEEE MultiMedia, 9(2):18–22, April/June 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2002/02/u2018.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/02/
u2018.pdf.

Dorai:2004:CBM

[DK04] Chitra Dorai and Martin Kienzle. Challenges of business media. IEEE MultiMedia, 11(4):18–21, October/December 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/04/u4018.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/04/
u4018.pdf.

Daras:2004:MTC

[DKRS04] Petros Daras, Ioannis Kompatsiaris, Theodoros Raptis, and Michael G. Strintzis. An MPEG-4 tool for composing 3D scenes. IEEE MultiMedia, 11(2): 58–71, April/June 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
02/u2058abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2004/02/u2058.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/02/
u2058.pdf.

Dong:2022:MMC

[DL22] Haiwei Dong and Jeannie S. A. Lee. The metaverse from a multimedia communications perspective. IEEE MultiMedia, 29(4):123–127, October/December 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

deLara:2001:CSB

[dLA01] Juan de Lara and Manuel Alfonseca. Constructing simulation-based Web documents. IEEE MultiMedia, 8(1):42–49, January/ March 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u1042.
pdf;   http://www.computer.
org/multimedia/mu2001/
u1042abs.htm.

Duan:2019:CDV

[DLB^{+}19] L. Duan, Y. Lou, Y. Bai, T. Huang, W. Gao, V. Chandrasekhar, J. Lin, S. Wang, and A. C. Kot. Compact descriptors for video analysis: The emerging MPEG standard. IEEE MultiMedia, 26(2):44–54, April/June 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Day:2001:TLS

[DLH01] Young Francis Day, Peiya Liu, and Liang H. Hsu. Transforming large-scale product documents into multimedia training manuals. IEEE MultiMedia, 8(3):39–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u3039.
pdf;   http://www.computer.
org/multimedia/mu2001/
u3039abs.htm.

Duan:2019:AOL

[DLW^{+}19] L. Duan, Y. Lou, S. Wang, W. Gao, and Y. Rui. AIoriented large-scale video management for smart city: Technologies, standards, and beyond. IEEE MultiMedia, 26(2):8–20, April/June 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Dalhoum:2012:DIS

[DMA^{+}12] Abdel Latif Abu Dalhoum, Basel A. Mahafzah, Aiman Ayyal Awwad, Ibraheem Aldhamari, Alfonso Ortega, and Manuel Alfonseca. Digital image scrambling using 2D cellular automata. IEEE Multi-Media, 19(4):28–36, October/December 2012. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Dittman:2000:MIC

[DN00] Jana Dittman and Frank Nack. Media impact: Copyright copywrong; Web resources. IEEE Multi-Media, 7(4):14–17, October/December 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u4014.
                                        pdf.

Donnelly:2010:MMT

[DNM^{+}10] Mark Donnelly, Chris Nugent, Sally McClean, Bryan Scotney, Sarah Mason, Peter Passmore, and David Craig. A mobile multimedia technology to aid those with Alzheimer’s Disease. IEEE MultiMedia, 17(2): 42–51, April/June 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Dong:2022:WVG

[DOD22] Jiong Dong, Kaoru Ota, and Mianxiong Dong. Why VR games sickness? An empirical study of capturing and analyzing VR games head movement dataset. IEEE MultiMedia, 29(2):74–82, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

deOliveira:2003:JMT

[dOHS^{+}03] Jauvane C. de Oliveira, Mojtaba Hosseini, Shervin Shirmohammadi, François Malric, Saeid Nourian, Abdulmotaleb El Saddik, and Nicolas D. Georganas. Java multimedia telecollaboration. IEEE Multi-Media, 10(3):18–29, July/ September 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
03/u3018abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/03/u3018.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/
u3018.pdf.

Dimitrova:1995:MRC

[DP95] Nevenka Dimitrova and Tina Porter. Media reviews: Children’s software enters new age; educational software is not for children only. IEEE MultiMedia, 2(1):100–102, Spring 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u10100.
                                        pdf.

Diepold:2005:MMA

[DPC05] Klaus Diepold, Fernando Pereira, and Wo Chang. MPEG-A: Multimedia application formats. IEEE MultiMedia, 12(4):34–41, October/December 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Davis:2009:USN

[DRB09] Stephen J. Davis, Christian H. Ritz, and Ian S. Burnett. Using social networking and collections to enable video semantics acquisition. IEEE MultiMedia, 16(4):52–61, October/ December 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Dimoulas:2015:SSM

[DS15] Charalampos A. Dimoulas and Andreas L. Symeonidis. Syncing shared multimedia through audiovisual bimodal segmentation. IEEE MultiMedia, 22(3):26–42, July/ September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/03/mmu2015030026-
abs.html.

DeFanti:1996:VVC

[DSB96] Tom DeFanti, Dan Sandin, and Maxine Brown. Visions and views: The coming defenestration: Immersive environments without windows. IEEE MultiMedia, 3(4):6–9, Winter 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1996/pdf/
u4006.pdf.

Dutta:2021:RTA

[DSGG21] Tanima Dutta, Aishwarya Soni, Prateek Gona, and Hari Prabhat Gupta. Real testbed for autonomous anomaly detection in power grid using low-cost unmanned aerial vehicles and aerial imaging. IEEE MultiMedia, 28(3):63–74, July/ September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Donderler:2003:BVD

[DSUG03] Mehmet Emin Dönderler, Ediz Saykol, Özgür Ulusoy, and Ugur Güdükbay. Bil-Video: a video database management system. IEEE MultiMedia, 10(1):66–70, January/March 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
01/extras/u1066.htm;
http://csdl.computer.
org/comp/mags/mu/2003/
01/u1066abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/01/u1066.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/01/
u1066.pdf.

Doganata:1994:MCE

[DT94] Yurdaer N. Doganata and Asser N. Tantawi. Making a cost-effective video server. IEEE MultiMedia, 1(4):22–30, Winter 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

DAronco:2017:PBC

[DTF17] Stefano D’Aronco, Laura Toni, and Pascal Frossard. Price-based controller for utility-aware HTTP adaptive streaming. IEEE MultiMedia, 24(2):20–29, April/June 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
02/mmu2017020020-abs.
html.

Doller:2008:SMQ

[DTG^{+}08] Mario Döller, Ruben Tous, Matthias Gruhne, Kyoungro Yoon, Masanori Sano, and Ian S. Burnett. Standards: The MPEG query format: Unifying access to multimedia retrieval systems. IEEE MultiMedia, 15(4):82–95, October/ December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Doller:2013:JJS

[DTT^{+}13] Mario Doller, Ruben Tous, Frederik Temmermans, Kyoungro Yoon, Je-Ho Park, Youngseop Kim, Florian Stegmaier, and Jaime Delgado. JPEG’s JPSearch standard: Harmonizing image management and search. IEEE MultiMedia, 20(4):38–48, October/ December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Dustdar:2000:GEI

[Dus00] Schahram Dustdar. Guest Editor’s introduction: Multimedia computer-supported cooperative work. IEEE MultiMedia, 7(4):24–25, October/December 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.

org/mu/books/mu2000/pdf/
u4024.pdf.

DelBimbo:1996:VPV

[DV96] Alberto Del Bimbo and Enrico Vicario. Visual programming of virtual worlds animation. IEEE MultiMedia, 3(1):40–49, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u1040.
pdf;   http://www.computer.
org/multimedia/mu1996/
u1040abs.htm.

Dorai:2001:CMA

[DV01] Chitra Dorai and Svetha Venkatesh. Computational media aesthetics: Finding meaning beautiful. IEEE MultiMedia, 8(4):10–12, October/ December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u4010.
pdf;   http://www.computer.
org/multimedia/mu2001/
u4010abs.htm.

Dorai:2003:GEI

[DV03] Chitra Dorai and Svetha Venkatesh. Guest Editors’ introduction: Bridging the semantic gap with computational media aesthetics. IEEE MultiMedia, 10(2):15–17, April/ June 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/02/u2015.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/02/
u2015.pdf.

Dittmann:2001:UCW

[DWN01] Jana Dittmann, Petra Wohlmacher, and Klara Nahrstedt. Using cryptographic and watermarking algorithms. IEEE Multi-Media, 8(4):54–65, October/December 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u4054.
pdf;   http://www.computer.
org/multimedia/mu2001/
u4054abs.htm.

Ding:2023:EMS

[DWX23] Yuqing Ding, Zhongcheng Wu, and Liyang Xie. Enabling manageable and secure hybrid P2P-CDN video-on-demand streaming services through coordinating blockchain and zero knowledge. IEEE MultiMedia, 30(1):36–51, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Duan:2008:DCC

[DWZ^{+}08] Ling-Yu Duan, Jinqiao Wang, Yan-Tao Zheng, Hanqing Lu, and Jesse S. Jin. Digesting commercial clips from TV streams. IEEE MultiMedia, 15(1):28–41, January/ March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Dimitrova:2000:VVE

[DYG^{+}00] Nevenka Dimitrova, Heather Yu, Francis Galliano, Rob Koenen, Avideh Zakhor, and Charles Bouman. Visions and views: Entry into the content forest: The role of multimedia portals; Web portals. IEEE MultiMedia, 7(3):14–20, July/ September 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u3014.
pdf.

Doller:2009:SMQ

[DYT^{+}09] Mario Doller, Armelle Ndjafa Natacha Yakou, Ruben Tous, Jaime Delgado, Matthias Gruhne, Miran Choi, and Tae-Beom Lim. Semantic MPEG query format validation and processing. IEEE MultiMedia, 16(4):22–33, October/December 2009. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Dimitrova:2002:AVC

[DZS^{+}02] Nevenka Dimitrova, Hong-Jiang Zhang, Behzad Shahraray, Ibrahim Sezan, Thomas Huang, and Avideh Zakhor. Applications of video-content analysis and retrieval. IEEE Multi-Media, 9(3):42–55, July/ September 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
03/u3042abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/03/u3042.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/03/
u3042.pdf.

Eakins:1998:SRT

[EBG98] John P. Eakins, Jago M.

Boardman, and Margaret E. Graham. Similarity retrieval of trademark images. IEEE MultiMedia, 5(2):53–63, April/ June 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u2053.
pdf;   http://www.computer.
org/multimedia/mu1998/
u2053abs.htm.

Effelsberg:1998:GEI

[Eff98] Wolfgang Effelsberg. Guest Editor’s introduction: Music in multimedia systems.

IEEE MultiMedia, 5(3):16, July/September 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u3016.
                                        pdf.

Effenberg:2005:MSE

[Eff05] Alfred O. Effenberg. Movement sonification: Effects on perception and action. IEEE MultiMedia, 12(2): 53–59, April/June 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ebrahimi:2016:JPT

[EFPS16] Touradj Ebrahimi, Siegfried Foessel, Fernando Pereira, and Peter Schelkens. JPEG Pleno: Toward an efficient representation of visual reality. IEEE Multi-Media, 23(4):14–20, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        04/mmu2016040014-abs.
                                        html.

ElSaddik:2001:RMC

[EFS01] Abdulmotaleb El Saddik, Stefan Fischer, and Ralf Steinmetz. Reusable multimedia content in Web-based learning systems. IEEE MultiMedia, 8(3):30–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u3030.
pdf;   http://www.computer.
org/multimedia/mu2001/
u3030abs.htm.

Ebner:2005:LUH

[EH05] Martin Ebner and Andreas Holzinger. Lurking: An underestimated humancomputer phenomenon. IEEE MultiMedia, 12(4): 70–75, October/December 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Evans:2009:NV

[EHH09] Helen Evans, Heiko Hansen, and Joey Hagedorn. Nuage vert. IEEE MultiMedia, 16 (3):13–15, July/September 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

ElSaddik:2018:DTC

[El 18] A. El Saddik. Digital twins: The convergence of multimedia technologies. IEEE MultiMedia, 25(2): 87–92, April/June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

ElSaddik:2020:MTI

[El 20] Abdulmotaleb El Saddik.

Multimedia and the tactile Internet. IEEE MultiMedia, 27(1):5–7, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ennis:2000:VPE

[EL00] Gareth Ennis and Malcolm Lindsay. VRML possibilities: The evolution of the Glasgow Model. IEEE MultiMedia, 7(2):48– 51, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2048.
pdf;   http://www.computer.
org/multimedia/mu2000/
u2048abs.htm.

Ellis:2006:AMI

[EL06] Daniel P. W. Ellis and Keansub Lee. Accessing minimal-impact personal audio archives. IEEE MultiMedia, 13(4):30–38, October/December 2006. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Escobar-Molano:1997:CDS

[EMG97] Martha L. Escobar-Molano and Shahram Ghandeharizadeh. On coordinated display of structured video. IEEE MultiMedia, 4(3):62–75, July/ September 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3062.
pdf;   http://www.computer.
org/multimedia/mu1997/
u3062abs.htm.

Echiffre:1998:SMA

[EMM^{+}98] Marica Echiffre, Claudio Marchisio, Pietro Marchisio, Paolo Panicciari, and Silvia Del Rossi. Standards: MHEC-5 — aims, concepts, and implementation issues. IEEE MultiMedia, 5(1):84–91, January/March 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u1084.
pdf.

Emond:2009:MWH

[Emo09] Bruno Emond. Multimedia at work: Harvesting resources for recording concurrent videoconferences. IEEE MultiMedia, 16(2): 92–95, April/June 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Egelhaaf:1996:PRD

[EMS96] Christian Egelhaaf, Eckhard Moeller, and Peter School. Project reports: Developing distributed multimedia telecommunication applications. IEEE MultiMedia, 3(4):76– 81, Winter 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u4076.
                                        pdf.

Ensor:1998:PRM

[Ens98] J. Robert Ensor. Project reports: Multimedia communications projects. IEEE MultiMedia, 5(1):97–101, January/March 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u1097.
                                        pdf.

Esteve:2006:MWF

[EP06] Manuel Esteve and Carlos E. Palau. Multimedia at work: a flexible video streaming system for urban traffic control. IEEE MultiMedia, 13(1):78–83, January/March 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Erzin:2006:MPR

[EYT^{+}06] Engin Erzin, Yücel Yemez, A. Murat Tekalp, Aytül Erçil, Hakan Erdogan, and Hüseyin Abut. Multimodal person recognition for human-vehicle interaction. IEEE MultiMedia, 13(2):18–31, April/June 2006. CODEN IEMUE4.

ISSN 1070-986X (print), 1941-0166 (electronic).

Falco:2007:AMC

[Fal07] Charles M. Falco. Artful media: Computer vision and art. IEEE MultiMedia, 14(2):8–11, April/June 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Farella:2005:PCI

[FBB^{+}05a] Elisabetta Farella, Davide Brunelli, Luca Benini, Bruno Ricco, and Maria Elena Bonfigli. Pervasive computing for interactive virtual heritage. IEEE Multi-Media, 12(3):46–58, July/ September 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fernstrom:2005:HDI

[FBB05b] Mikael Fernstrom, Eoin Brazil, and Liam Bannon. HCI design and interactive sonification for fingers and ears. IEEE MultiMedia, 12(2):36–44, April/June 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Farias:2008:SDT

[FCA08] Mylène C. Q. Farias, Marcelo M. Carvalho, and Marcelo S. Alencar. Standards: Digital television broadcasting in Brazil.

IEEE MultiMedia, 15(2): 64–70, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fu:2019:RBE

[FCL^{+}19] Y. Fu, W. Chen, K. Lai, Y. Zhou, and J. Tang. Rank-based encoding features for stereo matching. IEEE MultiMedia, 26(4): 28–42, October/December 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ferrara:2022:NFM

[FCM^{+}22] Simone Ferrara, Lorenzo Ciccarelli, Amaya Jiménez Moreno, Shiruo Zhao, Yetish Joshi, Guido Meardi, and Stefano Battista. The next frontier for MPEG-5 LCEVC: From HDR and immersive video to the metaverse. IEEE MultiMedia, 29(4):111–122, October/December 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fernandes:2015:GMS

[FDM^{+}15] Felix C. Fernandes, Xavier Ducloux, Zhan Ma, Esmaeil Faramarzi, Patrick Gendron, and Jiangtao Wen. The green metadata standard for energyefficient video consumption. IEEE MultiMedia, 22(1):80–87, January/ March 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/01/mmu2015010080-
abs.html.

Feijo:2009:MIV

[Fei09] Bruno Feijó. Media impact: Visualization, digital content, and simulation. IEEE MultiMedia, 16(1):8– 12, January/March 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ferretti:2018:CMP

[Fer18]  S. Ferretti.  Clustering

of musical pieces through complex networks: An assessment over guitar solos. IEEE MultiMedia, 25(4): 57–67, October/December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Fuchs:2005:IIC

[FF05] Harald Fuchs and Nikolaus Farber. ISMA interoperability and conformance. IEEE MultiMedia, 12(2): 96–102, April/June 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fu:2020:DRS

[FFW^{+}20] Bo Fu, Shilin Fu, Liyan Wang, Yuhan Dong, and Yonggong Ren. Deep residual split directed graph convolutional neural networks for action recognition. IEEE MultiMedia, 27(4):9–17, October/ December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fan:2014:OLH

[FGC^{+}14] Baojie Fan, Hao Gao, Yang Cong, Yingkui Du, and Yangdong Tang. Online learning a high-quality dictionary and classifier jointly for multitask object tracking. IEEE Multi-Media, 21(4):56–66, October/December 2014. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu2014040056-
abs.html.

Fridrich:2001:DLS

[FGD01] Jessica Fridrich, Miroslav Goljan, and Rui Du. Detecting LSB steganography in color and grayscale images. IEEE MultiMedia, 8(4):22–28, October/December 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u4022.
pdf;   http://www.computer.
org/multimedia/mu2001/
u4022abs.htm.

Fernandez:2001:MSD

[FGL^{+}01] David Fernández, Ana B.

García, David Larrabeiti, Arturo Azcorra, Piotr Pacyna, and Zdzislaw Papir. Multimedia services for distant work and education in an IP/ATM environment. IEEE Multi-Media, 8(3):68–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u3068.
pdf;   http://www.computer.
org/multimedia/mu2001/
u3068abs.htm.

Femminella:2012:EJC

[FGR12] Mauro Femminella, Francesco Giacinti, and Gianluca Reali. An extended Java call control for the Session Initiation Protocol. IEEE MultiMedia, 19(4):60–71, October/December 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fotouhi:2009:GEI

[FGS09] Farshad Fotouhi, William I.

Grosky, and Peter Stanchev. Guest Editors’ introduction: The many faces of semantics. IEEE Multi-Media, 16(2):16–17, April/ June 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Falchi:2008:EVS

[FGSS08] Fabrizio Falchi, Claudio Gennaro, Pasquale Savino, and Peter Stanchev. Efficient video-stream filtering. IEEE MultiMedia, 15 (1):52–62, January/March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Friedland:2008:EMS

[FHK08a] Gerald Friedland, Wolfgang Hürst, and Lars Knipping. Educational multimedia: Special section introduction: Educational multimedia. IEEE Multi-Media, 15(3):54–56, July/ September 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Friedland:2008:MIM

[FHK08b] Gerald Friedland, Wolfgang Hürst, and Lars Knipping. Media impact: Multimedia education in computer science: a little bit of everything is not enough. IEEE MultiMedia, 15(2): 78–82, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fong:2004:DLW

[FHL04] A. C. M. Fong, S. C. Hui, and C. T. Lui. Ondemand learning for a wireless campus. IEEE Multi-Media, 11(4):50–60, October/December 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/04/u4050.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/04/
u4050.pdf.

Friedland:2012:IDY

[FJ12] Gerald Friedland and Yohan Jin. Industry dares you: The ACM Multimedia Grand Challenge 2011. IEEE MultiMedia, 19 (1):12, 16, January/March 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fleckles:2006:AMP

[Fle06] Daniel Fleckles. Artful media: Picturing system. IEEE MultiMedia, 13(2):8– 14, April/June 2006. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fu:2017:MCD

[FLZ17] Silvery Fu, Jiangchuan Liu, and Wenwu Zhu. Multimedia content delivery with network function virtualization: The energy perspective. IEEE Multi-Media, 24(3):38–47, July/ September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        03/mmu2017030038-abs.
                                        html.

Francois:2005:VOF

[FNHB05] Alexandre R. J. Francois, Ram Nevatia, Jerry Hobbs, and Robert C. Bolles. VERL: An ontology framework for representing and annotating video events. IEEE MultiMedia, 12(4): 76–86, October/December 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fels:1998:MGM

[FNM98] Sidney Fels, Kazushi Nishimoto, and Kenji Mase. MusiKalscope: a graphical musical instrument. IEEE MultiMedia, 5(3):26–35, July/September 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u3026.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1998/
                                        u3026abs.htm.

Foote:2005:KOS

[Foo05] Jonathan Foote. Kooks, obsessives, Sturgeon’s Law, and the real meaning of search. IEEE MultiMedia, 12(3):4–7, July/September 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Foote:2006:MVW

[Foo06]  Jonathan Foote.  Me-

dia vision: Where’s my holodeck? The new frontier of media display. IEEE MultiMedia, 13(3): 104, 103, July/September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Foreman:2002:MDD

[For02] Dennis J. Foreman. Managing data in distributed multimedia conferencing applications. IEEE MultiMedia, 9(4):30–37, October/December 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
04/u4030abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/04/u4030.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/04/
u4030.pdf.

Fotouhi:2006:DNP

[Fot06a]  Farshad Fotouhi.  De-

partments: New products. IEEE MultiMedia, 13 (1):91–92, January/March 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fotouhi:2006:NP

[Fot06b] Farshad Fotouhi. New products. IEEE MultiMedia, 13(3):101–102, July/ September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fotouhi:2008:NPa

[Fot08a] Farshad Fotouhi. New products. IEEE Multi-Media, 15(3):88–89, July/ September 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fotouhi:2008:NPb

[Fot08b] Farshad Fotouhi. New products. IEEE MultiMedia, 15(4):C2–1, October/ December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fotouhi:2012:NP

[Fot12] Farshad Fotouhi. New products. IEEE Multi-Media, 19(3):88–c3, July/ September 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fu:2022:CKA

[FOW^{+}22] Yahui Fu, Shogo Okada, Longbiao Wang, Lili Guo, Yaodong Song, Jiaxing Liu, and Jianwu Dang. Context- and knowledgeaware graph convolutional network for multimodal emotion recognition. IEEE MultiMedia, 29(3):91–100, July/September 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Friedlander:1995:VVS

[Fri95] Larry Friedlander. Visions and views: Spaces for experience [multimedia presentation]. IEEE MultiMedia, 2(1):6–11, Spring 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u10006.
pdf.

Fortino:2007:CSC

[FRM^{+}07] Giancarlo Fortino, Wilma Russo, Carlo Mastroianni, Carlos E. Palau, and Manuel Esteve. CDNsupported collaborative media streaming control. IEEE MultiMedia, 14(2): 60–71, April/June 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Faver:2001:TDS

[FS01] Cheryl Faver and Getrude Stein. Toward a digital stage architecture: a long-term research agenda in digitally enabled theater. IEEE MultiMedia, 8(4):6–9, October/ December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u4006.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2001/
                                        u4006abs.htm.

Fernandez:2004:SDT

[FS04] Marcel Fernandez and Miguel Soriano. Softdecision tracing in fingerprinted multimedia content. IEEE MultiMedia, 11(2):38–46, April/ June 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/mu/2004/02/u2038abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/02/
                                        u2038.htm;   http://csdl.
                                        computer.org/dl/mags/
                                        mu/2004/02/u2038.pdf.

Finegan:2008:ABV

[FSC08] Ann Finegan, Josephine Starrs, and Leon Cmielewski. Art beat: Visualizing data: Seeker’s affective interaction. IEEE MultiMedia, 15(2):16–19, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Franchi:1999:MI

[FTCS99] Antonio Franchi, Eyal Trachtman, Louis Christodoulides, and Jay Sengupta. Multimedia via Inmarsat. IEEE MultiMedia, 6(4):15–19, October/December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1999/pdf/
u4015.pdf;   http://www.
computer.org/multimedia/
mu1999/u4015abs.htm.

Furht:1994:MSO

[Fur94] Borko Furht. Multimedia systems: An overview. IEEE MultiMedia, 1(1):47– 59, Spring 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Furht:1998:PRM

[FWI98] Borko Furht, Raymond Westwater, and Jeffrey Ice. Project reports: Multimedia broadcasting over the Internet. I. IEEE Multi-Media, 5(4):78–82, October/December 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u4078.
pdf.

Furht:1999:PRM

[FWI99] Borko Furht, Raymond Westwater, and Jeffrey Ice. Project reports: Multimedia broadcasting over the Internet. II. Video compression. IEEE Multi-Media, 6(1):85–89, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u1085.
                                        pdf.

Fu:2022:RID

[FWL22] Bo Fu, Liyan Wang, and Zhongxuan Luo. A robust image denoising method with multiview textureaware convolutional neural networks. IEEE Multi-Media, 29(3):80–90, July/ September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fan:2021:LBS

[FZH^{+}21] Chunling Fan, Yun Zhang, Raouf Hamzaoui, Djemel Ziou, and Qingshan Jiang. Learning-based satisfied user ratio prediction for symmetrically and asymmetrically compressed stereoscopic images. IEEE MultiMedia, 28(3):8–20, July/ September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fu:2024:DGA

[FZXL24] Dongxin Fu, Shaowu Zheng, Pengcheng Xie, and Weihua Li. Depth-guided aggregation for real-time binocular depth estimation network. IEEE MultiMedia, 31(2):36–47, April/ June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gayo-Avello:2015:SMD

[GA15] Daniel Gayo-Avello. Social media, democracy, and democratization. IEEE MultiMedia, 22(2):10–16, April/June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/02/mmu2015020010-
abs.html.

Grigoryan:2020:EGAa

[GA20a] A. M. Grigoryan and S. S.

Agaian. Evidence of golden and aesthetic proportions in colors of paintings of the prominent artists. IEEE MultiMedia, 27(1):8–16, January/March 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Grigoryan:2020:EGAb

[GA20b] Artyom M. Grigoryan and Sos S. Agaian. Evidence of golden and aesthetic proportions in colors of paintings of the prominent artists. IEEE MultiMedia, 27(1):8–16, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gibbs:1998:VSO

[GAB^{+}98] Simon Gibbs, Costas Arapis, Christian Breiteneder, Vali Lalioti, Sina Mostafawy, and Josef Speier. Virtual studios: An overview. IEEE MultiMedia, 5(1):18– 35, January/March 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1998/pdf/
u1018.pdf;   http://www.
computer.org/multimedia/
mu1998/u1018abs.htm.

Gorbova:2018:IVL

[GAL^{+}18] J. Gorbova, E. Avots, I. Lüsi, M. Fishel, S. Escalera, and G. Anbarjafari. Integrating vision and language for first-impression personality analysis. IEEE MultiMedia, 25(2):24–33, April/June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Glinert:1996:GEI

[GB96] Ephraim P. Glinert and Meera M. Blattner. Guest Editors’ introduction: Multimodal interaction. IEEE MultiMedia, 3(4):13, Winter 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1996/pdf/
u4013.pdf.

Golshani:1998:NDW

[GB98] Forouzan Golshani and Robin Baldwin. In the news: Digital-watermarking faces challenges; virtual human agents evolve; vendors urge, give SET a chance; radio on the net; media notes. IEEE MultiMedia, 5(3):6–9, July/ September 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u3006.
pdf.

Gunes:2015:SC

[GBH15] Hatice Gunes, Susanne Boll, and Alan Hanjalic. Scientific conferences. IEEE MultiMedia, 22(1):93–95, January/ March 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/01/mmu2015010093.
html.

Grosky:1997:NCC

[GBL97] William Grosky, Robin Baldwin, and Anne C. Lear. In the news: Customers cool to push-y technology; Taiwan’s telecom comes of age; shrinking images; streets paved with silicon: Malaysia’s multimedia super corridor. IEEE Multi-Media, 4(3):5–7, July/ September 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3005.
pdf.

Gerovac:1996:SSH

[GC96] Branko J. Gerovac and David C. Carver. Standards: Standardizing headers in the quest for interoperability. IEEE MultiMedia, 3(1):84–88, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u1084.

pdf. Comments of the use of Unicode for multimedia data storage.

Grachten:2017:TCA

[GCCGW17] Maarten Grachten, Carlos Eduardo Cancino-Chacon, Thassilo Gadermaier, and Gerhard Widmer. Toward computerassisted understanding of dynamics in symphonic music. IEEE MultiMedia, 24(1):36–46, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/

01/mmu2017010036-abs.
html.

Girod:2011:MVS

[GCGR11] Bernd Girod, Vijay Chandrasekhar, Radek Grzeszczuk, and Yuriy A. Reznik. Mobile visual search: Architectures, technologies, and the emerging MPEG standard. IEEE MultiMedia, 18 (3):86–94, July/September 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gupta:2012:CDF

[GCK12] Swati Gupta, Seongho Cho, and C.-C. Jay Kuo. Current developments and future trends in audio authentication. IEEE MultiMedia, 19(1):50–59, January/March 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gidney:1994:CFT

[GCM94] Eric Gidney, Annmarie Chandler, and Greg Mc-Farlane. CSCW for film and TV preproduction. IEEE MultiMedia, 1(2):16– 26, Summer 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gomez:2001:CMD

[GCP01] Jaime Gómez, Cristina Cachero, and Oscar Pastor. Conceptual modeling of device-independent Web applications. IEEE MultiMedia, 8(2):26–??, April/June 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u2026.
pdf;   http://www.computer.
org/multimedia/mu2001/
u2026abs.htm.

Gao:2014:VBO

[GD14] Yue Gao and Qionghai Dai. View-based 3D object retrieval: Challenges and approaches. IEEE Multi-Media, 21(3):52–57, July/ September 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gecsei:1997:ADM

[Gec97]  Jan Gecsei.  Adapta-

tion in distributed multimedia systems. IEEE MultiMedia, 4(2):58–66, April/June 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2058.
pdf;   http://www.computer.
org/multimedia/mu1997/
u2058abs.htm.

Gao:2023:IML

[GG23] Lei Gao and Ling Guan. Interpretability of machine learning: Recent advances and future prospects. IEEE MultiMedia, 30(4):105– 118, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gao:2017:CCS

[GGX^{+}17] Zhigang Gao, Hongyi Guo, Yunfeng Xie, Yanjun Luo, Huijuan Lu, and Ke Yan. ChildGuard: A child-safety monitoring system. IEEE MultiMedia, 24(4):48–57, October/December 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
04/mmu2017040048.html.

Gunes:2015:ESS

[GH15] Hatice Gunes and Hayley Hung. Emotional and social signals: A neglected frontier in multimedia computing? IEEE MultiMedia, 22(2):76–85, April/June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/02/mmu2015020076-
abs.html.

Gemmell:1994:DSM

[GHBC94] Jim Gemmell, Jiawei Han, Richard J. Beaton, and Stavros Christodoulakis. Delay-sensitive multimedia on disks. IEEE MultiMedia, 1(3):56–67, Autumn/Fall 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Guan:2014:EBG

[GHD^{+}14] Tao Guan, Yunfeng He, Liya Duan, Jianzhong Yang, Juan Gao, and Junqing Yu. Efficient BOF generation and compression for on-device mobile visual location recognition. IEEE MultiMedia, 21(2): 32–41, April/June 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Goose:2000:MTF

[GHR00] Stuart Goose, Wendy Hall, and Siegfried Reich. Microcosm TNG: a framework for distributed open hypermedia. IEEE MultiMedia, 7(3):52–60, July/ September 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u3052.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2000/
                                        u3052abs.htm.

Ghani:2011:PWK

[GI11] Dahlan Abdul Ghani and Mohd Sidin Bin Ahmad Ishak. Preserving Wayang Kulit for future generations. IEEE MultiMedia, 18(4):70–74, October/ December 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gibbs:1998:GEI

[Gib98] Simon Gibbs. Guest Editor’s introduction: Virtual studios. IEEE MultiMedia, 5(1):17, January/ March 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u1017.
pdf.

Gringeri:1998:TMV

[GKL^{+}98] Steven Gringeri, Bhumip Khasnabish, Arianne Lewis, Khaled Shuaib, Roman Egorov, and Bert Basch. Transmission of MPEG-2 video streams over ATM. IEEE MultiMedia, 5(1):58–71, January/March 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u1058.
pdf;   http://www.computer.
org/multimedia/mu1998/
u1058abs.htm.

Gay:1997:SMO

[GL97a] Valérie Gay and Peter Leydekkers. Standards: Multimedia in the ODP-RM standard. IEEE Multi-Media, 4(1):68–73, January/March 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u1068.
pdf.

Grosky:1997:NSV

[GL97b] William Grosky and Anne C. Lear. In the news: Smart but vulnerable; transitioning to digital TV; gaining voices; business briefs; technology notes; phoning the Web; converging competition; Intel presents MMX — and another upgrade dilemma; want more information? IEEE MultiMedia, 4(1):12–16, January/March 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u1012.
pdf.

Geigel:2003:UGA

[GL03] Joe Geigel and Alexander Loui. Using genetic algorithms for album page layouts. IEEE MultiMedia, 10(4):16–27, October/December 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
04/u4016abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/04/u4016.

htm;   http://csdl.computer.
org/dl/mags/mu/2003/04/
u4016.pdf.

Guo:2010:HMW

[GL10] Jing-Ming Guo and Yun-Fu Liu. Hiding multitone watermarks in halftone images. IEEE MultiMedia, 17(1):65, January/March 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Guan:2020:MTC

[GL20] Yuanyuan Guan and Ximing Li. Multilabel text classification with incomplete labels: a safe generative model with label manifold regularization and confidence constraint. IEEE MultiMedia, 27(4):38–47, October/December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Guo:2021:MAC

[GLB^{+}21] Yifei Guo, Bing Li, Xianye Ben, Yi Ren, Junping Zhang, Rui Yan, and Yujun Li. A magnitude and angle combined optical flow feature for microexpression spotting. IEEE MultiMedia, 28(2): 29–39, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Guo:2024:EIK

[GLLW24] Lanqing Guo, Yuxin Lin, Jian Li, and Bihan Wen. Exploiting illumination knowledge in the real world for low-light image enhancement. IEEE Multi-Media, 31(1):33–41, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Guerreiro:2008:TTM

[GLN^{+}08] Tiago Guerreiro, Paulo Lagoá, Hugo Nicolau, Daniel Gonçalves, and Joaquim A. Jorge. From tapping to touching: Making touch screens accessible to blind users. IEEE MultiMedia, 15(4):48–50, October/December 2008. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ginige:1995:HA

[GLR95] Athula Ginige, David B. Lowe, and John Robertson. Hypermedia authoring. IEEE MultiMedia, 2(4):24– 35, Winter 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u40024.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1995/
                                        u4024abs.htm.

Gao:2020:DDR

[GLZW20] Wang Gao, Lin Li, Xun Zhu, and Yuwei Wang. Detecting disaster-related tweets via multimodal adversarial neural network. IEEE MultiMedia, 27(4): 28–37, October/December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ginige:2001:GEIb

[GM01a] Athula Ginige and San Murugesan. Guest Editors’ introduction: The essence of Web engineering: Managing the diversity and complexity of Web application development. IEEE MultiMedia, 8(2):22– ??, April/June 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u2022.
pdf;   pdf/u2022.pdf.

Ginige:2001:GEIa

[GM01b] Athula Ginige and San Murugesan. Guest Editors’ introduction: Web engineering: An introduction. IEEE Multi-Media, 8(1):14–18, January/March 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u1014.
pdf.

Gopalakrishnan:2001:PWV

[GMV01] K. Gopalakrishnan, Nasir Memon, and Poorvi L. Vora. Protocols for watermark verification. IEEE MultiMedia, 8(4):66–70, October/December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://dlib.computer.
                                        org/mu/books/mu2001/pdf/
                                        u4066.pdf;   http://www.
                                        computer.org/multimedia/
                                        mu2001/u4066abs.htm.

Grosky:1996:NAH

[GO96] William L. Grosky and Michael Olan. In the news: Apple hopes to ride Internet, multimedia wave back to health; Forum promotes product standards; Asian perspectives on the Internet; focus: The Internet in Malaysia; digital watermarking stakes a claim on the Web; Chinese PC keyboard: a problem with a multimedia solution? Intel gets graphic with new chips. IEEE MultiMedia, 3(2):6– 9, Summer 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u2006.
                                        pdf.

Godbersen:2008:MIV

[God08] Heinrich Godbersen. Media impact: Virtual environments for anyone. IEEE MultiMedia, 15(3):90–95, July/September 2008. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Golshani:1994:IGM

[Gol94] Forouzan Golshani. Images galore: multimedia at Arizona State University. IEEE MultiMedia, 1(4):71–73, Winter 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Golshani:1995:MRS

[Gol95] Forouzan Golshani. Media reviews: On-the-shelf solutions. IEEE Multi-Media, 2(3):76, Fall 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Golshani:1996:MRS

[Gol96] Forouzan Golshani. Media reviews: Off-the-shelf solutions. IEEE Multi-Media, 3(2):86–87, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2083.
pdf.

Golshani:1997:MRSa

[Gol97a] Forouzan Golshani. Media reviews: On the shelf. IEEE MultiMedia, 4(1):83–85, January/ March 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u1083.
pdf.

Golshani:1997:MRSb

[Gol97b] Forouzan Golshani. Media reviews: On the shelf. IEEE MultiMedia, 4(3):93, July/September 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3091.
pdf.

Golshani:1997:MRT

[Gol97c] Forouzan Golshani. Media reviews: Teaching cyber style: C and C ++ Multimedia Cyber Classroom CD-ROM by Harvey Deitel and Paul Deitel (Prentice Hall PTR, 1995, $69.95, ISBN 0-13-231374-X); on the shelf. IEEE MultiMedia, 4(2):93– 94, April/June 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2093.
pdf.

Golshani:1998:NKM

[Gol98] Forouzan Golshani. In the news: Kiosks make their mark; remember when push was king?; multimedia notes; copyright update. IEEE Multi-Media, 5(2):9–13, April/ June 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u2009.
pdf.

Golshani:2002:EMU

[Gol02a] Forouzan Golshani. EIC’s message: Ushering in a new era. IEEE Multi-Media, 9(1):4, January/ March 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2002/pdf/u1004.
pdf;   http://www.computer.
org/multimedia/mu2002/
u1004abs.htm.

Golshani:2002:EDC

[Gol02b] Forouzan Golshani. Experiments in digital content. IEEE MultiMedia, 9(3):1– ??, July/September 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
03/u3001abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/03/u3001.
htm;   http://csdl.computer.

                                        org/dl/mags/mu/2002/03/
                                        u3001.pdf.

Golshani:2002:MTA

[Gol02c] Forouzan Golshani. From multimedia tools to artistic content. IEEE MultiMedia, 9(4):1–??, October/December 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2002/
                                        04/u4001abs.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mu/2002/04/u4001.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/04/
                                        u4001.pdf.

Goldman:2003:UTE

[Gol03a] Jerry Goldman. To undeaf their ears: The spoken word in a multimedia world. IEEE Multi-Media, 10(4):8–11, October/December 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2003/
                                        04/u4008abs.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mu/2003/04/u4008.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/04/
                                        u4008.pdf.

Golshani:2003:CBM

[Gol03b] Forouzan Golshani. Computational biometrics, multimedia analysis, and security. IEEE MultiMedia, 10(2):c2, 1, April/ June 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/02/u20c2.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/02/
u20c2.pdf.

Golshani:2003:EME

[Gol03c] Forouzan Golshani. EIC’s message: Expanding the scope of IEEE MultiMedia. IEEE MultiMedia, 10(1):c2, January/ March 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/01/u10c2.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/01/
u10c2.pdf.

Golshani:2003:EMR

[Gol03d] Forouzan Golshani. EIC’s message: Red, like the sound of a trumpet. IEEE MultiMedia, 10(3): c2, July/September 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
03/u30c2.pdf;   http:/
/csdl.computer.org/dl/

                                        mags/mu/2003/03/u30c2.
                                        htm.

Golshani:2003:EMS

[Gol03e] Forouzan Golshani. EIC’s message: Science, myths, and multimedia. IEEE MultiMedia, 10(4):c2, 1, October/December 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2003/
                                        04/u40c2.pdf;   http:/
                                        /csdl.computer.org/dl/
                                        mags/mu/2003/04/u40c2.
                                        htm.

Golshani:2004:EMD

[Gol04a] Forouzan Golshani. EIC’s message: Do you really know what you know? IEEE MultiMedia, 11 (4):c2, October/December 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2004/
                                        04/u40c2.pdf;   http:/
                                        /csdl.computer.org/dl/
                                        mags/mu/2004/04/u40c2.
                                        htm.

Golshani:2004:EMMb

[Gol04b] Forouzan Golshani. EIC’s message: Multimedia information lifecycle management. IEEE Multi-Media, 11(2):1–??, April/ June 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

csdl.computer.org/comp/
mags/mu/2004/02/u2001.
pdf;   http://csdl.computer.
org/dl/mags/mu/2004/02/
u2001.htm.

Golshani:2004:EMMa

[Gol04c] Forouzan Golshani. EIC’s message: Multimedia is correlated media. IEEE MultiMedia, 11(1):C2, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
01/u10c2.pdf;   http:/
/csdl.computer.org/dl/
mags/mu/2004/01/u10c2.
htm.

Golshani:2004:EMS

[Gol04d] Forouzan Golshani. EIC’s message: Still awaiting the killer app! IEEE MultiMedia, 11(3):c2, July/ September 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
03/u30c2.pdf;   http:/
/csdl.computer.org/dl/
mags/mu/2004/03/u30c2.
htm.

Golshani:2006:EMH

[Gol06a] Forouzan Golshani. EIC’s message: Happy trails. IEEE MultiMedia, 13(1):1, January/March 2006. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Golshani:2006:MVMa

[Gol06b] Forouzan Golshani. Media vision: Multimedia and reality. IEEE Multi-Media, 13(1):96, January/ March 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Golshani:2006:MVMb

[Gol06c] Forouzan Golshani. Media vision: Multimedia interfaces: Don’t move my cheese! IEEE MultiMedia, 13(2):96, 95, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Golshani:2007:MVH

[Gol07a] Forouzan Golshani. Media vision: Honeymoon on Mars. IEEE MultiMedia, 14(2):88, 87, April/June 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Golshani:2007:MVT

[Gol07b] Forouzan Golshani. Media vision: TUI or GUI– it’s a matter of somatics. IEEE MultiMedia, 14(1): 104, 103, January/March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Golshani:2008:MVD

[Gol08a] Forouzan Golshani. Media vision: Digital textbooks. IEEE MultiMedia, 15(2):c2, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Golshani:2008:MVA

[Gol08b] Forouzan Golshani. Media vision: The all-in-one box. IEEE MultiMedia, 15(3): 1, July/September 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gonzalez:2000:DM

[Gon00] Ruben Gonzalez. Disciplining multimedia. IEEE MultiMedia, 7(3):72–78, July/ September 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u3072.
pdf;   http://www.computer.
org/multimedia/mu2000/
u3072abs.htm.

Goodwin:2003:DAC

[Goo03] Michael Goodwin. Demystifying audio coding technologies. IEEE MultiMedia, 10(3):107–110, July/ September 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/03/u3107.

htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/
u3107.pdf.

Gordon:2005:FLL

[Gor05] Andrew S. Gordon. The fictionalization of lessons learned. IEEE MultiMedia, 12(4):12–14, October/ December 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gougher:1998:MRC

[Gou98a] Robert A. Gougher. Media reviews: Case studies in multimedia: Interactivity in Action: Case Studies of Multimedia Masterworks, Innovative Games, and Other Successful Interactive Products edited by Dominic Milano (Miller Freeman Books 1997, $39.95, ISBN 0-87930-481-2). IEEE MultiMedia, 5(2): 90, April/June 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u2090.
pdf.

Gougher:1998:MRMa

[Gou98b] Robert A. Gougher. Media reviews: Mastering 3D Web sites 3D Webmaster Version 5.50, (Superscape Virtual Reality Software, 1998, $995). IEEE MultiMedia, 5(3):70, July/ September 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u3069.
pdf.

Gougher:1998:MRMb

[Gou98c] Robert A. Gougher. Media reviews: Mastering 3D Web sites 3D Webmaster Version 5.50, (Superscape Virtual Reality Software, 1998, $995). IEEE MultiMedia, 5(3):70, July/ September 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u3069.
pdf.

Gougher:1999:MRPa

[Gou99a] R. A. Gougher. Media reviews: PC graphics programming (book review). IEEE MultiMedia, 6(3):93, July/September 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u3093.
pdf.

Gougher:1999:RCW

[Gou99b] R. A. Gougher. Reviewing a 3D creature workshop [media review]. IEEE MultiMedia, 6(4):92, October/ December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gougher:1999:MRH

[Gou99c] Robert A. Gougher. Media reviews: a handbook for the program formerly known as hash: The Animation:Master Handbook by Jeff Paries (Charles River Media, Rockland, Mass., $49.95, 450 pp, book and companion CD-ROM, ISBN 1-886801-71-1). IEEE MultiMedia, 6(2):92–93, April/ June 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u2091.
pdf.

Gougher:1999:MRPb

[Gou99d] Robert A. Gougher. Media reviews: 3D Photorealism Toolkit by Bill Fleming (John Wiley & Sons, New York, 1998, $39.99, 320 pp., ISBN 0-471-25346-4); workshop for computer animation: 3D Creature Workshop by Bill Fleming (Charles River Media, Rockland, Mass., 1998, $49.95, 400 pp., ISBN 1-886801-78-9). IEEE MultiMedia, 6(4):91–92, October/December 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u4091.
pdf.

Gougher:1999:MRS

[Gou99e] Robert A. Gougher. Media reviews: Sorting the Web by subject: Eastman Kodak NetSci-Fi and Net-Study edited by Michael Wolff (Dell Publishing 1997, $6.99 each, ISBN 0-440-22423-3). IEEE MultiMedia, 6(1):92–93, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u1090.
pdf.

Gougher:2000:MRC

[Gou00] Rob Gougher. Media reviews: Computer graphics explored: Reviewing Andrew Glassner’s Notebook: Recreational Computer Graphics by Andrew Glassner (Morgan Kaufmann Publishers, San Francisco, Calif., 1999, $44.95, 304 pp., ISBN 1-55860-598-3). IEEE MultiMedia, 7(1):85, January/ March 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u1085.
pdf.

Goudarzi:2015:DIA

[Gou15] Visda Goudarzi. Designing an interactive audio interface for climate science. IEEE MultiMedia, 22(1):41–47, January/ March 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/01/mmu2015010041-
                                        abs.html.

Guillemot:2015:KWR

[GOVI15] Mael Guillemot, Jean-Marc Odobez, Alessandro Vinciarelli, and Sandy Ingram. Klewel webcast: From research to growing company. IEEE Multi-Media, 22(4):94–99, October/December 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/04/mmu2015040094-
                                        abs.html.

Ghaleb:2020:MLBa

[GPA20a] E. Ghaleb, M. Popa, and S. Asteriadis. Metric learning-based multimodal audio-visual emotion recognition. IEEE MultiMedia, 27(1):37–48, January/March 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ghaleb:2020:MLBb

[GPA20b] Esam Ghaleb, Mirela Popa, and Stylianos Asteriadis. Metric learningbased multimodal audiovisual emotion recognition. IEEE MultiMedia, 27 (1):37–48, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Guha:1995:CPD

[GPLR95] Aloke Guha, Allalaghatta Pavan, Jonathan C. L. Liu, and Barry A. Roberts. Controlling the process with distributed multimedia. IEEE Multi-Media, 2(2):20–29, Summer 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u20020.
pdf;   http://www.computer.
org/multimedia/mu1995/
u2020abs.htm.

Goudarzi:2010:ORA

[GQ10] Pejman Goudarzi and Masoud Mohammadzadeh Qaratlu. Optimal rate allocation for video transmission over wireless ad hoc networks. IEEE MultiMedia, 17(1):142, January/March 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gao:2023:PAH

[GQFZ23] Guopeng Gao, Chuan Qin, Yaodong Fang, and Yuanding Zhou. Perceptual authentication hashing for digital images with contrastive unsupervised learning. IEEE MultiMedia, 30(3):129–140, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Grosky:2017:CRC

[GR17] William I. Grosky and Terry L. Ruas. The continuing reinvention of contentbased retrieval: Multimedia is not dead. IEEE MultiMedia, 24(1):6–11, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        01/mmu2017010006-abs.
                                        html.

Goose:2016:PVP

[GRF^{+}16] Stuart Goose, Larry Riddle, Chris Fuller, Tarun Gupta, and Aaron Marcus. PAZ: In-vehicle personalized audio zones. IEEE MultiMedia, 23(4):32–41, October/December 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/

04/mmu2016040032-abs.
html.

Gupta:2021:MPE

[GRG21] Varun Gupta, Luis Rubalcaba, and Chetna Gupta. Multimedia prototyping for early-stage startups endurance: Stage for new normal? IEEE MultiMedia, 28(4):107–116, October/December 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Grinstein:1996:NCB

[Gri96]  Georges Grinstein.  In

the news: Creating benchmarks for information exploration; NASA seeks tech partnets; A new dimension to data storage; VR leaps from arcade to boardroom. IEEE MultiMedia, 3(4):10– 12, Winter 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u4010.
pdf.

Grosky:1994:MIS

[Gro94] William I. Grosky. Multimedia information systems. IEEE MultiMedia, 1(1):12–24, Spring 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Grosky:1995:NEY

[Gro95a] William Grosky. In the news: Europeans yawn at interactive services; Oracle weaves a wider Web; unlikely partners seek programming; interactive trials test the market; Sprint serves up a Web site; DreamWorks and SGI plan studio; Microsoft geras up for network on-line service; newspapers link for the Web; virtual reality notes. IEEE MultiMedia, 2(3):4–7, Fall 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u30004.
pdf.

Grosky:1995:NT

[Gro95b]  William Grosky.  In

the news: Telemedicine. IEEE MultiMedia, 2(1): 4–5, Spring 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u10004.
pdf.

Grosky:1995:NVS

[Gro95c] William Grosky. In the news: Videodisk standards battle averted; copyright protection needed; interoperability standard sought; HTML revisited; MSN faces rivals; Internet voice no challenge yet; banking and checking on the net; transactions in turmoil. IEEE MultiMedia, 2 (4):5–7, Winter 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u40005.
pdf.

Grosky:1996:NLD

[Gro96a] William Grosky. In the news: Life and death on the Intranet; mapping systems watch crime; technology tested at the Olympics; browser battles; adding graphics on the Web. IEEE MultiMedia, 3(3):7–9, Fall 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u3007.
pdf.

Grosky:1996:NWH

[Gro96b] William Grosky. In the news: Whither the high definition in HDTV? TV on the Net; telecommunications overhauled; videoconferencing market grows; ITU videoconferencing standards; credit on the net; Bell Labs brews a rival to Sun’s Java; 3D multimedia comes on a chip; new print output. IEEE MultiMedia, 3(1):

6–9, Spring 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u1006.

pdf. Discusses the Inferno programming language.

Grosky:1997:GAN

[Gro97a] W. Grosky. Grosky appointed new EIC for IEEE MultiMedia. IEEE MultiMedia, 4(4):8, October/ December 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Grosky:1997:IVA

[Gro97b] W. Grosky. Indexing video archives. IEEE Multi-Media, 4(4):7, October/ December 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Grosky:1997:NPS

[Gro97c] William Grosky. In the news: Pushing streaming video. IEEE Multi-Media, 4(4):6, October/ December 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1997/pdf/u4006.
                                        pdf.

Grosky:1998:EMT

[Gro98a] William Grosky. EIC’s message: a time for change. IEEE MultiMedia, 5(1):16, January/March 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u1016.
pdf.

Grosky:1998:NFS

[Gro98b] William Grosky. In the news: Fingerprint sensors solve security problems; Web 3D graphics; protecting copyright on the net; DSL, IP poised to spread widely. IEEE MultiMedia, 5(4):6–7, October/December 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u4006.
pdf.

Grosky:1999:EMM

[Gro99a] William Grosky. EIC’s message: Managing content. IEEE MultiMedia, 6(3):(cover 2), July/ September 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u30c2.
pdf.

Grosky:1999:EMW

[Gro99b] William Grosky. EIC’s message: Welcome, new board members. IEEE MultiMedia, 6(1):10, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u1010.
                                        pdf.

Grosky:2000:EMNb

[Gro00a] William Grosky. EIC’s message: New coverage. IEEE MultiMedia, 7(4):5, October/ December 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u4005.
                                        pdf.

Grosky:2000:EMNa

[Gro00b] William Grosky. EIC’s message: New faces, new columns. IEEE MultiMedia, 7(2):4, April/ June 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u2004.
                                        pdf.

Grosky:2000:EMW

[Gro00c] William Grosky. EIC’s message: What’s new; tell us what you think; new board member. IEEE MultiMedia, 7(3):5, July/ September 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u3005.
pdf.

Grosky:2001:EF

[Gro01]  William Grosky.  EIC

farewell. IEEE Multi-Media, 8(4):5, October/ December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u4005.
pdf;   http://www.computer.
org/multimedia/mu2001/
u4005abs.htm.

Guarda:2021:NAL

[GRP21] André F. R. Guarda, Nuno M. M. Rodrigues, and Fernando Pereira. Neighborhood adaptive loss function for deep learning-based point cloud coding with implicit and explicit quantization. IEEE MultiMedia, 28(3):107–116, July/ September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Goldman-Segall:1994:WSI

[GS94] Ricki Goldman-Segall. Whose story is it, anyway? an ethnographic answer. IEEE MultiMedia, 1(4):7– 11, Winter 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Grosky:1995:NDI

[GS95] William Grosky and Dave Sims. In the news: Defining intellectual property in cyberspace; standards for the World Wide Web; deregulation is in the air; Internet audio tools; commerce stumbles onto the net; virtual money for real transactions. IEEE MultiMedia, 2(2):4–6, Summer 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u20004.
pdf.

Goeschka:2001:CLI

[GS01] Karl M. Goeschka and Markus W. Schranz. Client and legacy integration in object-oriented Web engineering. IEEE Multi-Media, 8(1):32–41, January/March 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u1032.
pdf;   http://www.computer.
org/multimedia/mu2001/
u1032abs.htm.

Gemmell:2006:GEI

[GS06] Jim Gemmell and Hari Sundaram. Guest Editors’ introduction: Capture, archival, and retrieval of personal experience. IEEE MultiMedia, 13(4):18–19, October/ December 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Grzegorzek:2010:LWF

[GSOI10] Marcin Grzegorzek, Sorin Sav, Noel E. O’Connor, and Ebroul Izquierdo. Local wavelet features for statistical object classification and localization. IEEE MultiMedia, 17(1): 118, January/March 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Golshani:1997:MRW

[GSW97] Forouzan Golshani, Christoph F.

Strnadl, and Robert Wilmes. Media reviews: Webcasting presentations: ActivePresenter (SPC Software Publishing, 1997. $49.95 three-attendee version, $99.95 unlimited attendee version); explaining video and imaging processing: Reviewing Video and Image Processing in Multimedia Systems by Borko Furht, Stephen W. Smoliar, and Hongjian Zhang (Kluwer Academic Publishers, 1995, 377 pp., ISBN 0-7923-9604-9); authoring multimedia: MMD — A Multimedia Database and Authoring Tool (Capella Computers, 1997, $6,750); on the shelf. IEEE Multi-Media, 4(4):83–87, October/December 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1997/pdf/u4083.
                                        pdf.

Golder:2008:CTG

[GT08] Scott Golder and Jelena Tesić. Collaborative tagging: Guest Editors’ introduction: Collaborative tagging of multimedia. IEEE MultiMedia, 15(3):12–13, July/September 2008. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Guo:2011:DHH

[GT11] Jing-Ming Guo and Jia-Jin Tsai. Data-hiding in halftone images using adaptive noise-balanced error diffusion. IEEE Multi-Media, 18(2):48–59, April/ June 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Grafl:2013:SMC

[GTH^{+}13] Michael Grafl, Christian Timmerer, Hermann Hellwagner, George Xilouris, Georgios Gardikis, Daniele Renzi, Stefano Battista, Eugen Borcoci, and Daniel Negru. Scalable media coding enabling contentaware networking. IEEE MultiMedia, 20(2):30–41, April/June 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Guo:2016:NSS

[GTQG16] Xin Guo, Yun Tie, Lin Qi, and Ling Guan. A novel semi-supervised dimensionality reduction framework. IEEE MultiMedia, 23(2):28–41, April/ June 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL https://

www.computer.org/csdl/
mags/mu/2016/02/mmu2016020028-
abs.html.

Gemmell:2000:GAV

[GTZ^{+}00] Jim Gemmell, Kentaro Toyama, C. Lawrence Zitnick, Thomas Kang, and Steven Seitz. Gaze awareness for videoconferencing:A software approach. IEEE MultiMedia, 7(4):26–35, October/ December 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u4026.

                                        pdf;   http://www.computer.
                                        org/multimedia/mu2000/
                                        u4026abs.htm.

Guglielmetti:2003:SAM

[Gug03] Lori Guglielmetti. Standardizing automotive multimedia interfaces. IEEE MultiMedia, 10(2):76–78, April/June 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2003/02/u2076.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/02/
                                        u2076.pdf.

Gomez:2007:MRN

[GV07] Jaime Gómez and Jose Luis Vicedo. Media reviews: Next-generation multimedia database retrieval. IEEE MultiMedia, 14(3): 106–107, July/September 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gao:2017:VSC

[GWC17] Guanyu Gao, Yonggang Wen, and Jianfei Cai. vCache: Supporting costefficient adaptive bitrate streaming. IEEE Multi-Media, 24(3):19–27, July/ September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/

03/mmu2017030019-abs.
html.

Guo:2022:ERM

[GWD^{+}22] Lili Guo, Longbiao Wang, Jianwu Dang, Yahui Fu, Jiaxing Liu, and Shifei Ding. Emotion recognition with multimodal transformer fusion framework based on acoustic and lexical information. IEEE MultiMedia, 29(2):94–103, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gu:2022:ELG

[GWL^{+}22] Ruibin Gu, Qiuxia Wu, Yuqiong Li, Wenxiong Kang, Wing W. Y. Ng, and Zhiyong Wang. Enhanced local and global learning for rotation-invariant point cloud representation. IEEE MultiMedia, 29(4):24–37, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Grosky:1997:NGG

[GY97] William Grosky and Robin Yeary. In the news: Gaming grows up; Webvertising; business notes; multimedia on the line; MTAC 97. IEEE Multi-Media, 4(2):10–13, April/ June 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1997/pdf/u2010.
                                        pdf.

Gao:2023:PND

[GYL^{+}23] Wei Gao, Hang Yuan, Guibiao Liao, Zixuan Guo, and Jianing Chen. PP8K: a new dataset for 8K UHD video compression and processing. IEEE MultiMedia, 30(3):100–109, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Grosky:2009:IPM

[GZC09] William I. Grosky, Chengcui Zhang, and Shu-Ching Chen. Intelligent and pervasive multimedia systems. IEEE MultiMedia, 16 (1):14–15, January/March 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Gong:2024:VRB

[GZP^{+}24] Rui Gong, Xue Zhang, Jianan Pan, Jie Guo, and Xiushan Nie. Vehicle reidentification based on convolution and vision transformer feature fusion. IEEE MultiMedia, 31(2): 61–68, April/June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hagsand:1996:IMV

[Hag96]  Olof Hagsand.  Inter-

active multiuser VEs in the DIVE system. IEEE MultiMedia, 3(1):30–39, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u1030.
pdf;   http://www.computer.
org/multimedia/mu1996/
u1030abs.htm.

Hakkila:2015:MUM

[Hak15] Jonna Hakkila. Mobile and ubiquitous multimedia. IEEE MultiMedia, 22(1):96, January/ March 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/01/mmu2015010096.
html.

Hall:1994:ETB

[Hal94] Wendy Hall. Ending the tyranny of the button. IEEE MultiMedia, 1(1):60– 68, Spring 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Halpine:2001:AMM

[Hal01] Susana Maria Halpine.

Artful media: Molecular visualization: a microcosm of the E-revolution. IEEE MultiMedia, 8(2):4– ??, April/June 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u2004.
                                        pdf.

Hanjalic:2015:MSR

[Han15] Alan Hanjalic. Multimedia search: From relevance to usefulness. IEEE MultiMedia, 22(1):2–3, January/March 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/01/mmu2015010002-
                                        abs.html.

Hanjalic:2017:MRW

[Han17]  Alan Hanjalic.  Multi-

media research: What is the right approach? IEEE MultiMedia, 24(2):4– 6, April/June 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        02/mmu2017020004.html.

Hasselbring:2002:WDI

[Has02] Wilhelm Hasselbring. Web data integration for Ecommerce applications. IEEE MultiMedia, 9(1):16– 25, January/March 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://dlib.computer.
                                        org/mu/books/mu2002/pdf/
                                        u1016.pdf;   http://www.

computer.org/multimedia/
mu2002/u1016abs.htm.

Hatefi:1996:MRE

[Hat96] Farid Hatefi. Media reviews: Encryption technology explained: Building in Big Brothers — The Cryptography Policy Debate, Lance J. Hoffman, Editor (Springer-Verlag, 1995, $29.95, ISBN 0-387-94441-9); on the shelf. IEEE MultiMedia, 3(4):87– 88, Winter 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u4087.
pdf.

Haval:2000:TDD

[Hav00] Nikhilesh Haval. Threedimensional documentation of complex heritage structures. IEEE Multi-Media, 7(2):52–55, April/ June 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2052.
pdf;   http://www.computer.
org/multimedia/mu2000/
u2052abs.htm.

Hayashi:1998:ICB

[Hay98] Masaki Hayashi. Image compositing based on virtual cameras. IEEE MultiMedia, 5(1):36–48, January/March 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u1036.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1998/
                                        u1036abs.htm.

Hu:2004:CIW

[HB04] Jianying Hu and Amit Bagga. Categorizing images in Web documents. IEEE Multi-Media, 11(1):22–30, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2004/
                                        01/u1022abs.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mu/2004/01/u1022.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/01/
                                        u1022.pdf.

Hinterberger:2005:POS

[HB05] Thilo Hinterberger and Gerold Baier. Parametric orchestral sonification of EEG in real time. IEEE MultiMedia, 12(2):70–79, April/June 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2005/
                                        02/extras/u2070x1.mp3.

Habermann:2017:ASC

[HCAMJ17] Philipp Habermann, Chi Ching Chi, Mauricio Alvarez-Mesa, and Ben Juurlink. Application-specific cache and prefetching for HEVC CABAC decoding. IEEE MultiMedia, 24(1):72–85, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
01/mmu2017010072-abs.
html.

Huet:2012:LSM

[HCH12] Benoit Huet, Tat-Seng Chua, and Alexander Hauptmann. Large-scale multimedia data collections. IEEE MultiMedia, 19(3):12–14, July/ September 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Han:2023:AFT

[HCLK23] Guang Han, Chen Cao, Jixin Liu, and Sam Kwong. Anchor-free tracker based on space-time memory network. IEEE MultiMedia, 30(1):73–83, January/ March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

He:2004:ISP

[He04]  Zhihai He.  An inte-

grated system perspective for multimedia communication. IEEE Multi-Media, 11(1):88–89, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
01/u1088abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2004/01/u1088.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/01/
u1088.pdf.

Han:2011:MRS

[HFdW11] Jungong Han, Dirk Farin, and Peter H. N. de With. A mixed-reality system for broadcasting sports video to mobile devices. IEEE MultiMedia, 18(2):72–84, April/June 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hirzalla:1995:TMI

[HFK95] Na’el Hirzalla, Benjamin Falchuk, and Ahmed Karmouch. A temporal model for interactive multimedia scenarios. IEEE MultiMedia, 2(3):24–31, Fall 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u30024.
pdf;   http://www.computer.
org/multimedia/mu1995/
u3024abs.htm.

Habib:2009:CSD

[HG09] Ahsan Habib and Stuart Goose. CommunityPVR: a service to deliver the long tail for on-demand TV. IEEE MultiMedia, 16 (3):28–38, July/September 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hou:2015:SGD

[HG15] Weilong Hou and Xinbo Gao. Saliency-guided deep framework for image quality assessment. IEEE MultiMedia, 22(2):46–55, April/June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/02/mmu2015020046-
abs.html.

Hong:2023:NLD

[HG23] Phuoc-Thanh Hong and Ling Guan. A novel learning dictionary for sparse coding-based key point detection. IEEE MultiMedia, 30(4):47–60, 2023. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hao:2016:SAS

[HGHW16] Shijie Hao, Yanrong Guo, Richang Hong, and Meng Wang. Scale-aware spatially guided mapping. IEEE MultiMedia, 23(3): 34–42, July/September 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
03/mmu2016030034-abs.
html.

Hermann:2005:GEI

[HH05] Thomas Hermann and Andy Hunt. Guest Editors’ introduction: An introduction to interactive sonification. IEEE MultiMedia, 12(2):20–24, April/June 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2005/
02/u2020.pdf.

Hornecker:2018:CPO

[HH18] E. Hornecker and M. Honauer. Creating performance-oriented multimedia projects as part of an interdisciplinary teaching practice. IEEE MultiMedia, 25(2):93–103, April/June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Harman:1995:LDI

[HHI^{+}95] Dale D. Harman, Gang Huang, Gi-Hong Im, Mai-Huong Nguyen, Jean-Jacques Werner, and Michael K. Wong. Local distribution for IMTV. IEEE MultiMedia, 2(3):14–23, Fall 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u30014.
pdf;   http://www.computer.
org/multimedia/mu1995/
u3014abs.htm.

Hsieh:2020:WBQ

[HHLW20] Jui-Hung Hsieh, King-Chu Hung, Je-Hung Liu, and Tsung-Ching Wu. Waveletbased quality-constrained ECG data compression system without decoding process. IEEE MultiMedia, 27(2):33–45, April/June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hight:2008:ABD

[Hig08] Jeremy Hight. Art beat:

Directions in locative media. IEEE MultiMedia, 15(3):4–5, July/September 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hirose:1997:IBV

[Hir97] Michitaka Hirose. Imagebased virtual world generation. IEEE MultiMedia, 4(1):27–33, January/ March 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u1027.

                                        pdf;   http://www.computer.
                                        org/multimedia/mu1997/
                                        u1027abs.htm.

Henze:2017:DSL

[HK17] Niels Henze and Kai Kunze. A Dagstuhl Seminar looks beyond virtual and augmented reality. IEEE MultiMedia, 24(2): 14–17, April/June 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        02/mmu2017020014-abs.
                                        html.

Hsu:2007:RMV

[HKC07] Winston H. Hsu, Lyndon S. Kennedy, and Shih-Fu Chang. Reranking methods for visual search. IEEE MultiMedia, 14(3):14–22, July/September 2007. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Harada:2010:APM

[HKM^{+}10] Noboru Harada, Yutaka Kamamoto, Takehiro Moriya, Hendry, Houari Sabirin, and Munchurl Kim. Archive and preservation of media content using MPEG-A. IEEE MultiMedia, 17(4): 94–99, October/December 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hunt:2004:MMI

[HKN04] Andy Hunt, Ross Kirk, and Matt Neighbour. Multiple media interfaces for music therapy. IEEE Multi-Media, 11(3):50–58, July/ September 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/03/u3050.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/03/
u3050.pdf.

Hosio:2010:SSB

[HKR10] Simo Hosio, Hannu Kukka, and Jukka Riekki. Social surroundings: Bridging the virtual and physical divide. IEEE MultiMedia, 17(2): 26–33, April/June 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hinojo-Lucena:2021:IUM

[HLADCR^{+}21] Francisco-Javier Hinojo-Lucena, Inmaculada Aznar-Díaz, María-Pilar Cáceres-Reche, Juan-Manuel Trujillo-Torres, and José-María Romero-Rodríguez. Instagram use as a multimedia platform for sharing images and videos: Links to smartphone addiction and self-esteem. IEEE Multi-Media, 28(1):48–55, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hile:2010:VNM

[HLB^{+}10] Harlan Hile, Alan Liu, Gaetano Borriello, Radek Grzeszczuk, Ramakrishna Vedantham, and Jana Kosecka. Visual navigation for mobile devices. IEEE MultiMedia, 17(2):16–25, April/June 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hess:1998:VIM

[HLN98] Christopher K. Hess, Daniel Lin, and Klara Nahrstedt. VistaMail: An integrated multimedia mailing system. IEEE Multi-Media, 5(4):13–23, October/December 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u4013.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1998/
                                        u4013abs.htm.

Hudson:2017:JSG

[HLNS17] Graham Hudson, Alain Leger, Birger Niss, and Istvan Sebestyen. JPEG at 25: Still going strong. IEEE MultiMedia, 24(2): 96–103, April/June 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
02/mmu2017020096-abs.
html.

Huang:2014:VMH

[HLP^{+}14] Thomas Huang, Vuong Le, Thomas Paine, Pooya Khorrami, and Usman Tariq. Visual media: History and perspectives. IEEE MultiMedia, 21(2):4– 10, April/June 2014. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

He:2015:ICM

[HLT^{+}15] Xiangjian He, Suhuai Luo, Dacheng Tao, Changsheng Xu, and Jie Yang. The 21st International Conference on MultiMedia Modeling. IEEE MultiMedia, 22(2):86–88, April/ June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/02/mmu2015020086.
html.

Hussain:2023:EIL

[HLT23] Muhammad Awais Hussain, Chun-Lin Lee, and Tsung-Han Tsai. An efficient incremental learning algorithm for sound classification. IEEE Multi-Media, 30(1):84–90, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hua:2016:EGF

[HLW16] Kai-Lung Hua, Kai-Han Lo, and Yu-Chiang Frank Frank Wang. Extended guided filtering for depth map upsampling. IEEE Multi-Media, 23(2):72–83, April/ June 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL https://

                                        www.computer.org/csdl/
                                        mags/mu/2016/02/mmu2016020072-
                                        abs.html.

Hung:2019:ASA

[HLW^{+}19] S. Hung, Y. Lai, S. Wong, C. Chiu, and C. Yao. Arbitrary screen-aware manga reading framework with parameter-optimized panel extraction. IEEE Multi-Media, 26(2):55–65, April/ June 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Huang:2010:MMV

[HLY10] Chung-Ming Huang, Chung-Wei Lin, and Chia-Ching Yang. Mobility management for video streaming on heterogeneous networks. IEEE MultiMedia, 17(1): 35, January/March 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Heller:1995:MT

[HM95] Rachelle S. Heller and C. Dianne Martin. A media taxonomy. IEEE MultiMedia, 2(4):36–45, Winter 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u40036.
pdf;   http://www.computer.
org/multimedia/mu1995/
u4036abs.htm.

Ma:1998:VBH

[hMLDM98] Wei hsiu Ma, Yen-Jen Lee, David H. C. Du, and Mark P. McCahill. Video-based hypermedia for education-on-demand. IEEE MultiMedia, 5(1):72– 83, January/March 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1998/pdf/
u1072.pdf;   http://www.
computer.org/multimedia/
mu1998/u1072abs.htm.

Hemmati:2017:QAB

[HMS17] Mahdi Hemmati, Bill Mc-Cormick, and Shervin Shirmohammadi. QoEaware bandwidth allocation for video traffic using sigmoidal programming. IEEE MultiMedia, 24(4):80–90, October/December 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        04/mmu2017040080-abs.
                                        html.

Hausenblaus:2007:VVI

[HN07] Michael Hausenblaus and Frank Nack. Visions and views: Interactivity = reflective expressiveness. IEEE MultiMedia, 14(2): 1, 4–7, April/June 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Herbaut:2017:DDO

[HNDP17] Nicolas Herbaut, Daniel Negru, David Dietrich, and Panagiotis Papadimitriou. Dynamic deployment and optimization of virtual content delivery networks. IEEE Multi-Media, 24(3):28–37, July/ September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        03/mmu2017030028-abs.
                                        html.

Heylighen:2003:WYS

[HNM03] Ann Heylighen, Herman Neuckermans, and Pieter Morisse. What you see is what you get. IEEE MultiMedia, 10(3):48–56, July/September 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
03/u3048abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/03/u3048.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/
u3048.pdf.

Hoces:1995:MRF

[Hoc95] Carlos Hoces. Media reviews: Fractal compression theory: Image Compression Using Fractals, edited by Yuval Fisher (Springer-Verlag, 1995, ISBN 0-387-94211-4, 324 pp., $49.50). IEEE MultiMedia, 2(3): 76, Fall 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hoces:1996:MRB

[Hoc96] Carlos Hoces. Media reviews: a broad guide to data compression: Introduction to Data Compression by Khalid Sayood (Morgan Kaufmann Publishers, 1995, 600 pp., $64.95, ISBN 1-55860-346-8). IEEE MultiMedia, 3 (3):91–92, Fall 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u3090.
pdf.

Hoffman:1994:NCM

[Hof94]  Helene Hoffman.  Net-

work compatible multimedia for medical training: Med Pics. IEEE MultiMedia, 1(3):71–73, Autumn/Fall 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hirayu:2000:CHV

[HOK00] Hidekazu Hirayu, Takeo Ojika, and Ryugo Kijima. Constructing the historic villages of Shirakawa-go in virtual reality. IEEE MultiMedia, 7(2):61–64, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u2061.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2000/
                                        u2061abs.htm.

Holthe:2010:MPF

[Hol10] Ole-Ivar Holthe. A multimedia presentation framework for Web 3.0 computer game experiences. IEEE MultiMedia, 17(4): 70–79, October/December 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Horowitz:1995:NPa

[Hor95a] Bradley Horowitz. New products. IEEE MultiMedia, 2(1):103, Spring 1995.

CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1995/pdf/
u10103.pdf.

Horowitz:1995:NPb

[Hor95b] Bradley Horowitz. New products. IEEE Multi-Media, 2(2):94–95, Summer 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u20094.
pdf.

Horowitz:1995:NPc

[Hor95c] Bradley Horowitz. New products. IEEE MultiMedia, 2(3):77–79, Fall 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1995/pdf/
u30077.pdf.

Horowitz:1995:NPd

[Hor95d] Bradley Horowitz. New products. IEEE Multi-Media, 2(4):97–98, Winter 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1995/pdf/
u40097.pdf.

Horowitz:1996:NPa

[Hor96a] Bradley Horowitz. New products. IEEE Multi-Media, 3(1):94–95, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1996/pdf/
u1094.pdf.

Horowitz:1996:NPb

[Hor96b] Bradley Horowitz. New products. IEEE Multi-Media, 3(2):94–95, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2094.
pdf.

Horowitz:1996:NPc

[Hor96c] Bradley Horowitz. New products. IEEE MultiMedia, 3(3):94–95, Fall 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1996/pdf/
u3094.pdf.

Horowitz:1996:NPF

[Hor96d] Bradley Horowitz. New products: FreeD links photographs and 3D; server views images on the Web; Mercury Mail delivers; Web tool provides database functions; building Javabased tables and forms; viewing interactive maps on the net; managing multimedia on demand; Internet video solution offered. IEEE MultiMedia, 3(4):89– 90, Winter 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u4089.
pdf.

Horowitz:1997:NPa

[Hor97a] Bradley Horowitz. New products. IEEE MultiMedia, 4(2):95, April/ June 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2095.
pdf.

Horowitz:1997:NPb

[Hor97b] Bradley Horowitz. New products. IEEE Multi-Media, 4(3):94–95, July/ September 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3094.
pdf.

Horowitz:1997:NPc

[Hor97c] Bradley Horowitz. New products. IEEE Multi-Media, 4(4):88–89, October/December 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1997/pdf/u4088.
                                        pdf.

Horowitz:1997:NPH

[Hor97d] Bradley Horowitz. New products: Hardware; software. IEEE MultiMedia, 4(1):86–87, January/ March 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1997/pdf/u1086.
                                        pdf.

Horowitz:1998:NPa

[Hor98a] Bradley Horowitz. New products. IEEE Multi-Media, 5(1):104, January/ March 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u1104.
                                        pdf.

Horowitz:1998:NPb

[Hor98b] Bradley Horowitz. New products. IEEE Multi-Media, 5(2):93–95, April/ June 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u2093.
                                        pdf.

Horowitz:1998:NPc

[Hor98c] Bradley Horowitz. New products. IEEE Multi-Media, 5(3):94–95, July/ September 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u3094.
pdf.

Horowitz:1998:NPd

[Hor98d] Bradley Horowitz. New products. IEEE Multi-Media, 5(4):89, October/ December 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u4089.
pdf.

Horowitz:1999:NP

[Hor99] Bradley Horowitz. New products. IEEE Multi-Media, 6(1):94–95, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u1094.
pdf.

Hoschka:1998:SIS

[Hos98] Philipp Hoschka. Standards: An introduction to the Synchronized Multimedia Integration Language; World Wide Web Consortium. IEEE MultiMedia, 5(4):84–88, October/December 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u4084.
pdf.

Hirose:1999:ILV

[HOY99] Michitaka Hirose, Tetsuro Ogi, and Toshio Yamada. Integrating live video for immersive environments. IEEE Multi-Media, 6(3):14–22, July/ September 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u3014.
pdf;   http://www.computer.
org/multimedia/mu1999/
u3014abs.htm.

Huang:1996:PRM

[HP96] Jiandong Huang and Allalaghatta Pavan. Project reports: The Mercuri Multimedia Laboratory at Honeywell. IEEE Multi-Media, 3(2):72–77, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2072.
pdf.

Hart:2005:RMR

[HPH05] Peter E. Hart, Kurt Pierson, and Jonathan J. Hull. Refocusing multimedia research on short clips. IEEE MultiMedia, 12(3):8–13, July/September 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hansen:2000:AMA

[HR00] Mark Hansen and Ben Rubin. Artful media: The audiences would be the artists and their life would be the arts. IEEE MultiMedia, 7(2):6–9, April/ June 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2006.
pdf.

Holbling:2008:MWI

[HRCK08] Günther Hölbling, Tilmann Rabl, David Coquil, and Harald Kosch. Multimedia at work: Interactive TV services on mobile devices. IEEE MultiMedia, 15(2): 72–76, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Huang:2021:EDC

[HRH^{+}21] Xiangdong Huang, Minjie Ren, Qiankun Han, Xiaoqi Shi, Jie Nie, Weizhi Nie, and An-An Liu. Emotion detection for conversations based on reinforcement learning framework. IEEE MultiMedia, 28(2): 76–85, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Herman:1996:SPEa

[HRvL96a] Ivan Herman, Graham J. Reynolds, and James van Loo. Standards: Premo: An emerging standard for multimedia presentation. Part I: Overview and framework. IEEE MultiMedia, 3(3):83–89, Fall 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u3083.
                                        pdf.

Herman:1996:SPEb

[HRvL96b] Ivan Herman, Graham J. Reynolds, and James van Loo. Standards: Premo: An emerging standard for multimedia presentation. Part II: Specification and applications. IEEE MultiMedia, 3(4):72–75, Winter 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u4072.
                                        pdf.

Holzbock:1999:AMS

[HS99] Matthias Holzbock and Christian Senninger. An aeronautical multimedia service demonstration at high frequencies. IEEE MultiMedia, 6(4):20–29, October/December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1999/pdf/
u4020.pdf;   http://www.
computer.org/multimedia/
mu1999/u4020abs.htm.

Hunter:2008:CAS

[HS08] Jane Hunter and Ronald Schroeter. Co-Annotea: a system for tagging relationships between multiple mixed-media objects. IEEE MultiMedia, 15(3): 42–53, July/September 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hinami:2017:ABM

[HS17] Ryota Hinami and Shin’ichi Satoh. Audience behavior mining: Integrating TV ratings with multimedia content. IEEE MultiMedia, 24(2):44–54, April/June 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
02/mmu2017020044-abs.
html.

Huang:2006:CSV

[HT06] Biao-Bing Huang and Shao-Xian Tang. A contrast-sensitive visible watermarking scheme. IEEE MultiMedia, 13(2):60–66, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hofmann:2008:VVD

[HT08] Günter Hofmann and Gerhard Thomas. Visions and views: Digital lifestyle 2020. IEEE MultiMedia, 15 (2):4–7, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hung:2022:UII

[HTH^{+}22] Yu-Hsiang Hung, Julianne Tan, Tai-Ming Huang, Shang-Che Hsu, Yi-Ling Chen, and Kai-Lung Hua. Unpaired image-to-image translation using negative learning for noisy patches. IEEE MultiMedia, 29(4): 59–68, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hemphill:1996:MWS

[HTL96] C. T. Hemphill, P. R. Thrift, and J. C. Linn. Multimedia at work: Speechaware multimedia. IEEE MultiMedia, 3(1):74–78, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u1074.
pdf.

Huang:1997:GEI

[Hua97] H. K. Huang. Guest Editors’ introduction: Multimedia applications in health care. IEEE MultiMedia, 4(2):23, April/ June 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2023.
pdf.

Huber:2004:MNG

[Hub04]  Josef F. Huber.  Mo-

bile next-generation networks. IEEE MultiMedia, 11(1):72–83, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
01/u1072abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2004/01/u1072.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/01/
u1072.pdf.

Hussain:2023:EIM

[HUU^{+}23] Tanveer Hussain, Samee Ullah Khan, Waseem Ullah, Ijaz Ul Haq, Min Je Kim, Mi Young Lee, and Sung Wook Baik. eCube-Land: an intelligent multiview video data modeling. IEEE MultiMedia, 30 (4):5–15, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Huang:1998:SIM

[HW98] Chung-Ming Huang and Chian Wang. Synchronization for interactive multimedia presentations. IEEE MultiMedia, 5(4):44–62, October/ December 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u4044.
pdf;   http://www.computer.
org/multimedia/mu1998/
u4044abs.htm.

Harrison:2010:AUN

[HWD10] Chris Harrison, Jason Wiese, and Anind K. Dey. Achieving ubiquity: The new third wave. IEEE MultiMedia, 17(3):8–12, July/ September 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hong:2012:MQA

[HWL^{+}12] Richang Hong, Meng Wang, Guangda Li, Liqiang Nie, Zheng-Jun Zha, and Tat-Seng Chua. Multimedia question answering.

IEEE MultiMedia, 19(4): 72–78, October/December 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hu:2014:TMS

[HWL^{+}14] Han Hu, Yonggang Wen, Huanbo Luan, Tat-Seng Chua, and Xuelong Li. Toward multiscreen social TV with geolocation-aware social sense. IEEE Multi-Media, 21(3):10–19, July/ September 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Helyer:2009:AMS

[HWV^{+}09] Nigel Helyer, Daniel Woo, Francesca Veronesi, Nigel Helyer, and Francesca Veronesi. Artful media: The sonic nomadic: Exploring mobile surroundsound interactions. IEEE MultiMedia, 16(2):12–15, April/June 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hu:2023:EAV

[HWZ^{+}23] Xinjue Hu, Chenchen Wang, Lin Zhang, Guo Chen, and Shervin Shirmohammadi. Edge-assisted virtual viewpoint generation for immersive light field. IEEE MultiMedia, 30(2):18–27, April/June 2023. CODEN IEMUE4.

ISSN 1070-986X (print), 1941-0166 (electronic).

Huang:2018:SRL

[HXXS18] X. Huang, C. Xu, L. Xiang, and J. Sang. Social relationship labeling based on multimodal behaviors and social interactions. IEEE MultiMedia, 25(4):68–78, October/ December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Hao:2019:EAV

[HYC^{+}19] Y. Hao, J. Yang, M. Chen, M. S. Hossain, and M. F. Alhamid. Emotion-aware video QoE assessment via transfer learning. IEEE MultiMedia, 26(1):31–40, January/March 2019. CO-DEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Huynh:2021:EEL

[HYLK21] Van Thong Huynh, Hyung-Jeong Yang, Guee-Sang Lee, and Soo-Hyung Kim. End-to-end learning for multimodal emotion recognition in video with adaptive loss. IEEE Multi-Media, 28(2):59–66, April/ June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Huang:2023:RTI

[HZK^{+}23] Zhe Huang, Fangfang Zhang, Lei Kou, Quande Yuan, and Yuanhong Liu. A real-time image encryption algorithm for a distributed energy system based on the 13-D complex chaotic sequence. IEEE MultiMedia, 30(4):71–81, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hsieh:2023:BAH

[HZSH23] Jui-Hung Hsieh, Zhi-Yu Zhang, Jing-Cheng Syu, and Mao-Cheng Hsieh. Bandwidth-aware highefficiency video coding design scheme on a multiprocessor system on chip. IEEE MultiMedia, 30(3): 37–47, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hu:2023:IHG

[HZY^{+}23] Fang Hu, Yin Zhang, Xingyu Yan, Mingfang Huang, and Xiong Zhang. An improved heterogeneous graph convolutional network for inter-relational medicine representation learning. IEEE Multi-Media, 30(1):52–61, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Issa:2008:VUA

[IGB^{+}08] Omneya Issa, Jean-Charles Grégoire, Yacine Belala, James Wong, and Marc Bage. 3G video uploading applications: Performance and improvements. IEEE MultiMedia, 15(4): 58–67, October/December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ionescu:2020:AIF

[IGR^{+}20] Bogdan Ionescu, Marian Ghenescu, Florin Răstoceanu, Răzvan Roman, and Marian Buric. Artificial intelligence fights crime and terrorism at a new level. IEEE MultiMedia, 27(2): 55–61, April/June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ikedo:2002:CRS

[Ike02] Tsuneo Ikedo. Creating realistic scenes in future multimedia systems. IEEE MultiMedia, 9(4):56–72, October/December 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
04/u4056abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/04/u4056.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/04/
u4056.pdf.

Ionescu:2016:RCI

[IMK^{+}16] Bogdan Ionescu, Henning Muller, Yiannis Kompatsiaris, Guillaume Gravier, and Anthony Vetro. Reports on CBMI 16 and ICME 16. IEEE Multi-Media, 23(4):88–93, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
04/mmu2016040088.html.

Iera:1999:QMA

[IMMP99] Antonio Iera, Antonella Molinaro, Salvatore Marano, and Miriam Petrone. QoS for multimedia applications in satellite systems. IEEE MultiMedia, 6(4):46–53, October/ December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u4046.
pdf;   http://www.computer.
org/multimedia/mu1999/
u4046abs.htm.

Ingalls:2013:AME

[Ing13] Todd Ingalls. Affect in media: Embodied media interaction in performance and public art. IEEE MultiMedia, 20(3):4–7, July/ September 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ilarri:2014:MDM

[ISDH14] Sergio Ilarri, Florence Sedes, Francesco De Natale, and Alan Hanjalic. Multimedia data management in mobile computing. IEEE MultiMedia, 21 (1):10–13, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Iqbal:2008:CDV

[ISSZ08] Razib Iqbal, Shervin Shirmohammadi, Abdulmotaleb El Saddik, and Jiying Zhao. Compresseddomain video processing for adaptation, encryption, and authentication. IEEE MultiMedia, 15(2):38–50, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2015:VMO

[iWD15] Cheng i Wang and Shlomo Dubnov. The variable Markov oracle: Algorithms for human gesture applications. IEEE MultiMedia, 22(4):52–67, October/ December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/04/mmu2015040052-
                                        abs.html.

Jain:1994:MC

[Jai94a] Ramesh Jain. Multimedia computing. IEEE MultiMedia, 1(1):3–6, Spring 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jain:1994:MP

[Jai94b] Ramesh Jain. Multimedia pollution. IEEE MultiMedia, 1(4):3–4, Winter 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jain:1994:SMS

[Jai94c]  Ramesh Jain.  Seman-

tics in multimedia systems. IEEE MultiMedia, 1(2):3– 4, Summer 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jain:1994:WMA

[Jai94d] Ramesh Jain. What is multimedia, anyway? IEEE MultiMedia, 1(3):3–??, Autumn/Fall 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jain:1995:ECMa

[Jai95a] Ramesh Jain. Editor-In-Chief’s message: Digital imaging. IEEE MultiMedia, 2(1):3, Spring 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1995/pdf/
u10003.pdf.

Jain:1995:ECMc

[Jai95b] Ramesh Jain. Editor-in-Chief’s message: Infoscopes. IEEE MultiMedia, 2(3):3, Fall 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u30003.
pdf.

Jain:1995:ECMd

[Jai95c]  Ramesh Jain.  Editor-

in-Chief’s message: Real telepresence. IEEE MultiMedia, 2(4):3–4, Winter 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1995/pdf/
u40003.pdf.

Jain:1995:ECMb

[Jai95d] Ramesh Jain. Editor-in-Chief’s message: Worldwide maze. IEEE MultiMedia, 2(2):3, Summer 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u20003.
pdf.

Jain:1996:ECMc

[Jai96a]  R. Jain.  Editor-in-

Chief’s message: Computing in natural environments. IEEE MultiMedia, 3(3):3–4, Fall 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u3003.
pdf.

Jain:1996:ECMa

[Jai96b] Ramesh Jain. Editor-in-Chief’s message: CAD for creative people. IEEE MultiMedia, 3(1):3–4, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1996/pdf/
u1003.pdf.

Jain:1996:ECMb

[Jai96c] Ramesh Jain. Editor-in-Chief’s message: Globalizing with the Web. IEEE MultiMedia, 3(2): 3, Summer 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2003.
pdf.

Jain:1996:ECMd

[Jai96d] Ramesh Jain. Editor-in-Chief’s message: Merging virtual and real.

IEEE MultiMedia, 3(4): 3, Winter 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u4003.
pdf.

Jain:1997:MVR

[Jai97a] Ramesh Jain. Media vision: a revolution in education. IEEE MultiMedia, 4(1):1–2, January/ March 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u1001.
pdf.

Jain:1997:MVC

[Jai97b] Ramesh Jain. Media vision: Couch potatoes and mouse clickers. IEEE MultiMedia, 4(3):1, July/ September 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3001.
pdf.

Jain:1997:MVE

[Jai97c] Ramesh Jain. Media vision: Erratum. IEEE MultiMedia, 4(1):1, January/March 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u1001.

pdf. Corrects a tableof-contents entry error for [MG96].

Jain:1997:MVI

[Jai97d] Ramesh Jain. Media vision: Immersive, interactive, navigational — needed for information access. IEEE MultiMedia, 4(4):1–2, October/ December 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u4001.
pdf.

Jain:1997:MVL

[Jai97e] Ramesh Jain. Media vision: Live art. IEEE MultiMedia, 4(2):1, April/ June 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2001.
pdf.

Jain:1998:MVC

[Jai98a] Ramesh Jain. Media vision: Communities 3000. IEEE MultiMedia, 5(2):1– 2, April/June 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/

                                        books/mu1998/pdf/u2001.
                                        pdf.

Jain:1998:MVM

[Jai98b] Ramesh Jain. Media vision: Multimedia documents. IEEE MultiMedia, 5(3):1, July/ September 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u3001.
                                        pdf.

Jain:1998:MVP

[Jai98c] Ramesh Jain. Media vision: Putting sensors on the Web. IEEE Multi-Media, 5(1):1, January/ March 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u1001.
                                        pdf.

Jain:1998:MVS

[Jai98d] Ramesh Jain. Media visions: Sensory overloading in documents. IEEE MultiMedia, 5(4):1–2, October/December 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u4001.
                                        pdf.

Jain:1999:MVE

[Jai99a] Ramesh Jain. Media vision: Event Web. IEEE MultiMedia, 6(2):1, April/ June 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u2001.
pdf.

Jain:1999:MVM

[Jai99b] Ramesh Jain. Media vision: Multimedia portals. IEEE MultiMedia, 6(1):1, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u1001.
pdf.

Jain:1999:MVC

[Jai99c] Ramesh Jain. Media vision: The convergence of PCs and TV. IEEE MultiMedia, 6(4):1, October/ December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u4001.
pdf.

Jain:2000:MVD

[Jai00a] Ramesh Jain. Media vision: Data, information, and experience. IEEE MultiMedia, 7(3):1, July/ September 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u3001.
pdf.

Jain:2000:MVI

[Jai00b]  Ramesh Jain.  Me-

dia vision: Informationitis. IEEE MultiMedia, 7(4):1, 5, October/ December 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u4001.
pdf.

Jain:2000:MVSa

[Jai00c] Ramesh Jain. Media vision: Searching for solutions. IEEE Multi-Media, 7(1):1, January/ March 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u1001.
pdf.

Jain:2000:MVSb

[Jai00d] Ramesh Jain. Media vision: Searching in context. IEEE MultiMedia, 7(2):1– 2, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/

books/mu2000/pdf/u2001.
pdf.

Jain:2001:MVK

[Jai01a] Ramesh Jain. Media vision: Knowledge and experience. IEEE Multi-Media, 8(4):4, October/ December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u4004.
pdf;   http://www.computer.
org/multimedia/mu2001/
u4004abs.htm.

Jain:2001:MVS

[Jai01b] Ramesh Jain. Media vision: Structuralizing multimedia data. IEEE MultiMedia, 8(2):1–??, April/ June 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u2001.
pdf.

Jain:2001:MVT

[Jai01c] Ramesh Jain. Media vision: Towards digital experience. IEEE Multi-Media, 8(1):1, January/ March 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u1001.
pdf.

Jain:2002:MVT

[Jai02a] Ramesh Jain. Media vision: Transformed experiences. IEEE Multi-Media, 9(1):1, January/ March 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2002/pdf/u1001.
pdf;   http://www.computer.
org/multimedia/mu2002/
u1001abs.htm.

Jain:2002:PTC

[Jai02b] Ramesh Jain. PC and TV convergence: Is it finally here? IEEE MultiMedia, 9(4):104, 103, October/December 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
04/u4104abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/04/u4104.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/04/
u4104.pdf.

Jain:2003:WDM

[Jai03a] Ramesh Jain. Are we doing multimedia? IEEE MultiMedia, 10(4):112, 111, October/December 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/

04/u4112abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/04/u4112.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/04/
u4112.pdf.

Jain:2003:BW

[Jai03b] Ramesh Jain. Beyond windows. IEEE MultiMedia, 10(2):88, 87, April/ June 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/02/u2088.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/02/
u2088.pdf.

Jain:2003:EIO

[Jai03c]  Ramesh Jain.  Expe-

rience isn’t only screen deep. IEEE MultiMedia, 10(1):80, 79, January/March 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/01/u1080.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/01/
u1080.pdf.

Jain:2003:MEC

[Jai03d]  Ramesh Jain.  Multi-

media electronic chronicles. IEEE MultiMedia, 10(3):112, 111, July/ September 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2003/03/u3112.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/03/
                                        u3112.pdf.

Jain:2005:WMI

[Jai05a] Ramesh Jain. I want my IPTV. IEEE MultiMedia, 12(3):96, 95, July/ September 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jain:2005:TMC

[Jai05b] Ramesh Jain. A true multimedia client. IEEE MultiMedia, 12(2):104, 103, April/June 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jaimes:2006:VVH

[Jai06]  Alejandro Jaimes.  Vi-

sions and views: Humancentered multimedia: Culture, deployment, and access. IEEE MultiMedia, 13 (1):12–19, January/March 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jain:2007:MVP

[Jai07] Ramesh Jain. Media vision: Photo retrieval: Multimedia’s chance to solve a real problem for real people. IEEE MultiMedia, 14(3):112, 111, July/ September 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jain:2015:LWV

[Jai15] Ramesh Jain. Let’s weave the Visual Web. IEEE MultiMedia, 22(3):66–72, July/September 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/03/mmu2015030066-
abs.html.

Jaimes:2016:CVS

[Jai16]  Alex Jaimes.  Com-

puter vision startups tackle AI. IEEE MultiMedia, 23(4):94–96, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
04/mmu2016040094-abs.
html.

Jain:2022:LHD

[Jai22] Ramesh Jain. Lifeblood of health is data. IEEE Multi-Media, 29(1):128–135, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jamieson:2007:AMU

[Jam07] Helen Varley Jamieson. Artful media: UpStage: a platform for creating and performing online. IEEE MultiMedia, 14(3):8–10, July/September 2007. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jayant:2003:VM

[Jay03] Nikil Jayant. Video and mobility. IEEE MultiMedia, 10(2):81–82, April/ June 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2003/02/u2081.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/02/
                                        u2081.pdf.

Jin:2023:DBC

[JCM^{+}23] Jianqiu Jin, Xiaoming Chen, Yu Meng, Xiangyang Gong, Bailin Yang, and Yuk Ying Chung. Deep blind chest X-ray image quality assessment with region-of-interest-guided attention. IEEE MultiMedia, 30(4): 36–46, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jeffay:1999:PRT

[Jef99] Kevin Jeffay. Project reports: Towards a betterthan-best-effort forwarding service for multimedia flows. IEEE Multi-Media, 6(4):84–87, October/December 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u4084.
pdf.

Junior:2013:SBS

[JGN^{+}13] Lucenildo Aquino Junior, Ruan Gomes, Manoel Silva Neto, Alexandre Duarte, Rostand Costa, and Guido Souza Filho. A software-based solution for distributing and displaying 3D UHD films. IEEE MultiMedia, 20(1):60–68, January/March 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jing:2013:MGC

[JIW13] Yushi Jing, Go Irie, and Marcel Worring. Multimedia Grand Challenge 2012. IEEE MultiMedia, 20(1): 2–6, January/March 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jain:2014:OS

[JJ14] Ramesh Jain and Laleh Jalali. Objective self. IEEE MultiMedia, 21(4):100– 110, October/December 2014. CODEN IEMUE4.

ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://www.computer.org/
csdl/mags/mu/2014/04/
mmu2014040100-abs.html.

Jaimes:2004:DBA

[JJO^{+}04] Alejandro Jaimes, Pamela Jennings, Hugo Ortega, Mark Tribe, and Christina Yang. Digital boundaries: The 2004 ACM Multimedia Interactive Art Program Exhibition. IEEE MultiMedia, 11(4):8–9, October/December 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/04/u4008.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/04/
u4008.pdf.

Jeon:2019:DLT

[JK19] J. Jeon and M. Kim. Discovering latent topics with saliency-weighted LDA for image scene understanding. IEEE MultiMedia, 26 (3):56–68, July/September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jain:2012:FMR

[JKP12] Anil K. Jain, Brendan Klare, and Unsang Park. Face matching and retrieval in forensics applications. IEEE MultiMedia, 19(1):20, 28, January/March 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jiang:2007:RPP

[JLD07] Hao Jiang, Ze-Nian Li, and Mark S. Drew. Recognizing posture in pictures with successive convexification and linear programming. IEEE MultiMedia, 14(2): 26–37, April/June 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Janzen:2017:EDR

[JM17] Ryan Janzen and Steve Mann. Extreme-dynamic-range sensing: Realtime adaptation to extreme signals. IEEE MultiMedia, 24(2):30–42, April/June 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
02/mmu2017020030-abs.
html.

Jin:2004:QSL

[JN04] Jingwen Jin and Klara Nahrstedt. QoS specification languages for distributed multimedia applications: a survey and taxonomy. IEEE Multi-Media, 11(3):74–87, July/ September 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2004/03/u3074.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/03/
                                        u3074.pdf.

Jadhav:2006:IIH

[JNK06] Chetan Jadhav, Pravin Nair, and Venkat Krovi. Individualized interactive home-based haptic telerehabilitation. IEEE Multi-Media, 13(3):32–39, July/ September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jain:2013:MRC

[JP13] Ramesh Jain and Balakrishnan Prabha Prabhakaran. Multimedia Retrieval Conference enjoys Texas hospitality. IEEE MultiMedia, 20(3): 8–9, July/September 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jain:2013:MSM

[JS13] Ramesh Jain and Malcolm Slaney. Micro stories and mega stories. IEEE MultiMedia, 20(1):86–90, January/March 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jin:2009:DIS

[JTCC09] Xing Jin, Ho-Shing Tang, S.-H. Gary Chan, and Kan-Leung Cheng. Deployment issues in scalable island multicast for peer-to-peer streaming. IEEE MultiMedia, 16(1):72–80, January/March 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Juhlin:2011:SMR

[Juh11] Oskar Juhlin. Social media on the road: Mobile technologies and future traffic research. IEEE MultiMedia, 18(1):8–10, January/March 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Jin:2017:WCM

[JW17] Yichao Jin and Yonggang Wen. When cloud media meet network function virtualization: Challenges and applications. IEEE MultiMedia, 24(3):72–82, July/September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
03/mmu2017030072-abs.
html.

Jiao:2014:LSM

[JWW^{+}14] Jianbo Jiao, Ronggang Wang, Wenmin Wang, Shengfu Dong, Zhenyu Wang, and Wen Gao. Local stereo matching with improved matching cost and disparity refinement. IEEE MultiMedia, 21(4):16–27, October/ December 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu2014040016-
abs.html.

Jin:2017:CCE

[JXS^{+}17] Qun Jin, Yong Xiang, Guozi Sun, Yao Liu, and Chin-Chen Chang. Cybersecurity for cyberenabled multimedia applications. IEEE Multi-Media, 24(4):10–13, October/December 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
04/mmu2017040010.html.

Jin:2008:LRA

[JYC08] Xing Jin, W.-P. Ken Yiu, and S.-H. Gary Chan. Loss recovery in applicationlayer multicast. IEEE MultiMedia, 15(1):18–27, January/March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ji:2014:LSG

[JYS^{+}14] Rongrong Ji, Yi Yang, Nicu Sebe, Kiyoharu Aizawa, and Liangliang Cao. Largescale geosocial multimedia [guest editorial]. IEEE MultiMedia, 21(3):7–9, July/September 2014. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ji:2011:VHO

[JYXT11] Rongrong Ji, Hongxun Yao, Xing Xie, and Qi Tian. Vocabulary hierarchy optimization and transfer for scalable image search. IEEE Multi-Media, 18(3):66–77, July/ September 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kara:2023:CHM

[KAB^{+}23] Burak Kara, Mehmet N.

Akcay, Ali C. Begen, Saba Ahsan, Igor D. D. Curcio, and Emre B. Aksu. Could head motions affect quality when viewing 360^{∘} videos? IEEE MultiMedia, 30(2): 28–37, April/June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kahaner:1996:SRT

[Kah96] David K. Kahaner. Special report: Traffic safety activities in Japan. IEEE MultiMedia, 3(3):16–17, Fall 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u3016.
                                        pdf.

Kazman:1996:FPI

[KAHHM96] Rick Kazman, Reem Al-Halimi, William Hunt, and Marilyn Mantei. Four paradigms for indexing video conferences. IEEE MultiMedia, 3(1):63–73, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u1063.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1996/
                                        u1063abs.htm.

Kalva:2006:SHV

[Kal06] Hari Kalva. Standards: The H.264 Video Coding Standard. IEEE MultiMedia, 13(4):86–90, October/ December 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kamara:1999:MWJ

[Kam99] Sheila Kamara. Multimedia at work: JavuNetwork: remote video production and storage. IEEE MultiMedia, 6(3):78–80, July/ September 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u3078.
pdf.

Kankanhalli:2020:MDP

[Kan20] Mohan Kankanhalli. Multimedia data privacy against machines. IEEE MultiMedia, 27(2):5–7, April/June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kasik:2004:SCI

[Kas04] David J. Kasik. Strategies for consistent image partitioning. IEEE MultiMedia, 11(1):32–41, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
01/u1032abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2004/01/u1032.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/01/
u1032.pdf.

Krosche:2004:MMM

[KBB04] Jens Krösche, Jörg Baldzer, and Susanne Boll. MobiDENK— mobile multimedia in monument conservation. IEEE MultiMedia, 11(2):72–77, April/June 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/02/u2072.

                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/02/
                                        u2072.pdf.

Kosch:2005:LCM

[KBD^{+}05] Harald Kosch, Laszlo Boszormenyi, Mario Doller, Mulugeta Libsie, Peter Schojer, and Andrea Kofler. The life cycle of multimedia metadata. IEEE MultiMedia, 12(1):80–86, January/March 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kankanhalli:2000:PRV

[KC00] Mohan S. Kankanhalli and Tat-Seng Chua. Project reports: Video modeling using strata-based annotation. IEEE MultiMedia, 7(1):68–74, January/ March 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u1068.
                                        pdf.

Kanaya:2000:TDM

[KCKC00] Ichiroh Kanaya, Qian Chen, Yuko Kanemoto, and Kunihiro Chihara. Three-dimensional modeling for virtual relic restoration. IEEE MultiMedia, 7(2):42–44, April/ June 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2042.
pdf;   http://www.computer.
org/multimedia/mu2000/
u2042abs.htm.

Kim:2010:TGD

[KCRO10] Yeongmi Kim, Jongeun Cha, Jeha Ryu, and Ian Oakley. A tactile glove design and authoring system for immersive multimedia. IEEE MultiMedia, 17 (3):34–44, July/September 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kong:2017:ODB

[KD17] Lingchao Kong and Rui Dai. Object-detection-based video compression for wireless surveillance systems. IEEE MultiMedia, 24(2):76–85, April/June 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
02/mmu2017020076-abs.
html.

Karmouch:1996:PSM

[KE96] Ahmed Karmouch and James Emery. A playback schedule model for multimedia documents. IEEE MultiMedia, 3(1):50–61, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u1050.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1996/
                                        u1050abs.htm.

Kelliher:2014:CM

[Kel14] Aisling Kelliher. Critical multimedia. IEEE MultiMedia, 21(4):4–7, October/December 2014. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2014/04/mmu2014040004-
                                        abs.html.

Kelliher:2015:MLC

[Kel15] Aisling Kelliher. Machines learning culture. IEEE MultiMedia, 22(2):18–22, April/June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/02/mmu2015020018-
                                        abs.html.

Kelliher:2016:TAE

[Kel16] Aisling Kelliher. Technology and the arts: Educational encounters of the third kind. IEEE Multi-Media, 23(3):8–11, July/ September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/

03/mmu2016030008-abs.
html.

Kelliher:2017:ACC

[Kel17] Aisling Kelliher. ACM Creativity and Cognition Conference 2017. IEEE MultiMedia, 24(4):6–9, October/December 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
04/mmu2017040006.html.

Kemp:1995:MSI

[Kem95]  Zarine Kemp.  Multi-

media and spatial information systems. IEEE MultiMedia, 2(4):68–76, Winter 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u40068.
pdf;   http://www.computer.
org/multimedia/mu1995/
u4068abs.htm.

Keramane:2000:VVW

[Ker00] Cherif Keramane. Visions and views: The Wireless World Web. IEEE MultiMedia, 7(2):10–14, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2010.
pdf.

Kervella:1997:MMM

[KG97] Brigitte Kervella and Valérie Gay. MHEGAM — A multimedia messaging system. IEEE MultiMedia, 4(4):22–29, October/ December 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1997/pdf/u4022.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1997/
                                        u4022abs.htm.

Kuzmanovic:2005:HSS

[KG05] Maja Kuzmanovic and Nik Gaffney. Human-scale systems in responsive environments. IEEE MultiMedia, 12(1):8–13, January/ March 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Karduck:1996:MWM

[KGG96] Achim P. Karduck, André Geiser, and Thomas Gutekunst. Multimedia at work: Multimedia technology in banking. IEEE Multi-Media, 3(4):82–86, Winter 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://dlib.computer.
                                        org/mu/books/mu1996/pdf/

u4082.pdf. Discusses the Portfolio Analysis System (PAS) using Java.

Kim:2002:WSF

[KGS^{+}02] Wooyoung Kim, Sven Graupner, Akhil Sahai, Dmitry Lenkov, Chetan Chudasama, Samuel Whedbee, Yuhua Luo, Bharati Desai, Howard Mullings, and Pui Wong. Web Espeak: Facilitating Webbased E-services. IEEE MultiMedia, 9(1):43–55, January/March 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2002/pdf/u1043.
pdf;   http://www.computer.
org/multimedia/mu2002/
u1043abs.htm.

Kucuktunc:2007:MWN

[KGU07] Onur Küçüktunç, Uğur Güdükbay, and Özgür Ulusoy. Multimedia at work: a natural language-based interface for querying a video database. IEEE Multi-Media, 14(1):83–89, January/March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kamboj:2021:MPD

[KHA^{+}21] Manvi Kamboj, Christian Hessler, Priyanka Asnani, Kais Riani, and Mohamed Abouelenien. Multimodal political deception detection. IEEE Multi-Media, 28(1):94–102, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Khut:2007:AMC

[Khu07] George Poonkhin Khut. Artful media: Cardiomorphologies: An inner journey through art. IEEE MultiMedia, 14(4):5–7, October/December 2007. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kirovski:2004:DWF

[KHY04] Darko Kirovski, Henrique, and Yacov Yacobi. A dual watermark-fingerprint system. IEEE MultiMedia, 11(3):59–73, July/ September 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2004/03/u3059.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/03/
                                        u3059.pdf.

Kim:1995:MWM

[Kim95] Michelle Y. Kim. Multimedia at work: a multimedia information system for home health-care support. IEEE MultiMedia, 2(4):83– 87, Winter 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/

books/mu1995/pdf/u40083.
pdf.

Kimura:1996:PBP

[Kim96] Takayuki Dan Kimura. A pen-based prosodic user interface for schoolchildren. IEEE MultiMedia, 3(4):48– 55, Winter 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u4048.
pdf;   http://www.computer.
org/multimedia/mu1996/
u4048abs.htm.

Kirda:2001:EEF

[KJKS01] Engin Kirda, Mehdi Jazayeri, Clemens Kerer, and Markus Schranz. Experiences in engineering flexible Web services. IEEE MultiMedia, 8(1):58–65, January/March 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u1058.
pdf;   http://www.computer.
org/multimedia/mu2001/
u1058abs.htm.

Kalva:2007:SVV

[KL07] Hari Kalva and Jae-Beom Lee. Standards: The VC-1 Video Coding Standard. IEEE MultiMedia, 14(4): 88–91, October/December 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Klatt:1996:EME

[Kla96] Garth Klatt. To the Editor: on Microsoft Explorer vs. Netscape Navigator. IEEE MultiMedia, 3(4):5, Winter 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u4004.
                                        pdf.

Kang:2009:SIM

[KLKW09] Hogab Kang, Keunyoung Lee, Taehyun Kim, and Xin Wang. Standards in industry: MPEG-21 and its interoperability with rights-information standards. IEEE MultiMedia, 16(1):94–99, January/ March 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kelly:1995:PRV

[KM95] Patrick H. Kelly and Saied Moezzi. Project reports: Visual computing laboratory. IEEE MultiMedia, 2(1):94–99, Spring 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u10094.
                                        pdf.

Kitson:1997:OVC

[KMB97] Frederick Lee Kitson, Tom Malzbender, and Vasudev Bhaskaran. Opportunities for visual computing in healthcare. IEEE MultiMedia, 4(2):46–57, April/June 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2046.
pdf;   http://www.computer.
org/multimedia/mu1997/
u2046abs.htm.

Karatzas:2001:MAE

[KMF^{+}01] Kostas Karatzas, Nicolas Moussiopoulos, Kurt Fedra, Achim Lohmeyer, and Christos Kouroumlis. A multimedia application for EIA studies. IEEE MultiMedia, 8(4):71–75, October/December 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u4071.
pdf;   http://www.computer.
org/multimedia/mu2001/
u4071abs.htm.

Kaiser:2017:RWI

[KMJ17] Rene Kaiser, Britta Meixner, and Joscha Jager. Reflecting on the Workshop on Interactive Content Consumption (WS-ICC) series. IEEE Multi-Media, 24(3):83–88, July/ September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        03/mmu2017030083.html.

Kernchen:2010:IMP

[KMM^{+}10] Ralf Kernchen, Stefan Meissner, Klaus Moessner, Pablo Cesar, Ishan Vaishnavi, Matthieu Boussard, and Cristian Hesselman. Intelligent multimedia presentation in ubiquitous multidevice scenarios. IEEE MultiMedia, 17(2): 52–63, April/June 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kumar:2016:ECI

[KMS16] Chandan Kumar, Raphael Menges, and Steffen Staab. Eye-controlled interfaces for multimedia interaction. IEEE MultiMedia, 23(4):6–13, October/ December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        04/mmu2016040006-abs.
                                        html.

Klapsing:2001:SWE

[KNC01] Reinhold Klapsing, Gustaf Neumann, and Wolfram Conen. Semantics in Web engineering: Applying the resource description framework. IEEE MultiMedia, 8(2):62–??, April/June 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u2062.
pdf;   http://www.computer.
org/multimedia/mu2001/
u2062abs.htm.

Kansal:2007:MIS

[KNLZ07] Aman Kansal, Suman Nath, Jie Liu, and Feng Zhao. Media impact: SenseWeb: An infrastructure for shared sensing. IEEE MultiMedia, 14(4): 8–13, October/December 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kleinholz:1994:SCM

[KO94] Lutz Kleinholz and Martin Ohly. Supporting cooperative medicine: the Bermed Project. IEEE MultiMedia, 1(4):44–53, Winter 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Konstantas:1999:DMR

[KOCG99] Dimitri Konstantas, Yann Orlarey, Olivier Carbonel, and Simon Gibbs. The distributed musical rehearsal environment. IEEE MultiMedia, 6(3):54–64, July/ September 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u3054.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1999/
                                        u3054abs.htm.

Kogler:1998:MWV

[Kog98] Rita Ann Kogler. Multimedia at work: Virtual set design. IEEE MultiMedia, 5(1):92–96, January/March 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u1092.
                                        pdf.

Kolli:1996:MRG

[Kol96]  R. Kolli.  Media re-

views: Good, if limited, advice for would-be multimedia authors. IEEE MultiMedia, 3(1):91–93, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u1091.
                                        pdf.

Kamal:2024:GAD

[KP24] Minhas Kamal and Balakrishnan Prabhakaran. Generative AI for 3-D point clouds. IEEE MultiMedia, 31(2):5–6, April/June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kraak:1996:IMG

[Kra96] Menno-Jan Kraak. Integrating multimedia in geographical information systems. IEEE Multi-Media, 3(2):59–65, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2059.
pdf;   http://www.computer.
org/multimedia/mu1996/
u2059abs.htm.

Kraemer:2000:MRI

[Kra00] Eileen Kraemer. Media reviews: Interaction in smart environments: Managing Interactions in Smart Environments, P. Nixon, G. Lacey, and S. Dobson, eds. (Springer-Verlag, Berlin, 2000, $84.95, 250 pp., ISBN 1-85233-228-X). IEEE MultiMedia, 7(3):91–92, July/ September 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u3091.
pdf.

Koelle:2019:UIC

[KRB19] M. Koelle, E. Rose, and S. Boll. Ubiquitous intelligent cameras between legal nightmare and social empowerment. IEEE MultiMedia, 26(2):76–86, April/June 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Krishnan:1999:MWT

[Kri99] Rajesh Krishnan. Multimedia at work: Timeshared video-on-demand: a workable solution. IEEE MultiMedia, 6(1):77–79, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u1077.
                                        pdf.

Kriechbaum:2003:SBD

[Kri03a] Werner Kriechbaum. Scores and beyond: The dark side of audio multimedia. IEEE MultiMedia, 10(3):15–17, July/September 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2003/03/u3015.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/03/
                                        u3015.pdf.

Krikke:2003:SKB

[Kri03b] Jan Krikke. South Korea beats the world in broadband gaming. IEEE MultiMedia, 10(2):12–14, April/June 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/02/u2012.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/02/
u2012.pdf.

Kanade:1997:VRC

[KRN97] Takeo Kanade, Peter Rander, and P. J. Narayanan. Virtualized reality: Constructing virtual worlds from real scenes. IEEE MultiMedia, 4(1):34–47, January/March 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u1034.
pdf;   http://www.computer.
org/multimedia/mu1997/
u1034abs.htm.

Kustanowitz:2006:HLP

[KS06] Jack Kustanowitz and Ben Shneiderman. Hierarchical layouts for photo libraries. IEEE MultiMedia, 13(4): 62–72, October/December 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kelliher:2012:TMS

[KS12] Aisling Kelliher and Malcolm Slaney. Tell me a story. IEEE MultiMedia, 19(1):4, 6, January/March 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Knees:2007:EMC

[KSPW07] Peter Knees, Markus Schedl, Tim Pohle, and Gerhard Widmer. Exploring music collections in virtual landscapes. IEEE MultiMedia, 14(3):46–54, July/September 2007. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kalker:2012:URM

[KSW12] Ton Kalker, Rajan Samtani, and Xin Wang. UltraViolet: Redefining the movie industry? IEEE MultiMedia, 19(1):7, 11, January/March 2012. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Keshary:2022:MDB

[KTB^{+}22] Shivom Keshary, Padmapritha. T, Korkut Bekiroglu, Subathra Seshadhri, and Seshadhri Srinivasan. Multimedia data-based artificial pancreas for type 2 diabetes. IEEE MultiMedia, 29(1):18–27, January/ March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kahol:2006:DMS

[KTP06] Kanav Kahol, Priyamvada Tripathi, and Sethuraman Panchanathan. Documenting motion sequences with a personalized annotation system. IEEE MultiMedia, 13(1):37–45, January/ March 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kiyokawa:2000:SOC

[KTY00] Kiyoshi Kiyokawa, Haruo Takemura, and Naokazu Yokoya. SeamlessDesign for 3D object creation. IEEE MultiMedia, 7(1):22–33, January/ March 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u1022.
pdf;   http://www.computer.
org/multimedia/mu2000/
u1022abs.htm.

Kapach:2019:CRO

[KUM^{+}19] Z. Kapach, A. Ulmer, D. Merrick, A. Alikhan, Y. Lu, A. Mohan, A. S. Kaseb, and G. K. Thiruvathukal. Cloud resource optimization for processing multiple streams of visual data. IEEE MultiMedia, 26 (3):31–41, July/September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kundur:2001:WDI

[Kun01]  Deepa Kundur.  Wa-

termarking with diversity: Insights and implications. IEEE MultiMedia, 8(4):46–52, October/ December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u4046.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2001/
                                        u4046abs.htm.

Koehler:2010:RHR

[KWG10] Christopher Koehler, Thomas Wischgoll, and Forouzan Golshani. Reconstructing the human ribcage in 3D with X-rays and geometric models. IEEE Multi-Media, 17(3):46–53, July/ September 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Kong:2022:BBI

[KWYK22] Kai Kong, Xiangjun Wu, Datao You, and Haibin Kan. 3D-BCNN-Based image encryption with finite computing precision. IEEE MultiMedia, 29(4):97–110, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Khan:2006:MRP

[KZ06] Muhammad Khurram Khan and Jiashu Zhang. Media reviews: Pointing a finger toward standardized and automated biometrics.

IEEE MultiMedia, 13(4): 73–75, October/December 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Larson:2003:VWS

[Lar03] James A. Larson. VoiceXML and the W3C speech interface framework. IEEE MultiMedia, 10(4):91–93, October/December 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
04/u4091abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/04/u4091.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/04/
u4091.pdf.

Liu:2022:BND

[LBZL22] Jingren Liu, Haoyue Bai, Haofeng Zhang, and Li Liu. Beyond normal distribution: More factual feature generation network for generalized zero-shot learning. IEEE MultiMedia, 29 (3):69–79, July/September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lu:2021:LAI

[LCCL21] Zhicheng Lu, Xiaoming Chen, Vera Yuk Ying Chung, and Sen Liu. LFI-Augmenter: Intelligent light field image editing with interleaved spatialangular convolution. IEEE MultiMedia, 28(4):84–95, October/December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Little:1994:NAM

[LCG^{+}94] Thomas D. C. Little, Michael G. Christel, Simon J. Gibbs, Christian Breiteneder, and Vicki de Mey. Novel applications of multimedia. IEEE MultiMedia, 1(1):82–83, Spring 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lim:2021:CGM

[LCL21] Yu-Quan Lim, Chee Seng Chan, and Fung Ying Loo. ClaviNet: Generate music with different musical styles. IEEE Multi-Media, 28(1):83–93, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Le:2023:SDR

[LCL23] The Van Le, Yongho Choi, and Jin Young Lee. Specular detection and rendering for immersive multimedia. IEEE MultiMedia, 30(2): 38–47, April/June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lin:2009:WBM

[LCM09] Hao-Tung Lin, Herng-Yow Chen, and Ting-Yi Ma. A Web-based music lecture database framework. IEEE MultiMedia, 16(4): 62–70, October/December 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ling:2012:EIC

[LCM^{+}12] Hefei Ling, Hongrui Cheng, Qingzhen Ma, Fuhao Zou, and WeiQi Yan. Efficient image copy detection using multiscale fingerprints. IEEE Multi-Media, 19(1):60–69, January/March 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lin:2011:WSF

[LCSC11] Lin Lin, Chao Chen, Mei-Ling Shyu, and Shu-Ching Chen. Weighted subspace filtering and ranking algorithms for video concept retrieval. IEEE Multi-Media, 18(3):32–43, July/ September 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Levy:2006:MWR

[LD06] Richard Levy and Peter Dawson. Multimedia at work: Reconstructing a Thule whalebone house using 3D imaging. IEEE MultiMedia, 13(2):78–83, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liu:2018:TEH

[LDZS18] Y. Liu, H. Dong, L. Zhang, and A. E. Saddik. Technical evaluation of HoloLens for multimedia: a first look. IEEE MultiMedia, 25(4):8–18, October/ December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Lee:1998:PVS

[Lee98] Jack Y. B. Lee. Parallel video servers: a tutorial. IEEE MultiMedia, 5(2):20– 28, April/June 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u2020.
pdf;   http://www.computer.
org/multimedia/mu1998/
u2020abs.htm.

Lin:2008:MWS

[LEGFD08] Ching-Yung Lin, Kate Ehrlich, Vicky Griffiths-Fisher, and Christopher Desforges. Multimedia at work: SmallBlue: People mining for expertise search. IEEE MultiMedia, 15(1):78–84, January/ March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2018:CIC

[LFX^{+}18] M. Li, H. Fan, Y. Xiang, Y. Li, and Y. Zhang. Cryptanalysis and improvement of a chaotic image encryption by firstorder time-delay system. IEEE MultiMedia, 25(3): 92–101, July/September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Lokki:2005:NAC

[LG05] Tapio Lokki and Matti Grohn. Navigation with auditory cues in a virtual environment. IEEE MultiMedia, 12(2):80–86, April/June 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/cg/2005/
04/extras/u2080x2.mpeg;
http://csdl.computer.
org/comp/mags/mu/2005/
02/extras/u2080x1.mpeg;
http://csdl.computer.
org/comp/mags/mu/2005/
02/extras/u2080x3.mpeg;
http://csdl.computer.
org/comp/mags/mu/2005/
02/extras/u2080x4.mpeg.

Liem:2017:MTE

[LGT17] Cynthia C. S. Liem, Emilia Gomez, and George Tzanetakis. Multimedia technologies for enriched music performance, production, and consumption. IEEE MultiMedia, 24(1):20–23, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
01/mmu2017010020-abs.
html.

Luke:1997:MWW

[LH97] Sean Luke and James Hendler. Multimedia at work: Web agents that work. IEEE MultiMedia, 4(3):76–80, July/ September 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3076.
pdf.

Lu:2018:DMI

[LH18] L. Lu and A. P. Harrison. Deep medical image computing in preventive and precision medicine. IEEE MultiMedia, 25(3): 109–113, July/September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Li:2023:BMS

[LH23] Fanjie Li and Xiao Hu. Background music for studying: a naturalistic experiment on music characteristics and user perception. IEEE MultiMedia, 30 (1):62–72, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lui:2006:MSM

[LHA06] Simon Lui, Andrew Horner, and Lydia Ayers. MIDI to SP-MIDI transcoding using phrase stealing. IEEE MultiMedia, 13(2):52–59, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lin:2020:KPS

[LHD^{+}20] Weiyao Lin, Xiaoyi He, Wenrui Dai, John See, Tushar Shinde, Hongkai Xiong, and Lingyu Duan. Key-point sequence lossless compression for intelligent video analysis. IEEE MultiMedia, 27(3):12–22, July/ September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2024:UGD

[LHH^{+}24] Hengfan Li, Xinwei Hong, Guohua Huang, Xuanbo Xu, and Qingfeng Xia. Uncertainty-guided different levels of pseudolabels for semisupervised medical image segmentation. IEEE MultiMedia, 31(1): 42–53, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lu:2011:VCI

[LHX11] Jian Lu, Xian-Sheng Hua, and Dong Xu. Visual content identification and search. IEEE Multi-Media, 18(3):8–10, July/ September 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ling:2017:ARR

[Lin17] Haibin Ling. Augmented reality in reality. IEEE MultiMedia, 24(3):10–15, July/September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        03/mmu2017030010.html.

Little:1994:CI

[Lit94] Thomas D. C. Little. Commerce on the Internet. IEEE MultiMedia, 1(4):74– 78, Winter 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liu:1994:ISI

[Liu94]  Peiya Liu.  ISO stan-

dards from an information infrastructure perspective. IEEE MultiMedia, 1(1):89– 90, Spring 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liu:2016:BSS

[Liu16] Qiong Liu. The BAMMF series in Silicon Valley. IEEE MultiMedia, 23(2): 95–96, April/June 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
02/mmu2016020095.html.

Licks:2005:GAI

[LJ05] Vinicius Licks and Ramiro Jordan. Geometric attacks on image watermarking systems. IEEE Multi-Media, 12(3):68–78, July/ September 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lee:2012:IRF

[LJJT12] Jung-Eun Lee, Rong Jin, Anil K. Jain, and Wei Tong. Image retrieval in forensics: Tattoo image database application. IEEE MultiMedia, 19 (1):40–49, January/March 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2017:CMD

[LJSP17] Yan Li, Young-Sik Jeong, Byeong-Seok Shin, and Jong Hyuk Park. Crowdsensing multimedia data:

Security and privacy issues. IEEE MultiMedia, 24(4):58–66, October/December 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
04/mmu2017040058.html.

Liu:2014:LSP

[LJSZ14] Risheng Liu, Wei Jin, Zhixun Su, and Changcheng Zhang. Latent subspace projection pursuit with online optimization for robust visual tracking. IEEE MultiMedia, 21(4):47–55, October/December 2014. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu2014040047-
abs.html.

Li:2013:PDI

[LJZ^{+}13] Liang Li, Shuqiang Jiang, Zheng-Jun Zha, Zhipeng Wu, and Qingming Huang. Partial-duplicate image retrieval via saliency-guided visual matching. IEEE MultiMedia, 20(3):13–23, July/September 2013. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lampi:2008:VCT

[LKBE08] Fleming Lampi, Stephan Kopf, Manuel Benz, and Wolfgang Effelsberg. A virtual camera team for lecture recording. IEEE MultiMedia, 15(3):58–61, July/ September 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lee:1994:MIV

[LKGR94] Woobin Lee, Yongmin Kim, Robert J. Gove, and Christopher R. Read. MediaStation 5000: Integrating video and audio. IEEE MultiMedia, 1(2):50– 61, Summer 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lin:2013:KRS

[LKLH13] Ying-Dar Lin, Chia-Yu Ku, Yuan-Cheng Lai, and Chia-Fon Hung. In-kernel relay for scalable one-to-many streaming. IEEE Multi-Media, 20(1):69–79, January/March 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liao:1997:SMP

[LL97] Wanjiun Liao and Victor O. K. Li. The split and merge protocol for interactive video-on-demand. IEEE Multi-Media, 4(4):51–62, October/December 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1997/pdf/u4051.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1997/
                                        u4051abs.htm.

Li:2008:CTA

[LL08] Qingfeng Li and Stephen C-Y. Lu. Collaborative tagging applications and approaches. IEEE Multi-Media, 15(3):14–21, July/ September 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lee:2014:SRI

[LL14] Jung-San Lee and Bo Li. Self-recognized image protection technique that resists large-scale cropping. IEEE MultiMedia, 21 (1):60–73, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liu:2016:PCC

[LLD^{+}16] Song Liu, Wanqing Li, Stephen Davis, Christian Ritz, and Hongda Tian. Planogram compliance checking based on detection of recurring patterns. IEEE MultiMedia, 23(2):54–63, April/June 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.

org/csdl/mags/mu/2016/
02/mmu2016020054-abs.
html.

Lu:2023:ALD

[LLEK23] Qianwen Lu, Jinho Lee, Yuki Endo, and Shunsuke Kamijo. Artistic line drawing rendering with priors of depth and edge density. IEEE MultiMedia, 30 (4):61–70, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liu:2022:CTI

[LLH22] Sheng Liu, Chengqing Li, and Qiao Hu. Cryptanalyzing two image encryption algorithms based on a first-order time-delay system. IEEE MultiMedia, 29 (1):74–84, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lin:2015:BMU

[LLJW15] Yin-Tzu Lin, Chuan-Lung Lee, Jyh-Shing Roger Jang, and Ja-Ling Wu. Bridging music using sound-effect insertion. IEEE Multi-Media, 22(4):30–38, October/December 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/04/mmu2015040030-
abs.html.

Li:2017:CIS

[LLL17a] Chengqing Li, Dongdong Lin, and Jinhu Lu. Cryptanalyzing an image-scrambling encryption algorithm of pixel bits. IEEE Multi-Media, 24(3):64–71, July/ September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
03/mmu2017030064-abs.
html.

Liu:2017:PBQ

[LLL^{+}17b] Maofu Liu, Ya Liu, Zhenguang Liu, Huijun Hu, and Wei Fang. Pooling-based quantitative approach to evaluating binarization algorithms. IEEE Multi-Media, 24(1):86–92, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
01/mmu2017010086-abs.
html.

Li:2018:CIE

[LLLH18] C. Li, D. Lin, J. Lü, and F. Hao. Cryptanalyzing an image encryption algorithm based on autoblocking and electrocardiography. IEEE MultiMedia, 25(4):46–56, October/ December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Li:2000:CRI

[LLP00] Xiangyang Li, Dongming Lu, and Yunhe Pan. Color restoration and image retrieval for Dunhuang fresco preservation. IEEE Multi-Media, 7(2):38–42, April/ June 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2038.
pdf;   http://www.computer.
org/multimedia/mu2000/
u2038abs.htm.

Li:2016:SMU

[LLP^{+}16] Qing Li, Rynson W. H.

Lau, Elvira Popescu, Yanghui Rao, Howard Leung, and Xinzhong Zhu. Social media for ubiquitous learning and adaptive tutoring [Guest editors’ introduction]. IEEE Multi-Media, 23(1):18–24, January/March 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
01/mmu2016010018-abs.
html.

Lu:2011:VST

[LLS11] Yan Lu, Shipeng Li, and Huifeng Shen. Virtualized screen: a third element for cloud-mobile convergence. IEEE MultiMedia, 18(2):4–11, April/June 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lin:2010:PDA

[LLT10] Qian Lin, Jerry Liu, and Daniel Tretter. Printing in a digital age. IEEE Multi-Media, 17(4):100–107, October/December 2010. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liu:2011:RTV

[LLY^{+}11] Bo Liu, Zhu Li, Linjun Yang, Meng Wang, and Xinmei Tian. Real-time video copy-location detection in large-scale repositories. IEEE MultiMedia, 18 (3):22–31, July/September 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liu:2014:FAC

[LLY14] Si Liu, Luoqi Liu, and Shuicheng Yan. Fashion analysis: Current techniques and future directions. IEEE MultiMedia, 21(2):72–79, April/June 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2018:WMH

[LLY^{+}18] L. Li, H. Li, W. Yuan, J. Lu, X. Feng, and C. Chang. A watermarking mechanism with high capacity for threedimensional mesh objects using integer planning. IEEE MultiMedia, 25(3): 49–64, July/September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Liu:2003:AVM

[LLZ03] Jiangchuan Liu, Bo Li, and Ya-Qin Zhang. Adaptive video multicast over the Internet. IEEE MultiMedia, 10(1):22–33, January/March 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
01/u1022abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/01/u1022.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/01/
u1022.pdf.

Liu:2022:TSA

[LLZ22] Hantang Liu, Wentong Li, and Jianke Zhu. Translational symmetry-aware facade parsing for 3-D building reconstruction. IEEE MultiMedia, 29(4):38–47, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Leonardi:2002:SIM

[LM02] Riccardo Leonardi and Pierangelo Migliorati. Semantic indexing of multimedia documents. IEEE MultiMedia, 9(2):44–51, April/June 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2002/
                                        02/u2044abs.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mu/2002/02/u2044.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2002/02/
                                        u2044.pdf.

Lederman:2018:RUM

[LMCP18] O. Lederman, A. Mohan, D. Calacci, and A. S. Pentland. Rhythm: A unified measurement platform for human organizations. IEEE MultiMedia, 25(1):26–38, January/ March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Li:2010:QAC

[LMH^{+}10] Guangda Li, Zhao-Yan Ming, Richang Hong, Tat-Seng Chua, Haojie Li, and Sheng TangChua. Question answering over communitycontributed Web videos. IEEE MultiMedia, 17(4): 46–57, October/December 2010. CODEN IEMUE4.

ISSN 1070-986X (print), 1941-0166 (electronic).

Laborie:2009:MQD

[LMS09] Sebastien Laborie, Ana-Maria Manzat, and Florence Sedes. Managing and querying distributed multimedia metadata. IEEE MultiMedia, 16(4):12–21, October/December 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liu:2022:PPV

[LMS^{+}22] Jixin Liu, Ru Meng, Ning Sun, Guang Han, and Sam Kwong. Privacy-preserving video fall detection via chaotic compressed sensing and GAN-based feature enhancement. IEEE Multi-Media, 29(4):14–23, October/December 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liu:2024:RGC

[LNM^{+}24] Chang Liu, Yi Niu, Mingming Ma, Fu Li, and Guangming Shi. Retinexguided channel groupingbased patch swap for arbitrary style transfer. IEEE MultiMedia, 31(1):7–18, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lowe:1995:PRH

[Low95] David Lowe. Project reports: HyVIS: The hypermedia and visual information systems group. IEEE MultiMedia, 2(3): 70–72, Fall 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u30070.
pdf.

Lim:2013:MEM

[LPL^{+}13] Youngkwon Lim, Kyungmo Park, Jin Young Lee, Shuichi Aoki, and Gerard Fernando. MMT: An emerging MPEG standard for multimedia delivery over the Internet. IEEE MultiMedia, 20(1):80–85, January/March 2013. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Leas:2008:DSA

[LPSZ08] Dennis Leas, Emilia Persoon, Neil Soiffer, and Michael Zacherle. Daisy 3: a standard for accessible multimedia books. IEEE MultiMedia, 15(4): 28–37, October/December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lacey:2007:MRS

[LRCY07] Gerard Lacey, Donncha Ryan, Derek Cassidy, and Derek Young. Mixedreality simulation of minimally invasive surgeries. IEEE MultiMedia, 14(4): 76–87, October/December 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Llorente:2013:SBA

[LRDTP13] Silvia Llorente, Eva Rodriguez, Jaime Delgado, and Victor Torres-Padrosa. Standards-based architectures for content management. IEEE MultiMedia, 20(4):62–72, October/ December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2000:LEC

[LS00] Ambrose Li and Keri Schreiner. Letter to the editor: Clearing up ClearType. IEEE MultiMedia, 7(3):4–5, July/ September 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u3004.

pdf. See [Sch00].

Larson:2017:BIM

[LSG^{+}17] Martha Larson, Mohammad Soleymani, Guillaume Gravier, Bogdan Ionescu, and Gareth J. F. Jones. The benchmarking initiative for multimedia evaluation: MediaEval 2016. IEEE Multi-Media, 24(1):93–96, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        01/mmu2017010093.html.

Liu:2022:ELC

[LSK^{+}22] Chao Liu, Heming Sun, Jiro Katto, Xiaoyang Zeng, and Yibo Fan. An efficient low-complexity convolutional neural network filter. IEEE MultiMedia, 29(2):83–93, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Licsar:2009:FSR

[LSKP09] Attila Licsar, Tamas Sziranyi, Laszlo Kovacs, and Balazs Pataki. A folk song retrieval system with a gesture-based interface. IEEE MultiMedia, 16(3): 48–59, July/September 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2024:CIE

[LSL24] Chengqing Li, Xianhui Shen, and Sheng Liu. Cryptanalyzing an image encryption algorithm underpinned by 2-D lagcomplex logistic map. IEEE MultiMedia, 31(1):

99–109, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lin:2021:SAE

[LSM^{+}21] Fuqiang Lin, Yiping Song, Xingkong Ma, Erxue Min, and Bo Liu. Sentimentaware emoji insertion via sequence tagging. IEEE MultiMedia, 28(2):40–48, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2000:OPR

[LSZM00] Du Li, Chengzheng Sun, Limin Zhou, and Richard R. Muntz. Operation propagation in real-time group editors. IEEE MultiMedia, 7(4):55–61, October/ December 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u4055.
pdf;   http://www.computer.
org/multimedia/mu2000/
u4055abs.htm.

Libeks:2011:YCJ

[LT11] Janis Libeks and Douglas Turnbull. You can judge an artist by an album cover: Using images for music annotation. IEEE MultiMedia, 18(4):30–37, October/ December 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2023:MLB

[LTB23] Pengju Li, Zhiyi Tan, and Bing-Kun Bao. Multiview language bias reduction for visual question answering. IEEE MultiMedia, 30(1):91–99, January/ March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lin:2020:DAF

[LTC^{+}20] Yong-Xiang Lin, Daniel Stanley Tan, Yung-Yao Chen, Ching-Chun Huang, and Kai-Lung Hua. Domain adaptation with foreground/background cues and gated discriminators. IEEE MultiMedia, 27(3): 44–53, July/September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lee:2005:GBO

[LTD05] Haeyoung Lee, Yiying Tong, and Mathieu Desbrun. Geodesics-based oneto-one parameterization of 3D triangle meshes. IEEE MultiMedia, 12(1):27–33, January/March 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lim:2003:HPC

[LTM03] Joo-Hwee Lim, Qi Tian, and Philippe Mulhem. Home photo content modeling for personalized eventbased retrieval. IEEE MultiMedia, 10(4):28–37, October/December 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
04/u4028abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/04/u4028.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/04/
u4028.pdf.

Lu:2007:LMG

[LTS^{+}07] Yijuan Lu, Qi Tian, Maribel Sanchez, Jennifer Neary, Feng Liu, and Yufeng Wang. Learning microarray gene expression data by hybrid discriminant analysis. IEEE Multi-Media, 14(4):22–31, October/December 2007. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2017:STG

[LTS17] Guangda Li and Chua Tat-Seng. Startup takes on growing demand for visual search technology. IEEE MultiMedia, 24(1):17–19, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        01/mmu2017010017-abs.
                                        html.

Luglio:1999:GEI

[Lug99] Michele Luglio. Guest Editor’s introduction: Mobile multimedia satellite communications. IEEE MultiMedia, 6(4):10–14, October/December 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u4010.
                                        pdf.

Luo:2004:MR

[Luo04] Jiebo Luo. Media reviews. IEEE MultiMedia, 11(2):78–79, April/ June 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2004/02/u2078.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/02/
                                        u2078.pdf.

Little:1994:PIV

[LV94] Thomas D. C. Little and Dinesh Venkatesh. Prospects for interactive video-on-demand. IEEE MultiMedia, 1(3):14–24, Autumn/Fall 1994. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2009:MSP

[LVL09] Xiaorong Li, Bharadwaj Veeravalli, and Hailong Li. Multimedia service provisioning and personalization on grid-based infrastructures: Now and the future. IEEE MultiMedia, 16 (1):36–45, January/March 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ludwig:2000:VGC

[LW00] Heiko Ludwig and Keith Whittingham. VEC: Gateways for cross-organizational document flow. IEEE MultiMedia, 7(4):62–72, October/December 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u4062.
pdf;   http://www.computer.
org/multimedia/mu2000/
u4062abs.htm.

Li:2022:SAI

[LW22] Gaozhe Li and Si Wu.

Scene-adaptive instance modification for semisupervised pedestrian detection. IEEE MultiMedia, 29(4): 69–79, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lo:2016:EBI

[LWH16] Kai-Han Lo, Yu-Chiang Frank Wang, and Kai-Lung Hua. Example-based image textural style transfer. IEEE MultiMedia, 23(4):60–66, October/December 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
04/mmu2016040060-abs.
html.

Li:2018:PPD

[LWH^{+}18] T. Li, Y. Wang, R. Hong, M. Wang, and X. Wu. pDisVPL: Probabilistic discriminative visual part learning for image classification. IEEE MultiMedia, 25(4):34–45, October/ December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Li:2017:SPP

[LWL^{+}17] Huining Li, Kun Wang, Xiulong Liu, Yanfei Sun, and Song Guo. A selective privacy-preserving approach for multimedia data. IEEE MultiMedia, 24(4):14–25, October/December 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/

04/mmu2017040014-abs.
html.

Li:2020:STU

[LWL^{+}20] Yaochen Li, Xiao Wu, Danhui Lu, Ling Li, Yuehu Liu, and Li Zhu. Style transfer of urban road images using generative adversarial networks with structural details. IEEE MultiMedia, 27 (3):54–65, July/September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2021:SID

[LWN^{+}21] Hui Li, Qingbo Wu, King Ngi Ngan, Hongliang Li, Meng, and Fanman. Single image dehazing via region adaptive two-shot network. IEEE MultiMedia, 28(3):97–106, July/ September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2022:AAI

[LWW^{+}22] Guanyue Li, Xiwen Wei, Si Wu, Zhiwen Yu, Sheng Qian, and Hau-San Wong. Adversarial adaptive interpolation in autoencoders for dually regularizing representation learning. IEEE MultiMedia, 29(3):57–68, July/September 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liu:2018:MCM

[LWZ^{+}18] R. Liu, S. Wei, Y. Zhao, Z. Zhu, and J. Wang. Multiview cross-media hashing with semantic consistency. IEEE MultiMedia, 25(2): 71–86, April/June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Lu:2022:PSD

[LXH^{+}22] Haifeng Lu, Shihao Xu, Xiping Hu, Edith Ngai, Yi Guo, Wei Wang, and Bin Hu. Postgraduate student depression assessment by multimedia gait analysis. IEEE MultiMedia, 29(2):56–65, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Luo:2019:DSM

[LY19] L. Luo and X. Yang. A 3-D scene management method based on the triangular mesh for large-scale Web3D scenes. IEEE MultiMedia, 26(3):69–78, July/ September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2019:HDC

[LYGT19] J. Li, P. Yuan, D. Gu, and Y. Tian. Hierarchical deep cosegmentation of primary objects in aerial videos. IEEE MultiMedia, 26(3):9– 18, July/September 2019.

CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liu:2015:DDS

[LYL^{+}15] Xionghao Liu, Wei Yang, Liang Lin, Qing Wang, Zhaoquan Cai, and Jianhuang Lai. Data-driven scene understanding with adaptively retrieved exemplars. IEEE Multi-Media, 22(3):82–92, July/ September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/03/mmu2015030082-
abs.html.

Liu:2016:MHN

[LZ16] Wei Liu and Tongtao Zhang. Multimedia hashing and networking. IEEE MultiMedia, 23(3):75–79, July/September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
03/mmu2016030075-abs.
html.

Liu:2022:DZW

[LZD^{+}22] Xiyao Liu, Yayun Zhang, Sibo Du, Jian Zhang, Ming Jiang, and Hui Fang. DIBR zero-watermarking based on invariant feature and geometric rectification. IEEE MultiMedia, 29(3):27–37, July/September 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liao:2019:PRD

[LZDC19] X. L. Liao, C. Zhang, M. Dong, and X. Chen. Person reidentification by deep structured prediction a fully parameterized approach. IEEE MultiMedia, 26(3):42–55, July/ September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Liang:2007:SDT

[LZH^{+}07] Weiqiang Liang, Wenjun Zhang, Dazhi He, Yunfeng Guan, Yao Wang, and Jun Sun. Standards: Digital terrestrial television broadcasting in China. IEEE MultiMedia, 14(3):92–97, July/September 2007. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2023:SLT

[LZH^{+}23] Shibo Li, Shuyuan Zhu, Yuzhou Huang, Shuaicheng Liu, Bing Zeng, Muhammad Ali Imran, Qammer H. Abbasi, and Jonathan Cooper. Shortlong-term propagationbased video inpainting. IEEE MultiMedia, 30(4): 91–104, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Li:2018:MPE

[LZZ^{+}18] B. Li, M. Zhou, Y. Zhang, X. Lin, and W. Guo. Model parameter estimation for CTU-level rate control in HEVC. IEEE MultiMedia, 25(3):79–91, July/ September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Li:2021:PMC

[LZZ^{+}21] Junru Li, Li Zhang, Kai Zhang, Hongbin Liu, Meng Wang, Shiqi Wang, Siwei Ma, and Wen Gao. Prediction with multicross component for future video coding. IEEE MultiMedia, 28(3):52–62, July/ September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ma:1996:MRB

[Ma96] Kai-Kuang Ma. Media reviews: Block truncation coding in perspective: Image Data Compression: Block Truncation Coding, by Belur V. Dasarathy (IEEE Computer Society Press, 1995, 284 pp., $50 list, $40 members, ISBN 0-89186-6876-4). IEEE MultiMedia, 3(2): 83–84, Summer 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2083.
pdf.

Maddage:2006:ASD

[Mad06] Namunu C. Maddage. Automatic structure detection for popular music. IEEE MultiMedia, 13(1):65–77, January/March 2006. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Minami:1998:VHM

[MAHT98] Kenichi Minami, Akihito Akutsu, Hiroshi Hamada, and Yoshinobu Tonomura. Video handling with music and speech detection. IEEE MultiMedia, 5(3):17–25, July/ September 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u3017.
pdf;   http://www.computer.
org/multimedia/mu1998/
u3017abs.htm.

Makela:2005:MSC

[Mak05] Tapio Makela. Multimedia software as culture: Towards critical interaction design. IEEE MultiMedia, 12(1):14–15, January/ March 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Malmud:2000:AMT

[Mal00] A. Deborah Malmud. Artful media: Tech(xt)s; interactive text games. IEEE MultiMedia, 7(4):6–9, October/December 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u4006.
pdf.

May:1996:MWM

[May96] G. S. May. Multimedia at work: Microelectronics fabrication training using multimedia and the Web. IEEE MultiMedia, 3(2):67– 71, Summer 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2067.
pdf.

Mehrotra:1999:NOA

[MB99] Rajiv Mehrotra and Robin Baldwin. In the news: Online auctions attract buyers; xDSL reaches out and touches someone; business trends; Lucent delivers IP telephony; MP3: Dividing communities; the net attracts top musicians; service-level agreements reassure small businesses; digital signature for GSM phones; voice XML forum created. IEEE MultiMedia, 6(2):4–8, April/ June 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u2004.
                                        pdf.

Mehrotra:2000:NDC

[MB00] Rajiv Mehrotra and Robin Baldwin. In the news: Digital copyright battles continue; Web notes; digital privacy; extending HTML; advancing the internet; tech news; portals on edge; Go Network changes direction. IEEE MultiMedia, 7(1):7–11, January/ March 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u1007.
                                        pdf.

Mateescu:2016:VAR

[MB16] Victor A. Mateescu and Ivan V. Bajic. Visual attention retargeting. IEEE MultiMedia, 23(1):82–91, January/March 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        01/mmu2016010082-abs.
                                        html.

Meyer-Boudnik:1995:ME

[MBE95] Thomas Meyer-Boudnik and Wolfgang Effelsberg.

MHEG explained. IEEE MultiMedia, 2(1):26–38, Spring 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u10026.
pdf;   http://www.computer.
org/multimedia/mu1995/
u1026abs.htm.

Mittal:2021:MCA

[MBM21] Trisha Mittal, Aniket Bera, and Dinesh Manocha. Multimodal and context-aware emotion perception model with multiplicative fusion. IEEE MultiMedia, 28(2): 67–75, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Mangla:1998:MWE

[MBPM98] Anamika Mangla, Huzaifah Basrai, Puneeth Punja, and Patricia Morreale. Multimedia at work: Enhancing call-detail reporting using multimedia technologies. IEEE Multi-Media, 5(4):74–77, October/December 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u4074.
pdf.

Min:2014:MST

[MBX14] Weiqing Min, Bing-Kun Bao, and Changsheng Xu. Multimodal spatiotemporal theme modeling for landmark analysis. IEEE MultiMedia, 21 (3):20–29, July/September 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

McGee:2004:TFC

[McG04] Kevin McGee. A touch of the future: Contactexpressive devices. IEEE MultiMedia, 11(1):8–11, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2004/
                                        01/u1008abs.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mu/2004/01/u1008.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/01/
                                        u1008.pdf.

McGandy:2006:AMM

[McG06] Michael J. McGandy. Artful media: Multimedia enlightenment: Transcendent experiences with the help of technology. IEEE MultiMedia, 13(1):4–8, January/March 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

McKerrow:2005:TCC

[McK05] Phillip John McKerrow. Teaching content creation with programming. IEEE MultiMedia, 12(3):36–45, July/September 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Mehrotra:2000:NPH

[MD00] Rajiv Mehrotra and Ben Delaney. In the news: Piracy hits entertainment nerve; getting ready for Bluetooth; tech notes; TV-Net convergence; digital cinema: Ready for opening night; the real advantage of digital cinema; Web resources. IEEE Multi-Media, 7(3):21–25, July/ September 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u3021.
pdf.

Morreale:1997:MWE

[MDK97] Patricia Morreale, Ranjit Deshpande, and Ameet Kamath. Multimedia at work: An electronic database delivery system. IEEE MultiMedia, 4(4):63–67, October/ December 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u4063.
pdf.

Mandilian:2010:IOC

[MDK10] Lauren Mandilian, Paul Diefenbach, and Youngmoo Kim. Information overload: a collaborative dance performance. IEEE MultiMedia, 17(1):8–13, January/March 2010. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Moreno:2015:ITB

[MdRCS15] Marcio Ferreira Moreno, Romualdo Monteiro de Resende Costa, and Luiz Fernando Gomes Soares. Interleaved time bases in hypermedia synchronization. IEEE MultiMedia, 22(4):68–78, October/ December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/04/mmu2015040068-
                                        abs.html.

Mehrotra:1998:MRS

[Meh98] Rajiv Mehrotra. Media reviews: On the shelf. IEEE MultiMedia, 5(2): 92, April/June 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u2090.
                                        pdf.

Mehrotra:1999:NMD

[Meh99] Rajiv Mehrotra. In the news: Mobility drives new markets; voice gains ground on the Internet; fractals model data traffic; DTV add-in cards versus set-top boxes versus convergence; HDTV (US) versus DTT (Europe); dueling over DVD and Divx; trend notes; tech news. IEEE MultiMedia, 6(1):14–19, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u1014.
pdf.

Mehta:2004:CDW

[Meh04] Nina Mehta. The curiously dissociative work of Spencer Finch. IEEE MultiMedia, 11(4):1, 4–7, October/December 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/04/u4001.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/04/
u4001.pdf.

Metz:2000:SDS

[Met00] Chris Metz. Standards:

Differentiated services; acronyms; Web resources. IEEE MultiMedia, 7(3):84–90, July/ September 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u3084.
                                        pdf.

Mich:2003:EDW

[MFG03] Luisa Mich, Mariangela Franch, and Loris Gaio. Evaluating and designing Web site quality. IEEE MultiMedia, 10(1):34–43, January/March 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2003/
                                        01/u1034abs.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mu/2003/01/u1034.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/01/
                                        u1034.pdf.

Muhlhauser:1996:SFP

[MG96] Max Mühlhäuser and Jan Gecsei. Services, frameworks, and paradigms for distributed multimedia applications. IEEE MultiMedia, 3(3):48–61, Fall 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u3048.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1996/
                                        u3048abs.htm.

Marinelli:1999:RPS

[MG99] Mauro Marinelli and Riccardo Giubilei. A regenerative payload for satellite multimedia communications. IEEE Multi-Media, 6(4):30–37, October/December 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u4030.
pdf;   http://www.computer.
org/multimedia/mu1999/
u4030abs.htm.

Martinez:2012:EMT

[MG12] Ignacio Martinez and Jose Garcia. Evaluation methodology for the technical viability of E-health services. IEEE MultiMedia, 19(4):38–47, October/ December 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Mei:2010:KDC

[MHL10] Tao Mei, Winston H. Hsu, and Jiebo Luo. Knowledge discovery from communitycontributed multimedia. IEEE MultiMedia, 17(4): 16–17, October/December 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Milenkovic:1998:DIS

[Mil98] Milan Milenkovic. Delivering interactive services via a digital TV infrastructure. IEEE Multi-Media, 5(4):34–43, October/December 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u4034.
pdf;   http://www.computer.
org/multimedia/mu1998/
u4034abs.htm.

Miranda:2003:FAN

[Mir03]  Maria Miranda.  Fic-

tive art in new media. IEEE MultiMedia, 10(2):4– 7, April/June 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/02/u2004.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/02/
u2004.pdf.

Miyajima:2017:DLT

[Miy17] Ryo Miyajima. Deep learning triggers a new era in industrial robotics. IEEE MultiMedia, 24(4):91–96, October/December 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
04/mmu2017040091-abs.
html.

Megret:2002:TSS

[MJ02] Rémi Mégret and Jean-Michel Jolion. Tracking scale-space blobs for video description. IEEE MultiMedia, 9(2):34–43, April/June 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
02/u2034abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/02/u2034.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/02/
u2034.pdf.

Maculewicz:2015:EEA

[MJSE15] Justyna Maculewicz, Antti Jylha, Stefania Serafin, and Cumhur Erkut. The effects of ecological auditory feedback on rhythmic walking interaction. IEEE Multi-Media, 22(1):24–31, January/March 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/01/mmu2015010024-
abs.html.

Muneesawang:2015:MIA

[MKK^{+}15] Paisarn Muneesawang, Naimul Mefraz Khan, Matthew Kyan, R. Bruce Elder, Nan Dong, Guoyu Sun, Haiyan Li, Ling Zhong, and Ling Guan. A machine intelligence approach to virtual ballet training. IEEE MultiMedia, 22(4):80–92, October/December 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/

                                        mags/mu/2015/04/mmu2015040080-
                                        abs.html.

Mulhem:2003:PVS

[MKYH03] Philippe Mulhem, Mohan S. Kankanhalli, Ji Yi, and Hadi Hassan. Pivot vector space approach for audio-video mixing. IEEE MultiMedia, 10(2):28–40, April/June 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2003/
                                        02/extras/u2028.htm;
                                        http://csdl.computer.
                                        org/comp/mags/mu/2003/
                                        02/u2028abs.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mu/2003/02/u2028.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/02/
                                        u2028.pdf.

Ma:2018:PSP

[ML18] H. Ma and W. Liu. A progressive search paradigm for the Internet of Things. IEEE MultiMedia, 25 (1):76–86, January/March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Ma:2017:NDC

[MLHK17] Dekui Ma, Jian Liang, Ran He, and Xiangwei Kong. Nonlinear discrete cross-modal hashing for visual-textual data. IEEE MultiMedia, 24(2):56–65, April/June 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
02/mmu2017020056-abs.
html.

Merabti:2006:DRM

[MLJ06] Madjid Merabti and David Llewellyn-Jones. Digital rights management in ubiquitous computing. IEEE MultiMedia, 13(2):32–42, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Mansouri:2007:TMS

[MLM^{+}07] Alamin Mansouri, Alexandra Lathuilière, Franck S. Marzani, Yvon Voisin, and Pierre Gouton. Toward a 3D multispectral scanner: An application to multimedia. IEEE MultiMedia, 14 (1):40–47, January/March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Martinez:2003:AWD

[MM03a] José M. Martínez and Francisco Morán. Authoring 744: Writing descriptions to create content. IEEE MultiMedia, 10(4):94–99, October/December 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2003/
                                        04/u4094abs.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mu/2003/04/u4094.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/04/
                                        u4094.pdf.

McGregor-Mento:2003:EWA

[MM03b] Matthew McGregor-Mento. The evolution of Web art—one pioneer’s perspective. IEEE Multi-Media, 10(3):4–8, July/ September 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2003/03/u3004.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/03/
                                        u3004.pdf.

Miller:2014:FNI

[MM14] Kieran Miller and Patricia Morreale. Finding the needle in the image stack: Performance metrics for big data image analysis. IEEE MultiMedia, 21 (1):84–89, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Mendes:2001:WME

[MMC01] Emilia Mendes, Nile Mosley, and Steve Counsell. Web metrics — estimating design and authoring effort. IEEE MultiMedia, 8(1):50–57, January/ March 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u1050.
pdf;   http://www.computer.
org/multimedia/mu2001/
u1050abs.htm.

Moreno:2008:DSM

[MMRM08] Lourdes Moreno, Paloma Martínez, and Belén Ruiz-Mezcua. Disability standards for multimedia on the Web. IEEE MultiMedia, 15(4):52–54, October/ December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Milovanovic:2013:WCH

[MMS13] Milos Milovanovic, Miroslav Minovic, and Dusan Starcevic. Walking in colors: Human gait recognition using Kinect and CBIR. IEEE MultiMedia, 20(4):28–36, October/December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Min:2015:BMB

[MMSB15] Kyung-Yeon Min, Jong-Hyun Ma, Dong-Gyu Sim, and Ivan V. Bajic. Bidirectional mesh-based frame rate up-conversion. IEEE MultiMedia, 22(2):36–45, April/June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/02/mmu2015020036-
                                        abs.html.

Miranda:2007:ABI

[MN07] Maria Miranda and Norie Neumark. Art beat: On the Internet: Turbulence and 1001 nights of networked performance. IEEE MultiMedia, 14(3):1, 4–5, July/September 2007. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Moustakas:2007:HRV

[MNK^{+}07] Konstantinos Moustakas, Georgios Nikolakis, Konstantinos Kostopoulos, Dimitrios Tzovaras, and Michael G. Strintzis. Haptic rendering of visual data for the visually impaired. IEEE MultiMedia, 14(1):62–72, January/March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

MacDorman:2007:MBA

[MNKI07] Karl F. MacDorman, Hiroshi Nobuta, Satoshi Koizumi, and Hiroshi Ishiguro. Memory-based attention control for activity recognition at a subway station. IEEE MultiMedia, 14(2):38–49, April/June 2007. CODEN IEMUE4.

ISSN 1070-986X (print), 1941-0166 (electronic).

Moezzi:1997:GEI

[Moe97] Saied Moezzi. Guest Editors’ introduction: Immersive telepresence. IEEE MultiMedia, 4(1):17, January/March 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u1017.
pdf.

Matsubara:2015:WAB

[MOK^{+}15] Masaki Matsubara, Takahiro Oba, Hideki Kadone, Hiroko Terasawa, Kenji Suzuki, and Masaki Iguchi. Wearable auditory biofeedback device for blind and sighted individuals. IEEE MultiMedia, 22(1):68–73, January/March 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/01/mmu2015010068-
abs.html.

Moskowitz:2003:BC

[Mos03] Scott Moskowitz. Bandwidth as currency. IEEE MultiMedia, 10(1):14–21, January/March 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/01/u1014.

                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/01/
                                        u1014.pdf.

Moussa:2006:PBM

[Mou06] Ahmed Shawky Moussa. Perception-based microtuning over MIDI networks. IEEE MultiMedia, 13(1):56–64, January/ March 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Mansilla:2017:MPM

[MP17] Wendy Ann Mansilla and Andrew Perkis. Multiuse playspaces: Mediating expressive community places. IEEE Multi-Media, 24(1):12–16, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        01/mmu2017010012-abs.
                                        html.

Mayer-Patel:2007:SAS

[MPG07] Ketan Mayer-Patel and David Gotz. Scalable, adaptive streaming for nonlinear media. IEEE MultiMedia, 14(3):68–83, July/September 2007. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Marinelli:2000:MWE

[MPL00] Donald Marinelli, Randy Pausch, and Janeen LaForce. Multimedia at work: Entertainment technology center; building virtual worlds; CUBE equipment: Make your own at home. IEEE Multi-Media, 7(4):78–81, October/December 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u4078.
pdf;   http://www.computer.
org/multimedia/mu2000/
../articles/entertain.
htm.

Maestre:2017:EMR

[MPM^{+}17] Esteban Maestre, Panagiotis Papiotis, Marco Marchini, Quim Llimona, Oscar Mayor, Alfonso Perez, and Marcelo M. Wanderley. Enriched multimodal representations of music performances: Online access and visualization. IEEE MultiMedia, 24(1):24–34, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
01/mmu2017010024-abs.
html.

Moens:2018:VLI

[MPST18] M. Moens, K. Pastra, K. Saenko, and T. Tuytelaars. Vision and language integration meets multimedia fusion. IEEE MultiMedia, 25(2):7–10, April/June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Morreale:2010:RTE

[MQC^{+}10] Patricia Morreale, Feng Qi, Paul Croft, Ryan Suleski, Brian Sinnicke, and Francis Kendall. Real-time environmental monitoring and notification for public safety. IEEE MultiMedia, 17(2):4–11, April/June 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Miralles:1995:MRU

[MS95] Montserrat Miralles and Juan Sancho. Media reviews: Understanding Hy-Time: Making Hypermedia Work: A User’s Guide to HyTime by Steven J. DeRose and David G. Durand (Kluwer Academic Publishers, 1994, ISBN 0-7923-9432-1, 384 pp., $68). IEEE MultiMedia, 2(3): 75–76, Fall 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u30075.
                                        pdf.

Manap:2016:NQA

[MSF16] Redzuan Abdul Manap, Ling Shao, and Alejandro F. Frangi. Nonparametric quality assessment of natural images. IEEE MultiMedia, 23(4):22–30, October/December 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
04/mmu2016040022-abs.
html.

Mu:2015:LSS

[MSR^{+}15] Mu Mu, Steven Simpson, Nicholas Race, Omar Niamut, Gijs Koot, Rianne Kaptein, Jacco Taal, and Luigi Mori. “Let’s Share a Story”: Socially enhanced multimedia storytelling. IEEE Multi-Media, 22(3):54–65, July/ September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/03/mmu2015030054-
abs.html.

Mase:2006:UEM

[MST^{+}06] Kenji Mase, Yasuyuki Sumi, Tomoji Toriyama, Megumu Tsuchikawa, Sadanori Ito, Shoichiro Iwasawa, Kiyoshi Kogure, and Norihiro Hagita. Ubiquitous experience media. IEEE MultiMedia, 13(4):20–29, October/December 2006. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Malloch:2018:GMI

[MSW18] J. Malloch, S. Sinclair, and M. M. Wanderley. Generalized multi-instance control mapping for interactive media systems. IEEE MultiMedia, 25(1):39–50, January/March 2018. CO-DEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Magnenat-Thalmann:2006:VVH

[MTB06] Nadia Magnenat-Thalmann and Ugo Bonanni. Visions and views: Haptics in virtual reality and multimedia. IEEE MultiMedia, 13 (3):6–11, July/September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Moustakas:2011:UMR

[MTD^{+}11] Konstantinos Moustakas, Dimitrios Tzovaras, Laila Dybkjaer, Niels Ole Bernsen, and Oya Aran. Using modality replacement to facilitate communication between visually and hearing-impaired people. IEEE MultiMedia, 18(2): 26–37, April/June 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Moezzi:1997:VVG

[MTG97] Saied Moezzi, Li-Cheng Tai, and Philippe Gerard. Virtual view generation for 3D digital video. IEEE MultiMedia, 4(1):18–26, January/March 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u1018.
pdf;   http://www.computer.
org/multimedia/mu1997/
u1018abs.htm.

Min:2020:UMC

[MTH^{+}20] Weiqing Min, Yonghong Tian, Zi Huang, Wen-Huang Cheng, and Abdulmotaleb El Saddik. Urban multimedia computing: Emerging methods in multimedia computing for urban data analysis and applications. IEEE MultiMedia, 27(3):8–11, July/ September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Moustakas:2006:MWM

[MTS^{+}06] Konstantinos Moustakas, Dimitrios Tzovaras, Michael G. Strintzis, Sebastien Carbini, Olivier Bernier, Jean Emmanuel Viallet, Stephan Raidt, Matei Mancas, Mariella Dimiccoli, Enver Yagci, Serdar Balci, and Eloisa Ibanez Leon. Multimedia at work: Masterpiece: Physical interaction and 3D contentbased search in VR applications. IEEE Multi-Media, 13(3):92–100, July/ September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Miller:1997:DCD

[MTW97] Renée J. Miller, Odysseas G. Tsatalos, and John H. Williams. DataWeb: Customizable database publishing for the Web. IEEE MultiMedia, 4(4):14–21, October/December 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://dlib.computer.
                                        org/mu/books/mu1997/pdf/
                                        u4014.pdf;   http://www.
                                        computer.org/multimedia/
                                        mu1997/u4014abs.htm.

Ma:2019:RBP

[MTXH19] L. Ma, Y. Tian, P. Xing, and T. Huang. Residualbased post-processing for HEVC. IEEE MultiMedia, 26(4):67–79, October/ December 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Murchu:2015:CMC

[Mur15] Nora O’ Murchu. Collaborative modes of curating software. IEEE MultiMedia, 22(1):88–92, January/March 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/01/mmu2015010088-
abs.html.

Muller:2011:MIC

[MW11] Christian Müller and Garrett Weinberg. Multimodal input in the car, today and tomorrow. IEEE Multi-Media, 18(1):98–103, January/March 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Miao:2021:MSB

[MXH21] Naiyang Miao, Feng Xue, and Richang Hong. Multimodal semantics-based supervised latent Dirichlet allocation for event classification. IEEE MultiMedia, 28(4):8–17, October/ December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Mallick:2022:DRS

[MYBP^{+}22] Rupayan Mallick, Thinhinane Yebda, Jenny Benois-Pineau, Akka Zemmari, Marion Pech, and Hélène Amieva. Detection of risky situations for frail adults with hybrid neural networks on multimodal health data. IEEE MultiMedia, 29(1):7–17, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Macedonia:1997:TNV

[MZ97] Michael R. Macedonia and Michael J. Zyda. A taxonomy for networked virtual environments. IEEE MultiMedia, 4(1):48–56, January/March 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1997/pdf/u1048.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1997/
                                        u1048abs.htm.

Mukti:2001:MWC

[MZSH01] Norhayati Abd. Mukti, Halimah Badioze Zaman, Tengku Mohd. Tengku Sembok, and Siew Pei Hwa. Multimedia at work: Courseware development to motivate life-long reading habits. IEEE Multi-Media, 8(4):76–81, October/December 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u4076.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2001/
                                        u4076abs.htm.

Ma:2016:NLF

[MZZ^{+}16] Siwei Ma, Xinfeng Zhang, Jian Zhang, Chuanmin Jia, Shiqi Wang, and Wen Gao. Nonlocal in-loop filter: The way toward nextgeneration video coding? IEEE MultiMedia, 23(2): 16–26, April/June 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
02/mmu2016020016-abs.
html.

Nack:2000:MIA

[Nac00] Frank Nack. Media impact: All content counts: The future in digital media computing is meta; welcome to media impact; Web resources. IEEE MultiMedia, 7(3):10–13, July/ September 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u3010.
pdf.

Nack:2001:MIP

[Nac01] Frank Nack. Media impact: Play the game; Web resources. IEEE MultiMedia, 8(1):8–10, January/March 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u1008.
pdf.

Nack:2003:AC

[Nac03a] Frank Nack. Aesthetics of contradiction. IEEE MultiMedia, 10(1):11–13, January/March 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2003/01/u1011.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/01/
                                        u1011.pdf.

Nack:2003:MMT

[Nac03b]  Frank Nack.  Migrat-

ing from mobile telephony to multipurpose gadgets. IEEE MultiMedia, 10(2):8– 11, April/June 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2003/02/u2008.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/02/
                                        u2008.pdf.

Nack:2004:FDM

[Nac04]  Frank Nack.  The fu-

ture in digital media computing is meta. IEEE MultiMedia, 11(2):10–13, April/June 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2004/02/u2010.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/02/
                                        u2010.pdf.

Nack:2005:YMR

[Nac05] Frank Nack. You must remember this. IEEE MultiMedia, 12(1):4–7, January/March 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nack:2007:MIF

[Nac07] Frank Nack. Media impact:

Faithfulness is not static. IEEE MultiMedia, 14(2): 12–15, April/June 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nack:2009:LB

[Nac09] Frank Nack. Let’s blend.

IEEE MultiMedia, 16(4): 2–4, October/December 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nack:2010:AR

[Nac10a] Frank Nack. Add to the real. IEEE MultiMedia, 17(1):4–7, January/March 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nack:2010:CIM

[Nac10b] Frank Nack. Color, interaction, media. IEEE Multi-Media, 17(4):4–7, October/ December 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nack:2010:SC

[Nac10c] Frank Nack. Social compatibility. IEEE MultiMedia, 17(3):4–7, July/ September 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nack:2014:CCB

[Nac14] Frank Nack. A closing circle: Back to authoring. IEEE MultiMedia, 21(3): 4–6, July/September 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nam:2022:DDM

[NAK^{+}22] Seung-Hun Nam, Wonhyuk Ahn, Myung-Joon Kwon, Jihyeon Kang, and In-Jae Yu. DHNet: Double MPEG-4 compression detection via multiple DCT histograms. IEEE Multi-Media, 29(2):11–22, April/ June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Naugle:1998:VVD

[Nau98]  Lisa Naugle.  Visions

and views: Digital dancing. IEEE MultiMedia, 5(4):8–12, October/ December 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/

books/mu1998/pdf/u4008.
pdf.

Nack:2004:AMA

[NAV04] Frank Nack, Brett Adams, and Svetha Venkatesh. Authoring multimedia authoring tools. IEEE Multi-Media, 11(3):1, 4–6, July/ September 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/03/u3001.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/03/
u3001.pdf.

Nwana:2017:QUO

[NC17] Amandianeze O. Nwana and Tsuhan Chen. Querying users as oracles in tag engines for personalized image tagging. IEEE MultiMedia, 24(2):66–75, April/June 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
02/mmu2017020066-abs.
html.

Neumark:2000:AMM

[Neu00] Norie Neumark. Artful media: Making contact with artful CD-ROMs. IEEE MultiMedia, 7(1):4–6, January/March 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u1004.
                                        pdf.

New:1994:DMN

[New94] Darren New. Delivering multimedia over narrowband networks. IEEE MultiMedia, 1(3):74–76, Autumn/Fall 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Newcomb:1995:SMIa

[New95a] Steven R. Newcomb. Standards: Multimedia interchange using SGML/ HyTime. Part I: Structures. IEEE MultiMedia, 2(2):86–89, Summer 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u20086.
                                        pdf.

Newcomb:1995:SMIb

[New95b] Steven R. Newcomb. Standards: Multimedia interchange using SGML/ HyTime. Part II: Principles and applications. IEEE MultiMedia, 2(3): 60–64, Fall 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/

books/mu1995/pdf/u30060.
pdf.

Newsam:2010:CWW

[New10] Shawn Newsam. Crowdsourcing: What is where: Community-contributed photos as volunteered geographic information. IEEE MultiMedia, 17(4):36–45, October/December 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nafaa:2011:IAP

[NG11] Abdelhamid Nafaa and Manon Gaucher. Implementation and analysis of a peer-to-peer retransmissions system for live video services. IEEE MultiMedia, 18(2):60–71, April/June 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Namitha:2023:IIE

[NGA23]  Kalakunnath  Namitha,

Madathilkulangara Geetha, and Narayanan Athi. An improved interaction estimation and optimization method for surveillance video synopsis. IEEE MultiMedia, 30(3):25–36, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nigten:2006:MIP

[NJ06] Anne Nigten and Simon Jones. Media impact:

Play! IEEE MultiMedia, 13(4):8–11, October/ December 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nie:2022:LLI

[NKZ^{+}22] Weizhi Nie, Yuqi Ke, Yue Zhao, Qi Liang, and Yuting Su. LIMAN: Local information-based multiattention network for 3D shape recognition. IEEE MultiMedia, 29(1):65–73, January/March 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nack:1999:EYWa

[NL99a] Frank Nack and Adam T. Lindsay. Everything you wanted to know about MPEG-7: Part 1. IEEE MultiMedia, 6(3):65–77, July/September 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u3065.
pdf;   http://www.computer.
org/multimedia/mu1999/
u3065abs.htm.

Nack:1999:EYWb

[NL99b] Frank Nack and Adam T. Lindsay. Everything you wanted to know about MPEG-7: Part 2. IEEE MultiMedia, 6(4):64–73, October/December 1999.

CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1999/pdf/
u4064.pdf;   http://www.
computer.org/multimedia/
mu1999/u4064abs.htm.

Nanni:2012:CFE

[NL12] Loris Nanni and Alessandra Lumini. Combining face and eye detectors in a high-performance facedetection system. IEEE MultiMedia, 19(4):20–27, October/December 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nikolovska:1998:PRS

[NM98] Lira Nikolovska and Jacquelyn Martino. Project reports: Spatial browsing to retrieve multimedia information. IEEE MultiMedia, 5(2):78–83, April/June 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u2078.
pdf.

Nakao:2006:TBK

[NMK^{+}06] Megumi Nakao, Kotaro Minato, Tomohiro Kuroda, Masaru Komori, Hiroshi Oyama, and Takashi Takahashi. Transferring bioelasticity knowledge through haptic interaction. IEEE MultiMedia, 13(3):50–60, July/September 2006. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Naaman:2008:ZCT

[NN08] Mor Naaman and Rahul Nair. ZoneTag’s collaborative tag suggestions: What is this person doing in my phone? IEEE Multi-Media, 15(3):34–40, July/ September 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nejdl:2015:PRP

[NN15] Wolfgang Nejdl and Claudia Niederee. Photos to remember, photos to forget. IEEE MultiMedia, 22(1):6–11, January/ March 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/01/mmu2015010006-
                                        abs.html.

Negrel:2013:WSI

[NPG13] Romain Negrel, David Picard, and Philippe-Henri Gosselin. Web-scale image retrieval using compact tensor aggregation of visual descriptors. IEEE Multi-Media, 20(3):24–33, July/ September 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nahrstedt:1995:QB

[NS95] Klara Nahrstedt and Jonathan M.

Smith. The QOS Broker. IEEE MultiMedia, 2(1): 53–67, Spring 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u10053.
pdf;   http://www.computer.
org/multimedia/mu1995/
u1053abs.htm.

Nagao:2001:SAT

[NSS01] Katashi Nagao, Yoshinari Shirrai, and Kevin Square. Semantic annotation and transcoding: Making Web content more accessible. IEEE MultiMedia, 8(2):69– ??, April/June 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u2069.
pdf;   http://www.computer.
org/multimedia/mu2001/
u2069abs.htm.

Naphade:2006:SLS

[NST^{+}06] Milind Naphade, John R.

Smith, Jelena Tesic, Shih-Fu Chang, Winston Hsu, Lyndon Kennedy, Alexander Hauptmann, and Jon Curtis. Standards: Largescale concept ontology for multimedia. IEEE Multi-Media, 13(3):86–91, July/ September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nwosu:1997:GEI

[NTB97] Kingsley C. Nwosu, Bhavani Thuraisingham, and P. Bruce Berra. Guest Editors’ introduction: Multimedia database systems — A new frontier. IEEE MultiMedia, 4(3):21–23, July/ September 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1997/pdf/u3021.
                                        pdf.

Nack:2005:OOD

[NvOH05] Frank Nack, Jacco van Ossenbruggen, and Lynda Hardman. That obscure object of desire: Multimedia metadata on the Web, Part 2. IEEE MultiMedia, 12(1):54–63, January/March 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ngo:2013:WSN

[NXKS13] Chong-Wah Ngo, Changsheng Xu, Wessel Kraaij, and Abdulmotaleb El Saddik. Web-scale nearduplicate search: Techniques and applications. IEEE MultiMedia, 20(3): 10–12, July/September 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Nakanishi:1999:FVS

[NYNI99] Hideyuki Nakanishi, Chikara Yoshida, Toshikazu Nishimura, and Toru Ishida. Free-Walk: a 3D virtual space for casual meetings. IEEE MultiMedia, 6(2):20–28, April/June 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u2020.
pdf;   http://www.computer.
org/multimedia/mu1999/
u2020abs.htm.

Ox:2000:AMC

[OB00] Jack Ox and David Britton. Artful media: The 21st Century virtual reality color organ; the technology behind the color organ. IEEE Multi-Media, 7(3):6–9, July/ September 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u3006.
pdf.

Obrist:2017:MEH

[OGM^{+}17] Marianna Obrist, Elia Gatti, Emanuela Maggioni, Chi Thanh Vi, and Carlos Velasco. Multisensory experiences in HCI.

IEEE MultiMedia, 24(2):9– 13, April/June 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
02/mmu2017020009-abs.
html.

OHare:2014:MGC

[OK14] Neil O’Hare and Yiannis Kompatsiaris. The Multimedia Grand Challenge 2013. IEEE MultiMedia, 21 (1):90–94, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ohya:1999:VM

[OKN^{+}99] Jun Ohya, Jun Kurumisawa, Ryohei Nakatsu, Kazuyuki Ebihara, Shoichiro Iwasawa, David Harwood, and Thanarat Horprasert. Virtual metamorphosis. IEEE MultiMedia, 6(2):29– 39, April/June 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u2029.
pdf;   http://www.computer.
org/multimedia/mu1999/
u2029abs.htm.

Over:2004:MRB

[OLIG04] Paul Over, Clement Leung, Horace Ip, and Michael Grubinger. Multimedia retrieval benchmarks. IEEE MultiMedia, 11(2):80–84, April/June 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/02/u2080.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/02/
u2080.pdf.

OGrady:2004:JTM

[OO04] Michael J. O’Grady and Gregory M. P. O’Hare. Just-in-time multimedia distribution in a mobile computing environment. IEEE MultiMedia, 11(4):62–74, October/ December 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/04/u4062.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/04/
u4062.pdf.

Oomen:2011:AAH

[Oom11] Johan Oomen. Accessing audiovisual heritage: a roadmap for collaborative innovation. IEEE Multi-Media, 18(4):4–10, October/December 2011. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Olsina:2002:MWA

[OR02] Luis Olsina and Gustavo Rossi. Measuring Web application quality with WebQEM. IEEE Multi-Media, 9(4):20–29, October/December 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
04/u4020abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/04/u4020.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/04/
u4020.pdf.

Obrenovic:2004:MDA

[OSS04] Zeljko Obrenovic, Dusan Starcevic, and Bran Selic. A model-driven approach to content repurposing. IEEE Multi-Media, 11(1):62–71, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
01/u1062abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2004/01/u1062.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/01/
u1062.pdf.

Oviatt:1996:UCM

[Ovi96]  Sharon Oviatt.  User-

centered modeling for spoken language and multimodal interfaces. IEEE MultiMedia, 3(4):26–35, Winter 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u4026.
pdf;   http://www.computer.
org/multimedia/mu1996/
u4026abs.htm.

Ou:2020:ABM

[OWZ^{+}20] Qiyuan Ou, Siwei Wang, Sihang Zhou, Miaomiao Li, Xifeng Guo, and En Zhu. Anchor-based multiview subspace clustering with diversity regularization. IEEE Multi-Media, 27(4):91–101, October/December 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Packer:1999:AMJ

[Pac99] Randall Packer. Artful media: Just what is multimedia, anyway? IEEE MultiMedia, 6(1):11–13, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u1011.
pdf.

Paine:2013:NMI

[Pai13] Garth Paine. New musical instrument design considerations. IEEE MultiMedia, 20(4):76–84, October/ December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Pan:1995:TMA

[Pan95] Davis Pan. A tutorial on MPEG/audio compression. IEEE MultiMedia, 2(2):60– 74, Summer 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u20060.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1995/
                                        u2060abs.htm.

Panchanathan:2002:VVU

[Pan02] Sethuraman Panchanathan. Visions and views: Ubiquitous multimedia: Bridging the digital divide. IEEE MultiMedia, 9(1):12–15, January/March 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2002/pdf/u1012.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2002/
                                        u1012abs.htm.

Panchanathan:2006:EMJ

[Pan06] Sethuraman Panchanathan. EIC’s message: The journey ahead. IEEE Multi-Media, 13(2):1, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Panchanathan:2007:EMC

[Pan07a] Sethuraman Panchanathan. EIC’s message: Creativity at the fusion of disciplines.

IEEE MultiMedia, 14(1):1, January/March 2007. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Panchanathan:2007:ECL

[Pan07b] Sethuraman Panchanathan.

From the Editor in Chief: Looking into the future. IEEE MultiMedia, 14(3): c2, July/September 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Panchanathan:2008:EML

[Pan08] Sethuraman Panchanathan.

EIC’s message: Looking forward: New ideas, new opportunities. IEEE MultiMedia, 15(1):c2, 1, January/March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Park:1998:MRD

[Par98] Youngchoon Park. Media reviews: Design paradigms for image databases. IEEE MultiMedia, 5(1):102, January/March 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u1102.
pdf.

Parker:1999:SDD

[Par99] Dana J. Parker. Standards:

Defining DVD. IEEE MultiMedia, 6(1):80–84, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u1080.
pdf.

Parekh:2012:UTA

[Par12] Ranjan Parekh. Using texture analysis for medical diagnosis. IEEE MultiMedia, 19(2):28–37, May/June 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Pande:2013:VDC

[PAS^{+}13] Amit Pande, Vishal Ahuja, Rajarajan Sivaraj, Eilwoo Baik, and Prasant Mohapatra. Video delivery challenges and opportunities in 4G networks. IEEE MultiMedia, 20(3):88–94, July/ September 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Popescu-Belis:2012:FIM

[PBLB12] Andrei Popescu-Belis, Denis Lalanne, and Herve Bourlard. Finding information in multimedia meeting records. IEEE MultiMedia, 19(2):48–57, May/June 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Prattichizzo:2007:PIH

[PBM^{+}07] Domenico Prattichizzo, Mauro Barni, Gloria Menegaz, Alessandro Formaglio, Hong Z. Tan, and Seungmoon Choi. Perceptual issues in haptic digital watermarking. IEEE MultiMedia, 14(3): 84–91, July/September 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Pletinckx:2000:VRH

[PCKS00] Daniel Pletinckx, Dirk Callebaut, Ann E. Killebrew, and Neil A. Silberman. Virtual-reality heritage presentation at Ename. IEEE MultiMedia, 7(2):45–48, April/ June 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2045.
pdf;   http://www.computer.
org/multimedia/mu2000/
u2045abs.htm.

Panchanathan:2016:PCM

[PCMT16] Sethuraman Panchanathan, Shayok Chakraborty, Troy McDaniel, and Ramin Tadayon. Person-centered multimedia computing: A new paradigm inspired by assistive and rehabilitative applications. IEEE MultiMedia, 23(3):12–19, July/September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        03/mmu2016030012-abs.
                                        html.

Papadopoulos:2015:SMS

[PCS^{+}15]  Symeon  Papadopoulos,

Pablo Cesar, David A. Shamma, Aisling Kelliher, and Ramesh Jain. Social multimedia and storytelling [Guest editors’ introduction]. IEEE Multi-Media, 22(3):10–13, July/ September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/03/mmu2015030010-
                                        abs.html.

Pierucci:2000:IMS

[PD00] Laura Pierucci and Enrico Del Re. An interactive multimedia satellite telemedicine service. IEEE MultiMedia, 7(2):76– 83, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u2076.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2000/
                                        u2076abs.htm.

Phelps:2009:GCR

[PEB09a] Andrew M. Phelps, Christopher A. Egert, and Jessica D. Bayliss. Games in the classroom at the Rochester Institute of Technology: a case study. IEEE MultiMedia, 16(3): 82, July/September 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Phelps:2009:GCU

[PEB09b] Andrew M. Phelps, Christopher A. Egert, and Jessica D. Bayliss. Games in the classroom: Using games as a motivator for the study of computing: Part 2. IEEE MultiMedia, 16(3):2–7, July/September 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Phelps:2009:MIG

[PEB09c] Andrew M. Phelps, Christopher A. Egert, and Jessica D. Bayliss. Media impact: Games in the classroom: Using games as a motivator for studying computing: Part 1. IEEE MultiMedia, 16(2):4– 8, April/June 2009. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Peterson:1996:MRC

[Pet96]  Steve Peterson.  Me-

dia reviews: a cornerstone for a new niche: A Technical Introduction to Digital Video by Charles A. Poynton (John Wiley & Sons, 1996, 352 pp., $42.95, ISBN 0-471-12253-X). IEEE MultiMedia, 3 (3):91, Fall 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u3090.
pdf.

Pourmohammadi-Fallah:2003:IDM

[PFAHA03] Yasser Pourmohammadi-Fallah, Kambiz Asrar-Haghighi, and Hussein Alnuweiri. Internet delivery of MPEG-4 objectbased multimedia. IEEE MultiMedia, 10(3):68–78, July/September 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
03/u3068abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/03/u3068.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/
u3068.pdf.

Perez:2023:DAE

[PFMMF23] Francisco Maciá Pérez, Iren Lorenzo Fonseca, José Vicente Berná Martínez, and Alex Maciá-Fiteni. Distributed architecture for an elderly accompaniment service based on IoT devices, AI, and cloud services. IEEE MultiMedia, 30(1):17–27, January/ March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Prockup:2013:OPC

[PGHK13] Matthew Prockup, David Grunberg, Alex Hrybyk, and Youngmoo E. Kim. Orchestral performance companion: Using realtime audio to score alignment. IEEE MultiMedia, 20(2):52–60, April/June 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Patel:1994:MFM

[PHG94] Sanjiv P. Patel, Grant Henderson, and Nicolas D. Georganas. The multimedia fax-MIME gateway. IEEE MultiMedia, 1(4):64– 70, Winter 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Podlaseck:2003:CPG

[PHS03] Mark Podlaseck, Robert Hoch, and Edith Schoenberg. This is not a catalog: Philip Glass’s works online. IEEE MultiMedia, 10(4):4–7, October/ December 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
04/u4004abs.htm;   http:

                                        //csdl.computer.org/dl/
                                        mags/mu/2003/04/u4004.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/04/
                                        u4004.pdf.

Picard:2016:ARS

[Pic16] Rosalind W. Picard. Automating the recognition of stress and emotion: From lab to real-world impact. IEEE MultiMedia, 23(3): 3–7, July/September 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        03/mmu2016030003-abs.
                                        html.

Panchanathan:2006:GEI

[PK06] Sethuraman Panchanathan and Kanav Kahol. Guest Editors’ introduction: Haptic user interfaces for multimedia systems. IEEE MultiMedia, 13(3):22–23, July/ September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Park:2015:EAF

[PKL15] Yongki Park, Jaehoon Kim, and Kyogu Lee. Effects of auditory feedback on menu selection in handgesture interfaces. IEEE MultiMedia, 22(1):32–40, January/March 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/01/mmu2015010032-
abs.html.

Papadopoulos:2004:RSI

[PKN^{+}04] Christos Papadopoulos, Chris Kyriakakis, Ulrich Neumann, Alexander A. Sawchuk, Cyrus Shahabi, and Roger Zimmermann. RMI system: Internet meets the future home theater. IEEE MultiMedia, 11(2):48–57, April/June 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
02/u2048abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2004/02/u2048.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/02/
u2048.pdf.

Platt:1996:SNS

[Pla96] Richard Platt. Standards:

New standard helps multimedia get off the ground. IEEE MultiMedia, 3(2):78– 82, Summer 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2078.
pdf.

Platt:2000:OFP

[Pla00] J. Platt. Optimal filtering for patterned displays. IEEE Signal Processing Letters, 7(7):179– 180, 2000. CODEN IS-PLEM. ISSN 1070-9908 (print), 1558-2361 (electronic). URL http:

//research.microsoft.
com/users/jplatt/clearType/
optFilt.pdf;   http://
research.microsoft.com/
users/jplatt/clearType/
optFilt.ps.gz.

Park:2011:OVR

[PLL^{+}11] Jonghun Park, Sang-Jin Lee, Sung-Jun Lee, Kwanho Kim, Beom-Suk Chung, and Yong-Ki Lee. Online video recommendation through tag-cloud aggregation. IEEE Multi-Media, 18(1):78–87, January/March 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Pierucci:2016:NNQ

[PM16] Laura Pierucci and Davide Micheli. A neural network for quality of experience estimation in mobile communications. IEEE MultiMedia, 23(4):42–49, October/December 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
04/mmu2016040042-abs.
html.

Paniagua-Martin:2011:SAA

[PMGCCPRM11] Fernando Paniagua-Martin, Angel Garcia-Crespo, Ricardo Colomo-Palacios, and Belen Ruiz-Mezcua. Semantic annotation architecture for accessible multimedia resources. IEEE MultiMedia, 18(2):16–25, April/June 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Pea:2004:DPI

[PMR^{+}04] Roy Pea, Michael Mills, Joseph Rosen, Kenneth Dauber, Wolfgang Effelsberg, and Eric Hoffert. The Diver Project: Interactive digital video repurposing. IEEE Multi-Media, 11(1):54–61, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
01/u1054abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2004/01/u1054.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/01/
u1054.pdf.

Park:2010:VWT

[PMV^{+}10] Soohong Park, Erik Mannens, Rik Van de Walle, Joakim Soderberg, Glenn Adams, Philippe Le Hegaret, and Choong Seon Hong.

Video in the Web: Technical challenges and standardization. IEEE Multi-Media, 17(4):90–93, October/December 2010. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Paracchini:2022:FSS

[PMV^{+}22] Marco Paracchini, Marco Marcon, Federica Villa, Iris Cusini, and Stefano Tubaro. Fast skin segmentation on low-resolution grayscale images for remote PhotoPlethysmoGraphy. IEEE MultiMedia, 29 (1):28–35, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Pande:2013:SMC

[PMZ13] Amit Pande, Prasant Mohapatra, and Joseph Zambreno. Securing multimedia content using joint compression and encryption. IEEE MultiMedia, 20(4):50–61, October/ December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Parker:2005:VBC

[PP05] Conrad Parker and Silvia Pfeiffer. Video blogging: Content to the max. IEEE MultiMedia, 12(2):4– 8, April/June 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Pentland:1994:VIS

[PPDH94] Alex Pentland, Rosalind W.

Picard, Glorianna Davenport, and Kenneth B. Haase. Video and image semantics: Advanced tools for telecommunications. IEEE MultiMedia, 1(2):73–75, Summer 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Pradhan:2008:HGC

[PPL08] Gaurav Pradhan, Balakrishnan Prabhakaran, and Chuanjun Li. Hand-gesture computing for the hearing and speech impaired. IEEE MultiMedia, 15(2): 20–27, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Patrikakis:2011:PCL

[PPP^{+}11] Charalampos Z. Patrikakis, Nikolaos Papaoulakis, Panagiotis Papageorgiou, Aristodemos Pnevmatikakis, Paul Chippendale, Mario S. Nunes, Rui Santos Cruz, Stefan Poslad, and Zhenchen Wang. Personalized coverage of large athletic events. IEEE MultiMedia, 18(4): 18–29, October/December 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Prabhakaran:2023:TDL

[Pra23a] Balakrishnan Prabhakaran. Taking a “deep” look at multimedia streaming. IEEE MultiMedia, 30(3):5– 7, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Prabhakaran:2023:WNT

[Pra23b] Balakrishnan Prabhakaran. Welcome to the new team members. IEEE MultiMedia, 30(1):5–6, January/ March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Prabhakaran:2024:MDS

[Pra24] Balakrishnan Prabhakaran. Multimedia data and security. IEEE MultiMedia, 31(1):5–6, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Pachet:2000:CAC

[PRC00] François Pachet, Pierre Roy, and Daniel Cazaly. A combinatorial approach to content-based music selection. IEEE MultiMedia, 7(1):44–51, January/ March 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u1044.
                                        pdf;   http://www.computer.

org/multimedia/mu2000/
u1044abs.htm.

Priscoli:1999:UAI

[Pri99] Francesco Delli Priscoli.

UMTS architecture for integrating terrestrial and satellite systems. IEEE MultiMedia, 6(4):38–45, October/December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1999/pdf/
u4038.pdf;   http://www.
computer.org/multimedia/
mu1999/u4038abs.htm.

Pazandak:1997:EOD

[PS97] Paul Pazandak and Jaideep Srivastava. Evaluating object DBMSs for multimedia. IEEE Multi-Media, 4(3):34–49, July/ September 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3034.
pdf;   http://www.computer.
org/multimedia/mu1997/
u3034abs.htm.

Pugliese:2015:STH

[PT15] Roberto Pugliese and Tapio Takala. Sonic trampoline: How audio feedback impacts the user’s experience of jumping. IEEE MultiMedia, 22(1):74–79, January/March 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/01/mmu2015010074-
                                        abs.html.

Pham:2011:NPN

[PTM11] Phi The Pham, Tinne Tuytelaars, and Marie-Francine Moens. Naming people in news videos with label propagation. IEEE MultiMedia, 18(3):44–55, July/September 2011. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Purcell:1997:VVD

[Pur97] Patrick Purcell. Visions and views: Digital media and the primacy of new content. IEEE MultiMedia, 4(4):9–13, October/December 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1997/pdf/u4009.
                                        pdf.

Purchase:1998:VVD

[Pur98] Helen Purchase. Visions and views: Defining multimedia. IEEE MultiMedia, 5(1):8–15, January/ March 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u1008.
                                        pdf.

Portales:2010:AIC

[PVAMM10] Cristina Portales, Maria J.

Vinals, Pau Alonso-Monasterio, and Maryland Morant. AR-immersive cinema at the Aula Natura Visitors Center. IEEE MultiMedia, 17(4):8–15, October/ December 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Palaiokrassas:2016:SMI

[PVK^{+}16] Georgios Palaiokrassas, Athanasios Voulodimos, Kleopatra Konstanteli, Nicholas Vretos, David Salama Osborne, Efstathia Chatzi, Petros Daras, and Theodora Varvarigou. Social media interaction and analytics for enhanced educational experiences. IEEE Multi-Media, 23(1):26–35, January/March 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
01/mmu2016010026-abs.
html.

Peng:2010:KBV

[PXL10] Jiang Peng and Qin Xiao-Lin. Keyframe-based video summary using visual attention clues. IEEE Multi-Media, 17(2):64–73, April/ June 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Papadopoulos:2011:CBL

[PZKV11]  Symeon  Papadopoulos,

Christos Zigkolis, Yiannis Kompatsiaris, and Athena Vakali. Cluster-based landmark and event detection for tagged photo collections. IEEE MultiMedia, 18(1):52–63, January/ March 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Peng:2016:MEF

[PZW^{+}16] Yang Peng, Xiaofeng Zhou, Daisy Zhe Wang, Ishan Patwa, Dihong Gong, and Chunsheng Victor Fang. Multimodal ensemble fusion for disambiguation and retrieval. IEEE MultiMedia, 23(2):42–52, April/June 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
02/mmu2016020042-abs.
html.

Qin:2023:CCB

[QGZF23] Chuan Qin, Shengyan Gao, Xinpeng Zhang, and Guorui Feng. CADW: CGAN-based attack on deep robust image watermarking. IEEE Multi-Media, 30(1):28–35, January/March 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Quackenbush:2005:MS

[QH05] Schuyler Quackenbush and Jurgen Herre. MPEG surround. IEEE MultiMedia, 12(4):18–23, October/ December 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Qin:2018:NUW

[QJC^{+}18] C. Qin, P. Ji, C. Chang, J. Dong, and X. Sun. Nonuniform watermark sharing based on optimal iterative BTC for image tampering recovery. IEEE MultiMedia, 25(3):36–48, July/ September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Qian:2021:NRN

[QPZ^{+}21] Lihui Qian, Tianxiang Pan, Yunfei Zheng, Jiajie Zhang, Mading Li, Bing Yu, and Bin Wang. Noreference nonuniform distorted video quality assessment based on deep multiple instance learning. IEEE MultiMedia, 28(1):28–37, January/March 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Qu:2019:RSM

[QSL^{+}19] J. Qu, P. Sun, X. Li, B. Wang, X. Lu, Z. Tang, and C. Zhang. A retrieval system of medicine molecules based on graph similarity. IEEE MultiMedia, 26(4):17–27, October/ December 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Quackenbush:2013:MUS

[Qua13]  Schuyler  Quackenbush.

MPEG unified speech and audio coding. IEEE Multi-Media, 20(2):72–78, April/ June 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Quek:1996:UGI

[Que96] Francis K. H. Quek. Unencumbered gestural interaction. IEEE MultiMedia, 3(4):36–47, Winter 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://dlib.computer.
                                        org/mu/books/mu1996/pdf/
                                        u4036.pdf;   http://www.
                                        computer.org/multimedia/
                                        mu1996/u4036abs.htm.

Quint:2003:SVG

[Qui03] Antoine Quint. Scalable vector graphics. IEEE MultiMedia, 10(3):99–102, July/September 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/03/u3099.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/
u3099.pdf.

Qing:2014:CAM

[QZ14] Linbo Qing and Wenjun Zeng. Context-adaptive modeling for waveletdomain distributed video coding. IEEE MultiMedia, 21(4):84–93, October/ December 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu2014040084-
abs.html.

Radenkovic:2005:NIS

[Rad05] Milena Radenkovic. Novel infrastructures for supporting mixed-reality experiences. IEEE MultiMedia, 12(2):12–19, April/June 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Richter:2016:JXN

[RAE16] Thomas Richter, Alessandro Artusi, and Touradj Ebrahimi. JPEG XT: A new family of JPEG backward-compatible standards. IEEE MultiMedia, 23(3):80–88, July/ September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        03/mmu2016030080-abs.
                                        html.

Randerson:2008:VVC

[Ran08] Janine Randerson. Visions and views: Circulating images: From Aotearoa to Shanghai. IEEE Multi-Media, 15(1):14–17, January/March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Rane:2014:SBT

[Ran14] Shantanu Rane. Standardization of biometric template protection. IEEE MultiMedia, 21(4):94–99, October/December 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://www.computer.org/
                                        csdl/mags/mu/2014/04/
                                        mmu2014040094-abs.html.

Rath:2000:VVB

[Rat00] Kamlesh Rath. Visions and views: Broadband networks. IEEE MultiMedia, 7(4):10–13, October/ December 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u4010.
                                        pdf.

Rocchesso:2003:SO

[RBF03] Davide Rocchesso, Roberto Bresin, and Mikael Fernström. Sounding objects. IEEE MultiMedia, 10(2): 42–52, April/June 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
02/extras/u2042.htm;
http://csdl.computer.
org/comp/mags/mu/2003/
02/u2042abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/02/u2042.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/02/
u2042.pdf.

RipperKos:2000:CDE

[RBK^{+}00] José Ripper Kós, Adriana Simeone Barbosa, Carlos Krykhtine, Erivelton Muniz da Silva, and Rodrigo Cury Paraizo. The city that doesn’t exist: Multimedia reconstruction of Latin American cities. IEEE MultiMedia, 7(2):56– 60, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2056.
pdf;   http://www.computer.
org/multimedia/mu2000/
u2056abs.htm.

Rooholamini:1995:ABM

[RC95] Reza Rooholamini and Vladimir Cherkassky. ATMbased multimedia servers. IEEE MultiMedia, 2(1): 39–52, Spring 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u10039.
pdf;   http://www.computer.
org/multimedia/mu1995/
u1039abs.htm.

Rowe:2006:MWC

[RC06] Lawrence A. Rowe and Vince Casalaina. Multimedia at work: Capturing conference presentations. IEEE MultiMedia, 13(4):76–84, October/ December 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ram:1999:PAF

[RCG^{+}99] Ashwin Ram, Richard Catrambone, Mark J. Guzdial, Colleen M. Kehoe, D. Scott McCrickard, and John T. Stasko. PML: Adding flexibility to multimedia presentations. IEEE MultiMedia, 6(2):40–52, April/June 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u2040.
pdf;   http://www.computer.
org/multimedia/mu1999/
u2040abs.htm.

Rodriguez:2015:MCF

[RDBRD15] Eva Rodriguez, Jaime Delgado, Laurent Boch, and Victor Rodriguez-Doncel. Media contract formalization using a standardized contract expression language. IEEE MultiMedia, 22(2):64–74, April/ June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/02/mmu2015020064-
abs.html.

Rodriguez-Doncel:2009:MVC

[RDD09] Victor Rodriguez-Doncel and Jaime Delgado. A media value chain ontology for MPEG-21. IEEE MultiMedia, 16(4):44–51, October/ December 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Robles-De-La-Torre:2006:IST

[RDLT06] Gabriel Robles-De-La-Torre.

The importance of the sense of touch in virtual and real environments. IEEE MultiMedia, 13(3):24–30, July/ September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ratcliffe:1999:RAS

[RDP99] Mark Ratcliffe, Tim Paul Davies, and Gillian Price.

Remote advisory services: a NEAT approach. IEEE MultiMedia, 6(1):36–48, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u1036.
pdf;   http://www.computer.
org/multimedia/mu1999/
u1036abs.htm.

Rebelsky:1995:MWD

[Reb95] Samuel A. Rebelsky. Multimedia at work: Designing interactive electronic conference proceedings. IEEE MultiMedia, 2(2):75–79, Summer 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u20075.
pdf.

Reisman:1994:FUA

[Rei94a] Sorel Reisman. Forum users ask for advice. IEEE MultiMedia, 1(3):4–??, Autumn/Fall 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Reisman:1994:NSW

[Rei94b] Sorel Reisman. A newer, simpler way to access the forum. IEEE MultiMedia, 1(2):6–??, Summer 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Reisman:1994:LIE

[Rei94c] Sorel Reisman. On-line information exchange ready. IEEE MultiMedia, 1(1):7– ??, Spring 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Reisman:1996:WMA

[Rei96a] S. Reisman. Web Sight:

Microsoft addresses the Web; Web Sight’s Web site. IEEE MultiMedia, 3 (3):5–6, Fall 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u3005.
pdf.

Reisman:1996:WTD

[Rei96b] Sorel Reisman. Web Sight:

Et tu, déjà vu? IEEE MultiMedia, 3(4):4–5, Winter 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1996/pdf/
u4004.pdf.

Reisman:1996:WKE

[Rei96c] Sorel Reisman. Web Sight:

Keeping an eye on the Web; what is Web Sight Web site? IEEE MultiMedia, 3(1):5, Spring 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u1005.
pdf.

Reisman:1996:WBG

[Rei96d] Sorol Reisman. Web Sight: Browsing — not groping — the Web: Web Sight’s Web site. IEEE MultiMedia, 3(2):4–5, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2004.
pdf.

Reisman:1997:WMH

[Rei97a] Sorel Reisman. Web Sight: Musings from hyperspace. IEEE MultiMedia, 4(1):4– 5, January/March 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1997/pdf/
u1004.pdf.

Reisman:1997:WTS

[Rei97b] Sorel Reisman. Web Sight: Taking stock of the Web. IEEE MultiMedia, 4(3):4, July/September 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3004.
pdf.

Reisman:1997:WTP

[Rei97c] Sorel Reisman. Web Sight:

Talking PCs. IEEE MultiMedia, 4(4):4–5, October/December 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u4004.
pdf.

Reisman:1997:WWB

[Rei97d] Sorel Reisman. Web Sight:

With Web browsers, who needs help? IEEE MultiMedia, 4(2):4, April/ June 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2004.
pdf.

Reisman:1998:WMK

[Rei98a] Sorel Reisman. Web Insight:

E-mail — killer app or social disaster? IEEE MultiMedia, 5(4):4–5, October/December 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u4004.
pdf.

Reisman:1998:WMD

[Rei98b] Sorel Reisman. Web Insight:

Multimedia is dead. IEEE MultiMedia, 5(1):4–5, January/March 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u1004.
pdf.

Reisman:1998:WUW

[Rei98c] Sorel Reisman. Web Insight: Uncluttering the Web maze. IEEE Multi-Media, 5(3):4–5, July/ September 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u3004.
pdf.

Reisman:1998:WWP

[Rei98d] Sorel Reisman. Web Insight: What’s with these PDAs, anyway? IEEE MultiMedia, 5(2):4–5, April/ June 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u2004.
pdf.

Reisman:1999:WGW

[Rei99a] Sorel Reisman. Web Insight: Out of the gallery, onto the Web. IEEE MultiMedia, 6(3):1, July/ September 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u3001.
pdf.

Reisman:1999:WUN

[Rei99b] Sorel Reisman. Web Insight:

Unlikely news flashes (for April First). IEEE MultiMedia, 6(1):4, January/ March 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u1004.
pdf.

Reisman:1999:WUL

[Rei99c] Sorel Reisman. Web Insight:

Utopia lost and soughtin vain. IEEE Multi-Media, 6(4):4, October/ December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u4004.
pdf.

Rodriguez:1995:SSL

[RFM95] Arturo A. Rodriguez, Martin Fisher, and Brian Markey. Standards: Scripting languages emerge in standards bodies. IEEE MultiMedia, 2(4):88–92, Winter 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u40088.

pdf. Discusses the proposed Standard Multimedia/Hypermedia Scripting Language (SMSL), a superset of Java.

Rossetto:2021:UCC

[RGH^{+}21] Luca Rossetto, Ralph Gasser, Silvan Heller, Mahnaz Parian-Scherb, Loris Sauter, Florian Spiess, Heiko Schuldt, Ladislav Pe ka, Tomá Sou ek, Miroslav Kratochvíl, Franti ek Mejzlík, Patrik Veselý, and Jakub Loko . On the user-centric comparative remote evaluation of interactive video search systems. IEEE MultiMedia, 28(4):18–28, October/December 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Raisamo:2006:TUM

[RHP^{+}06] Roope Raisamo, Arto Hippula, Saija Patomäki, Eva Tuominen, Virpi Pasto, and Matias Hasu. Testing usability of multimodal applications with visually impaired children. IEEE MultiMedia, 13(3):70–76, July/ September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Rueger:2014:AIC

[RJ14] Stefan Rueger and Joemon M. Jose. ACM International Conference on Multimedia Retrieval (ICMR 2014). IEEE MultiMedia, 21(3):72, July/ September 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Radenkovic:2006:MIE

[RL06] Milena Radenkovic and Tom Lodge. Media impact: Engaging the public through mass-scale multimedia networks. IEEE MultiMedia, 13(3):12–15, July/September 2006. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ren:2023:EDA

[RLJ^{+}23] Sucheng Ren, Wenxi Liu, Jianbo Jiao, Guoqiang Han, and Shengfeng He. Edge distraction-aware salient object detection. IEEE MultiMedia, 30(3): 63–73, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Rao:2021:FER

[RLW^{+}21] Tianrong Rao, Jie Li, Xiaoyu Wang, Yibo Sun, and Hong Chen. Facial expression recognition with multiscale graph convolutional networks. IEEE MultiMedia, 28(2):11–19, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Roy:2013:SMS

[RLZ13] Suman Deb Roy, Gilad Lotan, and Wenjun Zeng. Social multimedia signals: Sense, process, and put them to work. IEEE MultiMedia, 20(1):7–13, January/March 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Rodriguez:1994:EVC

[RM94] Arturo A. Rodriguez and Ken Morse. Evaluating video codecs. IEEE MultiMedia, 1(3):25–33, Autumn/Fall 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Rockwood:1999:VVT

[RM99] Alyn Rockwood and Janet McAndless. Visions and views: Through the looking glass: the synthesis of computer graphics and computer vision. IEEE MultiMedia, 6(3):8–11, July/September 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u3008.
                                        pdf.

Rui:2007:AMC

[RMPK07] Yong Rui, Ketan Mayer-Patel, and Wolfgang Klas. Advances in multimedia computing: Guest Editors’ introduction: Advances in multimedia computing. IEEE MultiMedia, 14(3):12–13, July/ September 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Refsland:2000:GEI

[ROAS00] Scot Thrane Refsland, Takeo Ojika, Alonzo C. Addison, and Robert Stone. Guest Editors’ introduction: Virtual heritage: Breathing new life into our ancient past. IEEE MultiMedia, 7(2):20–21, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2020.
pdf;   http://www.computer.
org/multimedia/mu2000/
u2020abs.htm.

Refsland:2000:LVK

[ROB00] Scot Thrane Refsland, Takeo Ojika, and Robert Berry, Jr. The living virtual Kinka Kuji Temple: a dynamic environment. IEEE MultiMedia, 7(2):65– 67, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u2065.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2000/
                                        u2065abs.htm.

Ross:2003:AVP

[Ros03] Keith W. Ross. Asynchronous voice: a personal account. IEEE MultiMedia, 10(2):70–74, April/June 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2003/02/u2070.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/02/
                                        u2070.pdf.

Roussel:1999:MWB

[Rou99] Nicolas Roussel. Mediascape: a Web-based media space. IEEE Multi-Media, 6(2):64–74, April/ June 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u2064.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1999/
                                        u2064abs.htm.

Rebhan:1997:MGM

[ROW^{+}97] Richard Rebhan, Sofie Olsson, Björn Wergeland, Pierre Karlsson, and Olle Franceschi. Multimedia goes mobile in broadcast networks. IEEE Multi-Media, 4(2):14–22, April/ June 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2014.
pdf;   http://www.computer.
org/multimedia/mu1997/
u2014abs.htm.

Ross:1997:MWD

[RP97] John Minor Ross and Garnet Price. Multimedia at work: Dante’s Divine Comedy: multimedia in an art gallery. IEEE MultiMedia, 4(2):88–92, April/June 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2088.
pdf.

Ramanathan:1994:APM

[RR94] Srinivas Ramanathan and P. Venkat Rangan. Architectures for personalized multimedia. IEEE Multi-Media, 1(1):37–46, Spring 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Rath:2005:CSF

[RR05] Matthias Rath and Davide Rocchesso. Continuous sonic feedback from a rolling ball. IEEE MultiMedia, 12(2):60–69, April/June 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2005/
                                        02/extras/u2060x1.mpg.

Rodriguez:2001:CHS

[RSR^{+}01] Fernando Gamboa Rodríguez, José Luis Pérez Silva, Felipe Lara Rosano, Francisco Cabiedes Contreras, and Andrea Miranda Vitela. Creating a high school physics video-based laboratory. IEEE Multi-Media, 8(3):78–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u3078.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2001/
                                        u3078abs.htm.

Raimond:2009:IMR

[RSS09] Yves Raimond, Christopher Sutton, and Mark Sandler. Interlinking music-related data on the Web. IEEE MultiMedia, 16(2):52–63, April/June 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ricca:2001:URW

[RT01] Filippo Ricca and Paolo Tonella. Understanding and restructuring Web sites with ReWeb. IEEE MultiMedia, 8(2):40–??, April/June 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u2040.
pdf;   http://www.computer.
org/multimedia/mu2001/
u2040abs.htm.

Ricca:2006:DAF

[RT06] Filippo Ricca and Paolo Tonella. Detecting anomaly and failure in Web applications. IEEE MultiMedia, 13(2):44–51, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Rui:2014:BDI

[Rui14a] Yong Rui. Big data and image search. IEEE MultiMedia, 21(3):2–3, July/ September 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Rui:2014:DNN

[Rui14b] Yong Rui. Deep neural networks: Another tool for multimedia computing. IEEE MultiMedia, 21(4):2–3, October/ December 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu2014040002-
abs.html.

Rui:2014:IMF

[Rui14c] Yong Rui. IEEE Multi-Media forges ahead. IEEE MultiMedia, 21(2):2–3, April/June 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Rui:2015:EBP

[Rui15a] Yong Rui. Establishing best papers for IEEE MultiMedia. IEEE Multi-Media, 22(3):2–3, July/ September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/03/mmu2015030002-
                                        abs.html.

Rui:2015:MGB

[Rui15b]  Yong Rui.  Multime-

dia goes beyond content. IEEE MultiMedia, 22(2): 2, April/June 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/02/mmu2015020002-
                                        abs.html.

Rui:2016:REC

[Rui16a] Yong Rui. Recognizing exceptional contributions. IEEE MultiMedia, 23(3): 2, July/September 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
03/mmu2016030002-abs.
html.

Rui:2016:UM

[Rui16b] Yong Rui. Understanding multimedia. IEEE MultiMedia, 23(2):2–3, April/June 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
02/mmu2016020002-abs.
html.

Rui:2016:WDE

[Rui16c]  Yong Rui.  Working

with the domain experts. IEEE MultiMedia, 23(1): 2–3, January/March 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
01/mmu2016010002-abs.
html.

Rui:2017:BPB

[Rui17a] Yong Rui. Best paper and best department article unveiled. IEEE MultiMedia, 24(3):4, July/ September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
03/mmu2017030004.html.

Rui:2017:CHM

[Rui17b] Yong Rui. Changes on the horizon for the multimedia community. IEEE MultiMedia, 24(4):4–5, October/December 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        04/mmu2017040004.html.

Rui:2017:AIA

[Rui17c] Yong Rui. From artificial intelligence to augmented intelligence. IEEE MultiMedia, 24(1):4–5, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        01/mmu2017010004-abs.
                                        html.

Rich:1994:ALC

[RWS^{+}94] Charles Rich, Richard C. Waters, Yves Schabes, William T. Freeman, Mark C. Torrance, Andrew R. Golding, and Michal Roth. An animated on-line community with artificial agents. IEEE MultiMedia, 1(4):32– 42, Winter 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Saint-Andre:2007:SJJ

[SA07a] Peter Saint-Andre. Standards: Jingle: Jabber does multimedia. IEEE MultiMedia, 14(1):90–94, January/March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Sarper:2007:IVP

[SA07b] Hasan Sarper and Isik Aybay. Improving VoD performance with LAN client back-end buffering. IEEE MultiMedia, 14(1):48–60, January/March 2007. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Singh:2020:LEC

[SAB^{+}20] Vivek K. Singh, Elisabeth André, Susanne Boll, Mireille Hildebrandt, and David A. Shamma. Legal and ethical challenges in multimedia research. IEEE MultiMedia, 27(2):46–54, April/June 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Saddik:2020:MTI

[Sad20] A. E. Saddik. Multimedia and the tactile Internet. IEEE MultiMedia, 27(1): 5–7, January/March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Shin:1999:PRM

[SAH^{+}99] Kang Shin, Tarek Abdelzaher, Seungjae Han, John Reumann, and Emmanuel Abram-Profeta. Project reports: Multimedia-friendly server and communication design. IEEE Multi-Media, 6(2):84–90, April/ June 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u2084.
pdf.

Sakauchi:1994:DVI

[Sak94] Masao Sakauchi. Database vision and image retrieval. IEEE MultiMedia, 1(1):79– 81, Spring 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Sanders:2000:MRT

[San00a] Clate Sanders. Media reviews: Tools and technologies of Internet video: e-Video, Producing Internet Video as Broadband Technologies Converge, H. Peter Alesso (Addison-Wesley, Reading, Mass., 2000, $44.95, 290 pp., includes CD-ROM, ISBN 0-201-70314-9). IEEE MultiMedia, 7(4):82, October/ December 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u4082.
pdf.

Santini:2000:LER

[San00b] Simone Santini. Letter to the editor: Redefining piracy in the Internet Age. IEEE Multi-Media, 7(4):4, October/ December 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u4004.
pdf.

Seckin-Baskomurcu:1997:MRU

[SB97] Gamze Seckin-Baskomurcu.

Media reviews: Understanding distributed multimedia: Communication and Computing for Distributed Multimedia Systems by Guojun Lu (Artech House, 1996, 416 pp., $69.00, ISBN 0-89006-884-4). IEEE MultiMedia, 4(3):92, July/ September 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3091.
pdf.

Stegmaier:2013:UAM

[SBB^{+}13] Florian Stegmaier, Werner Bailer, Tobias Burger, Mari Carmen Suarez-Figueroa, Erik Mannens, Jean-Pierre Evain, Martin Hoffernig, Pierre-Antoine Champin, Mario Doller, and Harald Kosch. Unified access to media metadata on the Web. IEEE MultiMedia, 20(2):22–29, April/June 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Shahabi:2010:GFV

[SBKK^{+}10] Cyrus Shahabi, Farnoush Banaei-Kashani, Ali Khoshgozaran, Luciano Nocera, and Songhua Xing. GeoDec: a framework to visualize and query geospatial data for decisionmaking. IEEE MultiMedia, 17(3):14–23, July/ September 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2000:PRA

[SBL00] Brian K. Smith, Erik Blankinship, and Tamara Lackner. Project reports: Annotation and education. IEEE MultiMedia, 7(2):84– 89, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2084.
pdf.

Soula:1997:VVS

[SBN97] Jean-Pierre Soula, Claude Baron, and Carmel Nestor. Visions and views: a shift from chalk: multimedia and instruction. IEEE MultiMedia, 4(2):5– 9, April/June 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2005.
pdf.

Sawhney:1997:ANV

[SBS97] Nitin Sawhney, David Balcom, and Ian Smith. Authoring and navigating video in space and time. IEEE MultiMedia, 4(4):30–39, October/ December 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u4030.
pdf;   http://www.computer.
org/multimedia/mu1997/
u4030abs.htm.

Smith:1997:VSW

[SC97] John R. Smith and Shih-Fu Chang. Visually searching the Web for content. IEEE MultiMedia, 4(3):12–20, July/September 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u3012.

                                        pdf;   http://www.computer.
                                        org/multimedia/mu1997/
                                        u3012abs.htm.

Shyu:2015:EMR

[SC15] Mei-Ling Shyu and Shu-Ching Chen. Emerging multimedia research and applications. IEEE Multi-Media, 22(4):11–13, October/December 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/04/mmu2015040011.
                                        html.

Shi:2016:LTU

[SC16] Lei Shi and Alexandra I. Cristea. Learners thrive using multifaceted open social learner modeling. IEEE Multi-Media, 23(1):36–47, January/March 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        01/mmu2016010036-abs.
                                        html.

Schank:1994:ALT

[Sch94] Roger C. Schank. Active learning through multimedia. IEEE MultiMedia, 1(1):69–78, Spring 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Schreiner:2000:NBA

[Sch00] Keri Schreiner. In the news: E-books: it’s all in the resolution; wireless woes; digital music evolution; voice portals. IEEE MultiMedia, 7(2):15– 19, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2015.

pdf. See letter [LS00].

Schiphorst:2006:MIA

[Sch06] Thecla Schiphorst. Media impact: Affectionate computing: Can we fall in love with a machine? IEEE MultiMedia, 13(1):20–23, January/March 2006. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Schoeffmann:2014:UCM

[Sch14] Klaus Schoeffmann. A user-centric media retrieval competition: The video browser showdown 2012– 2014. IEEE MultiMedia, 21(4):8–13, October/ December 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu2014040008-
abs.html.

Sartori:2016:CMA

[SCY^{+}16] Andreza Sartori, Dubravko Culibrk, Yan Yan, Remo Job, and Nicu Sebe. Computational modeling of affective qualities of abstract paintings. IEEE Multi-Media, 23(3):44–54, July/ September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
03/mmu2016030044-abs.
html.

Shih:1997:IMP

[SD97] Timothy K. Shih and Ruth E. Davis. IMMPS: a multimedia presentation design system. IEEE MultiMedia, 4(2):67–78, April/June 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2067.
pdf;   http://www.computer.
org/multimedia/mu1997/
u2067abs.htm.

Stork:2007:RRR

[SD07a] David Stork and Marco Duarte. Readers’ remarks: Revisiting computer image analysis and art. IEEE MultiMedia, 14(3):108– 109, July/September 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Stork:2007:AMC

[SD07b] David G. Stork and Marco F.

Duarte. Artful media: Computer vision, image analysis, and master art: Part 3. IEEE MultiMedia, 14(1):14–18, January/ March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Schreiner:2009:MOA

[SDEFK09] Florian Schreiner, Klaus Diepold, Mohamed Abo El-Fotouh, and Taehyun Kim. The MPEG open access application format. IEEE MultiMedia, 16(3):8– 12, July/September 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Shirmohammadi:1998:ABT

[SdOG98] Shervin Shirmohammadi, Jauvane C. de Oliveira, and Nicolas D. Georganas. Applet-based telecollaboration: a network-centric approach. IEEE Multi-Media, 5(2):64–73, April/ June 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u2064.
pdf;   http://www.computer.
org/multimedia/mu1998/
u2064abs.htm.

Serviss:2005:EWH

[Ser05] Ben Serviss. Escaping the world: High and low resolution in gaming. IEEE MultiMedia, 12(4):4–8, October/December 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Schwabe:2001:EWA

[SERL01] Daniel Schwabe, Luiselena Esmeraldo, Gustavo Rossi, and Fernando Lyardet. Engineering Web applications for reuse. IEEE Multi-Media, 8(1):20–31, January/March 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u1020.
pdf;   http://www.computer.
org/multimedia/mu2001/
u1020abs.htm.

Seinstra:2007:HPD

[SGK^{+}07] Frank J. Seinstra, Jan-Mark Geusebroek, Dennis Koelma, Cees G. M. Snoek, Marcel Worring, and Arnold W. M. Smeulders. High-performance distributed video content analysis with Parallel-Horus. IEEE MultiMedia, 14(4):64–75, October/ December 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Steinacker:2001:SMS

[SGS01] Achim Steinacker, Amir Ghavam, and Ralf Steinmetz. Standards: Metadata standards for Webbased resources. IEEE MultiMedia, 8(1):70–76, January/March 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u1070.
pdf.

Schertenleib:2004:CVO

[SGVT04]  Sebastien  Schertenleib,

Mario Gutiérrez, Frédéric Vexo, and Daniel Thalmann. Conducting a virtual orchestra. IEEE MultiMedia, 11(3):40–49, July/ September 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/03/u3040.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/03/
u3040.pdf.

Sun:2008:MWV

[SH08] Qibin Sun and Wolfgang Hürst. Multimedia at work: Video browsing on handheld devices-interface designs for the next generation of mobile video players. IEEE MultiMedia, 15 (3):76–83, July/September 2008. CODEN IEMUE4.

ISSN 1070-986X (print), 1941-0166 (electronic).

Smits:2010:SCS

[SH10] Ewine Smits and Alan Hanjalic. A system concept for socially enriched access to soccer video collections. IEEE MultiMedia, 17(4): 26–35, October/December 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Sharma:1995:PRI

[Sha95] Ravi S. Sharma. Project reports: “Infobahns” in Asia: promises and perils. IEEE MultiMedia, 2(2): 80–85, Summer 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u20080.
                                        pdf.

Shao:2004:MQS

[Sha04a] Huai-Rong Shao. Multimedia QoS support specified in various wireless network standards. IEEE Multi-Media, 11(4):75–77, October/December 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2004/04/u4075.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/04/
                                        u4075.pdf.

Sharda:2004:RR

[Sha04b] Nalin Sharda. Readers’ remarks. IEEE MultiMedia, 11(2):C3, April/June 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
02/u20c3.pdf;   http:/
/csdl.computer.org/dl/
mags/mu/2004/02/u20c3.
htm.

Sheth:1999:MRS

[She99] Amit Sheth. Media reviews: On the shelf. IEEE MultiMedia, 6(1):93, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u1090.
pdf.

Stenton:2007:MWM

[SHG^{+}07] Stuart P. Stenton, Richard Hull, Patrick M. Goddi, Josephine E. Reid, Ben J. C. Clayton, Tom J. Melamed, and Susie Wee. Multimedia at work: Mediascapes: Context-aware multimedia experiences. IEEE MultiMedia, 14(3): 98–105, July/September 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Shih:2009:IAR

[SHH09] Huang-Chia Shih, Chung-Lin Huang, and Jenq-Neng Hwang. An interactive attention-ranking system for video search. IEEE MultiMedia, 16(4):70–81, October/December 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Shneiderman:2002:MHN

[Shn02] Ben Shneiderman. Meeting human needs with new digital imaging technologies. IEEE Multi-Media, 9(4):8–14, October/December 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
04/u4008abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/04/u4008.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/04/
u4008.pdf.

Shu:1999:NPa

[Shu99a] Chiao-Fe Shu. New products. IEEE MultiMedia, 6(2):94–96, April/ June 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u2094.
pdf.

Shu:1999:NPb

[Shu99b] Chiao-Fe Shu. New products. IEEE MultiMedia, 6(3):95–96, July/ September 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u3095.
pdf.

Shu:1999:NPc

[Shu99c] Chiao-Fe Shu. New products. IEEE MultiMedia, 6(4):94–95, October/ December 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u4094.
pdf.

Shu:2000:NPa

[Shu00a] Chiao-Fe Shu. New products. IEEE MultiMedia, 7(1):88, January/ March 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u1088.
pdf.

Shu:2000:NPb

[Shu00b] Chiao-Fe Shu. New products. IEEE MultiMedia, 7(2):103–104, April/ June 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/

                                        books/mu2000/pdf/u2103.
                                        pdf.

Shu:2000:NPc

[Shu00c] Chiao-Fe Shu. New products. IEEE MultiMedia, 7(3):93–95, July/ September 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u3093.
                                        pdf.

Shu:2000:NPd

[Shu00d] Chiao-Fe Shu. New products. IEEE MultiMedia, 7(4):94–96, October/ December 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u4094.
                                        pdf.

Shu:2001:NP

[Shu01] Chiao-Fe Shu. New products. IEEE MultiMedia, 8(1):78–79, January/ March 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u1078.
                                        pdf.

Singh:1999:GEI

[Sin99] Gurminder Singh. Guest Editor’s introduction: Media spaces. IEEE Multi-Media, 6(2):18–19, April/ June 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u2018.
pdf.

Singh:2004:GEI

[Sin04] Gurminder Singh. Guest Editor’s introduction: Content repurposing. IEEE MultiMedia, 11(1):20–21, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
01/u1020.pdf;   http:/
/csdl.computer.org/comp/
mags/mu/2004/01/u1020abs.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/01/
u1020.htm.

Santini:2000:IBQ

[SJ00] Simone Santini and Ramesh Jain. Integrated browsing and querying for image databases. IEEE MultiMedia, 7(3):26–39, July/ September 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u3026.
pdf;   http://www.computer.
org/multimedia/mu2000/
u3026abs.htm.

Smith:2004:PCV

[SJ04] John R. Smith and Marwan A. Jabri. The 3G-324M protocol for conversational video telephony. IEEE MultiMedia, 11(3):102–105, July/ September 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2004/03/u3102.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/03/
                                        u3102.pdf.

Stork:2006:AMCb

[SJ06] David G. Stork and Micah K. Johnson. Artful media: Computer vision, image analysis, and master art: Part 2. IEEE MultiMedia, 13(4):12–17, October/ December 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Scherp:2009:ES

[SJ09] Ansgar Scherp and Ramesh Jain. An ecosystem for semantics. IEEE Multi-Media, 16(2):18–25, April/ June 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Senior:2006:ABP

[SJM06] Andrew Senior, Alejandro Jaimes, and Wolfgang Muench. Art beat: Presence/absence: The 2005 ACM Multimedia Interactive Art Exhibition. IEEE MultiMedia, 13(1):9–11, January/March 2006. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Shah:2007:AVS

[SJS07] Mubarak Shah, Omar Javed, and Khurram Shafique. Automated visual surveillance in realistic scenarios. IEEE MultiMedia, 14 (1):30–39, January/March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Schar:2000:UNL

[SK00] Sissel Guttormsen Schär and Helmut Krueger. Using new learning technologies with multimedia. IEEE MultiMedia, 7(3):40–51, July/ September 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u3040.
pdf;   http://www.computer.
org/multimedia/mu2000/
u3040abs.htm.

Sgouros:2003:AEE

[SK03] Nikitas M. Sgouros and Sophia Kousidou. Authoring and execution environments for robotic performance. IEEE Multi-Media, 10(3):38–47, July/ September 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
03/extras/u3038x1.wmv;
 http://csdl.computer.
org/comp/mags/mu/2003/
03/u3038abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/03/u3038.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/
u3038.pdf.

Sabirin:2012:DAF

[SKK^{+}12] Houari Sabirin, Munchurl Kim, Hui-Yong Kim, Han Kyu Lee, Minkyu Park, and Yong Han Kim. DMB application format for mobile multimedia services. IEEE MultiMedia, 19(2):38–47, May/June 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Schmidt:2008:VVM

[SKSH08] Albrecht Schmidt, Dagmar Kern, Sara Streng, and Paul Holleis. Visions and views: Magic beyond the screen. IEEE MultiMedia, 15(4):8–13, October/ December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Singh:2022:IMM

[SKWB22] Amit Kumar Singh, Deepa Kundur, Min Wu, and Mauro Barni. Integrity of multimedia and multimodal data: From capture to use. IEEE MultiMedia, 29(2):8–10, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Slaney:2011:PRW

[Sla11a] Malcolm Slaney. Precisionrecall is wrong for multimedia. IEEE MultiMedia, 18(3):4–7, July/September 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Slaney:2011:WSM

[Sla11b] Malcolm Slaney. Web-scale multimedia analysis: Does content matter? IEEE MultiMedia, 18(2):12–15, April/June 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Sul:1998:VSL

[SLW98] Chang Whan Sul, Kee Chang Lee, and Kwangyun Wohn. Virtual stage: a locationbased karaoke system. IEEE MultiMedia, 5(2):42– 52, April/June 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/

books/mu1998/pdf/u2042.
pdf;   http://www.computer.
org/multimedia/mu1998/
u2042abs.htm.

Sun:2022:DEF

[SLZ^{+}22] Xiao Sun, Junjie Lang, Shixing Zheng, Yating Huang, and Meng Wang. Dual expression fusion: a universal microexpression recognition framework. IEEE MultiMedia, 29(4):48–58, October/ December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Schaffert:2015:ISR

[SM15] Nina Schaffert and Klaus Mattes. Interactive sonification in rowing: Acoustic feedback for on-water training. IEEE Multi-Media, 22(1):58–67, January/March 2015. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/01/mmu2015010058-
abs.html.

Song:2018:WMM

[SM18] S. Song and T. Mei. When multimedia meets fashion. IEEE MultiMedia, 25(3): 102–108, July/September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

SilvadeAbreu:2021:TCD

[SMdS^{+}21] Raphael Silva de Abreu, Douglas Mattos, Joel dos Santos, Gheorghita Ghinea, and Débora Christina Muchaluat-Saade. Toward content-driven intelligent authoring of mulsemedia applications. IEEE MultiMedia, 28(1):7–16, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Stini:2006:VVD

[SMF06] Michael Stini, Martin Mauve, and Frank H. P. Fitzek. Visions and views: Digital ownership: From content consumers to owners and traders. IEEE MultiMedia, 13(4):1, 4–6, October/December 2006. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Sun:2017:BMN

[SMH^{+}17] Lifeng Sun, Ming Ma, Wen Hu, Haitian Pang, and Zhi Wang. Beyond 1 million nodes: A crowdsourced video content delivery network. IEEE MultiMedia, 24(3):54–63, July/ September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
03/mmu2017030054.html.

Smith:2005:GEI

[Smi05] John R. Smith. Guest Editor’s introduction: What’s new with MPEG? IEEE MultiMedia, 12(4):16–17, October/December 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2005/
                                        04/u4016.pdf.

Smith:2010:CT

[Smi10a] John R. Smith. Clicking on things. IEEE Multi-Media, 17(4):2–3, October/ December 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2010:NG

[Smi10b]  John R. Smith.  The

not so newlywed game. IEEE MultiMedia, 17(3): 2–3, July/September 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2010:RB

[Smi10c] John R. Smith. The red balloon. IEEE MultiMedia, 17(2):2–3, April/June 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2011:EY

[Smi11a] John R. Smith. Express yourself. IEEE MultiMedia, 18(1):2, 3, January/ March 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2011:GC

[Smi11b] John R. Smith. A great challenge. IEEE MultiMedia, 18(2):2–3, April/June 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2011:HME

[Smi11c]  John R. Smith.  His-

tory made every day. IEEE MultiMedia, 18(3): 2–3, July/September 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2011:OCN

[Smi11d] John R. Smith. Open channel for news. IEEE Multi-Media, 18(4):2–3, October/ December 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2012:MG

[Smi12a] John R. Smith. Minding the gap. IEEE MultiMedia, 19(1):2–3, January/March 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2012:PG

[Smi12b] John R. Smith. Problems gone by? IEEE Multi-Media, 19(4):2–4, October/ December 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2012:RMP

[Smi12c] John R. Smith. Rich media, poor media. IEEE MultiMedia, 19(3):2–3, July/September 2012. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2012:VO

[Smi12d] John R. Smith. A virtual opinion. IEEE MultiMedia, 19(2):2–3, May/June 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2013:HC

[Smi13a] John R. Smith. Herding cats. IEEE Multi-Media, 20(4):2–3, October/ December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2013:JF

[Smi13b]  John R. Smith.  Just

the facets. IEEE Multi-Media, 20(1):92, January/ March 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2013:LL

[Smi13c] John R. Smith. Lessons in learning. IEEE MultiMedia, 20(3):2–3, July/ September 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2013:RE

[Smi13d] John R. Smith. Reconstructing events. IEEE MultiMedia, 20(2):2–3, April/June 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2014:HMV

[Smi14]  John R. Smith.  How

many visual concepts? IEEE MultiMedia, 21(1): 2–3, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Strintzis:2010:MSF

[SMK^{+}10] Michael Gerassimos Strintzis, Athanasios Mademlis, Konstantinos Kostopoulos, Konstantinos Moustakas, and Dimitrios Tzovaras. A map-search framework based on attributed graph matching. IEEE MultiMedia, 17 (3):24–33, July/September 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smoliar:1994:RMM

[Smo94] Stephen W. Smoliar. The role of music in multimedia. IEEE MultiMedia, 1(1):9–11, Spring 1994.

CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Satoh:1999:NIN

[SNK99] Shin’ichi Satoh, Yuichi Nakamura, and Takeo Kanade. Name-It: Naming and detecting faces in news videos. IEEE MultiMedia, 6(1):22–35, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u1022.
pdf;   http://www.computer.
org/multimedia/mu1999/
u1022abs.htm.

Stone:2000:VHW

[SO00] Robert Stone and Takeo Ojika. Virtual heritage: What next? IEEE MultiMedia, 7(2):73–74, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2073.
pdf;   http://www.computer.
org/multimedia/mu2000/
u2073abs.htm.

Sodagar:2011:MDS

[Sod11] Iraj Sodagar. The MPEG– DASH standard for multimedia streaming over the Internet. IEEE MultiMedia, 18(4):62–67, October/ December 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Schwarz:2013:DSS

[SOS13] Sebastian Schwarz, Roger Olsson, and Marten Sjostrom. Depth sensing for 3DTV: A survey. IEEE MultiMedia, 20(4):10–17, October/ December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Subrahmanyam:2008:ISB

[SRA08] G. R. K. Sai Subrahmanyam, A. N. Rajagopalan, and Rangarajan Aravind. Importance sampling-based unscented Kalman filter for filmgrain noise removal. IEEE MultiMedia, 15(2):52–63, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Seligmann:2004:PIV

[SREH04] Doree Duncan Seligmann, Nicolas Roussel, Helen Evans, and Heiko Hansen. Proximity as an interface for video communication. IEEE MultiMedia, 11(3):12–16, July/ September 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/03/u3012.
htm;   http://csdl.computer.

                                        org/dl/mags/mu/2004/03/
                                        u3012.pdf.

Sukel:2024:MTF

[SRW24] Maarten Sukel, Stevan Rudinac, and Marcel Worring. Multimodal temporal fusion transformers are good product demand forecasters. IEEE MultiMedia, 31(2):48–60, April/June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Spohrer:2000:VVU

[SS00] Jim Spohrer and Mitch Stein. Visions and views: User experience in the pervasive computing age. IEEE Multi-Media, 7(1):12–17, January/March 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u1012.
                                        pdf.

Sarris:2001:CVH

[SS01] Nikolaos Sarris and Michael G. Strintzis. Constructing a videophone for the hearing impaired using MPEG-4 tools. IEEE Multi-Media, 8(3):56–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u3056.

pdf;   http://www.computer.
org/multimedia/mu2001/
u3056abs.htm.

Smith:2006:SMS

[SS06] John R. Smith and Peter Schirling. Standards: Metadata standards roundup. IEEE MultiMedia, 13(2):84–88, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Smith:2008:SSI

[SS08] John R. Smith and John R.

Smith. Standards: The search for interoperability. IEEE MultiMedia, 15 (3):84–87, July/September 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Snoek:2011:AMI

[SS11] Cees G. M. Snoek and Malcolm Slaney. Academia meets industry at the Multimedia Grand Challenge. IEEE MultiMedia, 18(1): 4–7, January/March 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Shamma:2012:DCH

[SS12] David A. Shamma and Malcolm Slaney. Don’t click here. IEEE MultiMedia, 19(3):4–6, July/ September 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Srivastava:2019:MGV

[SS19] G. Srivastava and R. Srivastava. Modification of gradient vector flow using directional contrast for salient object detection. IEEE MultiMedia, 26(4): 7–16, October/December 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Sukhmani:2019:ECC

[SSEKE19] S. Sukhmani, M. Sadeghi, M. Erol-Kantarci, and A. El Saddik. Edge caching and computing in 5G for mobile AR/VR and tactile Internet. IEEE MultiMedia, 26(1):21–30, January/March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Saleme:2019:CCD

[SSG19] E. B. Saleme, C. A. S. Santos, and G. Ghinea. Coping with the challenges of delivering multiple sensorial media. IEEE MultiMedia, 26(2):66–75, April/June 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Sun:2014:RTG

[SSLS14] Li Sun, Mingli Song, Zicheng Liu, and Ming-Ting Sun. Real-time gaze estimation with online calibration. IEEE MultiMedia, 21(4):28–37, October/ December 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu2014040028-
abs.html.

Shepherd:1996:PRQ

[SSR96] Doug Shepherd, Andrew Scott, and Tom Rodden. Project reports: Quality-of-service support for multimedia applications. IEEE MultiMedia, 3(3):78–82, Fall 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u3078.
pdf.

Svanera:2019:WFD

[SSS^{+}19] M. Svanera, M. Savardi, A. Signoroni, A. B. Kovács, and S. Benini. Who is the film’s director? Authorship recognition based on shot features. IEEE MultiMedia, 26(4):43–54, October/ December 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Subrahmanyam:1995:QAS

[SSSS95] P. A. Subrahmanyam, K. J.

Singh, Guy A. Story, and William Schell. Quality assurance in scripting.

IEEE MultiMedia, 2(2):50– 59, Summer 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u20050.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1995/
                                        u2050abs.htm.

Sun:2020:TSE

[SSW20] Xiao Sun, Yezhen Song, and Meng Wang. Toward sensing emotions with deep visual analysis: a longterm psychological modeling approach. IEEE Multi-Media, 27(4):18–27, October/December 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Shaw:2016:EMN

[ST16] Felix Shaw and Barry-John Theobald. Expressive modulation of neutral visual speech. IEEE MultiMedia, 23(4):68–78, October/December 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        04/mmu2016040068-abs.
                                        html.

Steinmetz:1995:AMO

[Ste95] Ralf Steinmetz. Analyzing the Multimedia Operating System. IEEE MultiMedia, 2(1):68–84, Spring 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u10068.
pdf;   http://www.computer.
org/multimedia/mu1995/
u1068abs.htm.

Steinmetz:2001:MIM

[Ste01]  Arnd Steinmetz.  Me-

dia impact: Media and distance: a learning experience. IEEE Multi-Media, 8(2):8–??, April/ June 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u2008.
pdf.

Shen:2008:DVT

[STH08] Bo Shen, Wai-Tian Tan, and Frederic Huve. Dynamic video transcoding in mobile environments. IEEE MultiMedia, 15(1):42–51, January/ March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Stiles:1995:VVJ

[Sti95] Chris Stiles. Visions and views: Japanese multimedia industry update. IEEE MultiMedia, 2(4):8– 11, Winter 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u40008.
                                        pdf.

Shetty:2018:IVC

[STL18] R. Shetty, H. R. Tavakoli, and J. Laaksonen. Image and video captioning with augmented neural architectures. IEEE MultiMedia, 25(2):34–46, April/June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Stork:2006:AMCa

[Sto06] David G. Stork. Artful media: Computer vision, image analysis, and master art: Part 1. IEEE Multi-Media, 13(3):16–20, July/ September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Stokman:2014:FSP

[Sto14] Harro Stokman. The future of smart photography. IEEE MultiMedia, 21 (3):66–71, July/September 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Strnadl:1996:MRS

[Str96] Christoph F. Strnadl. Media reviews: a successful guide to networked multimedia: Understanding Networked Multimedia:

Applications and Technology, by François Fluckiger (Prentice-Hall, 1995, 620 pp, ISBN 0-13-190992-4). IEEE MultiMedia, 3(2):84– 85, Summer 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2083.
pdf.

Strnadl:1999:MRE

[Str99a] Christoph F. Strnadl. Media reviews: Evaluating communication networks: Performance Evaluation of Communication Networks by Gary N. Higginbottom (Artech-House, Norwood, Mass., 1998, $83.00, 382 pp, ISBN 0-89006-870-4). IEEE MultiMedia, 6(2):91– 92, April/June 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u2091.
pdf.

Strnadl:1999:MRM

[Str99b] Christoph F. Strnadl. Media reviews: Mobile multimedia communications: Mobile Multimedia Communications by D. J. Goodman and D. Raychaudhuri (Plenum Publishing, New York, 1997, $95.00, 312 pp, ISBN 0-306- 45772-5). IEEE MultiMedia, 6(1):90–92, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1999/pdf/u1090.
                                        pdf.

Strnadl:2000:MRI

[Str00] Christoph F. Strnadl. Media reviews: Internetworked multimedia: Internetworking Multimedia by Jon Crowcroft, Mark Handley, and Ian Wakeman (Morgan Kaufmann Publishers, San Francisco, Calif., 1999, $54.95, 290 pp., ISBN 1-55860-584-3). IEEE MultiMedia, 7(2):101–102, April/ June 2000. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u2101.
                                        pdf.

Stuttgen:1995:NEM

[Stü95] Heinrich J. Stüttgen. Network evolution and multimedia communication. IEEE MultiMedia, 2(3): 42–59, Fall 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u30042.
                                        pdf;   http://www.computer.

org/multimedia/mu1995/
u3042abs.htm.

Subramanya:2006:RRM

[Sub06] S. E. Subramanya. Readers’ remarks: Mobile content production–major issues. IEEE MultiMedia, 13 (4):c3, October/December 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Sun:2005:UMS

[Sun05] Qibin Sun. Understanding multimedia security. IEEE MultiMedia, 12(4):88–89, October/December 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2005/
04/u4088.pdf.

Sussman:1999:SWP

[Sus99] David Sussman. Standards: WebDAV: a panacea for collaborative authoring? IEEE MultiMedia, 6(2):76–79, April/ June 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u2076.
pdf.

Schulze:2011:MGM

[SvdM11] Walter Schulze and Brink van der Merwe. Music generation with Markov models. IEEE MultiMedia, 18 (3):78–85, July/September 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Stamou:2006:SMA

[SvOPS06] Giorgos Stamou, Jacco van Ossenbruggen, Jeff Z. Pan, and Guus Schreiber. Standards: Multimedia annotations on the Semantic Web. IEEE MultiMedia, 13 (1):86–90, January/March 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Shu:2004:RRC

[SW04] Wei Shu and Min-You Wu. Resource requirements of closed-loop video delivery services. IEEE MultiMedia, 11(2):24–37, April/June 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
02/u2024abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2004/02/u2024.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/02/
u2024.pdf.

Sat:2009:AVQ

[SW09] Batu Sat and Benjamin W. Wah. Analyzing voice quality in popular VoIP applications. IEEE Multi-Media, 16(1):46–59, January/March 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Snoek:2008:SVR

[SWdR^{+}08] Cees G. M. Snoek, Marcel Worring, Ork de Rooij, Koen E. A. van de Sande, Rong Yan, and Alexander G. Hauptmann. Standards: VideOlympics: Real-time evaluation of multimedia retrieval systems. IEEE MultiMedia, 15 (1):86–91, January/March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Switzer:1996:MRI

[Swi96] Mary Kay Switzer. Media reviews: An informed citizens’ guide to cyberspace: Civilizing Cyberspace: Policy, Power, and the Information Highway by Steven E. Miller (ACM Press/Addison Wesley Publishing Company, 1996, 413 pages, $26.85, ISBN 0-201-84760-4). IEEE MultiMedia, 3(3):93, Fall 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1996/pdf/
u3090.pdf.

Shavitt:2011:MML

[SWW11] Yuval Shavitt, Ela Weinsberg, and Udi Weinsberg. Mining music from largescale, peer-to-peer networks. IEEE MultiMedia, 18(1):14–23, January/ March 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Shan:2016:FIM

[SXHH16] Zhenyu Shan, Yingjie Xia, Peipei Hou, and Jifeng He. Fusing incomplete multisensor heterogeneous data to estimate urban traffic. IEEE MultiMedia, 23(3):56–63, July/ September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
03/mmu2016030056-abs.
html.

Subramanya:2005:MCC

[SY05a] S. R. Subramanya and Byung K. Yi. Mobile content customization. IEEE MultiMedia, 12(4):104, 103, October/ December 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Subramanya:2005:UCM

[SY05b] S. R. Subramanya and Byung K. Yi. Usercontrolled, multimediaenhanced communication using prior knowledge and experience. IEEE Multi-Media, 12(2):90–95, April/ June 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Syed:2001:GEI

[Sye01] Mahbubur Rahman Syed.

Guest Editor’s introduction: Diminishing the distance in distance education. IEEE Multi-Media, 8(3):18–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u3018.
pdf;   pdf/u3018.pdf.

Shin:2006:MST

[SYJC06] Kwang-Sik Shin, Wan-Oh Yoon, Jin-Ha Jung, and Sang-Bang Choi. MPEG-4 stream transmission and synchronization schemes for parallel servers. IEEE MultiMedia, 13(1):24–36, January/March 2006. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Sabirin:2018:TRT

[SYN^{+}18] H. Sabirin, Q. Yao, K. Nonaka, H. Sankoh, and S. Naito. Toward realtime delivery of immersive sports content. IEEE MultiMedia, 25(2):61–70, April/June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Smoliar:1994:CBV

[SZ94] Stephen W. Smoliar and HongJiang Zhang. Contentbased video indexing and retrieval. IEEE MultiMedia, 1(2):62–72, Summer 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Srihari:2000:SSA

[SZ00] Rohini K. Srihari and Zhongfei Zhang. Show&Tell: a semi-automated image annotation system. IEEE MultiMedia, 7(3):61–71, July/September 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2000/pdf/u3061.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2000/
                                        u3061abs.htm.

Shirmohammadi:2017:LMR

[SZ17] Shervin Shirmohammadi and Chengcui Zhang. The latest multimedia research from ISM 2016. IEEE MultiMedia, 24(2):18–19, April/June 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        02/mmu2017020018.html.

Shen:2008:HET

[SZH^{+}08] Xiaojun Shen, Jilin Zhou, Abdelwahab Hamam, Saeid Nourian, Naim R. El-Far, François Malric, and Nicolas D. Georganas. Haptic-enabled telementoring surgery simulation. IEEE MultiMedia, 15 (1):64–76, January/March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Surman:2020:GFDa

[SZS^{+}20a] P. Surman, X. Zhang, W. Song, X. Xia, S. Wang, and Y. Zheng. Glassesfree 3-D and augmented reality display advances: From theory to implementation. IEEE MultiMedia, 27(1):17–26, January/ March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Surman:2020:GFDb

[SZS^{+}20b] Phil Surman, Xiangyu Zhang, Weitao Song, Xinxing Xia, Shizheng Wang, and Yuanjin Zheng. Glassesfree 3-D and augmented reality display advances: From theory to implementation. IEEE MultiMedia, 27(1):17–26, January/ March 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Su:2019:PBM

[SZW^{+}19] G. Su, C. Zhang, H. Wang, M. Chen, M. Kankanhalli, and R. Lienhart. Pushing the boundary of multimedia big data: An overview of IEEE MIPR. IEEE MultiMedia, 26(2): 87–91, April/June 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Tabor:1997:PRT

[Tab97]  John Tabor.  Project

reports: Talking heads. IEEE MultiMedia, 4(3):81– 83, July/September 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://dlib.computer.
                                        org/mu/books/mu1997/pdf/
                                        u3081.pdf.

Taima:2002:CWE

[Tai02] Kenji Taima. Can we ever charge Napster users? IEEE MultiMedia, 9(4): 76–81, October/December 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/dl/mags/mu/2002/04/
                                        u4076.htm;   http://csdl.
                                        computer.org/dl/mags/
                                        mu/2002/04/u4076.pdf.

Tate:1996:VVE

[Tat96] R. Grant Tate. Visions and views: European multimedia update; Web resources.

IEEE MultiMedia, 3(2):12– 17, Summer 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2012.
pdf.

Tonomura:1994:SVC

[TATS94] Yoshinobu Tonomura, Akihito Akutsu, Yukinobu Tangiguchi, and Gen Suzuki. Structured video computing. IEEE MultiMedia, 1(3):34–43, Autumn/Fall 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Taylor:1999:SDV

[Tay99] Jim Taylor. Standards:

DVD-video: multimedia for the masses. IEEE MultiMedia, 6(3):86–92, July/ September 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u3086.
pdf.

Tokuda:2001:OTS

[TC01] Naoyuki Tokuda and Liang Chen. An online tutoring system for language translation. IEEE Multi-Media, 8(3):46–??, July/ September 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u3046.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2001/
                                        u3046abs.htm.

Tan:2018:DVR

[TCM18] J. Tan, G. Cheung, and R. Ma. 360-degree virtualreality cameras for the masses. IEEE MultiMedia, 25(1):87–94, January/ March 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Tjondronegoro:2004:IHM

[TCP04] Dian Tjondronegoro, Yi-Ping Phoebe Chen, and Binh Pham. Integrating highlights for more complete sports video summarization. IEEE Multi-Media, 11(4):22–37, October/December 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2004/04/u4022.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/04/
                                        u4022.pdf.

Timmerer:2010:AMB

[TCPD10] Christian Timmerer, Filippo Chiariglione, Marius Preda, and Victor Rodriguez Doncel. Accelerating the media business with MPEG extensible middleware. IEEE MultiMedia, 17 (3):74–78, July/September 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Tian:2015:MBD

[TCS^{+}15] Yonghong Tian, Shu-Ching Chen, Mei-Ling Shyu, Tiejun Huang, Phillip Sheu, and Alberto Del Bimbo. Multimedia big data. IEEE MultiMedia, 22(3):93–95, July/ September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/03/mmu2015030093.
html.

Tian:2016:UME

[TCS16] Yonghong Tian, Min Chen, and Leonel Sousa. Ubiquitous multimedia: Emerging research on multimedia computing. IEEE MultiMedia, 23(2):12–15, April/June 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
02/mmu2016020012.html.

Tous:2012:AOB

[TDZ^{+}12] Ruben Tous, Jaime Delgado, Thomas Zinkl, Pere Toran, Gabriela Alcalde, Martin Goetz, and Olga Ferrer-Roca. The anatomy of an optical biopsy semantic retrieval system. IEEE MultiMedia, 19(2):16–27, May/June 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Teeter:1998:RMI

[Tee98a] B. R. Teeter. Reviewing Microsoft Internet information [book review]. IEEE MultiMedia, 5(3):69, July/ September 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Teeter:1998:MRI

[Tee98b] Brian R. Teeter. Media reviews: Internet information server basics: Microsoft Internet Information Server 4.0 Sourcebook by Michelle Petrovsky (John Wiley & Sons, 1998, $39.95, ISBN 0-471-17805-5). IEEE MultiMedia, 5(3):69, July/ September 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u3069.
                                        pdf.

Teeter:1999:MRC

[Tee99] Brian R. Teeter. Media reviews: a cookbook for JavaScript: The JavaScript CD Cookbook 2/e by Erica Sadun and J. Brook Monroe (Charles River Media, 1997, $44.95, ISBN 1-886801-65-7). IEEE MultiMedia, 6(1):90, January/ March 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Tesic:2005:MPC

[Tes05] Jelena Tesic. Metadata practices for consumer photos. IEEE MultiMedia, 12 (3):86–92, July/September 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Tung:2022:WAE

[TGB^{+}22] Caleb Tung, Abhinav Goel, Fischer Bordwell, Nick Eliopoulos, Xiao Hu, Yung-Hsiang Lu, and George K. Thiruvathukal. Why accuracy is not enough: The need for consistency in object detection. IEEE MultiMedia, 29(3):8–16, July/ September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Tsagarakis:2006:HEM

[TGC^{+}06] Nikolaos G. Tsagarakis, John O. Gray, Darwin G. Caldwell, Cinzia Zannoni, Marco Petrone, Debora Testi, and Marco Viceconti. Haptic-enabled multimodal interface for the planning of hip arthroplasty. IEEE MultiMedia, 13(3):40–48, July/ September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Tirkel:2001:UWE

[TH01] Andrew Z. Tirkel and Thomas E. Hall. A unique watermark for every image. IEEE Multi-Media, 8(4):30–37, October/December 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2001/pdf/u4030.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2001/
                                        u4030abs.htm.

Timmerer:2005:IAM

[TH05] Christian Timmerer and Hermann Hellwagner. Interoperable adaptive multimedia communication. IEEE MultiMedia, 12 (1):74–79, January/March 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Tian:2013:VCD

[THJG13] Yonghong Tian, Tiejun Huang, Menglin Jiang, and Wen Gao. Video copydetection and localization with a scalable cascading framework. IEEE Multi-Media, 20(3):72–86, July/ September 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Thompson:2000:SIC

[Tho00] David Thompson. Standards: IEEE 1394: Changing the way we do multimedia communications; integrating layers; resources. IEEE MultiMedia, 7(2):94– 100, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2094.
pdf.

Tian:2006:MRF

[Tia06] Qi Tian. Media reviews:

a foundational perspective for visual information retrieval. IEEE MultiMedia, 13(2):90–92, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Takeda:1996:UIA

[TIS96] Koji Takeda, Mitsuyuki Inaba, and Kazuo Sugihara. User interface and agent prototyping for flexible working. IEEE MultiMedia, 3(2):40–50, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u2040.
pdf;   http://www.computer.
org/multimedia/mu1996/
u2040abs.htm.

Titsworth:2002:WSM

[Tit02] Tammi Titsworth. Will streaming media yield a profit? IEEE Multi-Media, 9(4):15–18, October/December 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
04/u4015abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/04/u4015.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/04/
u4015.pdf.

Titsworth:2003:RMV

[Tit03a] Tammi Titsworth. A room with multiple views— multimedia in hotels. IEEE MultiMedia, 10(3):1, 9, July/September 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/03/u3001.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/
u3001.pdf.

Titsworth:2003:SBC

[Tit03b] Tammi Titsworth. Silver bullets and cybersecurity. IEEE MultiMedia, 10(1):64–65, January/ March 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/

mags/mu/2003/01/u1064.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/01/
u1064.pdf.

Titsworth:2006:TBL

[Tit06] Tammi Titsworth. Tech beat: Lost on your cell phone: Innovations from NAB 2006. IEEE Multi-Media, 13(3):1, 4–5, July/ September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Titsworth:2007:TBM

[Tit07] Tammi Titsworth. Tech beat: Mobile opportunities: Experiencing TV outside your living room. IEEE MultiMedia, 14(3): 6–7, July/September 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Tajadura-Jimenez:2015:SST

[TJBBFB15] Ana Tajadura-Jimenez, Nadia Bianchi-Berthouze, Enrico Furfaro, and Frederic Bevilacqua. Sonification of surface tapping changes behavior, surface perception, and emotion. IEEE MultiMedia, 22(1):48–57, January/ March 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/

mags/mu/2015/01/mmu2015010048-
abs.html.

Thomas-Kerr:2007:SFY

[TKBR^{+}07] Joseph A. I. Thomas-Kerr, Ian S. Burnett, Christian H. Ritz, Davy De Schrijver, Rik Van de Walle, and Sylvain Devillers. Standards: Is that a fish in your ear? A universal metalanguage for multimedia. IEEE MultiMedia, 14(2): 72–77, April/June 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Thoma:1997:CTV

[TL97] George R. Thoma and L. Rodney Long. Compressing and transmitting visible human images. IEEE MultiMedia, 4(2):36– 45, April/June 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2036.
pdf;   http://www.computer.
org/multimedia/mu1997/
u2036abs.htm.

Tao:2019:QOM

[TLJ^{+}19] X. Tao, Y. Liu, C. Jiang, Z. Wang, and X. Qin. QoE-oriented multimedia assessment: A facial expression recognition approach. IEEE MultiMedia, 26(1):41–50, January/ March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Tseng:2004:UMM

[TLS04] Belle L. Tseng, Ching-Yung Lin, and John R. Smith. Using MPEG-7 and MPEG-21 for personalizing video. IEEE Multi-Media, 11(1):42–53, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
01/u1042abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2004/01/u1042.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/01/
u1042.pdf.

Tu:2014:JVT

[TML^{+}14] Kewei Tu, Meng Meng, Mun Wai Lee, Tae Eun Choe, and Song-Chun Zhu. Joint video and text parsing for understanding events and answering queries. IEEE MultiMedia, 21(2):42–70, April/June 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Tadamura:1998:SCG

[TN98] Katsumi Tadamura and Eihachiro Nakamae. Synchronizing computer graphics animation and audio. IEEE MultiMedia, 5(4):63–73, October/ December 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u4063.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1998/
                                        u4063abs.htm.

Tokmakoff:2005:MER

[TNJ05] Andrew Tokmakoff, François-Xavier Nuttall, and Kyunghee Ji. MPEG-21 event reporting: Enabling multimedia E-commerce. IEEE Multi-Media, 12(4):50–59, October/December 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Tobagi:1995:MWD

[Tob95] Fouad A. Tobagi. Multimedia at work: Distance learning with digital video. IEEE MultiMedia, 2(1): 90–93, Spring 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u10090.
                                        pdf.

Turner:2001:CAC

[TR01] David A. Turner and Keith W. Ross. A comprehensive architecture for continuous media email. IEEE MultiMedia, 8(2):88– ??, April/June 2001. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u2088.
pdf;   http://www.computer.
org/multimedia/mu2001/
u2088abs.htm.

Torres:2008:ODF

[TSDD08] Víctor Torres, Carlos Serrão, Miguel Sales Dias, and Jaime Delgado. Open DRM and the future of media. IEEE MultiMedia, 15(2):28–36, April/June 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Tang:2021:ELL

[TSM21] Junshu Tang, Zhiwen Shao, and Lizhuang Ma. EGGAN: Learning latent space for fine-grained expression manipulation. IEEE MultiMedia, 28(3): 42–51, July/September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Tian:2011:VRO

[TT11] Xinmei Tian and Dacheng Tao. Visual reranking: From objectives to strategies. IEEE MultiMedia, 18 (3):12–21, July/September 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Teng:2022:MMS

[TWY^{+}22] Fei Teng, Dian Wang, Yue Yuan, Haibo Zhang, Amit Kumar Singh, and Zhihan Lv. Multimedia monitoring system of obstructive sleep apnea via a deep active learning model. IEEE Multi-Media, 29(3):48–56, July/ September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Takahashi:2000:HEG

[TY00] Kenji Takahashi and Eiji Yana. A hypermedia environment for global collaboration. IEEE Multi-Media, 7(4):36–47, October/December 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u4036.
pdf;   http://www.computer.
org/multimedia/mu2000/
u4036abs.htm.

Tang:2023:CDM

[TYW23] Rui Tang, Cheng Yang, and Yuxuan Wang. A cross-domain multimodal supervised latent topic model for item tagging and cold-start recommendation. IEEE MultiMedia, 30(3):48–62, 2023. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Uden:2002:DPW

[Ude02]  Lorna Uden.  Design

process for Web applications. IEEE MultiMedia, 9(4):47–55, October/ December 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
04/u4047abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/04/u4047.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/04/
u4047.pdf.

Ubik:2012:RAM

[UTZH12] Sven Ubik, Zdenek Travnicek, Petr Zejdl, and Jiri Halak. Remote access to 3D models for research, engineering, and art. IEEE Multi-Media, 19(4):12–19, October/December 2012. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Veksler:2024:CNN

[VAAI24] Maryna Veksler, Ramazan S.

Aygun, Kemal Akkaya, and Sitharama S. Iyengar. A convolutional neural network ensemble for video source camera forensics. IEEE MultiMedia, 31(2): 26–35, April/June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Vaitzblit:1995:MWH

[Vai95] Lev Vaitzblit. Multimedia at work: a highresolution video server for cinema of the future. IEEE MultiMedia, 2(3): 65–69, Fall 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u30065.
                                        pdf.

Viana:2014:MPR

[VBL^{+}14] Windson Viana, Reinaldo Braga, Fabricio D. A. Lemos, Joao M. O. de Souza, Rafael A. F. Carmo, Rossana M. C. Andrade, and Herve Martin. Mobile photo recommendation and logbook generation using context-tagged images. IEEE MultiMedia, 21 (1):24–34, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Valsesia:2019:TCB

[VCBM19] D. Valsesia, G. Coluccia, T. Bianchi, and E. Magli. ToothPic: Camera-based image retrieval on large scales. IEEE MultiMedia, 26(2):33–43, April/June 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Valente:2000:FTR

[VD00] Stéphane Valente and Jean-Luc Dugelay. Face tracking and realistic animations for telecommunicant clones. IEEE MultiMedia, 7(1):34–43, January/March 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u1034.
pdf;   http://www.computer.
org/multimedia/mu1998/
u1034abs.htm;   http:/
/www.computer.org/multimedia/
mu2000/u1034abs.htm#web_
extra_u1.

Vadakital:2022:MIV

[VDL^{+}22] Vinod Kumar Malamal Vadakital, Adrian Dziembowski, Gauthier Lafruit, Franck Thudor, Gwangsoon Lee, and Patrice Rondao Alface. The MPEG Immersive Video Standard current status and future outlook. IEEE MultiMedia, 29(3):101–111, July/ September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Vermillion:1998:MRL

[Ver98] Colleen Vermillion. Media reviews: Learning Java basics: Java 1.2 and JavaScript for C and C++ Programmers by Michael C. Daconta et al. (John Wiley & Sons,1998, $54.99, ISBN 0-471-18359-8). IEEE MultiMedia, 5(4):83, October/ December 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u4083.
pdf.

Vetro:2004:MDI

[Vet04] Anthony Vetro. MPEG-21 digital item adaptation: Enabling universal multimedia access. IEEE MultiMedia, 11(1):84–87, January/March 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
01/u1084abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2004/01/u1084.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/01/
u1084.pdf.

Vagliano:2018:OIB

[VGH^{+}18] I. Vagliano, F. Günther, M. Heinz, A. Apaolaza, I. Bienia, G. Breitfuss, T. Blume, C. Collyda, A. Fessl, S. Gottfried, P. Hasitschka, J. Kellermann, T. Köhler, A. Maas, V. Mezaris, A. Saleh, A. M. J. Skulimowski, S. Thalmann, M. Vigo, A. Wertner, M. Wiese, and A. Scherp. Open innovation in the big data era with the MOVING platform. IEEE MultiMedia, 25 (3):8–21, July/September 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Vin:1998:PRS

[Vin98] Harrick M. Vin. Project reports: Supporting nextgeneration distributed applications. IEEE Multi-Media, 5(3):78–83, July/ September 1998. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u3078.
pdf.

Viola:2023:VCS

[VJS^{+}23] Irene Viola, Jack Jansen, Shishir Subramanyam, Ignacio Reimat, and Pablo Cesar. VR2Gather: A collaborative, social virtual reality system for adaptive, multiparty real-time communication. IEEE MultiMedia, 30(2):48–59, April/June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Vazirgiannis:1999:SAM

[VKS99] Michalis Vazirgiannis, Ioannis Kostalas, and Timos Sellis. Specifying and authoring multimedia scenarios. IEEE Multi-Media, 6(3):24–37, July/ September 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u3024.
pdf;   http://www.computer.
org/multimedia/mu1999/
u3024abs.htm.

Voulodimos:2012:TDA

[VKV^{+}12] Athanasios Voulodimos, Dimitrios Kosmopoulos, Georgios Vasileiou, Emmanuel Sardis, Vasileios Anagnostopoulos, Constantinos Lalos, Anastasios Doulamis, and Theodora Varvarigou. A threefold dataset for activity and workflow recognition in complex industrial environments. IEEE Multi-Media, 19(3):42–52, July/ September 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Vogel:1995:DMQ

[VKvBG95] Andreas Vogel, Brigitte Kerhervé, Gregor von Bochmann, and Jan Gecsei. Distributed multimedia and QOS: a survey. IEEE MultiMedia, 2(2):10– 19, Summer 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u20010.

pdf;   http://www.computer.
org/multimedia/mu1995/
u2010abs.htm.

vanOssenbruggen:2004:OOD

[vONH04] Jacco van Ossenbruggen, Frank Nack, and Lynda Hardman. That obscure object of desire: Multimedia metadata on the Web, part 1. IEEE Multi-Media, 11(4):38–48, October/December 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/04/u4038.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/04/
u4038.pdf.

Vukoti:2018:CAM

[VRG18] V. Vukoti , C. Raymond, and G. Gravier. A crossmodal approach to multimodal fusion in video hyperlinking. IEEE Multi-Media, 25(2):11–23, April/ June 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Velonaki:2008:ABF

[VRSW08] Mari Velonaki, David Rye, Steve Scheding, and Stefan Williams. Art beat: Fishbird: Cross-disciplinary collaboration. IEEE MultiMedia, 15(1):10–12, January/March 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Valjamae:2008:HEU

[VTJL^{+}08] Aleksander Väljamäe, Ana Tajadura-Jiménez, Pontus Larsson, Daniel Västfjäll, and Mendel Kleiner. Handheld experiences: Using audio to enhance the illusion of self-motion. IEEE Multi-Media, 15(4):68–75, October/December 2008. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Verma:2021:AAM

[VVS21] Monu Verma, Santosh Kumar Vipparthi, and Girdhari Singh. AffectiveNet: Affective-motion feature learning for microexpression recognition. IEEE MultiMedia, 28(1):17–27, January/March 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Vendrig:2003:IAM

[VW03] Jeroen Vendrig and Marcel Worring. Interactive adaptive movie annotation. IEEE Multi-Media, 10(3):30–37, July/ September 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2003/
                                        03/u3030abs.htm;   http:

//csdl.computer.org/dl/
mags/mu/2003/03/u3030.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/
u3030.pdf.

Valente:2007:CAB

[VWA^{+}07] Dan Valente, Haibin Wang, Peter Andrews, Partha P. Mitra, Sigal Saar, Ofer Tchernichovski, Ilan Golani, and Yoav Benjamini. Characterizing animal behavior through audio and video signal processing. IEEE MultiMedia, 14(4):32–41, October/December 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Vetter:1995:UCT

[VWS95] Ronald Vetter, Charles Ward, and Susan Shapiro. Using color and text in multimedia projections. IEEE MultiMedia, 2(4):46– 54, Winter 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1995/pdf/u40046.
pdf;   http://www.computer.
org/multimedia/mu1995/
u4046abs.htm.

Woo:2016:USI

[WA16] Daniel N. Woo and Ramazan S. Aygun. Unsupervised speaker identification for TV news. IEEE MultiMedia, 23(4):50–58, October/December 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        04/mmu2016040050.html.

Walker:2003:IIA

[Wal03] Kevin Walker. Interactive and informative art. IEEE MultiMedia, 10(1):4–10, January/March 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2003/01/u1004.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/01/
                                        u1004.pdf.

Wang:2004:MRE

[Wan04]  Xin Wang.  MPEG-

21 rights expression language: Enabling interoperable digital rights management. IEEE Multi-Media, 11(4):84–87, October/December 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/mu/2004/04/u4084.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2004/04/
                                        u4084.pdf.

Wald:2008:USR

[WB08] Mike Wald and Keith Bain. Using speech recognition for real-time captioning of multiple speakers. IEEE MultiMedia, 15(4):56–57, October/December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Webster:2016:SDB

[WB16] Dane Webster and Ivica Ico Bukvic. Small data, big impact. IEEE Multi-Media, 23(1):6–9, January/March 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
01/mmu2016010006-abs.
html.

Wold:1996:CBC

[WBKW96] Erling Wold, Thom Blum, Douglas Keislar, and James Wheaton. Content-based classification, search, and retrieval of audio. IEEE MultiMedia, 3(3):27–36, Fall 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1996/pdf/u3027.
pdf;   http://www.computer.
org/multimedia/mu1996/
u3027abs.htm.

Wen:2022:FSA

[WCQ^{+}22] Yang Wen, Leiting Chen, Lifeng Qiao, Yu Deng, Haisheng Chen, Tian Zhang, and Chuan Zhou. FLeak-Seg: Automated fundus fluorescein leakage segmentation via crossmodal attention learning. IEEE MultiMedia, 29(2): 114–123, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2021:IER

[WCS^{+}21] Xinzhi Wang, Yudong Chang, Vijayan Sugumaran, Xiangfeng Luo, Peng Wang, and Hui Zhang. Implicit emotion relationship mining based on optimal and majority synthesis from multimodal data prediction. IEEE MultiMedia, 28(2):96–105, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Weiss:1995:CSV

[WDG95] Ron Weiss, Andrzej Duda, and David K. Gifford. Composition and search with a video algebra. IEEE MultiMedia, 2(1):12–25, Spring 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u10012.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu1995/
                                        u1012abs.htm.

Walther:2009:TME

[WdJ09] Joerg Walther and Claudia de Jong. Technology for more effective healthcare. IEEE MultiMedia, 16(4):5–7, October/ December 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2017:WMM

[WDZ^{+}17] Yufeng Wang, Wei Dai, Bo Zhang, Jianhua Ma, and Athanasios V. Vasilakos. Word of mouth mobile crowdsourcing: Increasing awareness of physical, cyber, and social interactions. IEEE Multi-Media, 24(4):26–37, October/December 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
04/mmu2017040026-abs.
html.

Weatherford:2000:NGP

[Wea00] Margaret Weatherford. In the news: Games people play online; wireless app roundup; streaming video for the masses; computing peer-to-peer. IEEE MultiMedia, 7(4):18–22, October/December 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u4018.
pdf.

Weight:2008:AML

[Wei08] Jenny Weight. Artful media: Locative systems using mobile phones. IEEE MultiMedia, 15(3):6–10, July/September 2008. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wassermann:2003:LSC

[WEV03] Klaus C. Wassermann, Kynan Eng, and Paul F. M. J. Verschure. Live soundscape composition based on synthetic emotions. IEEE MultiMedia, 10(4):82–90, October/December 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/mu/2003/
                                        04/u4082abs.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mu/2003/04/u4082.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/04/
                                        u4082.pdf.

Wang:2007:MDC

[WFI07] Guijin Wang, Satoshi Futemma, and Eisaburo Itakura. Multiple-description coding for overlay network streaming. IEEE Multi-Media, 14(1):74–82, January/March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2021:MSD

[WFRZ21] Zichi Wang, Guorui Feng, Yanli Ren, and Xinpeng Zhang. Multichannel steganography in digital images for multiple receivers. IEEE MultiMedia, 28(1):65–73, January/ March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2020:BCM

[WFZ^{+}20] Yang Wang, Meng Fang, Joey Tianyi Zhou, Tingting Mu, and Dacheng Tao. Big cross-modal social media data analytics with deep intelligence. IEEE MultiMedia, 27(4):6–8, October/December 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wray:1994:NMM

[WGH94] Stuart Wray, Tim Glauert, and Andy Hopper. Networked multimedia: The Medusa environment. IEEE MultiMedia, 1(4):54–63, Winter 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2021:MEA

[WGY^{+}21] Lijuan Wang, Wenya Guo, Xingxu Yao, Yuxiang Zhang, and Jufeng Yang. Multimodal event-aware network for sentiment analysis in tourism. IEEE MultiMedia, 28(2):49–58, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wong:1997:NMM

[WH97] Stephen T. C. Wong and H. K. Huang. Networked multimedia for medical imaging. IEEE Multi-Media, 4(2):24–35, April/ June 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u2024.
pdf;   http://www.computer.
org/multimedia/mu1997/
u2024abs.htm.

Wu:2011:EBW

[WH11] Lei Wu and Steven C. H. Hoi. Enhancing bagof-words models with semantics-preserving metric learning. IEEE MultiMedia, 18(1):24–37, January/March 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2015:VSR

[WHM15] Xueting Wang, Takatsugu Hirayama, and Kenji Mase. Viewpoint sequence recommendation based on contextual information for multiview video. IEEE MultiMedia, 22(4):40–50, October/ December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/04/mmu2015040040-
abs.html.

Wang:2024:SDN

[WHTC24] Ming-Hung Wang, Ting-Shen Hsieh, Yu-Yao Tseng, and Po-Wen Chi. A software-defined networkingdriven reliable transmission architecture for enhancing real-time video streaming quality. IEEE MultiMedia, 31(1):54–64, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wu:2022:DRC

[WHW^{+}22] Yulin Wu, Ruimin Hu, Xiaochen Wang, Chenhao Hu, and Shanfa Ke. Distortion reduction via CAE and DenseNet mixture network for low bitrate spatial audio object coding. IEEE MultiMedia, 29(1):55–64, January/March 2022. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wilmes:1995:MRR

[Wil95] R. Wilmes. Media reviews: Reviews of an image processing guide and graphing software. IEEE MultiMedia, 2(4):93–96, Winter 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u40093.
                                        pdf.

Wilmes:1996:MRV

[Wil96] Robert Wilmes. Media reviews: a VRML primer with depth: The VRML Sourcebook, by Andrea L. Ames, David R. Nadeau, and John L. Moreland (John Wiley and Sons, 1996, $29.95, ISBN 0-471-14159-3). IEEE MultiMedia, 3(3):90, Fall 1996. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1996/pdf/u3090.
                                        pdf.

Wilson:1998:MWU

[Wil98] Mark Wilson. Multimedia at work: Using MPEG-2 video at the Getty Museum. IEEE MultiMedia, 5(2):74–77, April/ June 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1998/pdf/u2074.
                                        pdf.

Wang:1999:IMW

[WJ99] Fang Ju Wang and Shaidah Jusoh. Integrating multiple Web-based geographic information systems. IEEE MultiMedia, 6(1):49–61, January/March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1999/pdf/u1049.
pdf;   http://www.computer.
org/multimedia/mu1999/
u1049abs.htm.

Westermann:2007:TCE

[WJ07] Utz Westermann and Ramesh Jain. Toward a common event model for multimedia applications. IEEE MultiMedia, 14(1):19–29, January/ March 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wu:2013:SMV

[WKC^{+}13] Guan-Long Wu, Yin-Hsi Kuo, Tzu-Hsuan Chiu, Winston H. Hsu, and Lexing Xie. Scalable mobile video retrieval with sparse projection learning and pseudo label mining. IEEE MultiMedia, 20 (3):47–57, July/September 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Williams:2015:EDT

[WKC^{+}15] Doug Williams, Ian Kegel, Pablo Cesar, Jack Jansen, Marian Ursu, and Erik Geelhoed. Experiments with distributed theatre. IEEE MultiMedia, 22(3): 4–9, July/September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://www.computer.org/
csdl/mags/mu/2015/03/
mmu2015030004-abs.html.

Wallace:2003:ERP

[WKS^{+}03] Manolis Wallace, Kostas Karpouzis, George Stamou, George Moschovitis, Stefanos Kollias, and Christos Schizas. The electronic road: Personalized content browsing. IEEE MultiMedia, 10(4):49–59, October/December 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
04/u4049abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/04/u4049.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/04/
u4049.pdf.

Windridge:2015:NML

[WKY^{+}15] David Windridge, Josef Kittler, Fei Yan, William Christmas, and Aftab Khan. A novel Markov logic rule induction strategy for characterizing sports video footage. IEEE MultiMedia, 22(2):24–35, April/June 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/02/mmu2015020024-
abs.html.

Wang:2021:CBT

[WLL^{+}21] Min Wang, Congyan Lang, Liqian Liang, Gengyu Lyu, Songhe Feng, Wang, and Tao. Class-balanced text to image synthesis with attentive generative adversarial network. IEEE MultiMedia, 28(3):21–31, July/ September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2021:SPP

[WLN^{+}21] Ning Wang, Boqian Liu, Muyao Niu, Kaidi Meng, Haojie Li, Bin Liu, and Zhihui Wang. Semantic place prediction with user attribute in social media. IEEE MultiMedia, 28(4): 29–37, October/December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wu:2011:MES

[WLPN11] Xiao Wu, Yi-Jie Lu, Qiang Peng, and Chong-Wah Ngo. Mining event structures from Web videos. IEEE MultiMedia, 18 (1):38–51, January/March 2011. CODEN IEMUE4.

ISSN 1070-986X (print), 1941-0166 (electronic).

Wu:2021:ISN

[WLWW21] Zongkai Wu, Zihan Liu, Ting Wang, and Donglin Wang. Improved speaker and navigator for visionand-language navigation. IEEE MultiMedia, 28(4): 55–63, October/December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2023:EIE

[WLZ23] Zhi Wang, Jiangchuan Liu, and Wenwu Zhu. Edge intelligence-empowered immersive media. IEEE MultiMedia, 30(2):8–17, April/June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2012:RTC

[WLZL12] Liyun Wang, Hefei Ling, Fuhao Zou, and Zhengding Lu. Real-time compresseddomain video watermarking resistance to geometric distortions. IEEE MultiMedia, 19(1):70–79, January/March 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wu:2020:MMV

[WLZL20] Qing Wu, Miaomiao Li, Xueling Zhu, and Bo Liu. MVIIDroid: a multiple view information integration approach for Android malware detection and family identification. IEEE MultiMedia, 27(4): 48–57, October/December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Williamson:2005:SPF

[WMS05] John Williamson and Roderick Murray-Smith. Sonification of probabilistic feedback through granular synthesis. IEEE MultiMedia, 12(2):45–52, April/June 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wu:1994:IFU

[WN94] Jian-Kang Wu and Arcot Desai Narasimhalu. Identifying faces using multiple retrievals. IEEE MultiMedia, 1(2):27–38, Summer 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2009:RDA

[WNY09] Xiang-Yang Wang, Pan-Pan Niu, and Hong-Ying Yang. A robust, digital-audio watermarking method. IEEE Multi-Media, 16(3):60–69, July/ September 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wu:2010:DDA

[WNZ10] Xiao Wu, Chong-Wah Ngo, and Wan-Lei Zhao. Data-driven approaches to community-contributed video applications. IEEE MultiMedia, 17(4):58–69, October/December 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Worsley:1997:IEA

[WO97] Debra J. Worsley and Tokunbo Ogunfunmi. Isochronous Ethernet — an ATM bridge for multimedia networking. IEEE Multi-Media, 4(1):58–67, January/March 1997. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1997/pdf/u1058.
pdf;   http://www.computer.
org/multimedia/mu1997/
u1058abs.htm.

Wongwirat:2006:HMS

[WO06] Olarn Wongwirat and Shigeyuki Ohara. Haptic media synchronization for remote surgery through simulation. IEEE Multi-Media, 13(3):62–69, July/ September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wojdala:1998:CVS

[Woj98] Andrew Wojdala. Challenges of virtual set technology. IEEE MultiMedia, 5(1):50–57, January/ March 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu1998/pdf/u1050.
pdf;   http://www.computer.
org/multimedia/mu1998/
u1050abs.htm.

Wang:2020:AGF

[WPC^{+}20] Huibing Wang, Jinjia Peng, Dongyan Chen, Guangqi Jiang, Tongtong Zhao, and Xianping Fu. Attribute-guided feature learning network for vehicle reidentification. IEEE MultiMedia, 27(4):112– 121, October/December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2002:SOC

[WS02] Fangju Wang and Yunli Sun. Spatial object clustering and buffering. IEEE MultiMedia, 9(1):26–42, January/March 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2002/pdf/u1026.
pdf;   http://www.computer.
org/multimedia/mu2002/
u1026abs.htm.

Welch:2005:IEB

[WSI^{+}05] Greg Welch, Andrei State, Adrian Ilie, Kok-Lim Low, Anselmo Lastra, Bruce Cairns, Herman Towles, Henry Fuchs, Ruigang Yang, Sascha Becker, Dan Russo, Jesse Funaro, and Andries van Dam. Immersive electronic books for surgical training. IEEE MultiMedia, 12(3):22–35, July/September 2005. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Whitston:1995:MRA

[WSR^{+}95] Howard Whitston, Vaclav Skala, Sorel Reisman, Marguerite C. Murphy, and Karen Whitehouse. Media reviews: Arts & letters takes you to the moon; practical computer graphics; evaluating the market; multimedia for experienced computer scientists; the computer as mystic. IEEE MultiMedia, 2(2):90– 93, Summer 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu1995/pdf/u20090.
                                        pdf.

Wong:2002:MAO

[WSS02] Steven H. Wong, Steven L. Swartz, and Dilip Sarkar. A middleware architecture for open and interoperable GISs. IEEE MultiMedia, 9(2):62–76, April/June 2002. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
02/u2062abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/02/u2062.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/02/
u2062.pdf.

Worring:2012:WUM

[WSS^{+}12] Marcel Worring, Paul Sajda, Simone Santini, David A. Shamma, Alan F. Smeaton, and Qiang Yang. Where is the user in multimedia retrieval? IEEE MultiMedia, 19(4):6–10, October/December 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wus:2014:TEM

[WTL^{+}14] John Wus, Jun Tian, Hongbing Li, Dong-Qing Zhang, Wei Jiang, Zhenyu Wu, and Heather Yu. Toward experiential mobile media processing. IEEE MultiMedia, 21(2):80–89, April/June 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wei:2023:CAL

[WTW23] Xiwen Wei, Yihan Tang, and Si Wu. Content-aware latent semantic direction fusion for multi-attribute editing. IEEE MultiMedia, 30(3):87–99, 2023. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wolfson:2014:NPQ

[WX14] Ouri Wolfson and Bo Xu. A new paradigm for querying blobs in vehicular networks. IEEE MultiMedia, 21(1):48–58, January/ March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wah:2023:OMP

[WX23] Benjamin W. Wah and Jingxi X. Xu. Optimizing multidimensional perceptual quality in online interactive multimedia. IEEE MultiMedia, 30(3): 119–128, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2017:FWA

[WXL^{+}17] Ran Wang, Guangquan Xu, Bin Liu, Yan Cao, and Xiaohong Li. Flow watermarking for antinoise and multistream tracing in anonymous networks. IEEE MultiMedia, 24(4):38–47, October/December 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
04/mmu2017040038-abs.
html.

Wang:2022:CFE

[WXY^{+}22] Yifan Wang, Xing Xu, Wei Yu, Ruicong Xu, Zuo Cao, and Heng Tao Shen. Comprehensive framework of early and late fusion for image sentence retrieval. IEEE MultiMedia, 29(3):38–47, July/ September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wan:2007:MWA

[WY07] Kongwah Wan and Xin Yan. Multimedia at work: Advertising insertion in sports Webcasts. IEEE MultiMedia, 14(2):78–82, April/June 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2016:CSC

[WYZ^{+}16] Wei Wang, Yan Yan, Luming Zhang, Richang Hong, and Nicu Sebe. Collaborative sparse coding for multiview action recognition. IEEE MultiMedia, 23(4):80–87, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        04/mmu2016040080-abs.
                                        html.

Wang:2008:ASM

[WZ08] Ye Wang and Bingjun Zhang. Applicationspecific music transcription for tutoring. IEEE Multi-Media, 15(3):70–74, July/ September 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wiberg:2010:LFI

[WZ10] Mikael Wiberg and Arkady Zaslavsky. Landscaping future interaction: Special issue on mobile and ubiquitous multimedia. IEEE MultiMedia, 17(2):12–15, April/June 2010. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wu:2017:OSC

[WZBKB17] Yongmeng Wu, Leshao Zhang, Nick Bryan-Kinns, and Mathieu Barthet. Open Symphony: Creative participation for audiences of live music performances. IEEE Multi-Media, 24(1):48–62, January/March 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.

org/csdl/mags/mu/2017/
01/mmu2017010048-abs.
html.

Wang:2012:MMA

[WZL12] Qia Wang, Wenjun Zeng, and Aleksandre G. Lobzhanidze. Mobile media in action: Remote target localization and tracking. IEEE Multi-Media, 19(3):74–80, July/ September 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wu:2021:MIB

[WZL^{+}21] Yang Wu, Yanyan Zhao, Xin Lu, Bing Qin, Yin Wu, Jian Sheng, and Jinlong Li. Modeling incongruity between modalities for multimodal sarcasm detection. IEEE MultiMedia, 28(2): 86–95, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Wang:2024:YOL

[WZL^{+}24] Xin Wang, Hongyan Zhang, Qianhe Liu, Wei Gong, Song Bai, and Hezhen You. You-only-look-once multiple-strategy printed circuit board defect detection model. IEEE Multi-Media, 31(1):76–87, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Xu:2019:SMT

[XCZH19] Y. Xu, H. Chen, W. Zhang, and J. Hwang. Smart media transport: a burgeoning intelligent system for next generation multimedia convergence service over heterogeneous networks in China. IEEE MultiMedia, 26(3):79–91, July/ September 2019. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Xu:2014:GBR

[XCZY14] Dan Xu, Peng Cui, Wenwu Zhu, and Shiqiang Yang. Graph-based residence location inference for social media users. IEEE Multi-Media, 21(4):76–83, October/December 2014. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu2014040076-
abs.html.

Xu:2004:CAM

[XDL04] Haifeng Xu, Joe Diamand, and Ajay Luthra. Client architecture for MPEG-4 streaming. IEEE MultiMedia, 11(2):16–23, April/June 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2004/
02/u2016abs.htm;   http:

//csdl.computer.org/dl/
mags/mu/2004/02/u2016.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/02/
u2016.pdf.

Xiao:2006:UTH

[XGN06] Yang Xiao, Max H. Garzon, and Michael Nolen. Using talking heads for real-time virtual videophone in wireless networks. IEEE MultiMedia, 13(3):78–84, July/ September 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Xie:2022:TDG

[XL22] Zhonghua Xie and Lingjun Liu. Transferring deep Gaussian denoiser for compressed sensing MRI reconstruction. IEEE MultiMedia, 29(4):5–13, October/ December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Xu:2022:ESI

[XLL^{+}22] Luoyuan Xu, Yawei Luo, Keyang Luo, Yuesong Wang, Tao Guan, Zhuo Chen, and Wenkai Liu. Exploiting the structure information of suppositional mesh for unsupervised multiview stereo. IEEE Multi-Media, 29(1):94–103, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Xia:2021:DRL

[XLT^{+}21] Haiying Xia, Changyuan Li, Yumei Tan, Lingyun Li, and Shuxiang Song. Destruction and reconstruction learning for facial expression recognition. IEEE MultiMedia, 28(2):20–28, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Xin:2013:LVR

[XNS^{+}13] Xin Xin, Abhishek Nagar, Gaurav Srivastava, Zhu Li, Felix Fernandes, and Aggelos K. Katsaggelos. Large visual repository search with hash collision design optimization. IEEE MultiMedia, 20(2): 62–71, April/June 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Xu:2013:CBP

[XPZ^{+}13] Mingliang Xu, Zhigeng Pan, Mingmin Zhang, Pei Lv, Pengyu Zhu, Yangdong Ye, and Wei Song. Character behavior planning and visual simulation in virtual 3D space. IEEE Multi-Media, 20(1):49–59, January/March 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Xie:2014:CDL

[XSS14] Lexing Xie, David A.

Shamma, and Cees Snoek. Content is dead … long live content: The new age of multimedia-hard problems. IEEE MultiMedia, 21(1):4–8, January/March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Xu:2015:OPQ

[XW15] Jingxi Xu and Benjamin W.

Wah. Optimizing the perceptual quality of realtime multimedia applications. IEEE MultiMedia, 22(4):14–28, October/ December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/04/mmu2015040014-
abs.html.

Xu:2016:FSU

[XWJ16] Baohan Xu, Xi Wang, and Yu-Gang Jiang. Fast summarization of usergenerated videos: Exploiting semantic, emotional, and quality clues. IEEE MultiMedia, 23(3):23–33, July/September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
03/mmu2016030023-abs.
html.

Xiao:2023:BEP

[XXZ^{+}23] Yao Xiao, Lei Xu, Can Zhang, Liehuang Zhu, and Yan Zhang. Blockchainempowered privacy-preserving digital object trading in the metaverse. IEEE MultiMedia, 30(2):81–90, April/June 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Xie:2016:GIW

[XZL^{+}16] Haoran Xie, Di Zou, Raymond Y. K. Lau, Fu Lee Wang, and Tak-Lam Wong. Generating incidental word-learning tasks via topic-based and load-based profiles. IEEE MultiMedia, 23(1):60–70, January/March 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2016/
                                        01/mmu2016010060-abs.
                                        html.

Xie:2021:FGS

[XZL21] Yun Xie, Shaowu Zheng, and Weihua Li. Featureguided spatial attention upsampling for real-time stereo matching network. IEEE MultiMedia, 28 (1):38–47, January/March 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Xu:2013:NSE

[XZYY13] Pengfei Xu, Lei Zhang, Kuiyuan Yang, and Hongxun Yao. Nested-SIFT for efficient image matching and retrieval. IEEE Multi-Media, 20(3):34–46, July/ September 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ye:2014:SEH

[YA14] Yan Ye and Pierre Andrivon. The scalable extensions of HEVC for ultrahigh-definition video delivery. IEEE MultiMedia, 21 (3):58–64, July/September 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Yan:2015:RMO

[Yan15]  Rong Yan.  The rise

of multimedia for online communication startups. IEEE MultiMedia, 22(4):100–104, October/ December 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/04/mmu2015040100-
abs.html.

Yang:2012:POA

[YCZ^{+}12] Lin Yang, Wei-Bang Chen, Chengcui Zhang, John K. Johnstone, Song Gao, and Gary Warner. Profiling online auction sellers using image-editing styles. IEEE MultiMedia, 19(1):29, 39, January/March 2012. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Yeh:2016:SII

[Yeh16] Mei-Chen Yeh. Selecting interesting image regions to automatically create cinemagraphs. IEEE Multi-Media, 23(1):72–81, January/March 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
01/mmu2016010072-abs.
html.

Yamada:2016:CSC

[YGM^{+}16] Masanori Yamada, Yoshiko Goda, Hideya Matsukawa, Kojiro Hata, and Seisuke Yasunami. A computersupported collaborative learning design for quality interaction. IEEE Multi-Media, 23(1):48–59, January/March 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
01/mmu2016010048-abs.
html.

Yuan:2018:DLA

[YGN18] B. Yuan, X. Gao, and Z. Niu. Discovering latent aspects for diversityinduced image retrieval. IEEE MultiMedia, 25(4): 19–33, October/December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Yang:2013:LRW

[YH13] Linjun Yang and Alan Hanjalic. Learning to rerank Web images. IEEE MultiMedia, 20(2):13–21, April/June 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ye:2016:IEA

[YH16] Guodong Ye and Xiaoling Huang. An image encryption algorithm based on autoblocking and electrocardiography. IEEE MultiMedia, 23(2):64–71, April/June 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
02/mmu2016020064-abs.
html.

Yang:2014:MSR

[YHFC14] Yimin Yang, Hsin-Yu Ha, Fausto C. Fleites, and Shu-Ching Chen. A multimedia semantic retrieval mobile system based on HCFGs. IEEE MultiMedia, 21(1):36–46, January/ March 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Yan:2011:LSM

[YHS11] Rong Yan, Benoit Huet, and Rahul Sukthankar. Large-scale multimedia retrieval and mining. IEEE MultiMedia, 18(1):11–13, January/March 2011. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ye:2015:MUH

[YHX15] Yan Ye, Yuwen He, and Xiaoyu Xiu. Manipulating ultra-high definition video traffic. IEEE Multi-Media, 22(3):73–81, July/ September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/mu/2015/03/mmu2015030073-
                                        abs.html.

Yiu:2007:CAL

[YJC07] W.-P. Ken Yiu, Xing Jin, and S.-H. Gary Chan. Challenges and approaches in large-scale P2P media streaming. IEEE MultiMedia, 14(2):50–59, April/June 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Yoshitaka:1994:KAC

[YKHI94] Atsuo Yoshitaka, Setsuko Kishida, Masahito Hirakawa, and Tadao Ichikawa. Knowledgeassisted content-based retrieval for multimedia databases. IEEE Multi-Media, 1(4):12–21, Winter 1994. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Yu:2001:TCB

[YKW01] Heather Yu, Xiangyang Kong, and Wayne Wolf. Techniques for contentbased graph authentication. IEEE MultiMedia, 8(4):38–45, October/ December 2001. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2001/pdf/u4038.
pdf;   http://www.computer.
org/multimedia/mu2001/
u4038abs.htm.

Yamamoto:2008:VSA

[YMON08] Daisuke Yamamoto, Tomoki Masuda, Shigeki Ohira, and Katashi Nagao. Video scene annotation based on Web social activities. IEEE MultiMedia, 15(3):22–32, July/September 2008. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Yan:2009:HTB

[YNC09] Rong Yan, Apostol Natsev, and Murray Campbell. Hybrid tagging and browsing approaches for efficient manual image annotation. IEEE MultiMedia, 16(2): 26–41, April/June 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Yetgin:2009:PDM

[YS09] Zeki Yetgin and Gamze Seckin. Progressive download for multimedia broadcast multicast service. IEEE MultiMedia, 16(2): 76–85, April/June 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Yu:2024:PHD

[YTL^{+}24] Mengzhu Yu, Zhenjun Tang, Xiaoping Liang, Xianquan Zhang, and Xinpeng Zhang. Perceptual hashing with deep and texture features. IEEE MultiMedia, 31(1):65–75, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Yu:2017:LFJ

[Yu17] Jingyi Yu. A light-field journey to virtual reality. IEEE MultiMedia, 24(2): 104–112, April/June 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/mu/2017/
                                        02/mmu2017020104-abs.
                                        html.

Yeh:2014:CFM

[YW14] Mei-Chen Yeh and Wen-Po Wu. Clustering faces in movies using an automatically constructed social network. IEEE Multi-Media, 21(2):22–31, April/ June 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Yang:2007:PCI

[YWT^{+}07] Qiong Yang, Chao Wang, Xiaoou Tang, Mo Chen, and Zhongfu Ye. Progressive cut: An image cutout algorithm that models user intentions. IEEE Multi-Media, 14(3):56–66, July/ September 2007. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Yang:2024:AAV

[YWT^{+}24] Jian Yang, Haixin Wang, Ittetsu Taniguchi, Yibo Fan, and Jinjia Zhou. aVCSR: Adaptive video compressive sensing using region-of-interest detection in the compressed domain. IEEE MultiMedia, 31(1): 19–32, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Yan:2019:MCT

[YWWW19] J. Yan, D. Wu, H. Wang, and R. Wang. Multipoint cooperative transmission for virtual reality in 5G new radio. IEEE Multi-Media, 26(1):51–58, January/March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Yang:2023:DDS

[YXH^{+}23] Yong Yang, Yukun Xia, Shuying Huang, Weiguo Wan, and Xuemei Sun. DSMGN: Dual-supervised mask generation network for infrared and visible image fusion. IEEE Multi-Media, 30(4):82–90, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Yang:2021:EQV

[YZSY21] Mingyi Yang, Wenjie Zou, Jiarun Song, and Fuzheng Yang. Enhancing QoE for viewport-adaptive 360degree video streaming: Perception analysis and implementation. IEEE MultiMedia, 28(4):64–73, October/December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zanbouri:2023:NFB

[ZAKNY23] Kouros Zanbouri, Hamza Mohammed Ridha Al-Khafaji, Nima Jafari Navimipour, and Şenay Yalçın. A new fog-based transmission scheduler on the Internet of Multimedia Things using a fuzzy-based quantum genetic algorithm. IEEE MultiMedia, 30(3): 74–86, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhang:2013:VDI

[ZCC^{+}13] Cha Zhang, Qin Cai, Philip A. Chou, Zhengyou Zhang, and Ricardo Martin-Brualla. Viewport: A distributed, immersive teleconferencing system with infrared dot pattern. IEEE MultiMedia, 20(1):17–27, January/March 2013. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhu:2019:MBV

[ZCH^{+}19] H. Zhu, Y. Cao, Q. Hu, W. Wang, T. Jiang, and Q. Zhang. Multi-bitrate video caching for D2Denabled cellular networks. IEEE MultiMedia, 26 (1):10–20, January/March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Zhou:2004:MME

[ZCL^{+}04] Zhiying Zhou, Adrian David Cheok, Wei Liu, Xiangdong Chen, Farzam Farbiz, Xubo Yang, and Michael Haller. Multisensory musical entertainment systems. IEEE MultiMedia, 11(3):88–101, July/ September 2004. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2004/03/u3088.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/03/
u3088.pdf.

Zhang:1998:NIL

[ZCSSM98] Aidong Zhang, Wendy Chang, Gholamhosein Sheikholeslami, and Tanveer F. Syeda-Mahmood. NetView: Integrating large-scale distributed visual databases. IEEE MultiMedia, 5(3):47– 59, July/September 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.
org/mu/books/mu1998/pdf/
u3047.pdf;   http://www.
computer.org/multimedia/
mu1998/u3047abs.htm.

Zhu:2015:MBD

[ZCWH15] Wenwu Zhu, Peng Cui, Zhi Wang, and Gang Hua. Multimedia big data computing. IEEE MultiMedia, 22(3):96–c3, July/ September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/03/mmu2015030096-
abs.html.

Zhang:2020:ICH

[ZCZ20] Jing Zhang, Tianzhen Chen, and Yong Zhang. Incremental clustering with hard centers. IEEE Multi-Media, 27(4):102–111, October/December 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhang:2019:TQM

[ZDS19] L. Zhang, H. Dong, and A. E. Saddik. Towards a QoE model to evaluate holographic augmented reality devices. IEEE MultiMedia, 26(2):21–32, April/June 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Zaharieva:2015:CPS

[ZDZ15] Maia Zaharieva, Manfred Del Fabro, and Matthias Zeppelzauer. Crossplatform social event detection. IEEE MultiMedia, 22(3):14–25, July/ September 2015. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2015/03/mmu2015030014-
abs.html.

Zeng:2016:UHM

[Zen16] Wenjun Kevin Zeng. Understanding humans in multimedia. IEEE MultiMedia, 23(4):4–5, October/December 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
04/mmu2016040004-abs.
html.

Zhou:2024:ADP

[ZF24] Jinjia Zhou and Chen Fu. Adaptive detachable partition-based reference frame recompression for video coding. IEEE MultiMedia, 31(2):17–25, April/June 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhao:2009:DIR

[ZG09] Zhong-Qiu Zhao and Herve Glotin. Diversifying image retrieval with affinitypropagation clustering on visual manifolds. IEEE MultiMedia, 16(4):34–43, October/December 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhang:2014:TQA

[ZGL^{+}14] Lin Zhang, Zhongyi Gu, Xiaoxu Liu, Hongyu Li, and Jianwei Lu. Training quality-aware filters for no-reference image quality assessment. IEEE Multi-Media, 21(4):67–75, October/December 2014. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu2014040067-
abs.html.

Zhou:2002:UKV

[ZH02] Xiang Sean Zhou and Thomas S. Huang. Unifying keywords and visual contents in image retrieval. IEEE MultiMedia, 9(2): 23–33, April/June 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2002/
02/u2023abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2002/02/u2023.
htm;   http://csdl.computer.
org/dl/mags/mu/2002/02/
u2023.pdf.

Zhang:2012:MKS

[Zha12]  Zhengyou Zhang.  Mi-

crosoft Kinect sensor and its effect. IEEE MultiMedia, 19(2):4–10, May/June 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhai:2020:LQL

[ZHC^{+}20] Qiang Zhai, Rui Huang, Hong Cheng, Huiqin Zhan, Jun Li, and Zicheng Liu. Learning quintuplet loss for large-scale visual geolocalization. IEEE Multi-Media, 27(3):34–43, July/ September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zheng:2000:VRE

[Zhe00] Jiang Yu Zheng. Virtual recovery and exhibition of heritage. IEEE MultiMedia, 7(2):31–34, April/June 2000. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/dlib.computer.org/mu/
books/mu2000/pdf/u2031.
pdf;   http://www.computer.
org/multimedia/mu2000/
u2031abs.htm.

Zheng:2003:DRP

[Zhe03] Jiang Yu Zheng. Digital route panoramas. IEEE MultiMedia, 10(3):57–67, July/September 2003. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.
org/comp/mags/mu/2003/
03/u3057abs.htm;   http:
//csdl.computer.org/dl/
mags/mu/2003/03/u3057.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/
u3057.pdf.

Zimmerman:2003:ECP

[Zim03] John Zimmerman. Exploiting celebrity to produce delight in multimedia tools. IEEE Multi-Media, 10(3):10–14, July/ September 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/dl/
mags/mu/2003/03/u3010.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/
u3010.pdf.

Zhang:2020:PPB

[ZJC^{+}20] Yue Zhang, Yi Jin, Jianqiang Chen, Shichao Kan, Yigang Cen, and Qi Cao. PGAN: Part-based nondirect coupling embedded GAN for person reidentification. IEEE Multi-Media, 27(3):23–33, July/ September 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhang:2018:BRF

[ZK18] X. P. S. Zhang and D. Kedmey. A budding romance: Finance and AI. IEEE MultiMedia, 25(4):79–83, October/December 2018. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Zhu:2017:FMD

[ZKWH17a] Xiaoqing Zhu, Harilaos Koumaras, Mea Wang, and David Hausheer. The future of multimedia distribution: An interview with Baochun Li, Diego R. Lopez, and Christian Timmerer. IEEE Multi-Media, 24(3):48–53, July/ September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
03/mmu2017030048-abs.
html.

Zhu:2017:NFV

[ZKWH17b] Xiaoqing Zhu, Harilaos Koumaras, Mea Wang, and David Hausheer. Network function virtualization and software-defined networking: Advancing multimedia distribution. IEEE MultiMedia, 24(3):16–18, July/September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2017/
03/mmu2017030016.html.

Zhao:2020:EEF

[ZLL^{+}20] Mingbo Zhao, Yu Liu, Xianrui Li, Zhao Zhang, and Yue Zhang. An end-to-end framework for clothing collocation based on semantic feature fusion. IEEE Multi-Media, 27(4):122–132, October/December 2020. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zha:2023:RMC

[ZLL^{+}23] Yufei Zha, Fan Li, Huanyu Li, Peng Zhang, and Wei Huang. Reversible modal conversion model for thermal infrared tracking. IEEE MultiMedia, 30 (3):8–24, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zeng:2014:FCR

[ZLS14] Wenjun Zeng, Zicheng Liu, and Eckehard Steinbach. Forging a close relationship with multimedia communities. IEEE MultiMedia, 21(4):14–15, October/ December 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

www.computer.org/csdl/
mags/mu/2014/04/mmu2014040014.
html.

Zhang:2012:BSC

[ZLT^{+}12] Chunjie Zhang, Jing Liu, Qi Tian, Yanjun Han, Hanqing Lu, and Songde Ma. A boosting, sparsity- constrained bilinear model for object recognition. IEEE MultiMedia, 19(2):58–68, May/June 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zeng:2006:SMA

[ZLZ06] Wenjun Zeng, Junqiang Lan, and Xinhua Zhuang. Security for multimedia adaptation: Architectures and solutions. IEEE MultiMedia, 13(2):68–76, April/June 2006. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhang:2021:VCC

[ZMFB21] Fan Zhang, Di Ma, Chen Feng, and David R. Bull. Video compression with CNN-based postprocessing. IEEE MultiMedia, 28(4):74–83, October/ December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zeleznik:2008:MIA

[ZMvD^{+}08] Robert Zeleznik, Timothy Miller, Andries van Dam, Chuanjun Li, Dana Tenneson, Christopher Maloney, and Joseph J. LaViola. Media impact: Applications and issues in pencentric computing. IEEE MultiMedia, 15(4):14–21, October/December 2008. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zaharieva:2011:FAA

[ZMZB11] Maia Zaharieva, Dalibor Mitrovic, Matthias Zeppelzauer, and Christian Breiteneder. Film analysis of archived documentaries. IEEE MultiMedia, 18(2): 38–47, April/June 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhou:2009:DTM

[ZOR^{+}09] Jiehan Zhou, Zhonghong Ou, Mika Rautiainen, Timo Koskela, and Mika Ylianttila. Digital television for mobile devices. IEEE MultiMedia, 16(1):60–71, January/ March 2009. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zichichi:2023:EMV

[ZRD23] Mirko Zichichi and Víctor Rodríguez-Doncel. Encoding of media value chain processes through blockchains and MPEG-21 smart contracts for media. IEEE MultiMedia, 30 (4):26–35, 2023. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhou:2019:MCT

[ZRW^{+}19] L. Zhou, J. J. P. C. Rodrigues, H. Wang, M. Martini, and V. C. M. Leung. 5G multimedia communications: Theory, technology, and application. IEEE MultiMedia, 26(1): 8–9, January/March 2019. CODEN IEMUE4. ISSN 1070-986x (print), 1941-0166 (electronic).

Zhang:2002:NSM

[ZSM02] Aidong Zhang, Yuqing Song, and Markus Mielke. NetMedia: Streaming multimedia presentations in distributed environments. IEEE MultiMedia, 9(1):56–73, January/ March 2002. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /dlib.computer.org/mu/
                                        books/mu2002/pdf/u1056.
                                        pdf;   http://www.computer.
                                        org/multimedia/mu2002/
                                        u1056abs.htm.

Zhao:2005:ISC

[ZSN^{+}05] Haixia Zhao, Benjamin K. Smith, Kent Norman, Catherine Plaisant, and Ben Shneiderman. Interactive sonification of choropleth maps. IEEE MultiMedia, 12(2):26–35, April/June 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhang:2022:VML

[ZSYW22] Yi Zhang, Jundong Shen, Cheng Yu, and Chongjun Wang. Views meet labels: Personalized relation refinement network for multiview multilabel learning. IEEE MultiMedia, 29(2): 104–113, April/June 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhu:2020:ALB

[ZSZ^{+}20] Lei Zhu, Jiayu Song, Xiaofeng Zhu, Chengyuan Zhang, Shichao Zhang, and Xinpan Yuan. Adversarial learning-based semantic correlation representation for cross-modal retrieval. IEEE MultiMedia, 27(4):79–90, October/ December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zha:2022:SSC

[ZSZ^{+}22] Yufei Zha, Jingxian Sun, Peng Zhang, Lichao Zhang, Abel Gonzalez-Garcia, and Wei Huang. Self-supervised cross-modal distillation for thermal infrared tracking. IEEE MultiMedia, 29(4): 80–96, October/December 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). See [ZZS^{+}24].

Zhu:2021:SSA

[ZWH^{+}21] Wenqian Zhu, Zhongyuan Wang, Hu, Ruimin, and Dengshi Li. From semantic to spatial awareness: Vehicle reidentification with multiple attention mechanisms. IEEE MultiMedia, 28(3):32–41, July/September 2021. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhang:2020:WSF

[ZWL^{+}20] Haijun Zhang, Xinghao Wang, Linlin Liu, Dongliang Zhou, and Zhao Zhang. WarpClothingOut: a stepwise framework for clothes translation from the human body to tiled images. IEEE MultiMedia, 27(4): 58–68, October/December 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhao:2021:ISI

[ZXHS21] Sicheng Zhao, Min Xu, Qingming Huang, and Björn W. Schuller. Introduction to the special issue on MMAC: Multimodal affective computing of large-scale multimedia data. IEEE MultiMedia, 28 (2):8–10, April/June 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhang:2013:NWE

[ZYJ^{+}13] Xin Zhang, Zhichao Ye, Lianwen Jin, Ziyong Feng, and Shaojie Xu. A new writing experience: Finger writing in the air using a Kinect sensor. IEEE Multi-Media, 20(4):85–93, October/December 2013. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhang:2016:GEI

[ZYNS16] Luming Zhang, Yi Yang, Feiping Nie, and Ling Shao. Guest Editors’ introduction: Perception, aesthetics, and emotion in multimedia quality modeling. IEEE MultiMedia, 23(3):20–22, July/ September 2016. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

https://www.computer.
org/csdl/mags/mu/2016/
03/mmu2016030020-abs.
html.

Zhang:2010:CMA

[ZYS^{+}10] Yuejie Zhang, Chunlei Yang, Yi Shen, Ning Zhou, Jianping Fan, and Hangzai Luo. A cross-modal approach to cleansing weakly tagged images. IEEE MultiMedia, 17(4):18–25, October/December 2010. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhao:2011:DTO

[ZYXW11] Gangqiang Zhao, Junsong Yuan, Jiang Xu, and Ying Wu. Discovering the thematic object in commercial videos. IEEE Multi-Media, 18(3):56–65, July/ September 2011. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhang:2021:CAL

[ZZC^{+}21] Fangfang Zhang, Xue Zhang, Maoyong Cao, Fengying Ma, and Zhengfeng Li. Characteristic analysis of 2D lag-complex logistic map and its application in image encryption. IEEE MultiMedia, 28(4):96–106, October/December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhang:2022:VSH

[ZZL^{+}22] Yinlong Zhang, Xiaoyan Zheng, Wei Liang, Sichao Zhang, and Xudong Yuan. Visual surveillance for human fall detection in healthcare IoT. IEEE MultiMedia, 29(1):36–46, January/March 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhao:2021:SFE

[ZZLG21] Xibin Zhao, Junjie Zhu, Bingjun Luo, and Yue Gao. Survey on facial expression recognition: History, applications, and challenges. IEEE MultiMedia, 28(4): 38–44, October/December 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zhu:2022:DMH

[ZZS^{+}22] Lei Zhu, Chengyuan Zhang, Jiayu Song, Shichao Zhang, Chunwei Tian, and Xinghui Zhu. Deep multigraph hierarchical enhanced semantic representation for crossmodal retrieval. IEEE MultiMedia, 29(3):17–26, July/ September 2022. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Zha:2024:ESS

[ZZS^{+}24] Yufei Zha, Lichao Zhang, Jingxian Sun, Abel Gonzalez-Garcia, Peng Zhang, and Wei Huang. Erratum to “Self-Supervised Cross-Modal Distillation for Thermal Infrared Tracking”. IEEE MultiMedia, 31(1):110, 2024. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). See [ZSZ^{+}22].

Zhao:2021:SRL

[ZZZ^{+}21] Tianyu Zhao, Jian Zhao, Wengang Zhou, Yun Zhou, and Houqiang Li. State representation learning with adjacent state consistency loss for deep reinforcement learning. IEEE MultiMedia, 28(3):117– 127, July/September 2021. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).